A Complete Bibliography of Scientific American: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

25 April 2024 Version 4.06

Title word cross-reference

1/f [1661].

10th [783]. 16th [1654]. 16th-Century [1654].

9 [609].

Abortion [1411]. Abrasive [1785]. Absolute [1735]. Absorption [190, 997, 1287]. Accelerator [835]. Accelerators [456, 121, 455]. Accuracy [1111]. Acetylcholine [1435]. Acid [1963]. Acoustic [560, 1965]. Acoustics [1448]. Acquires [2013]. Acquisition [1783]. Act [689]. Actinomycin [933]. Action [642, 834]. Activated [1274, 1273]. Active [1172]. Activity [173]. Adaptation [1755]. Additives [438]. Adjacency [1685]. Advanced [708]. Advances [260]. Adversary [893]. Advice [698, 697]. Advice-Taking [698, 697]. Aerodynamic [6]. aether [1681].

Afar [23]. Africa [397]. African [1986]. Aftereffects [1396]. Against [722, 1236]. Age [1570, 817, 1330, 1988, 900]. Ago [2, 18, 35, 51, 68, 85, 101, 118, 134, 153, 170, 187, 204, 220, 237, 253, 271, 287, 303, 319, 335, 355, 371, 387, 404, 420, 436, 452, 469, 485, 501, 519, 535, 554, 570, 586, 603, 620, 638, 654, 671, 687, 704, 720, 736, 757, 774, 792, 810, 826, 843, 859, 875, 891, 907, 923, 940, 959, 975, 991, 1008, 1024, 1042, 1059, 1076, 1092, 1109, 1125, 1142, 1163, 1180, 1198, 1217, 1246, 1263, 1279, 1294, 1309, 1218, 1326, 1341, 1360, 1377, 1392, 1409, 1425, 1440]. Ago [1456, 1471, 1488, 1503, 1518, 1536, 1557, 1575, 1591, 1608, 1624, 1642, 1658, 1675, 1691, 1709, 1726, 1742, 1759, 1775, 1791, 1809, 1826, 1842, 1858, 1875, 1892, 1908, 1924, 1940, 1959, 1975, 1992]. Agreement [1830]. Agricultural [124, 1354, 1349, 343, 457]. Agriculture [1350, 1355, 1352, 1353, 1412, 1351, 1094, 1343]. Air [1017, 127, 689]. Airborne [1567, 1566]. Aircraft [851, 42]. Airflow [1528]. Airglow [412]. Airplane [1916]. Akhenaten [1802]. Albatrosses [178]. Albinos [879]. Alcohol [1265]. Alfred [1036]. Algae [1527]. Algorithms [1462, 1618]. Alienation [929]. Alkali [1511]. Alloying [1298]. Alloys [1982, 1981]. Alternative [1713]. Altitude [25]. Amateur [14, 31, 47, 64, 81, 97, 114, 130, 149, 166, 183, 199, 216, 233, 249, 267, 283, 299, 315, 331, 351, 367, 383, 399, 416, 432, 448, 465, 481, 497, 515, 531, 550, 566, 582, 598, 616, 634, 650, 667, 683, 700, 716, 732, 753, 770, 788, 805, 822, 839, 855, 871, 887, 903, 919, 936, 955, 971, 987, 1003, 1020, 1038, 1055, 1072, 1088, 1105, 1121, 1138, 1159, 1176, 1194, 1211, 1239, 1259, 1515, 1532, 1554, 1572, 1588, 1604, 1621, 1639, 1655, 1672, 1688, 1706, 1723, 1739, 1756, 1772, 1788, 1805, 1823, 1839, 1855, 1872, 1889, 1905, 1921, 1937, 1956, 1972, 1989, 2006]. Ambiguity [916]. America [1237]. American [2009]. Amorphous [1475, 1476]. Amorphous-Semiconductor [1476, 1475]. AMP [529]. Amplification [1354]. Analysis [493, 111, 177, 1010, 1999, 1287, 579]. Anatomy [373]. Anchovy [690]. Ancient [460, 1854]. Andes [727, 258]. Andreas [377]. Andromeda [691]. Animal [1917, 1347, 1801, 696, 539]. Animals [1348, 1736, 1172, 1919, 1751, 1188]. Anions [1511]. Annihilation [1099, 1098]. Annual [201, 401, 600, 807, 1005, 1213, 1406, 1605, 1806, 2007, 261]. Antibodies [122]. Antibody [1415]. Antigens [1568]. Antisubmarine [503]. Ants [246, 1602, 577, 1096]. Ape [563]. Apollo [1847]. Appendicularians [1233]. Approach [509, 510]. Archaeology [878, 1832]. Archaic [94, 1387]. Architecture [1638, 884]. Arithmetic [243]. Arms [622, 4]. Army [577]. Aromatic [1266, 523]. Arrow [1205]. Artificial [240]. Assembly [1780, 1814, 1632, 1631]. Assessment [422, 20]. Assistance [454]. Asteroids [1011]. Astounding [1693]. Astrolabe [820]. Astronomy [428]. Astrophysics [632, 631]. Asymmetry [645]. Atherosclerosis [1432]. Atlantic [1133, 1134, 1903]. Atmosphere [1506, 898]. Atmospheric [1671, 208]. Atomic [1850]. Atoms [79, 580]. ATP [1652]. Attack [722]. Attitudes [389, 1696]. Attractant [910]. Auditory [766, 191]. Australia [1209]. Authors [3, 19, 36, 52, 69, 86, 102, 119, 135, 154, 171, 188, 205, 221, 238, 254, 272, 288, 304, 320, 336, 356, 372, 388, 405, 421, 437, 453, 470, 486, 502, 520, 536, 555, 571, 587, 604, 621, 639, 655, 672, 688, 705, 721, 737, 758, 775, 793, 811, 827, 844, 860, 876, 892, 908, 924, 941, 960, 976, 992, 1043, 1009, 1025, 1060, 1077, 1093, 1110, 1126, 1143, 1164, 1181, 1199, 1219, 1247, 1264, 1280, 1295, 1310, 1220, 1327, 1342, 1361, 1378, 1393, 1410, 1426, 1441, 1457, 1472, 1489, 1504, 1519]. Authors [1537, 1558, 1576, 1592, 1609, 1625, 1643, 1659, 1676, 1692, 1710, 1727, 1743, 1760, 1776, 1792, 1810, 1827, 1843, 1859, 1876, 1893, 1909, 1925, 1941, 1960, 1976, 1993]. Automata [232]. Automatic [177, 1898]. Automobile [479, 1713]. Automobiles [628, 1012]. Autonomic [5]. Available [1353].

Background [926, 1681]. Bacteria [1282, 1129, 1617, 10, 1314]. Bacterial [1928]. Bag [1918]. Ballistic [1979]. Ban [406, 572]. Bantu [1465]. Barr [1913]. Based [1979]. Basis [847, 1883]. Batavia [835]. Beam [574, 1862]. Beams [724, 1779]. Bear [443]. Bearing [1115]. Bearings [230]. Beats [766]. Beehive [463]. Bees [1719]. Beetle [1986]. Beetles [1335]. before [1236]. Began [1570]. Behavior [795, 1936, 847, 1667, 1705, 108, 207, 1335, 1753, 577]. between [796, 246, 79, 1225, 1316, 1459, 674, 1686, 797]. Bibliography [16, 33, 49, 66, 83, 99, 116, 132, 151, 168, 185, 202, 218, 235, 251, 269, 285, 301, 317, 333, 353, 369, 385, 402, 418, 434, 450, 467, 483, 499, 517, 533, 552, 568, 584, 601, 618, 636, 652, 669, 685, 702, 718, 734, 755, 772, 790, 808, 824, 841, 857, 873, 889, 905, 921, 938, 957, 973, 989, 1006, 1022, 1040, 1057, 1074, 1090, 1107, 1123, 1140, 1161, 1178, 1196, 1214, 1243, 1261, 1277, 1292, 1307, 1323, 1244, 1339, 1358, 1372, 1390, 1407, 1423, 1438, 1454, 1469, 1486, 1501, 1517, 1534]. Bibliography [1555, 1573, 1589, 1606, 1622, 1640, 1656, 1673, 1689, 1707, 1724, 1740, 1757, 1773, 1789, 1807, 1824, 1840, 1856, 1873, 1890, 1906, 1922, 1938, 1957, 1973, 1990, 2008]. Bicycle [646]. Binary [527]. binding [931]. Binds [933]. Binocular [528, 1270]. Biocrystals [1463]. Biological [1447, 1509, 1131, 995, 763, 70, 1034, 261, 125, 488, 1257, 869, 1770]. Biology [1079]. Biosphere [145, 147, 341, 136, 146, 138]. Bird [1136, 1831]. Birds [967, 58, 395, 728, 1771]. Birth [525, 1668]. BL [1521]. Black [1413, 473, 994, 1680]. Blacks [1851]. Blindness [1050]. Blood [177, 1764]. Blue [1527]. Blue-Green [1527]. Bodies [624, 1156]. Body [291, 1116]. Bok [1494]. Bomb [1174]. Bomber [722]. Bones [1869]. Book [506]. Books [1177, 635, 65, 1122, 167, 15, 32, 48, 82, 98, 115, 131, 150, 184, 217, 234, 250, 268, 284, 300, 316, 332, 352, 368, 384, 400, 433, 449, 466, 482, 498, 516, 532, 551, 567, 583, 617, 651, 668, 684, 701, 717, 733, 754, 771, 789, 823, 840, 856, 872, 888, 904, 920, 937, 956, 972, 988, 1021, 1039, 1056, 1073, 1089, 1106, 1139, 1160, 1195, 1241, 1260, 1276, 1291, 1306, 1322, 1242, 1338, 1357, 1375, 1389, 1422, 1437, 1453, 1468, 1485, 1500, 1516, 1533, 1539, 1560, 1578, 1611, 1627, 1645, 1662, 1678, 1695, 1712, 1729, 1745]. Books [1762, 1778, 1812, 1829, 1845, 1861, 1878, 1895, 1911, 1927, 1943, 1962, 1978, 200, 599, 806, 1004, 1212, 1405, 1594, 1794, 1995, 417]. Bottom [1832]. Bound [240]. Bowed [819]. Bracing [837]. Brain [1886, 1949, 1955, 714, 1952, 836, 461, 1953, 1384, 594, 309, 1944, 1951, 1950, 1954, 645, 1764, 41, 1530, 1948, 766, 2003, 423, 1786, 227, 1884, 1529]. Brain-Cooling [1886, 1884]. Brains [917]. Brasses [707]. Braun [853]. break [1531]. Breakers [212]. Breakup [156]. Breathe [1831, 395]. Breeder [172, 1443, 1442]. Brideprice [712]. Brightest [658]. Brines [55]. Britain [1429, 1001, 1603]. Brother [1787]. brown [1661]. Bruegel [1654]. Bruno [664]. Bubbles [1231, 294]. Building [440, 239]. Buildings [837]. Bumblebee [665]. Burgess [1919]. Burial [1387]. Bursts [1367, 1366]. Burying [1335]. Business [750]. Byzantine [322].

C [916]. Cable [358]. Calcitonin [157]. Calculator [1271]. Calefaction [72]. Can [606]. Canals [1237]. Cancer [1187, 1636, 1848, 1478, 760]. Candidates [2012]. Capture [1738]. Car [1528]. Carbenes [1256]. Carbon [142, 562, 1880, 1015, 362, 1612]. Carcinogens [1928]. Care [53, 255, 751, 861, 1127]. Carnivorous [1635]. Carroll [506]. Carrying [1916]. Carthage [1619]. Carthaginian [1035]. Case [1479]. Castings [998]. Casts [917]. Catalysis [393]. Catapults [1854]. Cataract [1206]. Catastrophe [1285]. Cathedrals [579]. Cathode [853]. Cats [1584]. Cause [606]. Causes [1131, 1980]. Cave [443]. Cell [29, 1401, 846, 1063, 1798, 424, 1274, 623, 1852, 1883, 1814, 1201, 816, 1297, 1796, 1273]. Cell-Surface [1297]. Cells [1066, 1665, 1384, 1652, 177, 108, 1715, 911, 1750, 229, 1686, 363, 227, 160]. Cellular [232, 538]. Celtic [1603]. Cement [1510]. Cemetery [94]. Census [305]. Center [1135, 477, 866, 1318]. Center-Pivot [1318]. central [1914]. Centrifuge [1730]. Century [1654]. Cerebellum [1014]. Ceremonial [477]. Chain [1964]. Chains [1051]. Change [363]. Changes [423]. Changing [950]. Channels [543, 1852]. Characteristics [1620]. Charge [829, 830]. Charge-Coupled [829, 830]. Charm [1564]. Checks [430]. Chemical [1748, 508, 745, 70, 1715, 281, 899]. Chemicals [1282]. Chemistry [562, 1950, 849, 1885, 57]. Chess [698, 697]. Child [1380]. Children [60]. Chimeras [1868]. Chimpanzee [607]. China [861, 588, 1094]. Chinese [1869, 626]. Chloroplasts [173]. Choice [1311]. Cholera [321]. Christ [611]. Chromosomal [1031]. Chromosomes [265]. Cilia [963]. cipher [1531]. Circuit [1542, 212]. Circuits [1544, 1546, 1543]. Circulation [12, 208]. Citizen [7, 24, 40, 56, 73, 90, 107, 123, 139, 158, 175, 192, 209, 225, 242, 259, 276, 292, 308, 324, 340, 360, 376, 392, 409, 425, 441, 459, 474, 490, 507, 524, 540, 559, 575, 591, 608, 625, 643, 660, 676, 692, 709, 725, 742, 762, 780, 798, 815, 832, 848, 864, 880, 896, 912, 928, 945, 964, 980, 996, 1013, 1030, 1048, 1064, 1081, 1097, 1114, 1130, 1147, 1169, 1186, 1204, 1226, 1252, 1268, 1284, 1299, 1315, 1227, 1331, 1346, 1365, 1382, 1397, 1414, 1430, 1446, 1461]. Citizen [1477, 1493, 1508, 1524, 1545, 1565, 1582, 1598, 1615, 1633, 1649, 1666, 1682, 1717, 1733, 1749, 1766, 1782, 1799, 1817, 1833, 1849, 1865, 1882, 1899, 1915, 1947, 1966, 2000, 1983, 1931, 1699, 659]. City [296]. Clams [1067]. Class [1222, 307, 1171]. Clay [1866]. Clean [1017, 557, 689]. Climate [478, 901]. Clock [95, 1257]. Clocks [1034, 261]. Cluster [104, 105]. clustering [1583]. Clusters [1562, 1784, 1563]. Coal [1200, 1296, 1813, 934, 845]. Coatings [194]. Code [527]. Coil [279]. Colicins [1201]. Collective [455, 456]. Collective-Effect [456, 455]. Colliculus [594]. collision [1459]. Collisions [768, 767]. Color [1569, 1132, 1168, 1600, 1050]. Columbian [1135]. Combinations [165]. Combinatorial [1653]. Combining [1415]. Comet [1888]. Comets [833]. Commerce [1584]. Communication [543, 1047, 777, 548, 547, 546, 1192, 246, 544, 541, 526, 545, 75, 1551, 537, 538, 539]. Communications [1523, 1431, 1522]. Community [241, 546]. Companions [1464]. Comparisons [1127]. Compartments [1917]. Complement [781]. Complexity [1703, 1700]. Composite [708]. Compound [1512]. Compounds [1266]. Computer [1027, 1028, 2013, 673, 1722, 979, 884, 1553, 1631, 1632, 228, 91, 57, 698, 1552, 697]. Computer-controlled [1631, 1632]. Computer-Managed [1027, 1028]. Computers [162]. Computing [1288]. Concepts [648]. Conduction [612]. Conductors [1964]. Confederacy [224]. Confinement [1918, 1381]. Confirmation [679]. Confusions [191]. Consequences [1587]. Constants [159]. Constructive [1971]. Consumer [1595]. Consumption [1012]. Continent [440]. Continent-Building [440]. Continental [576, 1036]. Continents [1818, 1449]. Contour [494]. Contours [1283]. Contraction [1183, 472]. Contrast [984, 494]. Control [1473, 328, 1936, 979, 622, 492, 1787, 706, 1898, 763, 163, 463, 961, 1399, 661, 869, 610, 472, 1550]. controlled [1631, 1632]. Controversies [1281]. Convection [1383]. Conversion [103, 345]. Conway [165]. Cooling [1638, 1886, 278, 1884]. Cooperative [834]. Coordination [968]. Copernican [1498]. Copernicus [801]. Coprolites [1018]. Coral [1932]. Corals [1932]. Corn [323]. Corona [764]. Cortex [1014]. Cosmic [711, 1701, 1716, 1366, 1367, 926, 1714, 1681]. Counterfeiting [1001]. Countries [950, 951, 1355, 953, 949, 948, 1795]. Coupled [829, 1820, 830]. Covenanters [378]. Crashworthiness [628]. Cratering [1419, 471]. Cricket [927]. Crisis [690, 1587]. Criteria [239]. Critters [1693]. Crocodile [1289]. Crocodilians [1970]. Crops [1253]. Cruise [1427]. Crust [898]. Cry [852]. Cryptobiosis [391]. Cryptography [673, 1934]. Crystal [61, 931, 62]. Crystals [613, 614]. Culture [578]. Currents [1000, 1383, 1738]. Curvature [1336]. curves [1661]. Cutting [915]. Cyanate [1063]. Cycle [1473, 142, 141, 143, 816, 137, 661, 140, 138]. Cycles [161, 144, 1347, 895]. Cyclic [529].

Data [1549]. Dating [1403, 1402]. Day [365]. Daylight [1879]. Daylight-Saving [1879]. Death [738]. Debate [1505, 1174, 572]. Decay [2004]. Decision [349]. Decision-making [349]. Deep [1172, 1492, 1818].

Deep-Sea [1172]. Defense [722]. Deformation [1070, 1069]. Deimos [1428]. Deinstitutionalization [1628]. Delayed [1912]. Delivery [1998, 53, 861]. Deposits [1630, 1629]. Deprivation [511]. Depth [1920]. Desert [381]. Deserted [1373]. Design [1118]. Detection [1000, 274]. Deuterium [885]. Developed [950, 949, 948, 1795]. Developing [1355]. Development [1380, 981, 1949, 1917, 995, 1355, 1999, 1770]. Device [1288]. Devices [1476, 829, 61, 1475, 830, 62]. Diabetes [1980]. Diagnosis [1683, 375]. Diamond [1191, 1190]. Did [1768, 446]. Differentiation [1715]. Difficult [1887]. Digest [413]. Dimensional [675, 1614]. Dimensions [2012, 2013, 966, 1344]. Dinosaur [1065]. Diode [307]. Diodes [422]. Dioxide [1612]. Direct [1229]. Directed [407]. Dirty [557]. Disarmament [961]. Disclinations [1601]. Discover [446]. Discovery [680, 1053, 595]. Discrimination [179]. Disease [1063, 375, 606, 1016, 495]. Diseases [730]. Disk [1864]. Disorders [1954]. Disoriented [818]. Display [61, 62]. Displays [91]. Disposal [925, 1490]. Dissection [795]. Distributed [1898]. Diversity [1131]. DNA [54, 1401, 1596, 1505, 10, 1228, 933, 407]. Do [444]. Doctor [526]. Doctor-Patient [526]. does [413]. Dolphins [1853]. dominoes [1483]. Double [1046, 1045]. Drained [1680]. Drawings [1769]. Dried [590]. Drift [576, 1036, 1681]. Drip [1581]. Drowned [969]. Drug [1998, 657]. Drug-Delivery [1998]. Drums [397]. Dual [1032, 1033]. Dual-Resonance [1033, 1032]. Duet [728]. Duet-Singing [728]. Dung [1986, 869]. Dust [1371]. Dwarfism [511]. Dying [741]. Dynamic [846]. Dynamics [982, 691].

Earlier [457]. Earliest [1451, 1697, 1750]. Early [296, 39, 79, 1587, 258]. Earth [641, 1332, 1836, 1718, 458, 339, 1383, 137, 1525, 898, 1151, 396, 1049]. Earth-Moon [458]. Earthquake [1078]. Earthquakes [1597]. East [1732]. Eastern [162]. Ecological [1752]. Ecology [1986]. Economic [37, 346, 1010, 454]. Economy [749, 1394, 605]. Ecosystem [312, 161, 1651]. Eden [232]. Effect [456, 455]. Effects [1912, 1116]. Efficiency [765, 1618]. Eggs [1333, 1831]. Eggshell [43]. Eggshells [126]. Elastic [291]. Elder [1654]. Electric [979]. Electric-Power [979]. Electrical [2003]. Electricity [1363]. Electrochemical [813]. Electromagnetic [759]. Electron [574, 256, 1098, 1099, 410, 767, 768, 1312, 257, 1313]. Electron-Hole [1312, 1313]. Electron-Positron [1099, 768, 1098, 767]. Electronic [1648, 1271, 694]. Electronics [21, 22, 1552]. Electrons [612]. Electrostatics [442]. Elementary [1224, 1033, 913, 1032]. Elementary-Particle [913]. Elements [508, 1542, 817]. Embryo [863]. Embryos [193, 1702]. Emission [962]. Emitting [1046, 1045]. Endemic [295]. Energetics [665, 1201]. Energies [778, 779]. Energy [344, 338, 341, 1651, 339, 1996, 1997, 342, 1015, 1716, 346, 445, 1587, 137, 632, 343, 812, 146, 337, 345, 347, 348, 138, 1714, 631]. Energy-Storage [1997, 1996]. Enforcing [689]. Engine [521, 729]. Engines [1713]. England [1373, 1988]. Enhanced [1679]. Enhanced-Radiation [1679].

Environment [547, 273]. Environmental [463]. Enzyme [1863]. Enzymes [240]. Epidemiology [1599]. Epstein [1913]. Eradication [1362]. Escape [513]. Escher [916]. Ethics [1248, 977]. Euphrates [1930]. Eurasia [1459]. Europe [162, 362]. European [1255]. Evoked [2003]. Evolution [1747, 1970, 1145, 1748, 785, 1232, 1984, 325, 1752, 1746, 678, 491, 277, 1750, 1753, 1864, 1751, 1754, 727]. Excavation [969]. Exchange [1207]. Exoelectrons [1418]. Exotic [1698, 580]. Expand [1269]. Experience [423]. Experiment [1044, 556]. Experimental [696]. Experimentation [1248]. Experiments [724, 1103, 1062, 512, 179]. Exploring [1480]. Expression [548]. Extending [406]. External [1896]. Extraterrestrial [1085]. Eye [1819, 968, 1512, 290]. Eye-Head [968]. Eyes [1801]. Eyewitness [993].

Fabrication [1546]. Faces [782]. Factor [1897, 1329]. Factors [454, 495]. Fall [680]. Families [1224]. Family [949, 915]. Fantastic [165, 786]. Farmer [1930]. Farming [39, 189]. Farmstead [1603]. Fast [1803, 172]. Faster [26]. Fastest [228]. Fat [730]. Fat-Metabolism [730]. Fault [377]. Feedback [163]. Feeding [476, 1705]. Ferdinand [853]. Fermat [1767]. Fertility [706, 1795]. Fertilization [1586, 1896]. Fiber [777]. Fibers [291]. Fibroplasia [1496]. Field [1836]. Fields [1157]. Figures [818]. Films [1231]. Final [1167, 1255]. Fine [641]. Finite [1336]. Fire [909]. Fireflies [1301]. First [1916, 1288, 446]. Fishes [624, 211, 1450, 281]. Fission [1221, 1324, 1402, 1403, 1490]. Fission-Track [1403, 1402]. Fixation [1447, 965]. Flashlight [1450]. Flavor [1168]. Flea [784]. Flew [343]. Flight [327, 1787, 759, 802]. Flight-Control [1787]. Flint [1901, 1585]. Floor [1133, 1172, 1134]. Flow [344, 341, 1651, 342, 1764, 1525]. fluctuations [1661]. Fluorescence [1274, 1273]. Fluorescence-Activated [1274, 1273]. Fly [110]. Flying [492, 784, 1188]. Flywheels [794]. Food [1051, 145, 1667, 438, 952, 1343]. Food-sharing [1667]. Footracing [1320]. Forage [1253]. Force [1846]. Forces [79]. Foreign [1095, 1026]. Forest [1651, 1086, 799]. Forever [1269]. Form [995]. Formation [1509, 1933, 1770, 1768]. Forming [1385]. Fortification [280]. Fortress [1035]. Fossil [576, 917]. Fossils [275]. Four [1569]. Four-Color-Map [1569]. fractal [1661]. Free [680, 1182, 125]. Freedom [548]. Frequency [984]. Frog [1333]. Frontier [1429]. Fruit [110]. Fuel [1798, 1796, 1012]. Fuel-Cell [1798, 1796]. Fuels [1395, 557]. Full [1443, 1442]. Full-Scale [1443, 1442]. Function [1568, 122, 1764]. Functional [41]. Functions [1585]. Fundamental [1564, 159]. Fusion [509, 894, 1929, 226, 289, 1779, 510]. Future [1320].

Galaxies [796, 1784, 1400, 1985, 88, 1128, 1864, 797, 1583]. Galaxy [866, 1914, 1880]. Galilean [1304]. Galileo [680, 1053, 1103, 1288]. Game [2012, 2013, 165, 232]. Games [231, 165, 13, 30, 46, 63, 80, 96, 113, 129, 148, 164, 182, 198, 232, 215, 248, 266, 282, 298, 314, 330, 350, 366, 382, 398, 415, 431, 447, 464, 480, 496, 514, 530, 549, 565, 581, 597, 786, 615, 633, 649, 666, 682, 699, 715, 731, 752, 769, 787, 804, 821, 838, 854, 870, 886, 902, 918, 935, 954, 970, 986, 1002, 1037, 1019, 1054, 1071, 1087, 1104, 1120, 1137, 1158, 1175, 1193, 1210, 1238, 1258, 1275, 1290, 1305, 1321, 1240, 1337, 1356, 1374, 1388, 1404, 1531, 1421, 1436, 1452, 1467, 1484, 1499, 1514, 1538]. Games [1559, 1577, 1593, 1693, 1610, 1626, 1644, 1660, 1677, 1694, 1711, 1728, 1744, 1761, 1777, 1793, 1811, 1828, 1844, 1860, 1877, 1894, 1910, 1926, 1942, 1961, 1977, 1994, 1661]. Gamma [1367, 1366]. Gamma-Ray [1367, 1366]. Garden [232]. Gas [1296, 1458, 1730, 1527]. Gaseous [1804]. Gasification [845]. Gauss [1513]. Gears [230]. Gene [1999, 1031]. Generation [1363, 1188]. Genes [723, 1112, 1837, 642, 911]. Genetic [795, 375, 173, 89, 44, 380]. Genetics [946, 1636, 414]. Geography [346]. Geometry [1231]. Geosynclines [440]. Geothermal [411]. Getting [1250, 740, 1251]. Giant [1128, 227, 1067]. Gigantopithecus [9]. Giordano [664]. Giotto [1888]. Giraffe [983]. Giving [977]. Gland [1863]. Global [1431, 208]. Globular [1562, 104, 105, 1563]. Globular-Cluster [104, 105]. Globules [1494]. Glory [914]. Glycoproteins [882]. go [26]. Goiter [295]. Gothic [579]. Government [255]. Graaff [121]. Gradient [1184, 1185]. Graduate [1026]. Grape [901]. Graphics [884]. Gravitation [978]. Gravitational [274]. Gravity [1250, 1251]. Grazing [312]. Great [479, 178, 572]. Greece [558]. Greek [1317, 969]. Green [1527]. Ground [1597, 1769]. Grouse [1684]. Growing [1066, 739]. Growth [37, 430, 1897, 1449]. Guests [246]. Guide [1654]. Guided [847, 1815, 1816]. Guided-Wave [1816, 1815]. Gum [390].

Hadrons [313]. Hailstones [264]. Hales [1303]. Halley [1888]. Hallucinations [1571]. Halos [1671]. Happens [1206]. Head [968, 1821]. Health [1628, 255]. Heartbeat [674]. Heat [1525]. Heavy [1166, 1647, 1165]. Heavy-Water [1166, 1165]. Helium [77, 1398, 76]. Hemoglobin [1333, 1800]. Hepatitis [1507]. Herbarium [1480]. Hereditary [730]. Herpes [760]. High [778, 765, 263, 256, 323, 25, 1184, 1185, 1015, 1070, 632, 588, 779, 262, 257, 1069, 631]. High-Efficiency [765]. High-Energy [1015, 632, 631]. High-Gradient [1184, 1185]. High-Lysine [323]. High-Resolution [256, 257]. Highest [1716, 1714]. Highest-Energy [1716, 1714]. Hilbert [783]. Hilltop [1988]. Histocompatibility [1568]. Historical [1319]. History [943, 460, 1528, 867, 1903]. Hole [1312, 1313]. Holes [1413, 473, 994]. Holograms [1369, 1368]. Homing [999]. Hominid [917]. Hominids [1667, 1732]. Hormonal [1936]. Hormones [573, 1225, 1249]. Hospital [747]. Host [1401]. Hot [1332, 55]. Housework [985]. Human [1248, 145, 147, 946, 1047, 943, 1636, 947, 414, 193, 942, 1953, 1896, 1954, 645, 696, 1344, 265, 852, 2003, 911, 1345, 944, 146, 1613, 1116, 1206]. Hunger [1344]. Hunters [900]. Hunting [342]. Hybrid [911]. Hydra [995]. Hydraulic [1687]. Hydrogen [605, 1850, 1174]. Hyperactive [60]. Hypodermic [214]. Hypothalamus [573]. hypothesis [1036].

Ice [900]. Ice-Age [900]. Ideology [408]. II [1830]. Ills [743]. Illusions [1173, 280, 191]. Image [542, 1170]. Images [1718]. Imaginary [1693]. Immune [981, 710, 606]. Immunization [1236]. Immunology [1478, 1297]. Implantable [1998]. Implantation [662]. Implosion [894]. Importation [1458]. Imprinting [522]. Improvement [380]. Inapparent [831]. Income [556]. Index [201, 401, 600, 807, 1005, 1213, 1406, 1605, 1806, 2007]. India [1352, 1459]. Indian [678, 112, 94, 1387]. Indicators [1010]. Individuality [488]. Indochina [471]. Induced [213]. Induction [306, 279]. Industrial [344]. Inertial [42]. Infant [663, 359, 852, 674]. Infants [444]. Infectious [657]. Inflation [373]. Influence [1838]. Influenza [1599]. Information [347, 348]. Infrared [681, 658]. Ink [1871, 1867]. Ink-Jet [1871, 1867]. Inorganic [850]. Insect [126]. Insects [1512, 1670, 1257, 1235]. Insert [1401]. instrumentation [1550]. Integrated [862]. Integration [389, 21, 1544, 1696, 22, 1543]. Intelligence [155, 1898, 1085]. Interaction [913]. Interactions [778, 1225, 779]. Interactive [1047]. Intercellular [75]. Intercontinental [1979, 428]. Interference [194]. Interferon [1460, 306]. Intergroup [179]. Interior [641, 1525]. Internal [1445]. International [961, 1663, 1127]. Interplanetary [1157]. Interstellar [1616, 644]. Intervention [744, 746, 745]. Intravenous [476]. Intrinsically [1887]. Invertebrates [513]. Investment [255]. Ion [1852, 662]. Iran [296]. Iranian [1638]. Iron [1570, 1229, 1988]. Iroquois [224]. Irrigation [1581, 1318]. Isolation [723]. Isotopes [1433]. Issue [1541]. itself [413].

Japan [37]. Japanese [1370]. Jenner [1236]. Jet [1871, 1867]. Job [1579]. John [165]. Joint [1613]. Journal [1115]. Judaism [611]. jump [1483]. Junctions [1686]. Jupiter [1304, 1267, 1154].

Kangaroos [1526]. Karate [1870]. Kepler [446]. Key [1934]. Kilauea [1969]. Kimberlite [1669]. kind [1531]. Kittens [589]. Kiwi [1721]. Knee [1613]. Knots [1900].

Labor [1846]. Laboratory [193, 522]. Labrador [1387]. Lacertae [1521]. Lactase [561]. Lactose [561]. Lake [1832]. Lake-Bottom [1832]. Lakes [1969]. Land [1979, 1771]. Land-Based [1979]. Language [461, 1783, 1465, 563, 626]. Languages [2001, 281]. Large [22, 21, 1544, 1543]. Large-Scale [21, 1544, 22, 1543]. Laser [427, 29, 894, 38, 800, 289, 1885, 1433]. Lasers [307, 629, 630]. Last [1767]. Lava [1969]. Law [680]. Laws [1637, 446]. Layers [45]. Lead [222]. Leap [784]. Learning [5, 1719, 589]. Lectins [1497]. Leech [814]. Left [247]. Legal [1411]. Legionellosis [1967]. Lek [1684]. Length [1935]. Lens [1334, 1206]. Leonardo [230]. Leptons [1647]. Lesson [1496]. Letters [1, 17, 34, 50, 67, 84, 100, 117, 133, 152, 169, 186, 203, 219, 236, 252, 270, 286, 302, 318, 334, 354, 370, 386, 403, 419, 435, 451, 468, 484, 500, 518, 534, 553, 569, 585, 602, 619, 637, 653, 670, 686, 703, 719, 735, 756, 773, 791, 809, 825, 842, 858, 874, 890, 906, 922, 939, 958, 974, 990, 1007, 1023, 1041, 1058, 1075, 1091, 1108, 1124, 1141, 1162, 1179, 1197, 1215, 1245, 1262, 1278, 1293, 1308, 1216, 1325, 1340, 1359, 1376, 1391, 1408, 1424, 1439, 1455, 1470, 1487, 1502, 1535]. Letters [1556, 1574, 1590, 1607, 1623, 1641, 1657, 1674, 1690, 1708, 1725, 1741, 1758, 1774, 1790, 1808, 1825, 1841, 1857, 1874, 1891, 1907, 1923, 1939, 1958, 1974, 1991]. Lewis [506]. Life [558, 695, 1748, 508, 1580, 1492, 867, 738, 2012, 2013, 165, 232]. Lift [1188]. Light [427, 1522, 1523, 1698, 45, 26, 997, 1369, 439, 1235, 1116, 1368]. Light-Wave [1523, 1522]. Limitation [4, 206]. Limited [1379]. Line [329]. Linear [1964]. Linear-Chain [1964]. Lines [190]. Lions [1083]. Liquefied [1458]. Liquid [61, 62, 1312, 1313]. Liquid-Crystal [61, 62]. Liquids [693, 593]. Lithosphere [1630, 1189, 1629]. Liver [1095]. Lives [408]. Living [1686, 363]. Lizard [1936]. Load [44]. Locusts [327]. Logic [506]. Long [1781]. Long-Wavelength [1781]. Looking [1567, 1566]. Lost [506]. Love [110]. loving [1314]. Low [1191, 1190]. Lunar [38, 361, 120]. Lung [663]. Lymphocyte [696]. Lysine [323].

Macaques [1370]. Machining [813, 1785]. made [43]. Magnetic [294, 1836, 245, 1184, 1185]. Magnetism [326]. Magnets [196]. Majorities [87]. Make [1652]. making [349]. Malaria [95]. Malpractice [1328]. Mammals [1886, 1884]. Man [1916, 743, 365, 1348, 1330, 1018, 258, 1754]. Man-Carrying [1916]. Man-Day [365]. Managed [1027, 1028]. Manipulation [1112]. Mantle [1383, 1049]. Manufacture [1027, 1028]. many [1935]. Map [1569]. Mapping [265]. Marine [513, 189]. Mariner [609]. Markers [488]. Market [1084]. Marriage [1207, 1795]. Mars [1650, 1223, 1580, 71, 1506, 609, 1153]. Masers [1701]. Mass [1302, 1771]. Massive [1668]. Materials [147, 930, 708, 1287, 1208, 931]. Mathematical [2011, 231, 1082, 165, 13, 30, 46, 63, 80, 96, 113, 129, 148, 164, 182, 198, 232, 215, 248, 266, 282, 298, 314, 330, 350, 366, 382, 398, 415, 431, 447, 464, 480, 496, 514, 530, 549, 565, 581, 597, 786, 615, 633, 649, 666, 682, 699, 715, 731, 752, 769, 787, 804, 821, 838, 854, 870, 886, 902, 918, 935, 954, 970, 986, 1002, 1037, 1019, 1054, 1071, 1087, 1104, 1120, 1137, 1158, 1175, 1193, 1210, 1238, 1258, 1275, 1290, 1305, 1321, 1240, 1337, 1356, 1374, 1388, 1404, 1531, 1421, 1436, 1452, 1467, 1484, 1499]. Mathematical [1514, 1538, 1559, 1577, 1593, 1693, 1661, 1610, 1626, 1644, 1660, 1677, 1694, 1711, 1728, 1744, 1761, 1777, 1793, 1811, 1828, 1844, 1860, 1877, 1894, 1910, 1926, 1942, 1961, 1977, 1994]. Mathematics [1971, 1653, 1934]. Mathemtical [1037]. Mating [1684]. Matrixes [240]. Matter [675, 489]. Maurits [916]. Maya [477, 1451, 1171]. Measurement [365]. Mechanical [1298, 1288, 1118]. Mechanics [1413]. Mechanism [1016]. Mechanisms [1747, 714, 1952, 1951, 1785, 1188]. Medical [1683, 748, 749, 53, 750, 1026, 1838, 1328, 751, 861, 1127]. Medication [214]. Medicine [656, 738]. Medieval [127, 1373]. Meditation [429]. Mediterranean [590]. Medium [1616]. Melanism [1017]. Membrane [1720, 1852, 1968, 1314]. Membranes [846, 424, 1814, 1201].

Memories [1547]. Memory [328, 1719, 1495, 1982, 1981]. Mendel [1837]. Mental [1628]. Merchant [1171]. Mercury [273, 1148]. Mesozoa [596]. messenger [1529]. Messengers [1530]. Metabolism [730, 1265]. Metabolizes [1095]. Metal [1630, 1385, 726, 630, 629, 1629]. Metal-Oxide [726]. Metal-Vapor [629, 630]. Metals [612, 1511, 1070, 1069]. Metastasis [1848]. Meteorites [711, 489]. Meteorology [1267]. Metric [103]. Mexico [1351]. Micro [1552]. Micro-electronics [1552]. Microbial [1492]. Microcircuits [574, 1634]. Microcomputers [1080]. Microelectronic [1547, 1544, 1542, 1546, 1543]. Microelectronics [1553, 1551, 1540, 1549, 1541, 1550]. Microprocessors [1548]. Microscope [256, 410, 1965, 257]. Microstructure [627, 1208]. Microvascular [1664]. Microwave [422]. Mid [1133, 1134]. Mid-Atlantic [1133, 1134]. Middle [1222]. Migraines [280]. Migration [1771]. Migrations [947, 647]. Migratory [1136]. Military [1763]. Millikan [1822]. Millimeter [881]. millions [1531]. Mimicry [967]. Mind [1232]. Mine [1901]. Mineral [144, 713]. Mining [1200]. Mirages [1234]. Mirror [1801]. Missile [487]. Missiles [1979, 1111, 1427, 776]. Missing [1479]. Mississippi [1135]. Mitochondria [173]. Model [846, 995, 1918]. Models [12, 1033, 329, 1032]. Molecular [1984, 1883, 1079]. Molecule [657]. Molecules [523, 1797, 8, 644]. Mollusks [227]. Monomolecular [45]. monoxide [1880]. Moon [326, 562, 458, 1152]. Mosquito [1119]. Mosquitoes [1705]. Mössbauer [364]. Most [1029, 1052]. Mother [674]. Moths [1017, 492, 910]. Motion [1597, 1101, 1920]. Motions [1820]. Mound [1387]. Mountains [440]. Move [92, 963]. Movement [714, 1952, 1883]. Movements [1819, 968, 290]. Moving [1665, 1417, 1416]. Mule [197]. Multicellular [1751]. Multiple [106, 776]. Multiple-Warhead [776]. Multistability [394]. Muscle [1183, 59, 834]. muscles [472]. Muscular [445]. Mushrooms [1052]. Music [1103, 1661]. Musical [1173, 982]. Mutations [213]. Mycenaean [558]. Mystery [999].

National [503, 487]. Natural [1324, 1458, 522, 1165, 1166]. Natural-Uranium [1166, 1165]. Nature [1266, 523, 1011, 210, 833]. Navigation [42, 1316, 1235]. Neanderthals [2002]. Nebula [390, 691]. Nebulas [1804, 962]. Necessity [1221]. Negative [1396, 556, 1735, 109]. Negative-Income-Tax [556]. Negotiations [1520]. Neoglaciation [93]. Neolithic [1901, 1988, 1930]. Nerve [623, 108, 1852, 1897, 1225, 1715]. Nerve-Cell [623, 1852]. Nerve-Growth [1897]. Nervous [5, 814, 1634]. Network [111]. Networks [544, 1084]. Neural [847]. Neurobiology [927]. Neuron [1945]. Neurons [1946]. Neurophysiology [528]. Neurotransmitters [897]. Neutral [1000, 1984]. Neutrino [724]. Neutron [1364, 293]. Neutron-scattering [1364]. Neutrons [1902]. Newborn [663, 589]. Newfoundland [94]. Nighttime [1718]. Nile [1330, 1289]. Nitrogen [1447, 143, 965]. Nonsocial [1434]. Nonstandard [493]. Nonvisual [439]. Nourish [1348]. Nuclear [1395, 1797, 877, 1379, 1182, 1061, 1912, 406, 1663, 564]. Nuclear-Free [1182]. Nuclear-Test [406]. Nuclei [1698]. Nucleotide [1596]. Number [648, 329]. Numbers [694]. Nutrient [161]. Nutrition [836, 1347, 1345].

Object [359]. Objects [190, 1521, 1029, 1847]. Ocean [925, 627, 881, 678]. Oceanic [12, 1051, 1630, 1771, 1629]. Oceans [898, 867]. Off [59]. Offensive [206]. Oil [1296, 1646]. Old [740]. Oldest [275, 1449]. Omnivorous [607]. Opals [1286]. Operator [1228]. Operator-Repressor [1228]. Opiate [1445]. Opiates [1445]. Optical [194, 777]. Optics [1801, 1781, 862, 1816, 1815]. Oracle [1869]. Order [1370, 297]. Ore [1229]. Organic [8, 489]. Organisms [1738]. Organization [41, 1948, 751]. Orientation [327, 1136]. Origin [1432, 1630, 1145, 1748, 88, 128, 1629]. Origins [648, 929, 1916, 414, 527, 214, 623, 163]. Other [1400]. Otherwise [1693]. our [472]. Outer [1155]. Oxide [726]. Oxygen [141].

Pacific [785]. Paleolithic [1585, 1255]. Paleoneurology [1232]. Pangaea [156]. Parabolic [1053]. Paradox [1482]. Parasite [95]. Parasitic [967]. parsec [1914]. Particle [1765, 1862, 913, 1779]. Particle-Beam [1862]. Particles [1224, 1099, 26, 1033, 1564, 883, 1157, 1098, 1032]. Parts [1027, 1028]. Passive [1638]. Pathways [879, 309]. Patient [526]. Pattern [1509, 260, 1770]. patterns [786]. Patterson [2010]. Peer [1561]. Pelagic [1102]. People [11]. Perception [394, 1132, 578, 1396, 1851, 1685, 1101, 1062, 868, 290, 1920, 818, 1270, 1417, 1416]. Performance [1320, 239]. Permanent [196]. Pernor [337]. Personal [1553]. Personality [128]. Peru [1769]. Pesticides [58]. Phenomena [693]. Phobos [1428]. Phosphenes [27]. Photographic [1334]. Photon [1302]. Photons [313]. Photosynthesis [765, 997]. Photosynthetic [1968]. Photovoltaic [1363]. Physical [159]. Physics [707, 1870, 819, 1935, 1637]. Physiological [244]. Physiology [25, 944, 429, 983, 357]. Piano [1820]. Pictorial [578]. Pieter [1654]. Pigeon [999]. Pigments [1050]. Pineal [1863]. Pines [380]. Pipes [1669]. Pivot [1318]. Placebos [977]. Plain [1255]. Plan [124]. Planets [1155, 1316]. Planning [477]. Plant [1066, 1347]. Plants [1798, 1348, 1635, 213, 1016, 1751, 1796]. Plate [475, 713, 867]. Plays [2011]. Pluralistic [1394]. Poetic [1498]. Poisoning [222]. Poisonous [1052]. Poker [1722]. Polarized [1235]. Polarized-Light [1235]. Policy [706, 812]. Poliovirus [1079]. Political [454]. Pollutants [208]. Polycyclic [1266]. Polymeric [1208]. Polymers [850]. Pons [1384]. Population [943, 414, 942, 640, 430, 895, 952, 112]. Populations [946, 947, 951, 948]. Portrait [1888]. Positron [1099, 768, 1098, 767]. Positrons [1113]. Post [1429]. Potential [1928]. Power [411, 1221, 894, 1798, 979, 226, 349, 1663, 462, 557, 1779, 1796]. Powered [1916]. Pre [1135, 1930]. Pre-Columbian [1135]. Pre-Neolithic [1930]. Precursor [1697]. Prediction [1078]. Prehistoric [1769]. Prehistory [1317, 362]. Prematurity [595]. Prenatal [375]. Preservation [1704].

Pressure [427, 263, 1191, 262, 1190]. Primeval [1985]. Primitive [1029]. Principle [1685]. Printing [1871, 1867]. Prints [916]. Privacy [673]. Probe [1113]. Problem [1569, 1187, 783, 106, 1579, 1737]. Problems [243, 1887, 1935]. Process [893, 145, 147, 146]. Processes [1380]. Processing [1549]. Product [1595]. Production [145, 147, 1646, 1813, 1354, 349, 146]. Professionalization [1846]. Program [1586]. programmed [786]. Programming [2001]. Progress [1929]. Projections [1170]. Proliferation [1061]. Prompt [1912]. Proof [1082, 1483]. Prospects [640, 226]. Prostaglandins [379]. Protection [909]. Protein [1183, 460, 763, 915]. Protein-Cutting [915]. Proteins [834, 1031, 915]. Protohuman [1667]. Proton [779, 778, 293, 1881]. Psychiatric [746]. Psychiatrists [893]. Psychological [495]. Public [1934, 706]. Public-Key [1934]. Pulp [865]. Pulsars [210]. Pulsating [1100]. Pupfish [381]. Pupil [1192]. Purple [1314].

Quantum [1413, 1987]. Quark [1918]. Quarks [1168, 1381]. Quasars [277]. Quasi [190]. Quasi-Stellar [190]. Question [1612]. Qumran [378].

Rabbit [1333]. Race [155, 479]. Racial [389, 1696]. Radar [1567, 1566]. Radiation [711, 1679, 1491, 926, 1681]. Radicals [125]. Radio [428, 1128]. radioactive [1490]. Rain [1963, 799]. Rainbow [1466]. Ramapithecus [1474]. Randomness [1082]. Rangelands [28]. Rat [1481]. Ray [1562, 1046, 504, 853, 1287, 1366, 1563, 1045, 505]. Rays [1202, 1716, 1781, 1714, 1203]. Razed [1802]. Reactions [1015, 899]. Reactor [1324, 1929, 1443, 1442]. Reactors [1166, 172, 1490, 1165]. Reading [512]. Real [329]. Real-Number [329]. Reality [1987]. Reception [439]. Receptor [910]. Receptors [681, 1249, 1445]. Recognition [260, 782, 10]. Recombinant [1505]. Recombinant-DNA [1505]. Reconnaissance [622]. Reconstruction [1170]. Record [576]. Recreations [2013]. Recurrent [831]. Red [229, 55]. Reduction [1229]. Reefs [1932, 491]. Reflector [38]. Reflectors [211]. Regeneration [1509]. Regulation [1031]. Relations [674]. Remember [78]. Remnants [1202, 311, 1203]. Renaissance [1065]. Repeated [54]. Repel [1119]. Repellents [1119]. Repetitive [1380]. Replacement [1613]. Repressor [1228]. Repressors [89]. Reprocessing [1395]. Reproduction [232, 58, 944]. Requirements [1345]. Research [509, 510]. Reservoirs [1473]. Resistance [657, 1016]. Resolution [256, 257]. Resonance [1033, 1032]. Resources [339, 1222, 1353, 713, 1270]. Respiratory [1800]. Response [1435, 606, 423]. Responses [1498, 513, 2003]. Retina [610]. Retinex [1600]. Retrolental [1496]. Review [1561]. Revolution [457, 1498]. Reward [1786]. Ribosome [1364]. Rich [1784]. Rickets [195]. Rift [1133, 1134]. Right [247]. Rings [478]. Rise [934, 1171]. River [72]. RNA [1614, 407]. RNA-Directed [407]. Robert [1822]. Robot [1254]. Rocks [361, 1449]. Rodents [895]. Role [1051, 1103, 1192, 1551, 674, 1549, 1550]. Roman [1429, 1001, 1619, 1687]. Roots [677]. Rotating [899]. Rotation [1068, 631, 632, 396]. row [1483]. Rules [1837]. Running [1803]. Rural [1084].

Saccadic [1819]. Sacred [310]. Sage [1684]. Salt [181, 1314, 180, 1830, 1520]. Salt-loving [1314]. Samaritans [1420]. San [377]. Sand [1209]. Sardinia [1035]. Satellite [1431]. Satellites [1304]. Satisfaction [1044]. Saving [1879]. Scale [21, 1544, 1443, 22, 1543, 1442]. Scales [1935]. Scanning [256, 410, 257]. scattering [1364]. Scheduling [1653]. School [748]. Science [7, 24, 40, 56, 73, 90, 107, 123, 139, 158, 175, 192, 209, 225, 242, 259, 276, 292, 308, 324, 340, 360, 376, 392, 409, 425, 441, 459, 474, 490, 507, 524, 540, 559, 575, 591, 659, 608, 625, 643, 660, 676, 692, 709, 725, 742, 762, 780, 798, 815, 832, 848, 864, 880, 896, 912, 928, 945, 964, 980, 996, 1013, 1030, 1048, 1064, 1081, 1097, 1114, 1130, 1147, 1169, 1186, 1204, 1226, 1252, 1268, 1284, 1299, 1315, 1227, 1331, 1346, 1365, 1382, 1397, 1414, 1430, 1446]. Science [1461, 1477, 1493, 1508, 1524, 1545, 1565, 1582, 1598, 1615, 1633, 1649, 1666, 1682, 1717, 1733, 1749, 1766, 1782, 1799, 1817, 1833, 1849, 1865, 1882, 1899, 1915, 1947, 1966, 2000, 1561, 1983, 1931, 1699, 1281, 1552]. Science-Textbook [1281]. Scientific [2009, 595]. Scientist [14, 31, 47, 64, 81, 97, 114, 130, 149, 166, 183, 199, 216, 233, 249, 267, 283, 299, 315, 331, 351, 367, 383, 399, 416, 432, 448, 465, 481, 497, 515, 531, 550, 566, 582, 598, 616, 634, 650, 667, 683, 700, 716, 732, 753, 770, 788, 805, 822, 839, 855, 871, 887, 903, 919, 936, 955, 971, 987, 1003, 1020, 1038, 1055, 1072, 1088, 1105, 1121, 1138, 1159, 1176, 1194, 1211, 1239, 1259, 1515, 1532, 1554, 1572, 1588, 1604, 1621, 1639, 1655, 1672, 1688, 1706, 1723, 1739, 1756, 1772, 1788, 1805, 1823, 1839, 1855, 1872, 1889, 1905, 1921, 1937, 1956, 1972, 1989, 2006]. Sclerosis [106]. Sea [1172, 1492, 181, 180, 55, 1680]. Search [1224, 1580, 1085, 994]. Sebei [712]. Second [76, 77, 1530, 1529]. Secretion [1167]. Security [503, 487]. See [78]. Segments [54]. Seine [310]. Self [232]. Self-Reproduction [232]. Semiconductor [1476, 726, 1475]. Sensing [1282]. Sensitivity [610]. Separation [1184, 1433, 1185]. Sequence [1596]. Serengeti [312]. Services [1628]. Settlements [1255]. Sex [910]. Sex-Attractant [910]. Shad [647]. Shale [1919]. Shape [763, 1981, 1982, 363]. Shape-Memory [1982, 1981]. Shapes [8, 408]. Shaping [1702]. sharing [1667]. Sheet [1385]. Short [328]. Short-Term [328]. Sickle [1063]. Sickle-Cell [1063]. Side [1566, 1567]. Side-Looking [1567, 1566]. Silent [87]. Silviculture [1834]. similarity [1483]. Singing [1448, 728]. Sister [1207]. Sister-Exchange [1207]. Site [1415, 1988]. Size [1192]. Sky [504, 505]. Slavery [1096]. Sleep [1329]. Sling [761]. Slips [803]. Slow [831]. Small [1946, 1271]. Smaller [1156]. Smallpox [1362, 1236]. Snakes [681, 92]. Snow [613, 614]. Soap [1231]. Soaring [802]. Social [1083, 1272, 1370, 547, 1434, 1335, 577, 297]. Societies [1481]. Society [344, 342, 343]. Soil [120]. Solar [1145, 1444, 1029, 1156, 1419, 849, 764, 1144, 1768]. Solid [77, 1113, 223, 76].

Solidification [1510, 998]. Solids [693]. Solitaire [165]. Solitons [1835]. Solution [1569]. some [2013]. Song [110, 927]. Sorting [1274, 1273]. Sound [77, 76]. Source [1836, 310]. Sources [445, 658, 916]. Southern [1988, 1603, 380]. Soybeans [828]. Space [1336, 1718]. Spatial [984, 1495]. Special [1541]. Specializations [1953]. Specificity [623]. Spectroscopy [800, 364]. Spectrum [1850, 412]. Speech [426, 1434]. Spending [1763]. Spent [985]. Sperm [1821]. Spiders [1272]. Spin [1881]. Spots [1332]. Spread [1465]. Stability [1663]. Stairs [966]. Standard [1879]. Star [1933]. Stars [1464, 525, 1562, 1046, 104, 1100, 223, 1668, 1563, 1045, 105]. Starvation [357]. State [1113, 898, 229]. Stationary [640]. Statistics [1482]. Status [950, 1460]. Steady [898]. Stein [1482]. Stellar [190, 1136, 1731]. Stellar-Orientation [1136]. Stephen [1303]. Steps [1167]. Steroid [1249]. Stethoscope [1838]. Stick [1617]. Stirling [729]. Stomach [413]. Stomatopods [1230]. Stone [1330, 1704]. Stone-Age [1330]. Storage [1997, 1996]. Storms [1371]. Strategic [4, 1111]. Strategy [877, 1763]. Strength [593]. Stress [207, 495]. String [819]. Strings [1820]. Strip [1200]. Strip-Mining [1200]. Stroke [1664]. Structural [579]. Structure [641, 1568, 460, 122, 245, 424, 1616, 1818, 293, 962, 1800, 1614]. Studies [1364]. Subduction [1189]. Subjective [1283]. Subjects [1248]. Submarines [487]. Substances [1095]. Succession [1086]. Successors [2013]. Sulfur [74]. Summer [1904]. Sun [1068, 1146]. Sunlike [1464]. Sunspots [1479]. Superconductivity [262, 263]. Superconductors [245, 374, 462]. Superdense [176]. Superfluid [1398]. Supergravity [1637]. Superhard [930, 931]. Superior [594]. Supernova [1202, 311, 1203, 1768]. Supernovas [1933, 1400, 1319]. Superphenix [1442, 1443]. Support [1561]. Surface [1650, 1132, 1332, 1851, 560, 71, 1670, 1297, 564]. Surgery [29, 1664]. Surgical [744, 1613]. Swim [1129]. Switch [1183]. Symbiosis [325]. Synchronous [1301]. Synchrotron [1491]. Synthesis [1191, 426, 407, 1190]. Synthetic [1720]. Synthetic-Membrane [1720]. System [1886, 1083, 1145, 981, 5, 1136, 1029, 1156, 1419, 1787, 710, 849, 1228, 781, 814, 103, 1786, 1144, 1634, 1684, 1884, 458, 1768]. Systems [1254, 1638, 1998, 979, 1997, 1946, 1349, 1752, 1311, 125, 1996].

take [1531]. Taking [698, 697]. Talking [397]. Tall [695]. Tandem [121]. Tar [1102]. Targets [1417, 1416]. Task [656]. Tax [556]. Teaching [563]. Technology [422, 20, 1720, 934, 726, 1654, 1595, 953, 1687, 588, 646]. Tectonics [475, 713, 867]. Tektite [1737]. Telephone [1648]. Television [358]. Telling [247]. Temperature [492]. Temperatures [1070, 1735, 1069]. Temple [969, 1802]. Tensile [593]. Tephra [2005]. Teratomas [1868]. Term [328]. Terminals [545]. Test [406, 572]. Test-Ban [572]. Testimony [993]. Tests [1928]. Textbook [1281]. Texture [1062, 564]. Their [246, 1401]. Theorem [1767]. Theories [913]. Theory [1984, 1600, 1900, 1466, 1703, 978, 1285, 1700, 1987, 348]. Therapeutic [241]. thermodynamics [348]. Thermostat [1736]. Think [444]. Thinking [1955]. Three [2012, 2013, 1614]. Three-dimensional [1614]. Thunder [1117]. Tidal [1034]. Tides [796, 458, 797]. Tillage [1412]. Time [1879, 1205, 611, 985]. Timekeeping [1863]. Tissue [1225]. Tissues [1702]. Tokamak [509, 510, 1929]. Tomato [1734]. Tongue [803]. Tools [1585]. Top [881]. toppling [1483]. Total [476]. Towers [278]. traced [786]. Track [1403, 1402]. Tracks [1665, 1803]. Trading [322]. Trajectories [1819]. Trajectory [1053]. Transfer [953, 1614]. Transmission [462]. Transparency [868]. Transplant [863]. Transplantation [1999]. Transport [1800]. Tree [478]. Trees [695, 1386, 1118]. Tremor [244]. Trends [112]. Triangle [23]. trigger [1768]. Triticale [932]. Tropical [799]. Tube [853]. Tumors [1300]. Turkana [1732]. Turkey [39]. Turkeys [297]. Turned [59]. Two [675, 446]. Two-Dimensional [675]. Tycho [801].

U.S. [1394, 1846, 1350, 28, 812]. Ukraine [900]. Ultracold [1902]. Ultrafast [693]. Ultrasound [1683]. Underdeveloped [951, 953]. Underground [1473]. Unification [1637]. Unified [913]. Uniqueness [595]. Universe [338, 885, 1336, 1269]. Unsolved [243]. Unusual [1188]. Upsilon [1765]. Uranium [1166, 1165]. Urban [1386, 1135]. Uses [1491, 127].

Vacuoles [1527]. Vacuum [2004]. Vapor [629, 630]. Vascularization [1300]. Venture [322]. Venus [1149, 1150]. Verbal [541]. Verification [1830]. Vertebrate [1736]. Vietnam [87]. Views [79]. Village [1373, 39, 1930, 112]. Vines [901]. Violate [1837]. Violent [796, 797]. Viral [1596, 1507]. Virus [1913, 1780, 606]. Viruses [1401, 831, 760]. Viscosity [109]. Vision [1951, 1600, 528]. Visual [1620, 1396, 1384, 1685, 542, 879, 1101, 1062, 290, 1920, 1050, 160]. Visualization [642]. Visually [847]. Voice [1448]. Volcanoes [1223]. Vortexes [851]. Voting [1311]. Vultures [802].

Wakes [851]. Walking [1399]. Walls [1066]. Wankel [521]. War [1379, 1912, 87]. Warfare [503]. Warhead [776]. Warm [624]. was [1680]. Wasps [1209]. Waste [925]. wastes [1490]. Water [1473, 176, 1166, 1670, 661, 140, 229, 1165]. Wave [1523, 1816, 1522, 1815]. Wavelength [1781]. Waves [1444, 560, 274]. Wax [1051]. Ways [2011]. Weak [1000]. Weaker [1250, 1251]. Weapon [761]. Weapons [877, 1061, 1679, 70, 1862, 1663, 206]. Wear [883]. Weaver [1602]. Wegener [1036]. Western [1200, 28]. Whale [1821]. Whistles [6]. White [1661, 1368, 1369]. White-Light [1369, 1368]. Whites [1851]. Will [1269]. Winchester [878]. Wind [837, 1444]. Winds [1731]. Wines [901]. Winning [2011]. without [1904, 1412]. Women [950, 408]. Wood [865]. Woodhenges [174]. Words [1620]. Work [1044]. World [124, 359, 1646, 640, 1813, 1222]. Worm [2010, 786]. would [1531]. Wright [1787]. Writing [1697].

X [1202, 1203, 1562, 1563, 1045, 1046, 504, 505, 1781, 1287]. X-ray [1563, 505, 1562, 504, 1287]. X-ray-emitting [1045, 1046]. X-rays [1203].

Yarn [592]. Year [1904]. Years [2, 18, 35, 51, 68, 85, 101, 118, 134, 153, 170, 187, 204, 220, 237, 253, 271, 287, 303, 319, 335, 355, 371, 387, 404, 420, 436, 452, 469, 485, 501, 519, 535, 554, 570, 586, 603, 620, 638, 654, 671, 687, 704, 720, 736, 757, 774, 792, 810, 826, 843, 859, 875, 891, 907, 923, 940, 959, 975, 991, 1008, 1024, 1042, 1059, 1076, 1092, 1109, 1125, 1142, 1163, 1180, 1198, 1217, 1246, 1263, 1279, 1294, 1309, 1218, 1326, 1341, 1360, 1377, 1392, 1409, 1425, 1440]. Years [1456, 1471, 1488, 1503, 1518, 1536, 1557, 1575, 1591, 1608, 1624, 1642, 1658, 1675, 1691, 1709, 1726, 1742, 1759, 1775, 1791, 1809, 1826, 1842, 1858, 1875, 1892, 1908, 1924, 1940, 1959, 1975, 1992, 1317, 1531]. York [11].

Zone [1034]. Zones [1182]. Zoo [1693].

References

Anonymous:1970:La

[1] Anonymous. Letters. Scientific American, 222(1):6, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

      URL http://www.nature.com/scientificamerican/journal/v222/
      n1/pdf/scientificamerican0170-6.pdf.

Anonymous:1970:YAa

[2] Anonymous. 50 and 100 years ago. Scientific American, 222(1):10, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

      (electronic). URL http://www.nature.com/scientificamerican/
      journal/v222/n1/pdf/scientificamerican0170-10.pdf.

Anonymous:1970:Aa

[3] Anonymous. The authors. Scientific American, 222(1):16, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

      URL http://www.nature.com/scientificamerican/journal/v222/
      n1/pdf/scientificamerican0170-16.pdf.

Rathjens:1970:LSA

[4] G. W. Rathjens and G. B. Kistiakowsky. The limitation of strategic arms. Scientific American, 222(1):19–29, January 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

      URL http://www.nature.com/scientificamerican/journal/v222/
      n1/pdf/scientificamerican0170-19.pdf.

DiCara:1970:LAN

[5] Leo V. DiCara. Learning in the autonomic nervous system. Scientific American, 222(1):30–39, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n1/pdf/scientificamerican0170-
30.pdf.

Chanaud:1970:AW

[6] Robert C. Chanaud. Aerodynamic whistles. Scientific American, 222(1):40–47, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n1/pdf/scientificamerican0170-
40.pdf.

Anonymous:1970:SCa

[7] Anonymous. Science and the citizen. Scientific American, 222(1):48–52, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n1/pdf/scientificamerican0170-48.pdf.

Lambert:1970:SOM

[8] Joseph B. Lambert. The shapes of organic molecules. Scientific American, 222(1):58–75, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n1/pdf/scientificamerican0170-
58.pdf.

Simons:1970:G

[9] Elwyn L. Simons and Peter C. Ettel. Gigantopithecus. Scientific American, 222(1):76–85, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n1/pdf/scientificamerican0170-
76.pdf.

Luria:1970:RDB

[10] Salvador E. Luria. The recognition of DNA in bacteria. Scientific American, 222(1):88–102, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n1/pdf/scientificamerican0170-
88.pdf.

Cowgill:1970:PY

[11] Ursula M. Cowgill. The people of York: 1538–1812. Scientific American, 222(1):104–112, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n1/pdf/scientificamerican0170-
104.pdf.

Baker:1970:MOC

[12] D. James Baker, Jr. Models of oceanic circulation. Scientific American, 222(1):114–121, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n1/pdf/scientificamerican0170-
114.pdf.

Gardner:1970:MGa

[13] Martin Gardner. Mathematical games. Scientific American, 222 (1):124–127, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n1/pdf/scientificamerican0170-
124.pdf.

Stong:1970:ASa

[14] C. L. Stong. The amateur scientist. Scientific American, 222(1):130–139, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n1/pdf/scientificamerican0170-130.pdf.

Morrison:1970:Ba

[15] Philip Morrison. Books. Scientific American, 222(1):141–144, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n1/pdf/scientificamerican0170-141.pdf.

Anonymous:1970:Ba

[16] Anonymous. Bibliography. Scientific American, 222(1):146, January 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n1/pdf/scientificamerican0170-146.pdf.

Anonymous:1970:Lb

[17] Anonymous. Letters. Scientific American, 222(2):6–7, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n2/pdf/scientificamerican0270-6.pdf.

Anonymous:1970:YAb

[18] Anonymous. 50 and 100 years ago. Scientific American, 222(2):9, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n2/pdf/scientificamerican0270-9.pdf.

Anonymous:1970:Ab

[19] Anonymous. The authors. Scientific American, 222(2):10, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n2/pdf/scientificamerican0270-10.pdf.

Brooks:1970:AT

[20] Harvey Brooks and Raymond Bowers. The assessment of technology. Scientific American, 222(2):13–21, February 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n2/pdf/scientificamerican0270-13.pdf.

Heath:1970:LSI

[21] F. G. Heath. Large-scale integration in electronics. Scientific American, 222(2):22–31, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n2/pdf/scientificamerican0270-
22.pdf.

Heath:1970:LIE

[22] F. G. Heath. Large-scale integration in electronics. Scientific American, 222(2):22–32, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Tazieff:1970:AT

[23] Haroun Tazieff. The afar triangle. Scientific American, 222(2):32–41, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n2/pdf/scientificamerican0270-32.pdf.

Anonymous:1970:SCb

[24] Anonymous. Science and the citizen. Scientific American, 222(2):42–46, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n2/pdf/scientificamerican0270-42.pdf.

Hock:1970:PHA

[25] Raymond J. Hock. The physiology of high altitude. Scientific American, 222(2):52–62, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n2/pdf/scientificamerican0270-
52.pdf.

Feinberg:1970:PGF

[26] Gerald Feinberg. Particles that go faster than light. Scientific American, 222(2):68–77, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n2/pdf/scientificamerican0270-
68.pdf.

Oster:1970:P

[27] Gerald Oster. Phosphenes. Scientific American, 222(2):82–87, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n2/pdf/scientificamerican0270-82.pdf.

Love:1970:RWU

[28] R. Merton Love. The rangelands of the Western U.S. Scientific American, 222(2):88–96, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n2/pdf/scientificamerican0270-
88.pdf.

Berns:1970:CSL

[29] Michael W. Berns and Donald E. Rounds. Cell surgery by laser. Scientific American, 222(2):98–110, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v222/n2/pdf/scientificamerican0270-
98.pdf.

Gardner:1970:MGb

[30] Martin Gardner. Mathematical games. Scientific American, 222 (2):112–114, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n2/pdf/scientificamerican0270-
112.pdf.

Stong:1970:ASb

[31] C. L. Stong. The amateur scientist. Scientific American, 222(2):116–120, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n2/pdf/scientificamerican0270-116.pdf.

Morrison:1970:Bb

[32] Philip Morrison. Books. Scientific American, 222(2):122–125, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n2/pdf/scientificamerican0270-122.pdf.

Anonymous:1970:Bb

[33] Anonymous. Bibliography. Scientific American, 222(2):126, February 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n2/pdf/scientificamerican0270-126.pdf.

Anonymous:1970:Lc

[34] Anonymous. Letters. Scientific American, 222(3):6–8, March 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n3/pdf/scientificamerican0370-6.pdf.

Anonymous:1970:YAc

[35] Anonymous. 50 and 100 years ago. Scientific American, 222(3):14–18, March 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n3/pdf/scientificamerican0370-14.pdf.

Anonymous:1970:Ac

[36] Anonymous. The authors. Scientific American, 222(3):25–26, March 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n3/pdf/scientificamerican0370-25.pdf.

Abegglen:1970:EGJ

[37] James C. Abegglen. The economic growth of Japan. Scientific American, 222(3):31–37, March 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n3/pdf/scientificamerican0370-
31.pdf.

Faller:1970:LLR

[38] James E. Faller and E. Joseph Wampler. The lunar laser reflector. Scientific American, 222(3):38–49, March 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n3/pdf/scientificamerican0370-38.pdf.

Cambel:1970:EFV

[39] Halet Cambel and Robert J. Braidwood. An early farming village in Turkey. Scientific American, 222(3):50–57, March 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n3/pdf/scientificamerican0370-50.pdf.

Anonymous:1970:SCc

[40] Anonymous. Science and the citizen. Scientific American, 222(3):58–64, March 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n3/pdf/scientificamerican0370-58.pdf.

Luria:1970:FOB

[41] A. R. Luria. The functional organization of the brain. Scientific American, 222(3):66–78, March 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n3/pdf/scientificamerican0370-
66.pdf.

Leondes:1970:INA

[42] Cornelius T. Leondes. Inertial navigation for aircraft. Scientific American, 222(3):80–86, March 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n3/pdf/scientificamerican0370-
80.pdf.

Taylor:1970:HEM

[43] T. G. Taylor. How an eggshell is made. Scientific American, 222(3):88– 95, March 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n3/pdf/scientificamerican0370-88.pdf.

Wills:1970:GL

[44] Christopher Wills. Genetic load. Scientific American, 222(3):98–107, March 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n3/pdf/scientificamerican0370-98.pdf.

Drexhage:1970:MLL

[45] Karl H. Drexhage. Monomolecular layers and light. Scientific American, 222(3):108–119, March 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n3/pdf/scientificamerican0370-
108.pdf.

Gardner:1970:MGc

[46] Martin Gardner. Mathematical games. Scientific American, 222(3):121– 124, March 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n3/pdf/scientificamerican0370-121.pdf.

Stong:1970:ASc

[47] C. L. Stong. The amateur scientist. Scientific American, 222(3):131–136, March 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n3/pdf/scientificamerican0370-131.pdf.

Morrison:1970:Bc

[48] Philip Morrison. Books. Scientific American, 222(3):141–144, March 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n3/pdf/scientificamerican0370-141.pdf.

Anonymous:1970:Bc

[49] Anonymous. Bibliography. Scientific American, 222(3):146, March 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n3/pdf/scientificamerican0370-146.pdf.

Anonymous:1970:Ld

[50] Anonymous. Letters. Scientific American, 222(4):6, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n4/pdf/scientificamerican0470-6.pdf.

Anonymous:1970:YAd

[51] Anonymous. 50 and 100 years ago. Scientific American, 222(4):8, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n4/pdf/scientificamerican0470-8.pdf.

Anonymous:1970:Ad

[52] Anonymous. The authors. Scientific American, 222(4):12, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n4/pdf/scientificamerican0470-12.pdf.

Garfield:1970:DMC

[53] Sidney R. Garfield. The delivery of medical care. Scientific American, 222(4):15–23, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n4/pdf/scientificamerican0470-
15.pdf.

Britten:1970:RSD

[54] Roy J. Britten and David E. Kohne. Repeated segments of DNA. Scientific American, 222(4):24–31, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n4/pdf/scientificamerican0470-
24.pdf.

Degens:1970:RSH

[55] Egon T. Degens and David A. Ross. The Red Sea hot brines. Scientific American, 222(4):32–42, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n4/pdf/scientificamerican0470-
32.pdf.

Anonymous:1970:SCd

[56] Anonymous. Science and the citizen. Scientific American, 222(4):44–48, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n4/pdf/scientificamerican0470-44.pdf.

Wahl:1970:CC

[57] Arnold C. Wahl. Chemistry by computer. Scientific American, 222(4):54–70, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n4/pdf/scientificamerican0470-
54.pdf.

Peakall:1970:PRB

[58] David B. Peakall. Pesticides and the reproduction of birds. Scientific American, 222(4):72–78, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n4/pdf/scientificamerican0470-
72.pdf.

Hoyle:1970:HMT

[59] Graham Hoyle. How is muscle turned on and off. Scientific American, 222(4):84–93, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n4/pdf/scientificamerican0470-
84.pdf.

Stewart:1970:HC

[60] Mark A. Stewart. Hyperactive children. Scientific American, 222(4):94– 98, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n4/pdf/scientificamerican0470-94.pdf.

Heilmeier:1970:LCD

[61] George H. Heilmeier. Liquid-crystal display devices. Scientific American, 222(4):100–106, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n4/pdf/scientificamerican0470-
100.pdf.

Heilmeier:1970:LDD

[62] G. H. Heilmeier. Liquid-crystal display devices. Scientific American, 222 (4):100–106, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gardner:1970:MGd

[63] Martin Gardner. Mathematical games. Scientific American, 222(4):108– 112, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n4/pdf/scientificamerican0470-108.pdf.

Stong:1970:ASd

[64] C. L. Stong. The amateur scientist. Scientific American, 222(4):114– 120, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n4/pdf/scientificamerican0470-114.pdf.

Gopal:1970:B

[65] S. Gopal. Books. Scientific American, 222(4):122–129, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n4/pdf/scientificamerican0470-122.pdf.

Anonymous:1970:Bd

[66] Anonymous. Bibliography. Scientific American, 222(4):130, April 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n4/pdf/scientificamerican0470-130.pdf.

Anonymous:1970:Le

[67] Anonymous. Letters. Scientific American, 222(5):6–7, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n5/pdf/scientificamerican0570-6.pdf.

Anonymous:1970:YAe

[68] Anonymous. 50 and 100 years ago. Scientific American, 222(5):8, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n5/pdf/scientificamerican0570-8.pdf.

Anonymous:1970:Ae

[69] Anonymous. The authors. Scientific American, 222(5):10, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n5/pdf/scientificamerican0570-10.pdf.

Meselson:1970:CBW

[70] Matthew S. Meselson. Chemical and biological weapons. Scientific American, 222(5):15–25, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n5/pdf/scientificamerican0570-
15.pdf.

Leighton:1970:SM

[71] Robert B. Leighton. The surface of Mars. Scientific American, 222(5): 26–41, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n5/pdf/scientificamerican0570-26.pdf.

Merriman:1970:CR

[72] Daniel Merriman. The calefaction of a river. Scientific American, 222(5): 42–52, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n5/pdf/scientificamerican0570-42.pdf.

Anonymous:1970:SCe

[73] Anonymous. Science and the citizen. Scientific American, 222(5):54– 58, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n5/pdf/scientificamerican0570-54.pdf.

Pratt:1970:S

[74] Christopher J. Pratt. Sulfur. Scientific American, 222(5):62–72, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n5/pdf/scientificamerican0570-62.pdf.

Loewenstein:1970:IC

[75] Werner R. Loewenstein. Intercellular communication. Scientific American, 222(5):78–86, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n5/pdf/scientificamerican0570-
78.pdf.

Bertman:1970:SSH

[76] B. Bertman and D. J. Sandiford. ‘second sound’ in solid helium. Scientific American, 222(5):92–101, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bertman:1970:SSS

[77] Bernard Bertman and David J. Sandiford. “Second sound” in solid helium. Scientific American, 222(5):92–101, May 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n5/pdf/scientificamerican0570-92.pdf.

Haber:1970:HWR

[78] Ralph Norman Haber. How we remember what we see. Scientific American, 222(5):104–112, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n5/pdf/scientificamerican0570-
104.pdf.

Holliday:1970:EVF

[79] Leslie Holliday. Early views on forces between atoms. Scientific American, 222(5):116–122, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n5/pdf/scientificamerican0570-
116.pdf.

Gardner:1970:MGe

[80] Martin Gardner. Mathematical games. Scientific American, 222(5):124– 127, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n5/pdf/scientificamerican0570-124.pdf.

Stong:1970:ASe

[81] C. L. Stong. The amateur scientist. Scientific American, 222(5):130– 134, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n5/pdf/scientificamerican0570-130.pdf.

Morrison:1970:Bd

[82] Philip Morrison. Books. Scientific American, 222(5):139–146, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n5/pdf/scientificamerican0570-139.pdf.

Anonymous:1970:Be

[83] Anonymous. Bibliography. Scientific American, 222(5):148, May 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n5/pdf/scientificamerican0570-148.pdf.

Anonymous:1970:Lf

[84] Anonymous. Letters. Scientific American, 222(6):6, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n6/pdf/scientificamerican0670-6.pdf.

Anonymous:1970:YAf

[85] Anonymous. 50 and 100 years ago. Scientific American, 222(6):8, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n6/pdf/scientificamerican0670-8.pdf.

Anonymous:1970:Af

[86] Anonymous. The authors. Scientific American, 222(6):12, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n6/pdf/scientificamerican0670-12.pdf.

Converse:1970:SMV

[87] Philip E. Converse and Howard Schuman. “Silent majorities” and the Vietnam War. Scientific American, 222(6):17–25, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n6/pdf/scientificamerican0670-17.pdf.

Rees:1970:OG

[88] Martin J. Rees and Joseph Silk. The origin of galaxies. Scientific American, 222(6):26–35, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n6/pdf/scientificamerican0670-
26.pdf.

Ptashne:1970:GR

[89] Mark Ptashne and Walter Gilbert. Genetic repressors. Scientific American, 222(6):36–44, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n6/pdf/scientificamerican0670-
36.pdf.

Anonymous:1970:SCf

[90] Anonymous. Science and the citizen. Scientific American, 222(6):46–52, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n6/pdf/scientificamerican0670-46.pdf.

Sutherland:1970:CD

[91] Ivan E. Sutherland. Computer displays. Scientific American, 222(6):56– 81, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n6/pdf/scientificamerican0670-56.pdf.

Gans:1970:HSM

[92] Carl Gans. How snakes move. Scientific American, 222(6):82–96, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n6/pdf/scientificamerican0670-82.pdf.

Denton:1970:N

[93] George H. Denton and Stephen C. Porter. Neoglaciation. Scientific American, 222(6):100–110, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n6/pdf/scientificamerican0670-
100.pdf.

Tuck:1970:AIC

[94] James A. Tuck. An archaic Indian cemetery in Newfoundland. Scientific American, 222(6):112–121, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n6/pdf/scientificamerican0670-
112.pdf.

Hawking:1970:CMP

[95] Frank Hawking. The clock of the malaria parasite. Scientific American, 222(6):123–131, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v222/n6/pdf/scientificamerican0670-
123.pdf.

Gardner:1970:MGf

[96] Martin Gardner. Mathematical games. Scientific American, 222(6):132– 136, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n6/pdf/scientificamerican0670-132.pdf.

Stong:1970:ASf

[97] C. L. Stong. The amateur scientist. Scientific American, 222(6):141– 145, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v222/n6/pdf/scientificamerican0670-141.pdf.

Morrison:1970:Be

[98] Philip Morrison. Books. Scientific American, 222(6):146–151, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n6/pdf/scientificamerican0670-146.pdf.

Anonymous:1970:Bf

[99] Anonymous. Bibliography. Scientific American, 222(6):152, June 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v222/
n6/pdf/scientificamerican0670-152.pdf.

Anonymous:1970:Lg

[100] Anonymous. Letters. Scientific American, 223(1):6–7, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n1/pdf/scientificamerican0770-6.pdf.

Anonymous:1970:YAg

[101] Anonymous. 50 and 100 years ago. Scientific American, 223(1):10, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n1/pdf/scientificamerican0770-10.pdf.

Anonymous:1970:Ag

[102] Anonymous. The authors. Scientific American, 223(1):15, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n1/pdf/scientificamerican0770-15.pdf.

Ritchie-Calder:1970:CMS

[103] Lord Ritchie-Calder. Conversion to the metric system. Scientific American, 223(1):17–25, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n1/pdf/scientificamerican0770-
17.pdf.

Iben:1970:GCS

[104] Icko Iben, Jr. Globular-cluster stars. Scientific American, 223(1):26– 39, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n1/pdf/scientificamerican0770-26.pdf.

Iben:1970:GS

[105] I. Iben, Jr. Globular-cluster stars. Scientific American, 223(1):26–39, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Dean:1970:MSP

[106] Geoffrey Dean. The multiple sclerosis problem. Scientific American, 223(1):40–46, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n1/pdf/scientificamerican0770-
40.pdf.

Anonymous:1970:SCg

[107] Anonymous. Science and the citizen. Scientific American, 223(1):48– 52, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n1/pdf/scientificamerican0770-48.pdf.

Kandel:1970:NCB

[108] Eric R. Kandel. Nerve cells and behavior. Scientific American, 223(1): 57–70, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n1/pdf/scientificamerican0770-57.pdf.

Starr:1970:NV

[109] Victor P. Starr and Norman E. Gaut. Negative viscosity. Scientific American, 223(1):72–80, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n1/pdf/scientificamerican0770-
72.pdf.

Bennet-Clark:1970:LSF

[110] H. C. Bennet-Clark and A. W. Ewing. The love song of the fruit fly. Scientific American, 223(1):84–92, July 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n1/pdf/scientificamerican0770-84.pdf.

Frank:1970:NA

[111] Howard Frank and Ivan T. Frisch. Network analysis. Scientific American, 223(1):94–103, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n1/pdf/scientificamerican0770-
94.pdf.

Taylor:1970:PTI

[112] Carl E. Taylor. Population trends in an Indian village. Scientific American, 223(1):106–114, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n1/pdf/scientificamerican0770-
106.pdf.

Gardner:1970:MGg

[113] Martin Gardner. Mathematical games. Scientific American, 223(1):117– 119, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n1/pdf/scientificamerican0770-117.pdf.

Stong:1970:ASg

[114] C. L. Stong. The amateur scientist. Scientific American, 223(1):120– 128, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n1/pdf/scientificamerican0770-120.pdf.

Morrison:1970:Bf

[115] Philip Morrison. Books. Scientific American, 223(1):131–135, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n1/pdf/scientificamerican0770-131.pdf.

Anonymous:1970:Bg

[116] Anonymous. Bibliography. Scientific American, 223(1):136, July 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n1/pdf/scientificamerican0770-136.pdf.

Anonymous:1970:Lh

[117] Anonymous. Letters. Scientific American, 223(2):6, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n2/pdf/scientificamerican0870-6.pdf.

Anonymous:1970:YAh

[118] Anonymous. 50 and 100 years ago. Scientific American, 223(2):10, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n2/pdf/scientificamerican0870-10.pdf.

Anonymous:1970:Ah

[119] Anonymous. The authors. Scientific American, 223(2):12, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n2/pdf/scientificamerican0870-12.pdf.

Wood:1970:LS

[120] John A. Wood. The lunar soil. Scientific American, 223(2):14–23, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n2/pdf/scientificamerican0870-14.pdf.

Rose:1970:TVG

[121] Peter H. Rose and Andrew B. Wittkower. Tandem Van De Graaff accelerators. Scientific American, 223(2):24–33, August 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n2/pdf/scientificamerican0870-24.pdf.

Edelman:1970:SFA

[122] Gerald M. Edelman. The structure and function of antibodies. Scientific American, 223(2):34–42, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n2/pdf/scientificamerican0870-
34.pdf.

Anonymous:1970:SCh

[123] Anonymous. Science and the citizen. Scientific American, 223(2):44–48, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n2/pdf/scientificamerican0870-44.pdf.

Boerma:1970:WAP

[124] Addeke H. Boerma. A world agricultural plan. Scientific American, 223(2):54–69, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n2/pdf/scientificamerican0870-
54.pdf.

Pryor:1970:FRB

[125] William A. Pryor. Free radicals in biological systems. Scientific American, 223(2):70–83, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n2/pdf/scientificamerican0870-
70.pdf.

Hinton:1970:IE

[126] H. E. Hinton. Insect eggshells. Scientific American, 223(2):84–91, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n2/pdf/scientificamerican0870-84.pdf.

White:1970:MUA

[127] Lynn White, Jr. Medieval uses of air. Scientific American, 223(2):92–100, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n2/pdf/scientificamerican0870-92.pdf.

Thomas:1970:OP

[128] Alexander Thomas, Stella Chess, and Herbert G. Birch. The origin of personality. Scientific American, 223(2):102–109, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n2/pdf/scientificamerican0870-102.pdf.

Gardner:1970:MGh

[129] Martin Gardner. Mathematical games. Scientific American, 223 (2):110–113, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n2/pdf/scientificamerican0870-
110.pdf.

Stong:1970:ASh

[130] C. L. Stong. The amateur scientist. Scientific American, 223(2):116–121, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n2/pdf/scientificamerican0870-116.pdf.

Morrison:1970:Bg

[131] Philip Morrison. Books. Scientific American, 223(2):123–127, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n2/pdf/scientificamerican0870-123.pdf.

Anonymous:1970:Bh

[132] Anonymous. Bibliography. Scientific American, 223(2):128, August 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n2/pdf/scientificamerican0870-128.pdf.

Anonymous:1970:Li

[133] Anonymous. Letters. Scientific American, 223(3):10–12, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n3/pdf/scientificamerican0970-10.pdf.

Anonymous:1970:YAi

[134] Anonymous. 50 and 100 years ago. Scientific American, 223(3):18–26, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n3/pdf/scientificamerican0970-18.pdf.

Anonymous:1970:Ai

[135] Anonymous. The authors. Scientific American, 223(3):33–38, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n3/pdf/scientificamerican0970-33.pdf.

Hutchinson:1970:B

[136] G. Evelyn Hutchinson. The Biosphere. Scientific American, 223 (3):44–53, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n3/pdf/scientificamerican0970-
44.pdf.

Oort:1970:ECE

[137] Abraham H. Oort. The energy cycle of the Earth. Scientific American, 223(3):54–63, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n3/pdf/scientificamerican0970-
54.pdf.

Woodwell:1970:ECB

[138] George M. Woodwell. The energy cycle of the Biosphere. Scientific American, 223(3):64–74, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n3/pdf/scientificamerican0970-
64.pdf.

Anonymous:1970:SCi

[139] Anonymous. Science and the citizen. Scientific American, 223(3):78–94, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n3/pdf/scientificamerican0970-78.pdf.

Penman:1970:WC

[140] H. L. Penman. The water cycle. Scientific American, 223(3):98–108, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n3/pdf/scientificamerican0970-98.pdf.

Cloud:1970:OC

[141] Preston Cloud and Aharon Gibor. The oxygen cycle. Scientific American, 223(3):110–123, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n3/pdf/scientificamerican0970-
110.pdf.

Bolin:1970:CC

[142] Bert Bolin. The carbon cycle. Scientific American, 223(3):124–132, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n3/pdf/scientificamerican0970-124.pdf.

Delwiche:1970:NC

[143] C. C. Delwiche. The nitrogen cycle. Scientific American, 223(3):136–146, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n3/pdf/scientificamerican0970-136.pdf.

Deevey:1970:MC

[144] Edward S. Deevey, Jr. Mineral cycles. Scientific American, 223 (3):148–158, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n3/pdf/scientificamerican0970-
148.pdf.

Brown:1970:HFP

[145] Lester R. Brown. Human food production as a process in the Biosphere. Scientific American, 223(3):160–170, September 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n3/pdf/scientificamerican0970-160.pdf.

Singer:1970:HEP

[146] S. Fred Singer. Human energy production as a process in the Biosphere. Scientific American, 223(3):174–190, September 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n3/pdf/scientificamerican0970-174.pdf.

Brown:1970:HMP

[147] Harrison Brown. Human materials production as a process in the Biosphere. Scientific American, 223(3):194–208, September 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n3/pdf/scientificamerican0970-194.pdf.

Gardner:1970:MGi

[148] Martin Gardner. Mathematical games. Scientific American, 223 (3):210–218, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n3/pdf/scientificamerican0970-
210.pdf.

Stong:1970:ASi

[149] C. L. Stong. The amateur scientist. Scientific American, 223(3):221–234, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n3/pdf/scientificamerican0970-221.pdf.

Morrison:1970:Bh

[150] Philip Morrison. Books. Scientific American, 223(3):239–261, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n3/pdf/scientificamerican0970-239.pdf.

Anonymous:1970:Bi

[151] Anonymous. Bibliography. Scientific American, 223(3):262–266, September 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n3/pdf/scientificamerican0970-262.pdf.

Anonymous:1970:Lj

[152] Anonymous. Letters. Scientific American, 223(4):6, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n4/pdf/scientificamerican1070-6.pdf.

Anonymous:1970:YAj

[153] Anonymous. 50 and 100 years ago. Scientific American, 223(4):8–9, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n4/pdf/scientificamerican1070-8.pdf.

Anonymous:1970:Aj

[154] Anonymous. The authors. Scientific American, 223(4):15, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n4/pdf/scientificamerican1070-15.pdf.

Bodmer:1970:IR

[155] Walter F. Bodmer and Luigi Luca Cavalli-Sforza. Intelligence and race. Scientific American, 223(4):19–29, October 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n4/pdf/scientificamerican1070-19.pdf.

Dietz:1970:BP

[156] Robert S. Dietz and John C. Holden. The breakup of Pangaea. Scientific American, 223(4):30–41, October 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n4/pdf/scientificamerican1070-30.pdf.

Rasmussen:1970:C

[157] Howard Rasmussen and Maurice M. Pechet. Calcitonin. Scientific American, 223(4):42–51, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n4/pdf/scientificamerican1070-
42.pdf.

Anonymous:1970:SCj

[158] Anonymous. Science and the citizen. Scientific American, 223(4):52–61, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n4/pdf/scientificamerican1070-52.pdf.

Taylor:1970:FPC

[159] Barry N. Taylor, Donald N. Langenberg, and William H. Parker. The fundamental physical constants. Scientific American, 223(4):62–70, 76–78, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n4/pdf/scientificamerican1070-62.pdf.

Young:1970:VC

[160] Richard W. Young. Visual cells. Scientific American, 223(4):80–91, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n4/pdf/scientificamerican1070-80.pdf.

Bormann:1970:NCE

[161] F. Herbert Bormann and Gene E. Likens. The nutrient cycles of an ecosystem. Scientific American, 223(4):92–101, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n4/pdf/scientificamerican1070-92.pdf.

Berenyi:1970:CEE

[162] Ivan Berenyi. Computers in Eastern Europe. Scientific American, 223(4):102–109, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n4/pdf/scientificamerican1070-
102.pdf.

Mayr:1970:OFC

[163] Otto Mayr. The origins of feedback control. Scientific American, 223(4):110–118, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n4/pdf/scientificamerican1070-
110.pdf.

Gardner:1970:MGj

[164] Martin Gardner. Mathematical games. Scientific American, 223 (4):120–123, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n4/pdf/scientificamerican1070-
120.pdf.

Gardner:1970:MGFa

[165] M. Gardner. Mathematical games: the fantastic combinations of John Conway’s new solitaire game ‘Life’. Scientific American, 223(4):120–123, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). The original description of Conway’s game of LIFE. See [2012].

Stong:1970:ASj

[166] C. L. Stong. The amateur scientist. Scientific American, 223(4):124–129, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n4/pdf/scientificamerican1070-124.pdf.

Mirsky:1970:B

[167] Alfred E. Mirsky. Books. Scientific American, 223(4):135–143, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n4/pdf/scientificamerican1070-135.pdf.

Anonymous:1970:Bj

[168] Anonymous. Bibliography. Scientific American, 223(4):144, October 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n4/pdf/scientificamerican1070-144.pdf.

Anonymous:1970:Lk

[169] Anonymous. Letters. Scientific American, 223(5):6, November 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n5/pdf/scientificamerican1170-6.pdf.

Anonymous:1970:YAk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n5/pdf/scientificamerican1170-8.pdf.

Anonymous:1970:Ak

[171] Anonymous. The authors. Scientific American, 223(5):10–11, November 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n5/pdf/scientificamerican1170-10.pdf.

Seaborg:1970:FBR

[172] Glenn T. Seaborg and Justin L. Bloom. Fast breeder reactors. Scientific American, 223(5):13–21, November 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n5/pdf/scientificamerican1170-13.pdf.

Goodenough:1970:GAM

[173] Ursula W. Goodenough and R. P. Levine. The genetic activity of mitochondria and chloroplasts. Scientific American, 223(5):22–29, November 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n5/pdf/scientificamerican1170-22.pdf.

Wainwright:1970:W

[174] Geoffrey Wainwright. Woodhenges. Scientific American, 223(5):30–41, November 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n5/pdf/scientificamerican1170-30.pdf.

Anonymous:1970:SCk

[175] Anonymous. Science and the citizen. Scientific American, 223(5):42–56, November 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n5/pdf/scientificamerican1170-42.pdf.

Derjaguin:1970:SW

[176] Boris V. Derjaguin. Superdense water. Scientific American, 223 (5):52–64, 69–71, November 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n5/pdf/scientificamerican1170-
52.pdf.

Ingram:1970:AAB

[177] Marylou Ingram and Kendall Preston, Jr. Automatic analysis of blood cells. Scientific American, 223(5):72–82, November 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n5/pdf/scientificamerican1170-72.pdf.

Tickell:1970:GA

[178] W. L. N. Tickell. The great albatrosses. Scientific American, 223(5):84–93, November 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n5/pdf/scientificamerican1170-
84.pdf.

Tajfel:1970:EID

[179] Henri Tajfel. Experiments in intergroup discrimination. Scientific American, 223(5):96–102, November 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n5/pdf/scientificamerican1170-
96.pdf.

MacIntyre:1970:WSS

[180] F. MacIntyre. Why the sea is salt. Scientific American, 223(5):104–115, November 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Maclntyre:1970:WSS

[181] Ferren Maclntyre. Why the sea is salt. Scientific American, 223 (5):104–115, November 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n5/pdf/scientificamerican1170-
104.pdf.

Gardner:1970:MGk

[182] Martin Gardner. Mathematical games. Scientific American, 223 (5):116–118, November 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n5/pdf/scientificamerican1170-
116.pdf.

Stong:1970:ASk

[183] C. L. Stong. The amateur scientist. Scientific American, 223(5):120–125, November 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n5/pdf/scientificamerican1170-120.pdf.

Morrison:1970:Bi

[184] Philip Morrison. Books. Scientific American, 223(5):126–131, November 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n5/pdf/scientificamerican1170-126.pdf.

Anonymous:1970:Bk

[185] Anonymous. Bibliography. Scientific American, 223(5):132, November 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n5/pdf/scientificamerican1170-132.pdf.

Anonymous:1970:Ll

[186] Anonymous. Letters. Scientific American, 223(6):6, December 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n6/pdf/scientificamerican1270-6.pdf.

Anonymous:1970:YAl

[187] Anonymous. 50 and 100 years ago. Scientific American, 223(6):8, December 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n6/pdf/scientificamerican1270-8.pdf.

Anonymous:1970:Al

[188] Anonymous. The authors. Scientific American, 223(6):10, December 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n6/pdf/scientificamerican1270-10.pdf.

Pinchot:1970:MF

[189] Gifford B. Pinchot. Marine farming. Scientific American, 223(6):14–21, December 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n6/pdf/scientificamerican1270-14.pdf.

Burbidge:1970:ALQ

[190] E. Margaret Burbidge and C. Roger Lynds. The absorption lines of quasi-stellar objects. Scientific American, 223(6):22–29, December 1970.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n6/pdf/scientificamerican1270-22.pdf.

Warren:1970:AIC

[191] Richard M. Warren and Roslyn P. Warren. Auditory illusions and confusions. Scientific American, 223(6):30–37, December 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n6/pdf/scientificamerican1270-30.pdf.

Anonymous:1970:SCl

[192] Anonymous. Science and the citizen. Scientific American, 223(6):38–41, December 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n6/pdf/scientificamerican1270-38.pdf.

Edwards:1970:HEL

[193] R. G. Edwards and Ruth E. Fowler. Human embryos in the laboratory. Scientific American, 223(6):44–54, December 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n6/pdf/scientificamerican1270-44.pdf.

Baumeister:1970:OIC

[194] Philip Baumeister and Gerald Pincus. Optical interference coatings. Scientific American, 223(6):58–75, December 1970. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v223/
n6/pdf/scientificamerican1270-58.pdf.

Loomis:1970:R

[195] W. F. Loomis. Rickets. Scientific American, 223(6):76–91, December 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n6/pdf/scientificamerican1270-76.pdf.

Becker:1970:PM

[196] Joseph J. Becker. Permanent magnets. Scientific American, 223 (6):92–100, December 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n6/pdf/scientificamerican1270-
92.pdf.

Savory:1970:M

[197] Thcodore H. Savory. The mule. Scientific American, 223(6):102–109, December 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n6/pdf/scientificamerican1270-102.pdf.

Gardner:1970:MGl

[198] Martin Gardner. Mathematical games. Scientific American, 223 (6):110–115, December 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v223/n6/pdf/scientificamerican1270-
110.pdf.

Stong:1970:ASl

[199] C. L. Stong. The amateur scientist. Scientific American, 223(6):116–121, December 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n6/pdf/scientificamerican1270-116.pdf.

Philip:1970:B

[200] Philip and Phylis Morrison. Books. Scientific American, 223(6):122–135, December 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n6/pdf/scientificamerican1270-122.pdf.

Anonymous:1970:AIa

[201] Anonymous. Annual index. Scientific American, 223(6):136–139, December 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n6/pdf/scientificamerican1270-136.pdf.

Anonymous:1970:Bl

[202] Anonymous. Bibliography. Scientific American, 223(6):140, December 1970. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v223/n6/pdf/scientificamerican1270-140.pdf.

Anonymous:1971:La

[203] Anonymous. Letters. Scientific American, 224(1):6–8, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n1/pdf/scientificamerican0171-6.pdf.

Anonymous:1971:YAa

[204] Anonymous. 50 and 100 years ago. Scientific American, 224(1):11, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n1/pdf/scientificamerican0171-11.pdf.

Anonymous:1971:Aa

[205] Anonymous. The authors. Scientific American, 224(1):12–13, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n1/pdf/scientificamerican0171-12.pdf.

Scoville:1971:LOW

[206] Herbert Scoville, Jr. The limitation of offensive weapons. Scientific American, 224(1):15–25, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n1/pdf/scientificamerican0171-
15.pdf.

Levine:1971:SB

[207] Seymour Levine. Stress and behavior. Scientific American, 224(1):26–31, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n1/pdf/scientificamerican0171-26.pdf.

Newell:1971:GCA

[208] Reginald E. Newell. The global circulation of atmospheric pollutants. Scientific American, 224(1):32–42, January 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n1/pdf/scientificamerican0171-32.pdf.

Anonymous:1971:SCa

[209] Anonymous. Science and the citizen. Scientific American, 224(1):44–47, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n1/pdf/scientificamerican0171-44.pdf.

Ostriker:1971:NP

[210] Jeremiah P. Ostriker. The nature of pulsars. Scientific American, 224(1):48–63, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n1/pdf/scientificamerican0171-
48.pdf.

Denton:1971:RF

[211] Eric Denton. Reflectors in fishes. Scientific American, 224(1):64–72, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n1/pdf/scientificamerican0171-64.pdf.

Rieder:1971:CB

[212] Werner Rieder. Circuit breakers. Scientific American, 224(1):76–85, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n1/pdf/scientificamerican0171-76.pdf.

Sigurbjornsson:1971:IMP

[213] Björn Sigurbjörnsson. Induced mutations in plants. Scientific American, 224(1):86–95, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n1/pdf/scientificamerican0171-
86.pdf.

Howard-Jones:1971:OHM

[214] Norman Howard-Jones. The origins of hypodermic medication. Scientific American, 224(1):96–102, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n1/pdf/scientificamerican0171-
96.pdf.

Gardner:1971:MGa

[215] Martin Gardner. Mathematical games. Scientific American, 224 (1):104–108, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n1/pdf/scientificamerican0171-
104.pdf.

Stong:1971:ASa

[216] C. L. Stong. The amateur scientist. Scientific American, 224(1):110–115, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n1/pdf/scientificamerican0171-110.pdf.

Morrison:1971:Ba

[217] Philip Morrison. Books. Scientific American, 224(1):117–121, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n1/pdf/scientificamerican0171-117.pdf.

Anonymous:1971:Ba

[218] Anonymous. Bibliography. Scientific American, 224(1):122, January 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n1/pdf/scientificamerican0171-122.pdf.

Anonymous:1971:Lb

[219] Anonymous. Letters. Scientific American, 224(2):6, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n2/pdf/scientificamerican0271-6.pdf.

Anonymous:1971:YAb

[220] Anonymous. 50 and 100 years ago. Scientific American, 224(2):10, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n2/pdf/scientificamerican0271-10.pdf.

Anonymous:1971:Ab

[221] Anonymous. The authors. Scientific American, 224(2):12, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n2/pdf/scientificamerican0271-12.pdf.

Chisolm:1971:LP

[222] J. Julian Chisolm, Jr. Lead poisoning. Scientific American, 224(2):15–23, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n2/pdf/scientificamerican0271-15.pdf.

Ruderman:1971:SS

[223] Malvin A. Ruderman. Solid stars. Scientific American, 224(2):24–31, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n2/pdf/scientificamerican0271-24.pdf.

Tuck:1971:IC

[224] James A. Tuck. The Iroquois Confederacy. Scientific American, 224(2):32–42, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n2/pdf/scientificamerican0271-
32.pdf.

Anonymous:1971:SCb

[225] Anonymous. Science and the citizen. Scientific American, 224(2):44–47, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n2/pdf/scientificamerican0271-44.pdf.

Gough:1971:PFP

[226] William C. Gough and Bernard J. Eastlund. The prospects of fusion power. Scientific American, 224(2):50–64, February 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n2/pdf/scientificamerican0271-50.pdf.

Willows:1971:GBC

[227] A. O. D. Willows. Giant brain cells in mollusks. Scientific American, 224(2):68–75, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n2/pdf/scientificamerican0271-
68.pdf.

Slotnick:1971:FC

[228] D. L. Slotnick. The fastest computer. Scientific American, 224(2):76–87, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n2/pdf/scientificamerican0271-76.pdf.

Solomon:1971:SWR

[229] Arthur K. Solomon. The state of water in red cells. Scientific American, 224(2):88–96, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n2/pdf/scientificamerican0271-
88.pdf.

Reti:1971:LBG

[230] Ladislao Reti. Leonardo on bearings and gears. Scientific American, 224(2):100–110, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n2/pdf/scientificamerican0271-
100.pdf.

Cardner:1971:MG

[231] Martin Cardner. Mathematical games. Scientific American, 224 (2):112–117, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n2/pdf/scientificamerican0271-
112.pdf.

Gardner:1971:MGO

[232] M. Gardner. Mathematical games: On cellular automata, selfreproduction, the Garden of Eden and the game of ‘Life’. Scientific American, 224(2):112–117, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). Follow-up article on LIFE and other CA rules.

Stong:1971:ASb

[233] C. L. Stong. The amateur scientist. Scientific American, 224(2):118–122, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n2/pdf/scientificamerican0271-118.pdf.

Morrison:1971:Bb

[234] Philip Morrison. Books. Scientific American, 224(2):125–129, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n2/pdf/scientificamerican0271-125.pdf.

Anonymous:1971:Bb

[235] Anonymous. Bibliography. Scientific American, 224(2):130, February 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n2/pdf/scientificamerican0271-130.pdf.

Anonymous:1971:Lc

[236] Anonymous. Letters. Scientific American, 224(3):8–10, March 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n3/pdf/scientificamerican0371-8.pdf.

Anonymous:1971:YAc

[237] Anonymous. 50 and 100 years ago. Scientific American, 224(3):12, March 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n3/pdf/scientificamerican0371-12.pdf.

Anonymous:1971:Ac

[238] Anonymous. The authors. Scientific American, 224(3):14, March 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n3/pdf/scientificamerican0371-14.pdf.

Wright:1971:PCB

[239] James R. Wright. Performance criteria in building. Scientific American, 224(3):16–24, March 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n3/pdf/scientificamerican0371-
16.pdf.

Mosbach:1971:EBA

[240] Klaus Mosbach. Enzymes bound to artificial matrixes. Scientific American, 224(3):26–33, March 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n3/pdf/scientificamerican0371-
26.pdf.

Almond:1971:TC

[241] Richard Almond. The therapeutic community. Scientific American, 224(3):34–42, March 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n3/pdf/scientificamerican0371-
34.pdf.

Anonymous:1971:SCc

[242] Anonymous. Science and the citizen. Scientific American, 224(3):44–48, March 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n3/pdf/scientificamerican0371-44.pdf.

DeLong:1971:UPA

[243] Howard DeLong. Unsolved problems in arithmetic. Scientific American, 224(3):50–60, March 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n3/pdf/scientificamerican0371-
50.pdf.

Lippold:1971:PT

[244] Olof Lippold. Physiological tremor. Scientific American, 224(3):65–73, March 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n3/pdf/scientificamerican0371-65.pdf.

Essmann:1971:MSS

[245] Uwe Essmann and Hermann Träuble. The magnetic structure of superconductors. Scientific American, 224(3):74–84, March 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n3/pdf/scientificamerican0371-74.pdf.

Holldobler:1971:CBA

[246] Bert Hölldobler. Communication between ants and their guests. Scientific American, 224(3):86–93, March 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v224/n3/pdf/scientificamerican0371-
86.pdf.

Corballis:1971:TLR

[247] Michael C. Corballis and Ivan L. Beale. On telling left from right. Scientific American, 224(3):96–104, March 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n3/pdf/scientificamerican0371-96.pdf.

Gardner:1971:MGb

[248] Martin Gardner. Mathematical games. Scientific American, 224(3):106– 109, March 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n3/pdf/scientificamerican0371-106.pdf.

Stong:1971:ASc

[249] C. L. Stong. The amateur scientist. Scientific American, 224(3):110–114, March 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n3/pdf/scientificamerican0371-110.pdf.

Morrison:1971:Bc

[250] Philip Morrison. Books. Scientific American, 224(3):117–123, March 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n3/pdf/scientificamerican0371-117.pdf.

Anonymous:1971:Bc

[251] Anonymous. Bibliography. Scientific American, 224(3):124, March 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n3/pdf/scientificamerican0371-124.pdf.

Anonymous:1971:Ld

[252] Anonymous. Letters. Scientific American, 224(4):8–9, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n4/pdf/scientificamerican0471-8.pdf.

Anonymous:1971:YAd

[253] Anonymous. 50 and 100 years ago. Scientific American, 224(4):10, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n4/pdf/scientificamerican0471-10.pdf.

Anonymous:1971:Ad

[254] Anonymous. The authors. Scientific American, 224(4):14, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n4/pdf/scientificamerican0471-14.pdf.

Lewis:1971:GIH

[255] Irving J. Lewis. Government investment in health care. Scientific American, 224(4):17–25, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n4/pdf/scientificamerican0471-
17.pdf.

Crewe:1971:HRS

[256] Albert V. Crewe. A high-resolution scanning electron microscope. Scientific American, 224(4):26–35, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n4/pdf/scientificamerican0471-
26.pdf.

Crewe:1971:HSE

[257] A. V. Crewe. A high-resolution scanning electron microscope. Scientific American, 224(4):26–35, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

MacNeish:1971:EMA

[258] Richard S. MacNeish. Early Man in the Andes. Scientific American, 224(4):36–46, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n4/pdf/scientificamerican0471-
36.pdf.

Anonymous:1971:SCd

[259] Anonymous. Science and the citizen. Scientific American, 224(4):48–53, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n4/pdf/scientificamerican0471-48.pdf.

Casey:1971:APR

[260] Richard G. Casey and George Nagy. Advances in pattern recognition. Scientific American, 224(4):56–64, 70–71, April 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n4/pdf/scientificamerican0471-56.pdf.

Pengelley:1971:ABC

[261] Eric T. Pengelley and Sally J. Asmundson. Annual biological clocks. Scientific American, 224(4):72–79, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v224/n4/pdf/scientificamerican0471-
72.pdf.

Brandt:1971:SAH

[262] N. B. Brandt and N. I. Ginzburg. Superconductivity at high pressure. Scientific American, 224(4):83–88, 90, 92, 94, April 1971. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Brandt:1971:SHP

[263] N. B. Brandt and N. I. Ginzburg. Superconductivity at high pressure. Scientific American, 224(4):83–94, April 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n4/pdf/scientificamerican0471-83.pdf.

Knight:1971:H

[264] Charles Knight and Nancy Knight. Hailstones. Scientific American, 224(4):96–103, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n4/pdf/scientificamerican0471-
96.pdf.

McKusick:1971:MHC

[265] Victor A. McKusick. The mapping of human chromosomes. Scientific American, 224(4):104–113, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n4/pdf/scientificamerican0471-
104.pdf.

Gardner:1971:MGc

[266] Martin Gardner. Mathematical games. Scientific American, 224(4):114– 117, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n4/pdf/scientificamerican0471-114.pdf.

Stong:1971:ASd

[267] C. L. Stong. The amateur scientist. Scientific American, 224(4):118– 122, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n4/pdf/scientificamerican0471-118.pdf.

Morrison:1971:Bd

[268] Philip Morrison. Books. Scientific American, 224(4):125–129, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n4/pdf/scientificamerican0471-125.pdf.

Anonymous:1971:Bd

[269] Anonymous. Bibliography. Scientific American, 224(4):130, April 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n4/pdf/scientificamerican0471-130.pdf.

Anonymous:1971:Le

[270] Anonymous. Letters. Scientific American, 224(5):6–9, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n5/pdf/scientificamerican0571-6.pdf.

Anonymous:1971:YAe

[271] Anonymous. 50 and 100 years ago. Scientific American, 224(5):10, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n5/pdf/scientificamerican0571-10.pdf.

Anonymous:1971:Ae

[272] Anonymous. The authors. Scientific American, 224(5):12, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n5/pdf/scientificamerican0571-12.pdf.

Goldwater:1971:ME

[273] Leonard J. Goldwater. Mercury in the environment. Scientific American, 224(5):15–21, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n5/pdf/scientificamerican0571-
15.pdf.

Weber:1971:DGW

[274] Joseph Weber. The detection of gravitational waves. Scientific American, 224(5):22–29, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n5/pdf/scientificamerican0571-
22.pdf.

Barghoorn:1971:OF

[275] Elso S. Barghoorn. The oldest fossils. Scientific American, 224(5):30– 42, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n5/pdf/scientificamerican0571-30.pdf.

Anonymous:1971:SCe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n5/pdf/scientificamerican0571-44.pdf.

Schmidt:1971:EQ

[277] Maarten Schmidt and Francis Bello. The evolution of quasars. Scientific American, 224(5):54–69, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n5/pdf/scientificamerican0571-
54.pdf.

Woodson:1971:CT

[278] Riley D. Woodson. Cooling towers. Scientific American, 224(5):70– 78, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n5/pdf/scientificamerican0571-70.pdf.

Shiers:1971:IC

[279] George Shiers. The induction coil. Scientific American, 224(5):80–87, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n5/pdf/scientificamerican0571-80.pdf.

Richards:1971:FIM

[280] Whitman Richards. The fortification illusions of migraines. Scientific American, 224(5):88–96, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n5/pdf/scientificamerican0571-
88.pdf.

Todd:1971:CLF

[281] John H. Todd. The chemical languages of fishes. Scientific American, 224(5):98–108, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n5/pdf/scientificamerican0571-
98.pdf.

Gardner:1971:MGd

[282] Martin Gardner. Mathematical games. Scientific American, 224(5):110– 116, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n5/pdf/scientificamerican0571-110.pdf.

Stong:1971:ASe

[283] C. L. Stong. The amateur scientist. Scientific American, 224(5):118– 124, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n5/pdf/scientificamerican0571-118.pdf.

Morrison:1971:Be

[284] Philip Morrison. Books. Scientific American, 224(5):127–131, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n5/pdf/scientificamerican0571-127.pdf.

Anonymous:1971:Be

[285] Anonymous. Bibliography. Scientific American, 224(5):132, May 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n5/pdf/scientificamerican0571-132.pdf.

Anonymous:1971:Lf

[286] Anonymous. Letters. Scientific American, 224(6):9, June 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n6/pdf/scientificamerican0671-9.pdf.

Anonymous:1971:YAf

[287] Anonymous. 50 and 100 years ago. Scientific American, 224(6):12, June 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n6/pdf/scientificamerican0671-12.pdf.

Anonymous:1971:Af

[288] Anonymous. The authors. Scientific American, 224(6):16–17, June 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n6/pdf/scientificamerican0671-16.pdf.

Lubin:1971:FL

[289] Moshe J. Lubin and Arthur P. Fraas. Fusion by laser. Scientific American, 224(6):21–33, June 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n6/pdf/scientificamerican0671-
21.pdf.

Noton:1971:EMV

[290] David Noton and Lawrence Stark. Eye movements and visual perception. Scientific American, 224(6):34–43, June 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n6/pdf/scientificamerican0671-34.pdf.

Ross:1971:EFB

[291] Russell Ross and Paul Bornstein. Elastic fibers in the body. Scientific American, 224(6):44–52, June 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n6/pdf/scientificamerican0671-
44.pdf.

Anonymous:1971:SCf

[292] Anonymous. Science and the citizen. Scientific American, 224(6):54–59, June 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n6/pdf/scientificamerican0671-54.pdf.

Kendall:1971:SPN

[293] Henry W. Kendall and Wolfgang K. H. Panofsky. The structure of the proton and the neutron. Scientific American, 224(6):60–77, June 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n6/pdf/scientificamerican0671-60.pdf.

Bobeck:1971:MB

[294] Andrew H. Bobeck and H. E. D. Scovil. Magnetic bubbles. Scientific American, 224(6):78–90, June 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n6/pdf/scientificamerican0671-
78.pdf.

Gillie:1971:EG

[295] R. Bruce Gillie. Endemic goiter. Scientific American, 224(6):92–101, June 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n6/pdf/scientificamerican0671-92.pdf.

C:1971:ECI

[296] C. C. and Martha Lamberg-Karlovsky. An early city in Iran. Scientific American, 224(6):102–111, June 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v224/n6/pdf/scientificamerican0671-
102.pdf.

Watts:1971:SOT

[297] C. Robert Watts and Allen W. Stokes. The social order of turkeys. Scientific American, 224(6):112–118, June 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n6/pdf/scientificamerican0671-112.pdf.

Gardner:1971:MGe

[298] Martin Gardner. Mathematical games. Scientific American, 224(6):120– 123, June 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n6/pdf/scientificamerican0671-120.pdf.

Stong:1971:ASf

[299] C. L. Stong. The amateur scientist. Scientific American, 224(6):124– 128, June 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v224/n6/pdf/scientificamerican0671-124.pdf.

Morrison:1971:Bf

[300] Philip Morrison. Books. Scientific American, 224(6):131–135, June 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n6/pdf/scientificamerican0671-131.pdf.

Anonymous:1971:Bf

[301] Anonymous. Bibliography. Scientific American, 224(6):136, June 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v224/
n6/pdf/scientificamerican0671-136.pdf.

Anonymous:1971:Lg

[302] Anonymous. Letters. Scientific American, 225(1):6–7, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n1/pdf/scientificamerican0771-6.pdf.

Anonymous:1971:YAg

[303] Anonymous. 50 and 100 years ago. Scientific American, 225(1):10– 11, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n1/pdf/scientificamerican0771-10.pdf.

Anonymous:1971:Ag

[304] Anonymous. The authors. Scientific American, 225(1):14, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n1/pdf/scientificamerican0771-14.pdf.

Hauser:1971:C

[305] Philip M. Hauser. The census of 1970. Scientific American, 225(1):17– 25, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n1/pdf/scientificamerican0771-17.pdf.

Hilleman:1971:II

[306] Maurice R. Hilleman and Alfred A. Tytell. The induction of Interferon. Scientific American, 225(1):26–31, July 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n1/pdf/scientificamerican0771-26.pdf.

Panish:1971:NCD

[307] Morton B. Panish and Izuo Hayashi. A new class of diode lasers. Scientific American, 225(1):32–40, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n1/pdf/scientificamerican0771-
32.pdf.

Anonymous:1971:SCg

[308] Anonymous. Science and the citizen. Scientific American, 225(1):42– 45, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n1/pdf/scientificamerican0771-42.pdf.

Heimer:1971:PB

[309] Lennart Heimer. Pathways in the brain. Scientific American, 225(1):48– 60, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n1/pdf/scientificamerican0771-48.pdf.

Deyts:1971:SSS

[310] Simone-Antoinette Deyts. The sacred source of the Seine. Scientific American, 225(1):65–73, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n1/pdf/scientificamerican0771-
65.pdf.

Gorenstein:1971:SR

[311] Paul Gorenstein and Wallace Tucker. Supernova remnants. Scientific American, 225(1):74–85, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n1/pdf/scientificamerican0771-
74.pdf.

Bell:1971:GES

[312] Richard H. V. Bell. A grazing ecosystem in the Serengeti. Scientific American, 225(1):86–93, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n1/pdf/scientificamerican0771-
86.pdf.

Murphy:1971:PH

[313] Frederick V. Murphy and David E. Yount. Photons as hadrons. Scientific American, 225(1):94–104, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n1/pdf/scientificamerican0771-
94.pdf.

Gardner:1971:MGf

[314] Martin Gardner. Mathematical games. Scientific American, 225(1):106– 109, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n1/pdf/scientificamerican0771-106.pdf.

Stong:1971:ASg

[315] C. L. Stong. The amateur scientist. Scientific American, 225(1):110– 114, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n1/pdf/scientificamerican0771-110.pdf.

Morrison:1971:Bg

[316] Philip Morrison. Books. Scientific American, 225(1):117–121, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n1/pdf/scientificamerican0771-117.pdf.

Anonymous:1971:Bg

[317] Anonymous. Bibliography. Scientific American, 225(1):122, July 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n1/pdf/scientificamerican0771-122.pdf.

Anonymous:1971:Lh

[318] Anonymous. Letters. Scientific American, 225(2):6–7, August 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n2/pdf/scientificamerican0871-6.pdf.

Anonymous:1971:YAh

[319] Anonymous. 50 and 100 years ago. Scientific American, 225(2):8, August 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n2/pdf/scientificamerican0871-8.pdf.

Anonymous:1971:Ah

[320] Anonymous. The authors. Scientific American, 225(2):13, August 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n2/pdf/scientificamerican0871-13.pdf.

Hirschhorn:1971:C

[321] Norbert Hirschhorn and William B. Greenough III. Cholera. Scientific American, 225(2):15–21, August 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n2/pdf/scientificamerican0871-
15.pdf.

Bass:1971:BTV

[322] George F. Bass. A Byzantine trading venture. Scientific American, 225(2):22–33, August 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n2/pdf/scientificamerican0871-
22.pdf.

Harpstead:1971:HLC

[323] Dale D. Harpstead. High-lysine corn. Scientific American, 225(2):34–42, August 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n2/pdf/scientificamerican0871-34.pdf.

Anonymous:1971:SCh

[324] Anonymous. Science and the citizen. Scientific American, 225(2):44–47, August 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n2/pdf/scientificamerican0871-44.pdf.

Margulis:1971:SE

[325] Lynn Margulis. Symbiosis and evolution. Scientific American, 225(2):48– 57, August 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n2/pdf/scientificamerican0871-48.pdf.

Dyal:1971:MM

[326] Palmer Dyal and Curtis W. Parkin. The magnetism of the Moon. Scientific American, 225(2):62–73, August 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n2/pdf/scientificamerican0871-62.pdf.

Camhi:1971:FOL

[327] Jeffrey M. Camhi. Flight orientation in locusts. Scientific American, 225(2):74–81, August 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n2/pdf/scientificamerican0871-
74.pdf.

Atkinson:1971:CST

[328] Richard C. Atkinson and Richard M. Shiffrin. The control of shortterm memory. Scientific American, 225(2):82–90, August 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n2/pdf/scientificamerican0871-82.pdf.

Steen:1971:NMR

[329] Lynn Arthur Steen. New models of the real-number line. Scientific American, 225(2):92–99, August 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n2/pdf/scientificamerican0871-
92.pdf.

Gardner:1971:MGg

[330] Martin Gardner. Mathematical games. Scientific American, 225 (2):102–105, August 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n2/pdf/scientificamerican0871-
102.pdf.

Stong:1971:ASh

[331] C. L. Stong. The amateur scientist. Scientific American, 225(2):106–115, August 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n2/pdf/scientificamerican0871-106.pdf.

Morrison:1971:Bh

[332] Philip Morrison. Books. Scientific American, 225(2):116–119, August 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n2/pdf/scientificamerican0871-116.pdf.

Anonymous:1971:Bh

[333] Anonymous. Bibliography. Scientific American, 225(2):120, August 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n2/pdf/scientificamerican0871-120.pdf.

Anonymous:1971:Li

[334] Anonymous. Letters. Scientific American, 225(3):9–14, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n3/pdf/scientificamerican0971-9.pdf.

Anonymous:1971:YAi

[335] Anonymous. 50 and 100 years ago. Scientific American, 225(3):18–20, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n3/pdf/scientificamerican0971-18.pdf.

Anonymous:1971:Ai

[336] Anonymous. The authors. Scientific American, 225(3):25–32, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n3/pdf/scientificamerican0971-25.pdf.

Starr:1971:EP

[337] Chauncey Starr. Energy and pernor. Scientific American, 225(3):36–49, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n3/pdf/scientificamerican0971-36.pdf.

Dyson:1971:EU

[338] Freeman J. Dyson. Energy in the universe. Scientific American, 225(3):50–59, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n3/pdf/scientificamerican0971-
50.pdf.

Hubbert:1971:ERE

[339] M. King Hubbert. The energy resources of the Earth. Scientific American, 225(3):60–70, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n3/pdf/scientificamerican0971-
60.pdf.

Anonymous:1971:SCi

[340] Anonymous. Science and the citizen. Scientific American, 225(3):74–84, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n3/pdf/scientificamerican0971-74.pdf.

Gates:1971:FEB

[341] David M. Gates. The flow of energy in the Biosphere. Scientific American, 225(3):88–100, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n3/pdf/scientificamerican0971-
88.pdf.

Kemp:1971:FEH

[342] William B. Kemp. The flow of energy in a hunting society. Scientific American, 225(3):104–115, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v225/n3/pdf/scientificamerican0971-
104.pdf.

Rappaport:1971:FEA

[343] Roy A. Rappaport. The flew of energy in an agricultural society. Scientific American, 225(3):116–132, September 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n3/pdf/scientificamerican0971-116.pdf.

Cook:1971:FEI

[344] Earl Cook. The flow of energy in an industrial society. Scientific American, 225(3):134–144, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n3/pdf/scientificamerican0971-
134.pdf.

Summers:1971:CE

[345] Claude M. Summers. The conversion of energy. Scientific American, 225(3):148–160, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n3/pdf/scientificamerican0971-
148.pdf.

Luten:1971:EGE

[346] Daniel B. Luten. The economic geography of energy. Scientific American, 225(3):164–175, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n3/pdf/scientificamerican0971-
164.pdf.

Tribus:1971:EI

[347] Myron Tribus and Edward C. McIrvine. Energy and information. Scientific American, 225(3):179–188, September 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n3/pdf/scientificamerican0971-179.pdf.

Tribus:1971:EIT

[348] M. Tribus and E. C. McIrvine. Energy and information (thermodynamics and information theory). Scientific American, 225(3):179–188, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Katz:1971:DMP

[349] Milton Katz. Decision-making in the production of power. Scientific American, 225(3):191–200, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v225/n3/pdf/scientificamerican0971-
191.pdf.

Gardner:1971:MGh

[350] Martin Gardner. Mathematical games. Scientific American, 225 (3):204–212, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n3/pdf/scientificamerican0971-
204.pdf.

Stong:1971:ASi

[351] C. L. Stong. The amateur scientist. Scientific American, 225(3):218–224, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n3/pdf/scientificamerican0971-218.pdf.

Morrison:1971:Bi

[352] Philip Morrison. Books. Scientific American, 225(3):229–240, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n3/pdf/scientificamerican0971-229.pdf.

Anonymous:1971:Bi

[353] Anonymous. Bibliography. Scientific American, 225(3):244–246, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n3/pdf/scientificamerican0971-244.pdf.

Anonymous:1971:Lj

[354] Anonymous. Letters. Scientific American, 225(4):8, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n4/pdf/scientificamerican1071-8.pdf.

Anonymous:1971:YAj

[355] Anonymous. 50 and 100 years ago. Scientific American, 225(4):10, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n4/pdf/scientificamerican1071-10.pdf.

Anonymous:1971:Aj

[356] Anonymous. The authors. Scientific American, 225(4):12, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n4/pdf/scientificamerican1071-12.pdf.

Young:1971:PS

[357] Vernon R. Young and Nevin S. Scrimshaw. The physiology of starvation. Scientific American, 225(4):14–21, October 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n4/pdf/scientificamerican1071-14.pdf.

Knox:1971:CT

[358] William T. Knox. Cable television. Scientific American, 225(4):22–29, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n4/pdf/scientificamerican1071-22.pdf.

Bower:1971:OWI

[359] T. G. R. Bower. The object in the world of the infant. Scientific American, 225(4):30–38, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n4/pdf/scientificamerican1071-
30.pdf.

Anonymous:1971:SCj

[360] Anonymous. Science and the citizen. Scientific American, 225(4):40–44, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n4/pdf/scientificamerican1071-40.pdf.

Mason:1971:LR

[361] Brian Mason. The lunar rocks. Scientific American, 225(4):48–58, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n4/pdf/scientificamerican1071-48.pdf.

Renfrew:1971:CPE

[362] Colin Renfrew. Carbon 14 and the prehistory of Europe. Scientific American, 225(4):63–72, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n4/pdf/scientificamerican1071-
63.pdf.

Wessells:1971:HLC

[363] Norman K. Wessells. How living cells change shape. Scientific American, 225(4):76–82, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n4/pdf/scientificamerican1071-
76.pdf.

Herber:1971:MS

[364] R. H. Herber. Mössbauer spectroscopy. Scientific American, 225(4):86– 95, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n4/pdf/scientificamerican1071-86.pdf.

Ferguson:1971:MMD

[365] Eugene S. Ferguson. The measurement of “Man-Day”. Scientific American, 225(4):96–103, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n4/pdf/scientificamerican1071-
96.pdf.

Gardner:1971:MGi

[366] Martin Gardner. Mathematical games. Scientific American, 225 (4):104–107, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n4/pdf/scientificamerican1071-
104.pdf.

Stong:1971:ASj

[367] C. L. Stong. The amateur scientist. Scientific American, 225(4):108–112, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n4/pdf/scientificamerican1071-108.pdf.

Morrison:1971:Bj

[368] Philip Morrison. Books. Scientific American, 225(4):115–119, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n4/pdf/scientificamerican1071-115.pdf.

Anonymous:1971:Bj

[369] Anonymous. Bibliography. Scientific American, 225(4):120, October 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n4/pdf/scientificamerican1071-120.pdf.

Anonymous:1971:Lk

[370] Anonymous. Letters. Scientific American, 225(5):8, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n5/pdf/scientificamerican1171-8.pdf.

Anonymous:1971:YAk

[371] Anonymous. 50 and 100 years ago. Scientific American, 225(5):10, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n5/pdf/scientificamerican1171-10.pdf.

Anonymous:1971:Ak

[372] Anonymous. The authors. Scientific American, 225(5):12, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n5/pdf/scientificamerican1171-12.pdf.

Fisher:1971:AI

[373] W. Halder Fisher. The anatomy of inflation: 1953–1975. Scientific American, 225(5):15–21, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n5/pdf/scientificamerican1171-
15.pdf.

Geballe:1971:NS

[374] T. H. Geballe. New superconductors. Scientific American, 225(5):22–33, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n5/pdf/scientificamerican1171-22.pdf.

Friedmann:1971:PDG

[375] Theodore Friedmann. Prenatal diagnosis of genetic disease. Scientific American, 225(5):34–42, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v225/n5/pdf/scientificamerican1171-
34.pdf.

Anonymous:1971:SCk

[376] Anonymous. Science and the citizen. Scientific American, 225(5):46–50, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n5/pdf/scientificamerican1171-46.pdf.

Anderson:1971:SAF

[377] Don L. Anderson. The San Andreas Fault. Scientific American, 225(5):52–68, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n5/pdf/scientificamerican1171-
52.pdf.

Talmon:1971:NCQ

[378] Shemaryahu Talmon. The new covenanters of Qumran. Scientific American, 225(5):72–81, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n5/pdf/scientificamerican1171-
72.pdf.

Pike:1971:P

[379] John E. Pike. Prostaglandins. Scientific American, 225(5):84–92, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n5/pdf/scientificamerican1171-84.pdf.

Zobel:1971:GIS

[380] Bruce J. Zobel. The genetic improvement of Southern Pines. Scientific American, 225(5):94–103, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v225/n5/pdf/scientificamerican1171-
94.pdf.

Brown:1971:DP

[381] James H. Brown. The desert pupfish. Scientific American, 225 (5):104–110, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n5/pdf/scientificamerican1171-
104.pdf.

Gardner:1971:MGj

[382] Martin Gardner. Mathematical games. Scientific American, 225 (5):114–121, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n5/pdf/scientificamerican1171-
114.pdf.

Stong:1971:ASk

[383] C. L. Stong. The amateur scientist. Scientific American, 225(5):122–127, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n5/pdf/scientificamerican1171-122.pdf.

Morrison:1971:Bk

[384] Philip Morrison. Books. Scientific American, 225(5):129–134, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n5/pdf/scientificamerican1171-129.pdf.

Anonymous:1971:Bk

[385] Anonymous. Bibliography. Scientific American, 225(5):136–138, November 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n5/pdf/scientificamerican1171-136.pdf.

Anonymous:1971:Ll

[386] Anonymous. Letters. Scientific American, 225(6):6–7, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n6/pdf/scientificamerican1271-6.pdf.

Anonymous:1971:YAl

[387] Anonymous. 50 and 100 years ago. Scientific American, 225(6):8, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n6/pdf/scientificamerican1271-8.pdf.

Anonymous:1971:Al

[388] Anonymous. The authors. Scientific American, 225(6):10, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n6/pdf/scientificamerican1271-10.pdf.

Greeley:1971:ATR

[389] Andrew M. Greeley and Paul B. Sheatsley. Attitudes toward racial integration. Scientific American, 225(6):13–19, December 1971. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v225/
n6/pdf/scientificamerican1271-13.pdf.

Maran:1971:GN

[390] Stephen P. Maran. The Gum Nebula. Scientific American, 225(6):20–29, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n6/pdf/scientificamerican1271-20.pdf.

Crowe:1971:C

[391] John H. Crowe and Alan F. Cooper, Jr. Cryptobiosis. Scientific American, 225(6):30–36, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n6/pdf/scientificamerican1271-
30.pdf.

Anonymous:1971:SCl

[392] Anonymous. Science and the citizen. Scientific American, 225(6):38–42, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n6/pdf/scientificamerican1271-38.pdf.

Haensel:1971:C

[393] Vladimir Haensel and Robert L. Burwell, Jr. Catalysis. Scientific American, 225(6):46–58, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n6/pdf/scientificamerican1271-
46.pdf.

Attneave:1971:MP

[394] Fred Attneave. Multistability in perception. Scientific American, 225(6):62–71, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n6/pdf/scientificamerican1271-
62.pdf.

Schmidt-Nielsen:1971:HBB

[395] Knut Schmidt-Nielsen. How birds breathe. Scientific American, 225(6):72–79, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n6/pdf/scientificamerican1271-
72.pdf.

Smylie:1971:RE

[396] D. E. Smylie and L. Mansinha. The rotation of the Earth. Scientific American, 225(6):80–88, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v225/n6/pdf/scientificamerican1271-
80.pdf.

Carrington:1971:TDA

[397] John F. Carrington. The talking drums of Africa. Scientific American, 225(6):90–94, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n6/pdf/scientificamerican1271-
90.pdf.

Gardner:1971:MGk

[398] Martin Gardner. Mathematical games. Scientific American, 225 (6):96–99, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n6/pdf/scientificamerican1271-
96.pdf.

Stong:1971:ASl

[399] C. L. Stong. The amateur scientist. Scientific American, 225(6):100–104, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n6/pdf/scientificamerican1271-100.pdf.

Morrison:1971:Bl

[400] Philip Morrison and Phylis Morrison. Books. Scientific American, 225(6):106–115, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v225/n6/pdf/scientificamerican1271-
106.pdf.

Anonymous:1971:AIa

[401] Anonymous. Annual index. Scientific American, 225(6):116–118, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n6/pdf/scientificamerican1271-116.pdf.

Anonymous:1971:Bl

[402] Anonymous. Bibliography. Scientific American, 225(6):120, December 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v225/n6/pdf/scientificamerican1271-120.pdf.

Anonymous:1972:La

[403] Anonymous. Letters. Scientific American, 226(1):6, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n1/pdf/scientificamerican0172-6.pdf.

Anonymous:1972:YAa

[404] Anonymous. 50 and 100 years ago. Scientific American, 226(1):8–9, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n1/pdf/scientificamerican0172-8.pdf.

Anonymous:1972:Aa

[405] Anonymous. The authors. Scientific American, 226(1):10–11, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n1/pdf/scientificamerican0172-10.pdf.

Myers:1972:ENT

[406] Henry R. Myers. Extending the nuclear-test ban. Scientific American, 226(1):13–23, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n1/pdf/scientificamerican0172-
13.pdf.

Temin:1972:RDD

[407] Howard M. Temin. RNA-directed DNA synthesis. Scientific American, 226(1):24–33, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n1/pdf/scientificamerican0172-
24.pdf.

Lipman-Blumen:1972:HIS

[408] Jean Lipman-Blumen. How ideology shapes Women’s lives. Scientific American, 226(1):34–42, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n1/pdf/scientificamerican0172-
34.pdf.

Anonymous:1972:SCa

[409] Anonymous. Science and the citizen. Scientific American, 226(1):44–52, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n1/pdf/scientificamerican0172-44.pdf.

Everhart:1972:SEM

[410] Thomas E. Everhart and Thomas L. Hayes. The scanning electron microscope. Scientific American, 226(1):54–61, 64–67, 69, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n1/pdf/scientificamerican0172-54.pdf.

Barnea:1972:GP

[411] Joseph Barnea. Geothermal power. Scientific American, 226(1):70–77, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n1/pdf/scientificamerican0172-70.pdf.

Ingham:1972:SA

[412] M. F. Ingham. The spectrum of the airglow. Scientific American, 226(1):78–85, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n1/pdf/scientificamerican0172-
78.pdf.

Davenport:1972:WSD

[413] Horace W. Davenport. Why the stomach does not digest itself. Scientific American, 226(1):86–93, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n1/pdf/scientificamerican0172-
86.pdf.

Eckhardt:1972:PGH

[414] Robert B. Eckhardt. Population genetics and human origins. Scientific American, 226(1):94–103, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n1/pdf/scientificamerican0172-
94.pdf.

Gardner:1972:MGa

[415] Martin Gardner. Mathematical games. Scientific American, 226 (1):104–107, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n1/pdf/scientificamerican0172-
104.pdf.

Stong:1972:ASa

[416] C. L. Stong. The amateur scientist. Scientific American, 226(1):108–111, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n1/pdf/scientificamerican0172-108.pdf.

Sivin:1972:B

[417] N. Sivin. Books. Scientific American, 226(1):113–121, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n1/pdf/scientificamerican0172-113.pdf.

Anonymous:1972:Ba

[418] Anonymous. Bibliography. Scientific American, 226(1):122, January 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n1/pdf/scientificamerican0172-122.pdf.

Anonymous:1972:Lb

[419] Anonymous. Letters. Scientific American, 226(2):6–7, February 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n2/pdf/scientificamerican0272-6.pdf.

Anonymous:1972:YAb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n2/pdf/scientificamerican0272-8.pdf.

Anonymous:1972:Ab

[421] Anonymous. The authors. Scientific American, 226(2):10, February 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n2/pdf/scientificamerican0272-10.pdf.

Bowers:1972:TAM

[422] Raymond Bowers and Jeffrey Frey. Technology assessment and microwave diodes. Scientific American, 226(2):13–21, February 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n2/pdf/scientificamerican0272-13.pdf.

Rosenzweig:1972:BCR

[423] Mark R. Rosenzweig, Edward L. Bennett, and Marian Cleeves Diamond. Brain changes in response to experience. Scientific American, 226(2):22–29, February 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n2/pdf/scientificamerican0272-
22.pdf.

Fox:1972:SCM

[424] C. Fred Fox. The structure of cell membranes. Scientific American, 226(2):30–38, February 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n2/pdf/scientificamerican0272-
30.pdf.

Anonymous:1972:SCb

[425] Anonymous. Science and the citizen. Scientific American, 226(2):40–44, February 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n2/pdf/scientificamerican0272-40.pdf.

Flanagan:1972:SS

[426] James L. Flanagan. The synthesis of speech. Scientific American, 226(2):48–56, 58, February 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n2/pdf/scientificamerican0272-
48.pdf.

Ashkin:1972:PLL

[427] Arthur Ashkin. The pressure of laser light. Scientific American, 226(2):62–71, February 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n2/pdf/scientificamerican0272-
62.pdf.

Kellermann:1972:IRA

[428] K. I. Kellermann. Intercontinental radio astronomy. Scientific American, 226(2):72–83, February 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n2/pdf/scientificamerican0272-
72.pdf.

Wallace:1972:PM

[429] Robert Keith Wallace and Herbert Benson. The physiology of meditation. Scientific American, 226(2):84–90, February 1972. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n2/pdf/scientificamerican0272-84.pdf.

Langer:1972:CPG

[430] William L. Langer. Checks on population growth: 1750–1850. Scientific American, 226(2):92–99, February 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n2/pdf/scientificamerican0272-
92.pdf.

Gardner:1972:MGb

[431] Martin Gardner. Mathematical games. Scientific American, 226 (2):100–104, February 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n2/pdf/scientificamerican0272-
100.pdf.

Stong:1972:ASb

[432] C. L. Stong. The amateur scientist. Scientific American, 226(2):106–111, February 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n2/pdf/scientificamerican0272-106.pdf.

Morrison:1972:Ba

[433] Philip Morrison. Books. Scientific American, 226(2):113–117, February 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n2/pdf/scientificamerican0272-113.pdf.

Anonymous:1972:Bb

[434] Anonymous. Bibliography. Scientific American, 226(2):118, February 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n2/pdf/scientificamerican0272-118.pdf.

Anonymous:1972:Lc

[435] Anonymous. Letters. Scientific American, 226(3):8, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n3/pdf/scientificamerican0372-8.pdf.

Anonymous:1972:YAc

[436] Anonymous. 50 and 100 years ago. Scientific American, 226(3):10, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n3/pdf/scientificamerican0372-10.pdf.

Anonymous:1972:Ac

[437] Anonymous. The authors. Scientific American, 226(3):12, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n3/pdf/scientificamerican0372-12.pdf.

Kermode:1972:FA

[438] G. O. Kermode. Food additives. Scientific American, 226(3):15–21, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n3/pdf/scientificamerican0372-15.pdf.

Menaker:1972:NLR

[439] Michael Menaker. Nonvisual light reception. Scientific American, 226(3):22–29, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n3/pdf/scientificamerican0372-
22.pdf.

Dietz:1972:GMC

[440] Robert S. Dietz. Geosynclines, mountains and continent-building. Scientific American, 226(3):30–38, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v226/n3/pdf/scientificamerican0372-
30.pdf.

Anonymous:1972:SCc

[441] Anonymous. Science and the citizen. Scientific American, 226(3):40–44, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n3/pdf/scientificamerican0372-40.pdf.

Moore:1972:E

[442] A. D. Moore. Electrostatics. Scientific American, 226(3):46–58, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n3/pdf/scientificamerican0372-46.pdf.

Kurten:1972:CB

[443] Björn Kurtén. The cave bear. Scientific American, 226(3):60–72, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n3/pdf/scientificamerican0372-60.pdf.

Kagan:1972:DIT

[444] Jerome Kagan. Do infants think? Scientific American, 226(3):74–82, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n3/pdf/scientificamerican0372-74.pdf.

Margaria:1972:SME

[445] Rodolfo Margaria. The sources of muscular energy. Scientific American, 226(3):84–91, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n3/pdf/scientificamerican0372-
84.pdf.

Wilson:1972:HDK

[446] Curtis Wilson. How did Kepler discover his first two laws? Scientific American, 226(3):92–106, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n3/pdf/scientificamerican0372-
92.pdf.

Gardner:1972:MGc

[447] Martin Gardner. Mathematical games. Scientific American, 226(3):108– 112, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n3/pdf/scientificamerican0372-108.pdf.

Stong:1972:ASc

[448] C. L. Stong. The amateur scientist. Scientific American, 226(3):114–119, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n3/pdf/scientificamerican0372-114.pdf.

Morrison:1972:Bb

[449] Philip Morrison. Books. Scientific American, 226(3):121–125, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n3/pdf/scientificamerican0372-121.pdf.

Anonymous:1972:Bc

[450] Anonymous. Bibliography. Scientific American, 226(3):126, March 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n3/pdf/scientificamerican0372-126.pdf.

Anonymous:1972:Ld

[451] Anonymous. Letters. Scientific American, 226(4):6, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n4/pdf/scientificamerican0472-6.pdf.

Anonymous:1972:YAd

[452] Anonymous. 50 and 100 years ago. Scientific American, 226(4):8, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n4/pdf/scientificamerican0472-8.pdf.

Anonymous:1972:Ad

[453] Anonymous. The authors. Scientific American, 226(4):10, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n4/pdf/scientificamerican0472-10.pdf.

Myrdal:1972:PFE

[454] Gunnar Myrdal. Political factors in economic assistance. Scientific American, 226(4):15–21, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n4/pdf/scientificamerican0472-
15.pdf.

Keefe:1972:CA

[455] D. Keefe. Collective-effect accelerators. Scientific American, 226(4):22– 33, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Keefe:1972:CEA

[456] Denis Keefe. Collective-effect accelerators. Scientific American, 226(4):22–33, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n4/pdf/scientificamerican0472-
22.pdf.

Solheim:1972:EAR

[457] Wilhelm G. Solheim II. An earlier agricultural revolution. Scientific American, 226(4):34–41, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n4/pdf/scientificamerican0472-
34.pdf.

Goldreich:1972:TEM

[458] Peter Goldreich. Tides and the Earth-Moon system. Scientific American, 226(4):42–52, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n4/pdf/scientificamerican0472-
42.pdf.

Anonymous:1972:SCd

[459] Anonymous. Science and the citizen. Scientific American, 226(4):54–57, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n4/pdf/scientificamerican0472-54.pdf.

Dickerson:1972:SHA

[460] Richard E. Dickerson. The structure and history of an ancient protein. Scientific American, 226(4):58–72, April 1972. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n4/pdf/scientificamerican0472-58.pdf.

Geschwind:1972:LB

[461] Norman Geschwind. Language and the brain. Scientific American, 226(4):76–83, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n4/pdf/scientificamerican0472-
76.pdf.

Snowden:1972:SPT

[462] Donald P. Snowden. Superconductors for power transmission. Scientific American, 226(4):84–91, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n4/pdf/scientificamerican0472-
84.pdf.

Morse:1972:ECB

[463] Roger A. Morse. Environmental control in the beehive. Scientific American, 226(4):92–98, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n4/pdf/scientificamerican0472-
92.pdf.

Gardner:1972:MGd

[464] Martin Gardner. Mathematical games. Scientific American, 226(4):100– 104, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n4/pdf/scientificamerican0472-100.pdf.

Stong:1972:ASd

[465] C. L. Stong. The amateur scientist. Scientific American, 226(4):106– 111, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n4/pdf/scientificamerican0472-106.pdf.

Morrison:1972:Bc

[466] Philip Morrison. Books. Scientific American, 226(4):113–117, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n4/pdf/scientificamerican0472-113.pdf.

Anonymous:1972:Bd

[467] Anonymous. Bibliography. Scientific American, 226(4):118, April 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n4/pdf/scientificamerican0472-118.pdf.

Anonymous:1972:Le

[468] Anonymous. Letters. Scientific American, 226(5):8–10, May 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n5/pdf/scientificamerican0572-8.pdf.

Anonymous:1972:YAe

[469] Anonymous. 50 and 100 years ago. Scientific American, 226(5):12, May 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n5/pdf/scientificamerican0572-12.pdf.

Anonymous:1972:Ae

[470] Anonymous. The authors. Scientific American, 226(5):16, May 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n5/pdf/scientificamerican0572-16.pdf.

Westing:1972:CI

[471] Arthur H. Westing and E. W. Pfeiffer. The cratering of Indochina. Scientific American, 226(5):20–29, May 1972. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n5/pdf/scientificamerican0572-20.pdf.

Merton:1972:HWC

[472] P. A. Merton. How we control the contraction of our muscles. Scientific American, 226(5):30–37, May 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n5/pdf/scientificamerican0572-
30.pdf.

Penrose:1972:BH

[473] Roger Penrose. Black holes. Scientific American, 226(5):38–46, May 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n5/pdf/scientificamerican0572-38.pdf.

Anonymous:1972:SCe

[474] Anonymous. Science and the citizen. Scientific American, 226(5):48– 54, May 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n5/pdf/scientificamerican0572-48.pdf.

Dewey:1972:PT

[475] John F. Dewey. Plate tectonics. Scientific American, 226(5):56–68, May 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n5/pdf/scientificamerican0572-56.pdf.

Dudrick:1972:TIF

[476] Stanley J. Dudrick and Jonathan E. Rhoads. Total intravenous feeding. Scientific American, 226(5):73–80, May 1972. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n5/pdf/scientificamerican0572-73.pdf.

Hammond:1972:PMC

[477] Norman Hammond. The planning of a Maya ceremonial center. Scientific American, 226(5):82–91, May 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n5/pdf/scientificamerican0572-
82.pdf.

Fritts:1972:TRC

[478] Harold C. Fritts. Tree rings and climate. Scientific American, 226(5):92– 100, May 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n5/pdf/scientificamerican0572-92.pdf.

Ickx:1972:GAR

[479] Jacques Ickx. The great automobile race of 1895. Scientific American, 226(5):102–111, May 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n5/pdf/scientificamerican0572-
102.pdf.

Gardner:1972:MGe

[480] Martin Gardner. Mathematical games. Scientific American, 226(5):112– 117, May 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n5/pdf/scientificamerican0572-112.pdf.

Stong:1972:ASe

[481] C. L. Stong. The amateur scientist. Scientific American, 226(5):118– 126, May 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n5/pdf/scientificamerican0572-118.pdf.

Morrison:1972:Bd

[482] Philip Morrison. Books. Scientific American, 226(5):128–137, May 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n5/pdf/scientificamerican0572-128.pdf.

Anonymous:1972:Be

[483] Anonymous. Bibliography. Scientific American, 226(5):138–140, May 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n5/pdf/scientificamerican0572-138.pdf.

Anonymous:1972:Lf

[484] Anonymous. Letters. Scientific American, 226(6):6–7, June 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n6/pdf/scientificamerican0672-6.pdf.

Anonymous:1972:YAf

[485] Anonymous. 50 and 100 years ago. Scientific American, 226(6):8, June 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n6/pdf/scientificamerican0672-8.pdf.

Anonymous:1972:Af

[486] Anonymous. The authors. Scientific American, 226(6):12, June 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n6/pdf/scientificamerican0672-12.pdf.

Scoville:1972:MSN

[487] Herbert Scoville, Jr. Missile submarines and national security. Scientific American, 226(6):15–27, June 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n6/pdf/scientificamerican0672-
15.pdf.

Reisfeld:1972:MBI

[488] Ralph A. Reisfeld and Barry D. Kahan. Markers of biological individuality. Scientific American, 226(6):28–37, June 1972. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n6/pdf/scientificamerican0672-28.pdf.

Lawless:1972:OMM

[489] James G. Lawless, Clair E. Folsome, and Keith A. Kvenvolden. Organic matter in meteorites. Scientific American, 226(6):38–46, June 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n6/pdf/scientificamerican0672-38.pdf.

Anonymous:1972:SCf

[490] Anonymous. Science and the citizen. Scientific American, 226(6):51–53, June 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n6/pdf/scientificamerican0672-51.pdf.

Newell:1972:ER

[491] Norman D. Newell. The evolution of reefs. Scientific American, 226(6):54–65, June 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n6/pdf/scientificamerican0672-
54.pdf.

Heinrich:1972:TCF

[492] Bernd Heinrich and George A. Bartholomew. Temperature control in flying moths. Scientific American, 226(6):70–77, June 1972. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n6/pdf/scientificamerican0672-70.pdf.

Davis:1972:NA

[493] Martin Davis and Reuben Hersh. Nonstandard analysis. Scientific American, 226(6):78–86, June 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n6/pdf/scientificamerican0672-
78.pdf.

Ratliff:1972:CC

[494] Floyd Ratliff. Contour and contrast. Scientific American, 226(6):90– 101, June 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n6/pdf/scientificamerican0672-90.pdf.

Weiss:1972:PFS

[495] Jay M. Weiss. Psychological factors in stress and disease. Scientific American, 226(6):104–113, June 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v226/n6/pdf/scientificamerican0672-
104.pdf.

Gardner:1972:MGf

[496] Martin Gardner. Mathematical games. Scientific American, 226(6):114– 118, June 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n6/pdf/scientificamerican0672-114.pdf.

Stong:1972:ASf

[497] C. L. Stong. The amateur scientist. Scientific American, 226(6):122– 128, June 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v226/n6/pdf/scientificamerican0672-122.pdf.

Morrison:1972:Be

[498] Philip Morrison. Books. Scientific American, 226(6):131–135, June 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n6/pdf/scientificamerican0672-131.pdf.

Anonymous:1972:Bf

[499] Anonymous. Bibliography. Scientific American, 226(6):136, June 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v226/
n6/pdf/scientificamerican0672-136.pdf.

Anonymous:1972:Lg

[500] Anonymous. Letters. Scientific American, 227(1):6, July 1972. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n1/pdf/scientificamerican0772-6.pdf.

Anonymous:1972:YAg

[501] Anonymous. 50 and 100 years ago. Scientific American, 227(1):8, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n1/pdf/scientificamerican0772-8.pdf.

Anonymous:1972:Ag

[502] Anonymous. The authors. Scientific American, 227(1):13, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n1/pdf/scientificamerican0772-13.pdf.

Garwin:1972:AWN

[503] Richard L. Garwin. Antisubmarine warfare and national security. Scientific American, 227(1):14–25, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n1/pdf/scientificamerican0772-
14.pdf.

Schnopper:1972:XRS

[504] Herbert W. Schnopper and John P. Delvaille. The X-ray sky. Scientific American, 227(1):26–37, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n1/pdf/scientificamerican0772-
26.pdf.

Schnopper:1972:XS

[505] H. W. Schnopper and J. P. Delvaille. The X-ray sky. Scientific American, 227(1):26–37, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bartley:1972:LCL

[506] W. W. Bartley III. Lewis Carroll’s lost book on logic. Scientific American, 227(1):38–46, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n1/pdf/scientificamerican0772-
38.pdf.

Anonymous:1972:SCg

[507] Anonymous. Science and the citizen. Scientific American, 227(1):48– 51, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n1/pdf/scientificamerican0772-48.pdf.

Frieden:1972:CEL

[508] Earl Frieden. The chemical elements of life. Scientific American, 227(1): 52–60, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n1/pdf/scientificamerican0772-52.pdf.

Coppi:1972:TAF

[509] Bruno Coppi and Jan Rem. The Tokamak approach in fusion research. Scientific American, 227(1):65–75, July 1972. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n1/pdf/scientificamerican0772-65.pdf.

Coppie:1972:TAF

[510] B. Coppie and J. Rem. The Tokamak approach in fusion research. Scientific American, 227(1):65–75, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gardner:1972:DD

[511] Lytt I. Gardner. Deprivation dwarfism. Scientific American, 227(1):76– 82, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n1/pdf/scientificamerican0772-76.pdf.

Kolers:1972:ER

[512] Paul A. Kolers. Experiments in reading. Scientific American, 227(1):84– 91, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n1/pdf/scientificamerican0772-84.pdf.

Feder:1972:ERM

[513] Howard M. Feder. Escape responses in marine invertebrates. Scientific American, 227(1):92–100, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n1/pdf/scientificamerican0772-
92.pdf.

Gardner:1972:MGg

[514] Martin Gardner. Mathematical games. Scientific American, 227(1):102– 105, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n1/pdf/scientificamerican0772-102.pdf.

Stong:1972:ASg

[515] C. L. Stong. The amateur scientist. Scientific American, 227(1):106– 112, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n1/pdf/scientificamerican0772-106.pdf.

Morrison:1972:Bf

[516] Philip Morrison. Books. Scientific American, 227(1):117–121, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n1/pdf/scientificamerican0772-117.pdf.

Anonymous:1972:Bg

[517] Anonymous. Bibliography. Scientific American, 227(1):122, July 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n1/pdf/scientificamerican0772-122.pdf.

Anonymous:1972:Lh

[518] Anonymous. Letters. Scientific American, 227(2):8, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n2/pdf/scientificamerican0872-8.pdf.

Anonymous:1972:YAh

[519] Anonymous. 50 and 100 years ago. Scientific American, 227(2):11, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n2/pdf/scientificamerican0872-11.pdf.

Anonymous:1972:Ah

[520] Anonymous. The authors. Scientific American, 227(2):13, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n2/pdf/scientificamerican0872-13.pdf.

Cole:1972:WE

[521] David E. Cole. The Wankel engine. Scientific American, 227(2):14–23, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n2/pdf/scientificamerican0872-14.pdf.

Hess:1972:INL

[522] Eckhard H. Hess. “Imprinting” in a natural laboratory. Scientific American, 227(2):24–31, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n2/pdf/scientificamerican0872-
24.pdf.

Breslow:1972:NAM

[523] Ronald Breslow. The nature of aromatic molecules. Scientific American, 227(2):32–40, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n2/pdf/scientificamerican0872-
32.pdf.

Anonymous:1972:SCh

[524] Anonymous. Science and the citizen. Scientific American, 227(2):42–46, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n2/pdf/scientificamerican0872-42.pdf.

Bok:1972:BS

[525] Bart J. Bok. The birth of stars. Scientific American, 227(2):48–61, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n2/pdf/scientificamerican0872-48.pdf.

Korsch:1972:DPC

[526] Barbara M. Korsch and Vida Francis Negrete. Doctor-patient communication. Scientific American, 227(2):66–74, August 1972. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n2/pdf/scientificamerican0872-66.pdf.

Heath:1972:OBC

[527] F. G. Heath. Origins of the binary code. Scientific American, 227(2):76– 83, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n2/pdf/scientificamerican0872-76.pdf.

Pettigrew:1972:NBV

[528] John D. Pettigrew. The neurophysiology of binocular vision. Scientific American, 227(2):84–95, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n2/pdf/scientificamerican0872-
84.pdf.

Pastan:1972:CA

[529] Ira Pastan. Cyclic AMP. Scientific American, 227(2):97–105, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n2/pdf/scientificamerican0872-97.pdf.

Gardner:1972:MGh

[530] Martin Gardner. Mathematical games. Scientific American, 227 (2):106–109, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n2/pdf/scientificamerican0872-
106.pdf.

Stong:1972:ASh

[531] C. L. Stong. The amateur scientist. Scientific American, 227(2):110–117, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n2/pdf/scientificamerican0872-110.pdf.

Morrison:1972:Bg

[532] Philip Morrison. Books. Scientific American, 227(2):118–123, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n2/pdf/scientificamerican0872-118.pdf.

Anonymous:1972:Bh

[533] Anonymous. Bibliography. Scientific American, 227(2):124, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n2/pdf/scientificamerican0872-124.pdf.

Anonymous:1972:Li

[534] Anonymous. Letters. Scientific American, 227(3):8–12, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n3/pdf/scientificamerican0972-8.pdf.

Anonymous:1972:YAi

[535] Anonymous. 50 and 100 years ago. Scientific American, 227(3):16–20, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n3/pdf/scientificamerican0972-16.pdf.

Anonymous:1972:Ai

[536] Anonymous. The authors. Scientific American, 227(3):22–26, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n3/pdf/scientificamerican0972-22.pdf.

Pierce:1972:C

[537] John R. Pierce. Communication. Scientific American, 227(3):30–41, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n3/pdf/scientificamerican0972-30.pdf.

Stent:1972:CC

[538] Gunther S. Stent. Cellular communication. Scientific American, 227(3):42–51, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n3/pdf/scientificamerican0972-
42.pdf.

Wilson:1972:AC

[539] Edward O. Wilson. Animal communication. Scientific American, 227(3):52–60, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n3/pdf/scientificamerican0972-
52.pdf.

Anonymous:1972:SCi

[540] Anonymous. Science and the citizen. Scientific American, 227(3):64–69, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n3/pdf/scientificamerican0972-64.pdf.

Jakobson:1972:VC

[541] Roman Jakobson. Verbal communication. Scientific American, 227(3):72–80, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n3/pdf/scientificamerican0972-
72.pdf.

Gombrich:1972:VI

[542] E. H. Gombrich. The visual image. Scientific American, 227(3):82–96, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n3/pdf/scientificamerican0972-82.pdf.

Busignies:1972:CC

[543] Henri Busignies. Communication channels. Scientific American, 227(3):98–113, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n3/pdf/scientificamerican0972-
98.pdf.

Inose:1972:CN

[544] Hiroshi Inose. Communication networks. Scientific American, 227 (3):116–128, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n3/pdf/scientificamerican0972-
116.pdf.

Kretzmer:1972:CT

[545] Ernest R. Kretzmer. Communication terminals. Scientific American, 227(3):130–140, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n3/pdf/scientificamerican0972-
130.pdf.

Goldmark:1972:CC

[546] Peter C. Goldmark. Communication and the community. Scientific American, 227(3):142–150, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v227/n3/pdf/scientificamerican0972-
142.pdf.

Gerbner:1972:CSE

[547] George Gerbner. Communication and social environment. Scientific American, 227(3):152–160, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v227/n3/pdf/scientificamerican0972-
152.pdf.

Emerson:1972:CFE

[548] Thomas I. Emerson. Communication and freedom of expression. Scientific American, 227(3):163–172, September 1972. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n3/pdf/scientificamerican0972-163.pdf.

Gardner:1972:MGi

[549] Martin Gardner. Mathematical games. Scientific American, 227 (3):176–182, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n3/pdf/scientificamerican0972-
176.pdf.

Stong:1972:ASi

[550] C. L. Stong. The amateur scientist. Scientific American, 227(3):185–190, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n3/pdf/scientificamerican0972-185.pdf.

Morrison:1972:Bh

[551] Philip Morrison. Books. Scientific American, 227(3):193–206, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n3/pdf/scientificamerican0972-193.pdf.

Anonymous:1972:Bi

[552] Anonymous. Bibliography. Scientific American, 227(3):208–210, September 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n3/pdf/scientificamerican0972-208.pdf.

Anonymous:1972:Lj

[553] Anonymous. Letters. Scientific American, 227(4):8, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n4/pdf/scientificamerican1072-8.pdf.

Anonymous:1972:YAj

[554] Anonymous. 50 and 100 years ago. Scientific American, 227(4):11, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n4/pdf/scientificamerican1072-11.pdf.

Anonymous:1972:Aj

[555] Anonymous. The authors. Scientific American, 227(4):14, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n4/pdf/scientificamerican1072-14.pdf.

Kershaw:1972:NIT

[556] David N. Kershaw. A negative-income-tax experiment. Scientific American, 227(4):19–25, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n4/pdf/scientificamerican1072-
19.pdf.

Squires:1972:CPD

[557] Arthur M. Squires. Clean power from dirty fuels. Scientific American, 227(4):26–35, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n4/pdf/scientificamerican1072-
26.pdf.

Chadwick:1972:LMG

[558] John Chadwick. Life in Mycenaean Greece. Scientific American, 227(4):36–44, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n4/pdf/scientificamerican1072-
36.pdf.

Anonymous:1972:SCj

[559] Anonymous. Science and the citizen. Scientific American, 227(4):46–48, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n4/pdf/scientificamerican1072-46.pdf.

Kino:1972:ASW

[560] Gordon S. Kino and John Shaw. Acoustic surface waves. Scientific American, 227(4):50–68, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n4/pdf/scientificamerican1072-
50.pdf.

Kretchmer:1972:LL

[561] Norman Kretchmer. Lactose and lactase. Scientific American, 227(4):70– 78, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n4/pdf/scientificamerican1072-70.pdf.

Eglinton:1972:CCM

[562] Geoffrey Eglinton, James R. Maxwell, and Colin T. Pillinger. The carbon chemistry of the Moon. Scientific American, 227(4):80–90, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n4/pdf/scientificamerican1072-80.pdf.

Premack:1972:TLA

[563] Ann James Premack and David Premack. Teaching language to an ape. Scientific American, 227(4):92–99, October 1972. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n4/pdf/scientificamerican1072-92.pdf.

Zafiratos:1972:TNS

[564] Chris D. Zafiratos. The texture of the nuclear surface. Scientific American, 227(4):100–108, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n4/pdf/scientificamerican1072-
100.pdf.

Gardner:1972:MGj

[565] Martin Gardner. Mathematical games. Scientific American, 227 (4):110–112B, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n4/pdf/scientificamerican1072-
110.pdf.

Stong:1972:ASj

[566] C. L. Stong. The amateur scientist. Scientific American, 227(4):114–118, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n4/pdf/scientificamerican1072-114.pdf.

Morrison:1972:Bi

[567] Philip Morrison. Books. Scientific American, 227(4):121–127, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n4/pdf/scientificamerican1072-121.pdf.

Anonymous:1972:Bj

[568] Anonymous. Bibliography. Scientific American, 227(4):128, October 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n4/pdf/scientificamerican1072-128.pdf.

Anonymous:1972:Lk

[569] Anonymous. Letters. Scientific American, 227(5):6–8, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n5/pdf/scientificamerican1172-6.pdf.

Anonymous:1972:YAk

[570] Anonymous. 50 and 100 years ago. Scientific American, 227(5):10, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n5/pdf/scientificamerican1172-10.pdf.

Anonymous:1972:Ak

[571] Anonymous. The authors. Scientific American, 227(5):13, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n5/pdf/scientificamerican1172-13.pdf.

York:1972:GTB

[572] Herbert F. York. The Great Test-Ban Debate. Scientific American, 227(5):15–23, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n5/pdf/scientificamerican1172-
15.pdf.

Guillemin:1972:HH

[573] Roger Guillemin and Roger Burgus. The hormones of the hypothalamus. Scientific American, 227(5):24–33, November 1972. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n5/pdf/scientificamerican1172-24.pdf.

Broers:1972:MEB

[574] A. N. Broers and M. Hatzakis. Microcircuits by electron beam. Scientific American, 227(5):34–44, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v227/n5/pdf/scientificamerican1172-
34.pdf.

Anonymous:1972:SCk

[575] Anonymous. Science and the citizen. Scientific American, 227(5):49–54, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n5/pdf/scientificamerican1172-49.pdf.

Hallam:1972:CDF

[576] A. Hallam. Continental drift and the fossil record. Scientific American, 227(5):56–66, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n5/pdf/scientificamerican1172-
56.pdf.

Topoff:1972:SBA

[577] Howard R. Topoff. The social behavior of army ants. Scientific American, 227(5):70–79, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n5/pdf/scientificamerican1172-
70.pdf.

Deregowski:1972:PPC

[578] Jan B. Deregowski. Pictorial perception and culture. Scientific American, 227(5):82–88, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n5/pdf/scientificamerican1172-
82.pdf.

Mark:1972:SAG

[579] Robert Mark. The structural analysis of Gothic cathedrals. Scientific American, 227(5):90–99, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v227/n5/pdf/scientificamerican1172-
90.pdf.

Wiegand:1972:EA

[580] Clyde E. Wiegand. Exotic atoms. Scientific American, 227(5): 102–108, 110, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n5/pdf/scientificamerican1172-
102.pdf.

Gardner:1972:MGk

[581] Martin Gardner. Mathematical games. Scientific American, 227 (5):114–118, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n5/pdf/scientificamerican1172-
114.pdf.

Stong:1972:ASk

[582] C. L. Stong. The amateur scientist. Scientific American, 227(5):120–124, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n5/pdf/scientificamerican1172-120.pdf.

Morrison:1972:Bj

[583] Philip Morrison. Books. Scientific American, 227(5):127–133, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n5/pdf/scientificamerican1172-127.pdf.

Anonymous:1972:Bk

[584] Anonymous. Bibliography. Scientific American, 227(5):134–136, November 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n5/pdf/scientificamerican1172-134.pdf.

Anonymous:1972:Ll

[585] Anonymous. Letters. Scientific American, 227(6):6, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n6/pdf/scientificamerican1272-6.pdf.

Anonymous:1972:YAl

[586] Anonymous. 50 and 100 years ago. Scientific American, 227(6):8–9, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n6/pdf/scientificamerican1272-8.pdf.

Anonymous:1972:Al

[587] Anonymous. The authors. Scientific American, 227(6):10–11, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n6/pdf/scientificamerican1272-10.pdf.

Tsu:1972:HTC

[588] Raphael Tsu. High technology in China. Scientific American, 227(6):13–17, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n6/pdf/scientificamerican1272-
13.pdf.

Rosenblatt:1972:LNK

[589] Jay S. Rosenblatt. Learning in newborn kittens. Scientific American, 227(6):18–25, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n6/pdf/scientificamerican1272-
18.pdf.

Hsu:1972:WMD

[590] Kenneth J. Hsü. When the Mediterranean dried UP. Scientific American, 227(6):26–36, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n6/pdf/scientificamerican1272-
26.pdf.

Anonymous:1972:SCl

[591] Anonymous. Science and the citizen. Scientific American, 227(6):40–43, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n6/pdf/scientificamerican1272-40.pdf.

Backer:1972:Y

[592] Stanley Backer. Yarn. Scientific American, 227(6):46–56, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n6/pdf/scientificamerican1272-46.pdf.

Apfel:1972:TSL

[593] Robert E. Apfel. The tensile strength of liquids. Scientific American, 227(6):58–62, 67–71, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n6/pdf/scientificamerican1272-
58.pdf.

Gordon:1972:SCB

[594] Barbara Gordon. The superior colliculus of the brain. Scientific American, 227(6):72–82, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n6/pdf/scientificamerican1272-
72.pdf.

Stent:1972:PUS

[595] Gunther S. Stent. Prematurity and uniqueness in scientific discovery. Scientific American, 227(6):84–93, December 1972. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v227/
n6/pdf/scientificamerican1272-84.pdf.

Lapan:1972:M

[596] Elliot A. Lapan and Harold Morowitz. The Mesozoa. Scientific American, 227(6):94–101, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n6/pdf/scientificamerican1272-
94.pdf.

Gardner:1972:MGl

[597] Martin Gardner. Mathematical games. Scientific American, 227 (6):102–106, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v227/n6/pdf/scientificamerican1272-
102.pdf.

Stong:1972:ASl

[598] C. L. Stong. The amateur scientist. Scientific American, 227(6):108–111, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n6/pdf/scientificamerican1272-108.pdf.

Philip:1972:B

[599] Philip and Phylis Morrison. Books. Scientific American, 227(6):112–123, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n6/pdf/scientificamerican1272-112.pdf.

Anonymous:1972:AIa

[600] Anonymous. Annual index. Scientific American, 227(6):124–126, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n6/pdf/scientificamerican1272-124.pdf.

Anonymous:1972:Bl

[601] Anonymous. Bibliography. Scientific American, 227(6):128, December 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v227/n6/pdf/scientificamerican1272-128.pdf.

Anonymous:1973:La

[602] Anonymous. Letters. Scientific American, 228(1):8, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n1/pdf/scientificamerican0173-8.pdf.

Anonymous:1973:YAYa

[603] Anonymous. 50 and 100 years ago: 50 and 100 years ago. Scientific American, 228(1):10, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n1/pdf/scientificamerican0173-
10.pdf.

Anonymous:1973:Aa

[604] Anonymous. The authors. Scientific American, 228(1):11, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n1/pdf/scientificamerican0173-11.pdf.

Gregory:1973:HE

[605] Derek P. Gregory. The hydrogen economy. Scientific American, 228(1):13–21, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n1/pdf/scientificamerican0173-
13.pdf.

Notkins:1973:HIR

[606] Abner Louis Notkins and Hilary Koprowski. How the immune response to a virus can cause disease. Scientific American, 228(1):22–31, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n1/pdf/scientificamerican0173-22.pdf.

Teleki:1973:OC

[607] Geza Teleki. The omnivorous chimpanzee. Scientific American, 228(1):32–42, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n1/pdf/scientificamerican0173-
32.pdf.

Anonymous:1973:SCa

[608] Anonymous. Science and the citizen. Scientific American, 228(1):44–46, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n1/pdf/scientificamerican0173-44.pdf.

Murray:1973:MM

[609] Bruce C. Murray. Mars from Mariner 9. Scientific American, 228 (1):48–63, 6–9, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n1/pdf/scientificamerican0173-
48.pdf.

Werblin:1973:CSR

[610] Frank S. Werblin. The control of sensitivity in the retina. Scientific American, 228(1):70–79, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n1/pdf/scientificamerican0173-
70.pdf.

Stone:1973:JTC

[611] Michael E. Stone. Judaism at the time of Christ. Scientific American, 228(1):80–87, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n1/pdf/scientificamerican0173-
80.pdf.

Azbel:1973:CEM

[612] M. Ya. Azbel’, M. I. Kaganov, and I. M. Lifshitz. Conduction electrons in metals. Scientific American, 228(1):88–98, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n1/pdf/scientificamerican0173-88.pdf.

Charles:1973:SC

[613] Charles and Nancy Knight. Snow crystals. Scientific American, 228(1):100–107, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n1/pdf/scientificamerican0173-
100.pdf.

Knight:1973:SC

[614] C. Knight and N. Knight. Snow crystals. Scientific American, 228(1): 100–107, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gardner:1973:MGa

[615] Martin Gardner. Mathematical games. Scientific American, 228 (1):108–115, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n1/pdf/scientificamerican0173-
108.pdf.

Stong:1973:ASa

[616] C. L. Stong. The amateur scientist. Scientific American, 228(1):116–121, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n1/pdf/scientificamerican0173-116.pdf.

Morrison:1973:Ba

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n1/pdf/scientificamerican0173-123.pdf.

Anonymous:1973:Ba

[618] Anonymous. Bibliography. Scientific American, 228(1):128, January 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n1/pdf/scientificamerican0173-128.pdf.

Anonymous:1973:Lb

[619] Anonymous. Letters. Scientific American, 228(2):6–7, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n2/pdf/scientificamerican0273-6.pdf.

Anonymous:1973:YAYb

[620] Anonymous. 50 and 100 years ago: 50 and 100 years ago. Scientific American, 228(2):8–9, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n2/pdf/scientificamerican0273-
8.pdf.

Anonymous:1973:Ab

[621] Anonymous. The authors. Scientific American, 228(2):12–13, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n2/pdf/scientificamerican0273-12.pdf.

Greenwood:1973:RAC

[622] Ted Greenwood. Reconnaissance and arms control. Scientific American, 228(2):14–25, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n2/pdf/scientificamerican0273-
14.pdf.

Jacobson:1973:ONC

[623] Marcus Jacobson and R. Kevin Hunt. The origins of nerve-cell specificity. Scientific American, 228(2):26–35, February 1973. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n2/pdf/scientificamerican0273-26.pdf.

Carey:1973:FWB

[624] Francis G. Carey. Fishes with warm bodies. Scientific American, 228(2):36–44, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n2/pdf/scientificamerican0273-
36.pdf.

Anonymous:1973:SCb

[625] Anonymous. Science and the citizen. Scientific American, 228(2):46–49, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n2/pdf/scientificamerican0273-46.pdf.

Wang:1973:CL

[626] William S-Y. Wang. The Chinese language. Scientific American, 228(2):50–60, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n2/pdf/scientificamerican0273-
50.pdf.

Gregg:1973:MO

[627] Michael C. Gregg. The microstructure of the ocean. Scientific American, 228(2):64–77, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n2/pdf/scientificamerican0273-
64.pdf.

Miller:1973:CA

[628] Patrick M. Miller. The crashworthiness of automobiles. Scientific American, 228(2):78–86, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n2/pdf/scientificamerican0273-
78.pdf.

Silfvast:1973:MVL

[629] William T. Silfvast. Metal-vapor lasers. Scientific American, 228(2):88–97, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n2/pdf/scientificamerican0273-
88.pdf.

Silfvast:1973:ML

[630] W. T. Silfvast. Metal-vapor lasers. Scientific American, 228(2):89–97, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Pacini:1973:RHA

[631] F. Pacini and M. J. Rees. Rotation in high-energy astrophysics. Scientific American, 228(2):98–105, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Pacini:1973:RHE

[632] Franco Pacini and Martin J. Rees. Rotation in high-energy astrophysics. Scientific American, 228(2):98–105, February 1973. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n2/pdf/scientificamerican0273-98.pdf.

Gardner:1973:MGb

[633] Martin Gardner. Mathematical games. Scientific American, 228 (2):106–109, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n2/pdf/scientificamerican0273-
106.pdf.

Stong:1973:ASb

[634] C. L. Stong. The amateur scientist. Scientific American, 228(2):110–115, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n2/pdf/scientificamerican0273-110.pdf.

Cohen:1973:B

[635] I. Bernard Cohen. Books. Scientific American, 228(2):117–123, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n2/pdf/scientificamerican0273-117.pdf.

Anonymous:1973:Bb

[636] Anonymous. Bibliography. Scientific American, 228(2):124, February 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n2/pdf/scientificamerican0273-124.pdf.

Anonymous:1973:Lc

[637] Anonymous. Letters. Scientific American, 228(3):8–9, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n3/pdf/scientificamerican0373-8.pdf.

Anonymous:1973:YAYc

[638] Anonymous. 50 and 100 years ago: 50 and 100 years ago. Scientific American, 228(3):10, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n3/pdf/scientificamerican0373-
10.pdf.

Anonymous:1973:Ac

[639] Anonymous. The authors. Scientific American, 228(3):12–13, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n3/pdf/scientificamerican0373-12.pdf.

Frejka:1973:PSW

[640] Tomas Frejka. The prospects for a stationary world population. Scientific American, 228(3):15–23, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n3/pdf/scientificamerican0373-
15.pdf.

Bolt:1973:FSE

[641] Bruce A. Bolt. The fine structure of the Earth’s interior. Scientific American, 228(3):24–33, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n3/pdf/scientificamerican0373-
24.pdf.

Miller:1973:VGA

[642] O. L. Miller, Jr. The visualization of genes in action. Scientific American, 228(3):34–42, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n3/pdf/scientificamerican0373-
34.pdf.

Anonymous:1973:SCc

[643] Anonymous. Science and the citizen. Scientific American, 228(3):44–48, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n3/pdf/scientificamerican0373-44.pdf.

Turner:1973:IM

[644] Barry K. Turner. Interstellar molecules. Scientific American, 228 (3):50–63, 67–69, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n3/pdf/scientificamerican0373-
50.pdf.

Kimura:1973:AHB

[645] Doreen Kimura. The asymmetry of the human brain. Scientific American, 228(3):70–78, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n3/pdf/scientificamerican0373-
70.pdf.

Wilson:1973:BT

[646] S. S. Wilson. Bicycle technology. Scientific American, 228(3):81–91, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n3/pdf/scientificamerican0373-81.pdf.

Leggett:1973:MS

[647] William C. Leggett. The migrations of the shad. Scientific American, 228(3):92–98, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n3/pdf/scientificamerican0373-
92.pdf.

Brainerd:1973:ONC

[648] Charles J. Brainerd. The origins of number concepts. Scientific American, 228(3):101–109, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n3/pdf/scientificamerican0373-
101.pdf.

Gardner:1973:MGc

[649] Martin Gardner. Mathematical games. Scientific American, 228(3):110– 113, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n3/pdf/scientificamerican0373-110.pdf.

Stong:1973:ASc

[650] C. L. Stong. The amateur scientist. Scientific American, 228(3):114–119, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n3/pdf/scientificamerican0373-114.pdf.

Morrison:1973:Bb

[651] Philip Morrison. Books. Scientific American, 228(3):121–127, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n3/pdf/scientificamerican0373-121.pdf.

Anonymous:1973:Bc

[652] Anonymous. Bibliography. Scientific American, 228(3):128, March 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n3/pdf/scientificamerican0373-128.pdf.

Anonymous:1973:Ld

[653] Anonymous. Letters. Scientific American, 228(4):8–9, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n4/pdf/scientificamerican0473-8.pdf.

Anonymous:1973:YAYd

[654] Anonymous. 50 and 100 years ago: 50 and 100 years ago. Scientific American, 228(4):10, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n4/pdf/scientificamerican0473-
10.pdf.

Anonymous:1973:Ad

[655] Anonymous. The authors. Scientific American, 228(4):11, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n4/pdf/scientificamerican0473-11.pdf.

Glazier:1973:TM

[656] William H. Glazier. The task of medicine. Scientific American, 228(4):13–17, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n4/pdf/scientificamerican0473-
13.pdf.

Clowes:1973:MID

[657] Royston C. Clowes. The molecule of infectious drug resistance. Scientific American, 228(4):18–27, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n4/pdf/scientificamerican0473-
18.pdf.

Neugebauer:1973:BIS

[658] G. Neugebauer and Eric E. Becklin. The brightest infrared sources. Scientific American, 228(4):28–40, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v228/n4/pdf/scientificamerican0473-
28.pdf.

Anonymous:1973:SC

[659] Anonymous. Science and the citizen. Scientific American, 228(4):43– 44, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1973:SCd

[660] Anonymous. Science and the citizen. Scientific American, 228(4):43–45, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n4/pdf/scientificamerican0473-43.pdf.

Peixoto:1973:CWC

[661] José P. Peixoto and M. Ali Kettani. The control of the water cycle. Scientific American, 228(4):46–61, April 1973. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n4/pdf/scientificamerican0473-46.pdf.

Morehead:1973:II

[662] Frederick F. Morehead, Jr. and Billy L. Crowder. Ion implantation. Scientific American, 228(4):64–71, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v228/n4/pdf/scientificamerican0473-
64.pdf.

Avery:1973:LNI

[663] Mary Ellen Avery, Nai-San Wang, and H. William Taeusch, Jr. The lung of the newborn infant. Scientific American, 228(4):74–85, April 1973.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n4/pdf/scientificamerican0473-74.pdf.

Lerner:1973:GB

[664] Lawrence S. Lerner and Edward A. Gosselin. Giordano Bruno. Scientific American, 228(4):86–94, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n4/pdf/scientificamerican0473-
86.pdf.

Heinrich:1973:EB

[665] Bernd Heinrich. The energetics of the bumblebee. Scientific American, 228(4):96–102, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n4/pdf/scientificamerican0473-
96.pdf.

Gardner:1973:MGd

[666] Martin Gardner. Mathematical games. Scientific American, 228(4):106– 111, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n4/pdf/scientificamerican0473-106.pdf.

Stong:1973:ASd

[667] C. L. Stong. The amateur scientist. Scientific American, 228(4):112– 117, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n4/pdf/scientificamerican0473-112.pdf.

Morrison:1973:Bc

[668] Philip Morrison. Books. Scientific American, 228(4):119–123, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n4/pdf/scientificamerican0473-119.pdf.

Anonymous:1973:Bd

[669] Anonymous. Bibliography. Scientific American, 228(4):124, April 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n4/pdf/scientificamerican0473-124.pdf.

Anonymous:1973:Le

[670] Anonymous. Letters. Scientific American, 228(5):8–9, May 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n5/pdf/scientificamerican0573-8.pdf.

Anonymous:1973:YAYe

[671] Anonymous. 50 and 100 years ago: 50 and 100 years ago. Scientific American, 228(5):10–11, May 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n5/pdf/scientificamerican0573-
10.pdf.

Anonymous:1973:Ae

[672] Anonymous. The authors. Scientific American, 228(5):12–13, May 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n5/pdf/scientificamerican0573-12.pdf.

Feistel:1973:CCP

[673] Horst Feistel. Cryptography and computer privacy. Scientific American, 228(5):15–23, May 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n5/pdf/scientificamerican0573-
15.pdf.

Salk:1973:RHR

[674] Lee Salk. The role of the heartbeat in the relations between mother and infant. Scientific American, 228(5):24–29, May 1973. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n5/pdf/scientificamerican0573-24.pdf.

Dash:1973:TDM

[675] J. C. Dash. Two-dimensional matter. Scientific American, 228(5):30– 40, May 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n5/pdf/scientificamerican0573-30.pdf.

Anonymous:1973:SCe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n5/pdf/scientificamerican0573-42.pdf.

Epstein:1973:R

[677] Emanuel Epstein. Roots. Scientific American, 228(5):48–58, May 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n5/pdf/scientificamerican0573-48.pdf.

McKenzie:1973:EIO

[678] D. P. McKenzie and J. G. Sclater. The evolution of the Indian Ocean. Scientific American, 228(5):62–72, May 1973. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n5/pdf/scientificamerican0573-62.pdf.

Salmon:1973:C

[679] Wesley C. Salmon. Confirmation. Scientific American, 228(5):75–83, May 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n5/pdf/scientificamerican0573-75.pdf.

Drake:1973:GDL

[680] Stillman Drake. Galileo’s discovery of the Law of Free Fall. Scientific American, 228(5):84–92, May 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n5/pdf/scientificamerican0573-
84.pdf.

Gamow:1973:IRS

[681] R. Igor Gamow and John F. Harris. The infrared receptors of snakes. Scientific American, 228(5):94–100, May 1973. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n5/pdf/scientificamerican0573-94.pdf.

Gardner:1973:MGe

[682] Martin Gardner. Mathematical games. Scientific American, 228(5):102– 107, May 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n5/pdf/scientificamerican0573-102.pdf.

Stong:1973:ASe

[683] C. L. Stong. The amateur scientist. Scientific American, 228(5):108– 113, May 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n5/pdf/scientificamerican0573-108.pdf.

Morrison:1973:Bd

[684] Philip Morrison. Books. Scientific American, 228(5):115–119, May 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n5/pdf/scientificamerican0573-115.pdf.

Anonymous:1973:Be

[685] Anonymous. Bibliography. Scientific American, 228(5):120, May 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n5/pdf/scientificamerican0573-120.pdf.

Anonymous:1973:Lf

[686] Anonymous. Letters. Scientific American, 228(6):6, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n6/pdf/scientificamerican0673-6.pdf.

Anonymous:1973:YAYf

[687] Anonymous. 50 and 100 years ago: 50 and 100 years ago. Scientific American, 228(6):8–9, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n6/pdf/scientificamerican0673-
8.pdf.

Anonymous:1973:Af

[688] Anonymous. The authors. Scientific American, 228(6):10–12, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n6/pdf/scientificamerican0673-10.pdf.

deNevers:1973:ECA

[689] Noel de Nevers. Enforcing the clean air act of 1970. Scientific American, 228(6):14–21, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n6/pdf/scientificamerican0673-
14.pdf.

Idyll:1973:AC

[690] C. P. Idyll. The anchovy crisis. Scientific American, 228(6):22–29, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n6/pdf/scientificamerican0673-22.pdf.

Rubin:1973:DAN

[691] Vera C. Rubin. The dynamics of the Andromeda Nebula. Scientific American, 228(6):30–37, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v228/n6/pdf/scientificamerican0673-
30.pdf.

Anonymous:1973:SCf

[692] Anonymous. Science and the citizen. Scientific American, 228(6):38–40, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n6/pdf/scientificamerican0673-38.pdf.

Alfano:1973:UPL

[693] R. R. Alfano and S. L. Shapiro. Ultrafast phenomena in liquids and solids. Scientific American, 228(6):42–55, 58, 60, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n6/pdf/scientificamerican0673-42.pdf.

Sobel:1973:EN

[694] Alan Sobel. Electronic numbers. Scientific American, 228(6):64–73, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n6/pdf/scientificamerican0673-64.pdf.

Denison:1973:LTT

[695] William C. Denison. Life in tall trees. Scientific American, 228(6):74– 80, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n6/pdf/scientificamerican0673-74.pdf.

Lerner:1973:HLE

[696] Richard A. Lerner and Frank J. Dixon. The human lymphocyte as an experimental animal. Scientific American, 228(6):82–91, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n6/pdf/scientificamerican0673-82.pdf.

Zobrist:1973:ACC

[697] A. L. Zobrist and F. R. Carlson, Jr. An advice-taking chess computer. Scientific American, 228(6):92–105, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zobrist:1973:ATC

[698] Albert L. Zobrist and Frederic R. Carlson, Jr. An advice-taking chess computer. Scientific American, 228(6):92–105, June 1973. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n6/pdf/scientificamerican0673-92.pdf.

Gardner:1973:MGf

[699] Martin Gardner. Mathematical games. Scientific American, 228(6):106– 110, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n6/pdf/scientificamerican0673-106.pdf.

Stong:1973:ASf

[700] C. L. Stong. The amateur scientist. Scientific American, 228(6):112– 115, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n6/pdf/scientificamerican0673-112.pdf.

Morrison:1973:Be

[701] Philip Morrison. Books. Scientific American, 228(6):117–121, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v228/
n6/pdf/scientificamerican0673-117.pdf.

Anonymous:1973:Bf

[702] Anonymous. Bibliography. Scientific American, 228(6):122–124, June 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v228/n6/pdf/scientificamerican0673-122.pdf.

Anonymous:1973:Lg

[703] Anonymous. Letters. Scientific American, 229(1):8, July 1973. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v229/
n1/pdf/scientificamerican0773-8.pdf.

Anonymous:1973:YAYg

[704] Anonymous. 50 and 100 years ago: 50 and 100 years ago. Scientific American, 229(1):12, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n1/pdf/scientificamerican0773-
12.pdf.

Anonymous:1973:Ag

[705] Anonymous. The authors. Scientific American, 229(1):14, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v229/
n1/pdf/scientificamerican0773-14.pdf.

Jaffe:1973:PPF

[706] Frederick S. Jaffe. Public policy on fertility control. Scientific American, 229(1):17–23, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n1/pdf/scientificamerican0773-
17.pdf.

Benade:1973:PB

[707] Arthur H. Benade. The physics of brasses. Scientific American, 229(1): 24–35, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n1/pdf/scientificamerican0773-24.pdf.

Clauser:1973:ACM

[708] Henry R. Clauser. Advanced composite materials. Scientific American, 229(1):36–44, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n1/pdf/scientificamerican0773-
36.pdf.

Anonymous:1973:SCg

[709] Anonymous. Science and the citizen. Scientific American, 229(1):46– 51, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n1/pdf/scientificamerican0773-46.pdf.

Jerne:1973:IS

[710] Niels Kaj Jerne. The immune system. Scientific American, 229(1):52– 60, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n1/pdf/scientificamerican0773-52.pdf.

Cameron:1973:MCR

[711] I. R. Cameron. Meteorites and cosmic radiation. Scientific American, 229(1):64–73, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n1/pdf/scientificamerican0773-
64.pdf.

Goldschmidt:1973:BS

[712] Walter Goldschmidt. The brideprice of the Sebei. Scientific American, 229(1):74–85, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n1/pdf/scientificamerican0773-
74.pdf.

Rona:1973:PTM

[713] Peter A. Rona. Plate tectonics and mineral resources. Scientific American, 229(1):86–95, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n1/pdf/scientificamerican0773-
86.pdf.

Evarts:1973:BMM

[714] Edward V. Evarts. Brain mechanisms in movement. Scientific American, 229(1):96–103, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n1/pdf/scientificamerican0773-
96.pdf.

Gardner:1973:MGg

[715] Martin Gardner. Mathematical games. Scientific American, 229(1):104– 109, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n1/pdf/scientificamerican0773-104.pdf.

Stong:1973:ASg

[716] C. L. Stong. The amateur scientist. Scientific American, 229(1):110– 115, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n1/pdf/scientificamerican0773-110.pdf.

Morrison:1973:Bf

[717] Philip Morrison. Books. Scientific American, 229(1):117–121, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v229/
n1/pdf/scientificamerican0773-117.pdf.

Anonymous:1973:Bg

[718] Anonymous. Bibliography. Scientific American, 229(1):122, July 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v229/
n1/pdf/scientificamerican0773-122.pdf.

Anonymous:1973:Lh

[719] Anonymous. Letters. Scientific American, 229(2):6–7, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v229/
n2/pdf/scientificamerican0873-6.pdf.

Anonymous:1973:YAYh

[720] Anonymous. 50 and 100 years ago: 50 and 100 years ago. Scientific American, 229(2):8, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n2/pdf/scientificamerican0873-
8.pdf.

Anonymous:1973:Ah

[721] Anonymous. The authors. Scientific American, 229(2):9, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v229/
n2/pdf/scientificamerican0873-9.pdf.

English:1973:DAB

[722] Richard D. English and Dan I. Bolef. Defense against bomber attack. Scientific American, 229(2):11–19, August 1973. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v229/
n2/pdf/scientificamerican0873-11.pdf.

Brown:1973:IG

[723] Donald D. Brown. The isolation of genes. Scientific American, 229(2):20– 29, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n2/pdf/scientificamerican0873-20.pdf.

Barish:1973:ENB

[724] Barry C. Barish. Experiments with neutrino beams. Scientific American, 229(2):30–38, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n2/pdf/scientificamerican0873-
30.pdf.

Anonymous:1973:SCh

[725] Anonymous. Science and the citizen. Scientific American, 229(2):42–47, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n2/pdf/scientificamerican0873-42.pdf.

Hittinger:1973:MOS

[726] William C. Hittinger. Metal-oxide semiconductor technology. Scientific American, 229(2):48–59, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n2/pdf/scientificamerican0873-
48.pdf.

James:1973:EA

[727] David E. James. The evolution of the Andes. Scientific American, 229(2):60–69, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n2/pdf/scientificamerican0873-
60.pdf.

Thorpe:1973:DSB

[728] W. H. Thorpe. Duet-singing birds. Scientific American, 229(2):70–79, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n2/pdf/scientificamerican0873-70.pdf.

Walker:1973:SE

[729] Graham Walker. The Stirling engine. Scientific American, 229(2):80–87, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n2/pdf/scientificamerican0873-80.pdf.

Brady:1973:HFM

[730] Roscoe O. Brady. Hereditary fat-metabolism diseases. Scientific American, 229(2):88–97, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n2/pdf/scientificamerican0873-
88.pdf.

Gardner:1973:MGh

[731] Martin Gardner. Mathematical games. Scientific American, 229 (2):98–101, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n2/pdf/scientificamerican0873-
98.pdf.

Stong:1973:ASh

[732] C. L. Stong. The amateur scientist. Scientific American, 229(2):104–109, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n2/pdf/scientificamerican0873-104.pdf.

Morrison:1973:Bg

[733] Philip Morrison. Books. Scientific American, 229(2):111–115, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n2/pdf/scientificamerican0873-111.pdf.

Anonymous:1973:Bh

[734] Anonymous. Bibliography. Scientific American, 229(2):116, August 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n2/pdf/scientificamerican0873-116.pdf.

Anonymous:1973:Li

[735] Anonymous. Letters. Scientific American, 229(3):8–13, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v229/
n3/pdf/scientificamerican0973-8.pdf.

Anonymous:1973:YAYi

[736] Anonymous. 50 and 100 years ago: 50 and 100 years ago. Scientific American, 229(3):14–15, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n3/pdf/scientificamerican0973-
14.pdf.

Anonymous:1973:Ai

[737] Anonymous. The authors. Scientific American, 229(3):16–18, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n3/pdf/scientificamerican0973-16.pdf.

White:1973:LDM

[738] Kerr L. White. Life and death and medicine. Scientific American, 229(3):22–33, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n3/pdf/scientificamerican0973-
22.pdf.

Tanner:1973:G

[739] J. M. Tanner. Growing up. Scientific American, 229(3):34–43, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n3/pdf/scientificamerican0973-34.pdf.

Leaf:1973:GO

[740] Alexander Leaf. Getting old. Scientific American, 229(3):44–52, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n3/pdf/scientificamerican0973-44.pdf.

Morison:1973:D

[741] Robert S. Morison. Dying. Scientific American, 229(3):54–62, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n3/pdf/scientificamerican0973-54.pdf.

Anonymous:1973:SCi

[742] Anonymous. Science and the citizen. Scientific American, 229(3):64–74, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n3/pdf/scientificamerican0973-64.pdf.

Dingle:1973:IM

[743] John H. Dingle. The ills of man. Scientific American, 229(3):76–84, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n3/pdf/scientificamerican0973-76.pdf.

Child:1973:SI

[744] Charles G. Child III. Surgical intervention. Scientific American, 229(3):90–98, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n3/pdf/scientificamerican0973-
90.pdf.

Mellinkoff:1973:CI

[745] Sherman M. Mellinkoff. Chemical intervention. Scientific American, 229(3):102–112, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n3/pdf/scientificamerican0973-
102.pdf.

Eisenberg:1973:PI

[746] Leon Eisenberg. Psychiatric intervention. Scientific American, 229 (3):116–127, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n3/pdf/scientificamerican0973-
116.pdf.

Knowles:1973:H

[747] John H. Knowles. The hospital. Scientific American, 229(3):128–137, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n3/pdf/scientificamerican0973-128.pdf.

Ebert:1973:MS

[748] Robert H. Ebert. The medical school. Scientific American, 229 (3):138–148, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n3/pdf/scientificamerican0973-
138.pdf.

Feldstein:1973:ME

[749] Martin S. Feldstein. The medical economy. Scientific American, 229(3):151–159, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n3/pdf/scientificamerican0973-
151.pdf.

Goddard:1973:MB

[750] James L. Goddard. The medical business. Scientific American, 229(3):161–166, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n3/pdf/scientificamerican0973-
161.pdf.

Saward:1973:OMC

[751] Ernest W. Saward. The organization of medical care. Scientific American, 229(3):169–175, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n3/pdf/scientificamerican0973-
169.pdf.

Gardner:1973:MGi

[752] Martin Gardner. Mathematical games. Scientific American, 229 (3):176–183, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n3/pdf/scientificamerican0973-
176.pdf.

Stong:1973:ASi

[753] C. L. Stong. The amateur scientist. Scientific American, 229(3):184–189, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n3/pdf/scientificamerican0973-184.pdf.

Morrison:1973:Bh

[754] Philip Morrison. Books. Scientific American, 229(3):191–196, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n3/pdf/scientificamerican0973-191.pdf.

Anonymous:1973:Bi

[755] Anonymous. Bibliography. Scientific American, 229(3):198–200, September 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n3/pdf/scientificamerican0973-198.pdf.

Anonymous:1973:Lj

[756] Anonymous. Letters. Scientific American, 229(4):8, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v229/
n4/pdf/scientificamerican1073-8.pdf.

Anonymous:1973:YAYj

[757] Anonymous. 50 and 100 years ago: 50 and 100 years ago. Scientific American, 229(4):10–13, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n4/pdf/scientificamerican1073-
10.pdf.

Anonymous:1973:Aj

[758] Anonymous. The authors. Scientific American, 229(4):14–15, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n4/pdf/scientificamerican1073-14.pdf.

Kolm:1973:EF

[759] Henry H. Kolm and Richard D. Thornton. Electromagnetic flight. Scientific American, 229(4):17–25, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v229/n4/pdf/scientificamerican1073-
17.pdf.

Rafferty:1973:HVC

[760] Keen A. Rafferty, Jr. Herpes viruses and cancer. Scientific American, 229(4):26–33, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n4/pdf/scientificamerican1073-
26.pdf.

Korfmann:1973:SW

[761] Manfred Korfmann. The sling as a weapon. Scientific American, 229(4):34–42, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n4/pdf/scientificamerican1073-
34.pdf.

Anonymous:1973:SCj

[762] Anonymous. Science and the citizen. Scientific American, 229(4):47–51, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n4/pdf/scientificamerican1073-47.pdf.

Koshland:1973:PSB

[763] Daniel E. Koshland, Jr. Protein shape and biological control. Scientific American, 229(4):52–64, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n4/pdf/scientificamerican1073-
52.pdf.

Pasachoff:1973:SC

[764] Jay M. Pasachoff. The solar corona. Scientific American, 229(4):68–79, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n4/pdf/scientificamerican1073-68.pdf.

Bjorkman:1973:HEP

[765] Olle Björkman and Joseph Berry. High-efficiency photosynthesis. Scientific American, 229(4):80–93, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v229/n4/pdf/scientificamerican1073-
80.pdf.

Oster:1973:ABB

[766] Gerald Oster. Auditory beats in the brain. Scientific American, 229(4):94–102, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n4/pdf/scientificamerican1073-
94.pdf.

Litke:1973:EC

[767] A. M. Litke and R. Wilson. Electron-positron collisions. Scientific American, 229(4):104–113, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Litke:1973:EPC

[768] Alan M. Litke and Richard Wilson. Electron-positron collisions. Scientific American, 229(4):104–113, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v229/n4/pdf/scientificamerican1073-
104.pdf.

Gardner:1973:MGj

[769] Martin Gardner. Mathematical games. Scientific American, 229 (4):114–118, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n4/pdf/scientificamerican1073-
114.pdf.

Stong:1973:ASj

[770] C. L. Stong. The amateur scientist. Scientific American, 229(4):120–125, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n4/pdf/scientificamerican1073-120.pdf.

Morrison:1973:Bi

[771] Philip Morrison. Books. Scientific American, 229(4):127–131, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n4/pdf/scientificamerican1073-127.pdf.

Anonymous:1973:Bj

[772] Anonymous. Bibliography. Scientific American, 229(4):132, October 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n4/pdf/scientificamerican1073-132.pdf.

Anonymous:1973:Lk

[773] Anonymous. Letters. Scientific American, 229(5):8–9, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v229/
n5/pdf/scientificamerican1173-8.pdf.

Anonymous:1973:YAYk

[774] Anonymous. 50 and 100 years ago: 50 and 100 years ago. Scientific American, 229(5):12–13, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n5/pdf/scientificamerican1173-
12.pdf.

Anonymous:1973:Ak

[775] Anonymous. The authors. Scientific American, 229(5):16–17, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n5/pdf/scientificamerican1173-16.pdf.

York:1973:MWM

[776] Herbert F. York. Multiple-warhead missiles. Scientific American, 229(5):18–27, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n5/pdf/scientificamerican1173-
18.pdf.

Cook:1973:COF

[777] J. S. Cook. Communication by optical fiber. Scientific American, 229(5):28–35, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n5/pdf/scientificamerican1173-
28.pdf.

Amaldi:1973:PIH

[778] Ugo Amaldi. Proton interactions at high energies. Scientific American, 229(5):36–44, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n5/pdf/scientificamerican1173-
36.pdf.

Amaldi:1973:PIA

[779] U. Amaldi. Proton interactions at high energies. Scientific American, 229 (5):36–50, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1973:SCk

[780] Anonymous. Science and the citizen. Scientific American, 229(5):46–50, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n5/pdf/scientificamerican1173-46.pdf.

Mayer:1973:CS

[781] Manfred M. Mayer. The complement system. Scientific American, 229(5):54–66, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n5/pdf/scientificamerican1173-
54.pdf.

Harmon:1973:RF

[782] Leon D. Harmon. The recognition of faces. Scientific American, 229(5):70–82, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n5/pdf/scientificamerican1173-
70.pdf.

Davis:1973:HP

[783] Martin Davis and Reuben Hersh. Hilbert’s 10th problem. Scientific American, 229(5):84–91, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v229/n5/pdf/scientificamerican1173-
84.pdf.

Rothschild:1973:FLF

[784] Miriam Rothschild, Y. Schlein, K. Parker, C. Neville, and S. Sternberg. The flying leap of the flea. Scientific American, 229(5):92–100, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n5/pdf/scientificamerican1173-92.pdf.

Heezen:1973:EP

[785] Bruce C. Heezen and Ian D. MacGregor. The evolution of the Pacific. Scientific American, 229(5):102–112, November 1973. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v229/
n5/pdf/scientificamerican1173-102.pdf.

Gardner:1973:MGFa

[786] M. Gardner. Mathematical games: Fantastic patterns traced by programmed ‘worms’. Scientific American, 229(5):116–123, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). Description of Patterson’s Worm system. Also see [2010].

Gardner:1973:MGk

[787] Martin Gardner. Mathematical games. Scientific American, 229 (5):116–123, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n5/pdf/scientificamerican1173-
116.pdf.

Stong:1973:ASk

[788] C. L. Stong. The amateur scientist. Scientific American, 229(5):124–139, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n5/pdf/scientificamerican1173-124.pdf.

Morrison:1973:Bj

[789] Philip Morrison. Books. Scientific American, 229(5):131–135, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n5/pdf/scientificamerican1173-131.pdf.

Anonymous:1973:Bk

[790] Anonymous. Bibliography. Scientific American, 229(5):136, November 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n5/pdf/scientificamerican1173-136.pdf.

Anonymous:1973:Ll

[791] Anonymous. Letters. Scientific American, 229(6):8, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v229/
n6/pdf/scientificamerican1273-8.pdf.

Anonymous:1973:YAYl

[792] Anonymous. 50 and 100 years ago: 50 and 100 years ago. Scientific American, 229(6):9, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n6/pdf/scientificamerican1273-
9.pdf.

Anonymous:1973:Al

[793] Anonymous. The authors. Scientific American, 229(6):10–14, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n6/pdf/scientificamerican1273-10.pdf.

Post:1973:F

[794] Richard F. Post and Stephen F. Post. Flywheels. Scientific American, 229(6):17–23, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n6/pdf/scientificamerican1273-
17.pdf.

Benzer:1973:GDB

[795] Seymour Benzer. Genetic dissection of behavior. Scientific American, 229(6):24–37, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n6/pdf/scientificamerican1273-
24.pdf.

Alar:1973:VTB

[796] Alar and Juri Toomre. Violent tides between galaxies. Scientific American, 229(6):38–48, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n6/pdf/scientificamerican1273-
38.pdf.

Toomre:1973:VTB

[797] A. Toomre and J. Toomre. Violent tides between galaxies. Scientific American, 229(6):38–48, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1973:SCl

[798] Anonymous. Science and the citizen. Scientific American, 229(6):50–56, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n6/pdf/scientificamerican1273-50.pdf.

Richards:1973:TRF

[799] Paul W. Richards. The tropical rain forest. Scientific American, 229(6):58–67, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n6/pdf/scientificamerican1273-
58.pdf.

Feld:1973:LS

[800] M. S. Feld and V. S. Letokhov. Laser spectroscopy. Scientific American, 229(6):69–72, 78–85, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n6/pdf/scientificamerican1273-
69.pdf.

Gingerich:1973:CT

[801] Owen Gingerich. Copernicus and Tycho. Scientific American, 229(6):86–101, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n6/pdf/scientificamerican1273-
86.pdf.

Pennycuick:1973:SFV

[802] C. J. Pennycuick. The soaring flight of vultures. Scientific American, 229(6):102–109, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n6/pdf/scientificamerican1273-
102.pdf.

Fromkin:1973:ST

[803] Victoria A. Fromkin. Slips of the tongue. Scientific American, 229(6):110–117, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n6/pdf/scientificamerican1273-
110.pdf.

Gardner:1973:MGl

[804] Martin Gardner. Mathematical games. Scientific American, 229 (6):118–121, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v229/n6/pdf/scientificamerican1273-
118.pdf.

Stong:1973:ASl

[805] C. L. Stong. The amateur scientist. Scientific American, 229(6):122–128, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n6/pdf/scientificamerican1273-122.pdf.

Philip:1973:B

[806] Philip and Phylis Morrison. Books. Scientific American, 229(6):131–143, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n6/pdf/scientificamerican1273-131.pdf.

Anonymous:1973:AIa

[807] Anonymous. Annual index. Scientific American, 229(6):144–146, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n6/pdf/scientificamerican1273-144.pdf.

Anonymous:1973:Bl

[808] Anonymous. Bibliography. Scientific American, 229(6):148, December 1973. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v229/n6/pdf/scientificamerican1273-148.pdf.

Anonymous:1974:La

[809] Anonymous. Letters. Scientific American, 230(1):8, January 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n1/pdf/scientificamerican0174-8.pdf.

Anonymous:1974:YAa

[810] Anonymous. 50 and 100 years ago. Scientific American, 230(1):12, January 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n1/pdf/scientificamerican0174-12.pdf.

Anonymous:1974:Aa

[811] Anonymous. The authors. Scientific American, 230(1):16, January 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n1/pdf/scientificamerican0174-16.pdf.

Rose:1974:EPU

[812] David J. Rose. Energy policy in the U.S. Scientific American, 230(1):20– 29, January 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n1/pdf/scientificamerican0174-20.pdf.

Hoare:1974:EM

[813] James P. Hoare and Mitchell A. LaBoda. Electrochemical machining. Scientific American, 230(1):30–37, January 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n1/pdf/scientificamerican0174-30.pdf.

Nicholls:1974:NSL

[814] John G. Nicholls and David Van Essen. The nervous system of the leech. Scientific American, 230(1):38–48, January 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n1/pdf/scientificamerican0174-38.pdf.

Anonymous:1974:SCa

[815] Anonymous. Science and the citizen. Scientific American, 230(1):50–52, January 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n1/pdf/scientificamerican0174-50.pdf.

Mazia:1974:CC

[816] Daniel Mazia. The cell cycle. Scientific American, 230(1):54–64, January 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n1/pdf/scientificamerican0174-54.pdf.

Schramm:1974:AE

[817] David N. Schramm. The age of the elements. Scientific American, 230(1):69–77, January 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n1/pdf/scientificamerican0174-
69.pdf.

Rock:1974:PDF

[818] Irvin Rock. The perception of disoriented figures. Scientific American, 230(1):78–85, January 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n1/pdf/scientificamerican0174-
78.pdf.

Schelleng:1974:PBS

[819] John C. Schelleng. The physics of the bowed string. Scientific American, 230(1):87–95, January 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n1/pdf/scientificamerican0174-
87.pdf.

North:1974:A

[820] J. D. North. The astrolabe. Scientific American, 230(1):96–106, January 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n1/pdf/scientificamerican0174-96.pdf.

Gardner:1974:MGa

[821] Martin Gardner. Mathematical games. Scientific American, 230 (1):108–113, January 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n1/pdf/scientificamerican0174-
108.pdf.

Stong:1974:ASa

[822] C. L. Stong. The amateur scientist. Scientific American, 230(1):114–120, January 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n1/pdf/scientificamerican0174-114.pdf.

Morrison:1974:Ba

[823] Philip Morrison. Books. Scientific American, 230(1):123–129, January 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n1/pdf/scientificamerican0174-123.pdf.

Anonymous:1974:Ba

[824] Anonymous. Bibliography. Scientific American, 230(1):130, January 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n1/pdf/scientificamerican0174-130.pdf.

Anonymous:1974:Lb

[825] Anonymous. Letters. Scientific American, 230(2):6, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n2/pdf/scientificamerican0274-6.pdf.

Anonymous:1974:YAb

[826] Anonymous. 50 and 100 years ago. Scientific American, 230(2):10–11, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n2/pdf/scientificamerican0274-10.pdf.

Anonymous:1974:Ab

[827] Anonymous. The authors. Scientific American, 230(2):12–13, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n2/pdf/scientificamerican0274-12.pdf.

Dovring:1974:S

[828] Folke Dovring. Soybeans. Scientific American, 230(2):14–21, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n2/pdf/scientificamerican0274-14.pdf.

Amelio:1974:CCD

[829] Gilbert F. Amelio. Charge-coupled devices. Scientific American, 230(2):22–31, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n2/pdf/scientificamerican0274-
22.pdf.

Amelio:1974:CD

[830] G. F. Amelio. Charge-coupled devices. Scientific American, 230(2): 22–31, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holland:1974:SIR

[831] John J. Holland. Slow, inapparent and recurrent viruses. Scientific American, 230(2):32–40, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n2/pdf/scientificamerican0274-
32.pdf.

Anonymous:1974:SCb

[832] Anonymous. Science and the citizen. Scientific American, 230(2):42–45, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n2/pdf/scientificamerican0274-42.pdf.

Whipple:1974:NC

[833] Fred L. Whipple. The nature of comets. Scientific American, 230(2):48–57, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n2/pdf/scientificamerican0274-
48.pdf.

Murray:1974:CAM

[834] John M. Murray and Annemarie Weber. The cooperative action of muscle proteins. Scientific American, 230(2):58–71, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n2/pdf/scientificamerican0274-58.pdf.

Wilson:1974:BA

[835] R. R. Wilson. The Batavia Accelerator. Scientific American, 230(2):72–83, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n2/pdf/scientificamerican0274-
72.pdf.

Fernstrom:1974:NB

[836] John D. Fernstrom and Richard J. Wurtman. Nutrition and the brain. Scientific American, 230(2):84–91, February 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n2/pdf/scientificamerican0274-84.pdf.

Condit:1974:WBB

[837] Carl W. Condit. The wind bracing of buildings. Scientific American, 230(2):92–105, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n2/pdf/scientificamerican0274-
92.pdf.

Gardner:1974:MGb

[838] Martin Gardner. Mathematical games. Scientific American, 230 (2):106–108, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n2/pdf/scientificamerican0274-
106.pdf.

Stong:1974:ASb

[839] C. L. Stong. The amateur scientist. Scientific American, 230(2):110–115, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n2/pdf/scientificamerican0274-110.pdf.

Morrison:1974:Bb

[840] Philip Morrison. Books. Scientific American, 230(2):117–121, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n2/pdf/scientificamerican0274-117.pdf.

Anonymous:1974:Bb

[841] Anonymous. Bibliography. Scientific American, 230(2):122, February 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n2/pdf/scientificamerican0274-122.pdf.

Anonymous:1974:Lc

[842] Anonymous. Letters. Scientific American, 230(3):8–10, March 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n3/pdf/scientificamerican0374-8.pdf.

Anonymous:1974:YAc

[843] Anonymous. 50 and 100 years ago. Scientific American, 230(3):12, March 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n3/pdf/scientificamerican0374-12.pdf.

Anonymous:1974:Ac

[844] Anonymous. The authors. Scientific American, 230(3):16, March 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n3/pdf/scientificamerican0374-16.pdf.

Perry:1974:GC

[845] Harry Perry. The gasification of coal. Scientific American, 230(3):19–25, March 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n3/pdf/scientificamerican0374-19.pdf.

Capaldi:1974:DMC

[846] Roderick A. Capaldi. A dynamic model of cell membranes. Scientific American, 230(3):26–33, March 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n3/pdf/scientificamerican0374-
26.pdf.

Ewert:1974:NBV

[847] Jörg-Peter Ewert. The neural basis of visually guided behavior. Scientific American, 230(3):34–42, March 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n3/pdf/scientificamerican0374-
34.pdf.

Anonymous:1974:SCc

[848] Anonymous. Science and the citizen. Scientific American, 230(3):44–46, March 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n3/pdf/scientificamerican0374-44.pdf.

Lewis:1974:CSS

[849] John S. Lewis. The chemistry of the solar system. Scientific American, 230(3):50–60, 65, March 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n3/pdf/scientificamerican0374-
50.pdf.

Allcock:1974:IP

[850] Harry R. Allcock. Inorganic polymers. Scientific American, 230(3):66– 74, March 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n3/pdf/scientificamerican0374-66.pdf.

Chigier:1974:VAW

[851] Norman A. Chigier. Vortexes in aircraft wakes. Scientific American, 230(3):76–83, March 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n3/pdf/scientificamerican0374-
76.pdf.

Ostwald:1974:CHI

[852] Peter F. Ostwald and Philip Peltzman. The cry of the human infant. Scientific American, 230(3):84–90, March 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n3/pdf/scientificamerican0374-84.pdf.

Shiers:1974:FBC

[853] George Shiers. Ferdinand Braun and the cathode ray tube. Scientific American, 230(3):92–101, March 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n3/pdf/scientificamerican0374-
92.pdf.

Gardner:1974:MGc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n3/pdf/scientificamerican0374-102.pdf.

Stong:1974:ASc

[855] C. L. Stong. The amateur scientist. Scientific American, 230(3):110–115, March 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n3/pdf/scientificamerican0374-110.pdf.

Morrison:1974:Bc

[856] Philip Morrison. Books. Scientific American, 230(3):117–123, March 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n3/pdf/scientificamerican0374-117.pdf.

Anonymous:1974:Bc

[857] Anonymous. Bibliography. Scientific American, 230(3):124, March 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n3/pdf/scientificamerican0374-124.pdf.

Anonymous:1974:Ld

[858] Anonymous. Letters. Scientific American, 230(4):9–11, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n4/pdf/scientificamerican0474-9.pdf.

Anonymous:1974:YAd

[859] Anonymous. 50 and 100 years ago. Scientific American, 230(4):14, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n4/pdf/scientificamerican0474-14.pdf.

Anonymous:1974:Ad

[860] Anonymous. The authors. Scientific American, 230(4):16, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n4/pdf/scientificamerican0474-16.pdf.

Sidel:1974:DMC

[861] Victor W. Sidel and Ruth Sidel. The delivery of medical care in China. Scientific American, 230(4):19–27, April 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n4/pdf/scientificamerican0474-19.pdf.

Tien:1974:IO

[862] P. K. Tien. Integrated optics. Scientific American, 230(4):28–35, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n4/pdf/scientificamerican0474-28.pdf.

Beer:1974:ET

[863] Alan E. Beer and R. E. Billingham. The embryo as a transplant. Scientific American, 230(4):36–47, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n4/pdf/scientificamerican0474-
36.pdf.

Anonymous:1974:SCd

[864] Anonymous. Science and the citizen. Scientific American, 230(4):48–51, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n4/pdf/scientificamerican0474-48.pdf.

Hall:1974:WP

[865] F. Keith Hall. Wood pulp. Scientific American, 230(4):52–62, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n4/pdf/scientificamerican0474-52.pdf.

Sanders:1974:CG

[866] R. H. Sanders and G. T. Wrixon. The center of the galaxy. Scientific American, 230(4):66–77, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n4/pdf/scientificamerican0474-
66.pdf.

Valentine:1974:PTH

[867] James W. Valentine and Eldridge M. Moores. Plate tectonics and the history of life in the oceans. Scientific American, 230(4):80–89, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n4/pdf/scientificamerican0474-80.pdf.

Metelli:1974:PT

[868] Fabio Metelli. The perception of transparency. Scientific American, 230(4):90–98, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n4/pdf/scientificamerican0474-
90.pdf.

Waterhouse:1974:BCD

[869] D. F. Waterhouse. The biological control of dung. Scientific American, 230(4):100–109, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n4/pdf/scientificamerican0474-
100.pdf.

Gardner:1974:MGd

[870] Martin Gardner. Mathematical games. Scientific American, 230(4):110– 114, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n4/pdf/scientificamerican0474-110.pdf.

Stong:1974:ASd

[871] C. L. Stong. The amateur scientist. Scientific American, 230(4):116– 121, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n4/pdf/scientificamerican0474-116.pdf.

Morrison:1974:Bd

[872] Philip Morrison. Books. Scientific American, 230(4):123–127, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n4/pdf/scientificamerican0474-123.pdf.

Anonymous:1974:Bd

[873] Anonymous. Bibliography. Scientific American, 230(4):128, April 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n4/pdf/scientificamerican0474-128.pdf.

Anonymous:1974:Le

[874] Anonymous. Letters. Scientific American, 230(5):8, May 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n5/pdf/scientificamerican0574-8.pdf.

Anonymous:1974:YAe

[875] Anonymous. 50 and 100 years ago. Scientific American, 230(5):12– 13, May 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n5/pdf/scientificamerican0574-12.pdf.

Anonymous:1974:Ae

[876] Anonymous. The authors. Scientific American, 230(5):16–17, May 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n5/pdf/scientificamerican0574-16.pdf.

Carter:1974:NSN

[877] Barry Carter. Nuclear strategy and nuclear weapons. Scientific American, 230(5):20–31, May 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n5/pdf/scientificamerican0574-
20.pdf.

Biddle:1974:AW

[878] Martin Biddle. The archaeology of Winchester. Scientific American, 230(5):32–43, May 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n5/pdf/scientificamerican0574-
32.pdf.

Guillery:1974:VPA

[879] R. W. Guillery. Visual pathways in albinos. Scientific American, 230(5): 44–54, May 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n5/pdf/scientificamerican0574-44.pdf.

Anonymous:1974:SCe

[880] Anonymous. Science and the citizen. Scientific American, 230(5):59– 61, May 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n5/pdf/scientificamerican0574-59.pdf.

MacIntyre:1974:TMO

[881] Ferren MacIntyre. The top millimeter of the ocean. Scientific American, 230(5):62–77, May 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n5/pdf/scientificamerican0574-
62.pdf.

Sharon:1974:G

[882] Nathan Sharon. Glycoproteins. Scientific American, 230(5):78–86, May 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n5/pdf/scientificamerican0574-78.pdf.

Scott:1974:PW

[883] Douglas Scott, William W. Seifert, and Vernon C. Westcott. The particles of wear. Scientific American, 230(5):88–97, May 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n5/pdf/scientificamerican0574-88.pdf.

Greenberg:1974:CGA

[884] Donald P. Greenberg. Computer graphics in architecture. Scientific American, 230(5):98–106, May 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n5/pdf/scientificamerican0574-
98.pdf.

Pasachoff:1974:DU

[885] Jay M. Pasachoff and William A. Fowler. Deuterium in the universe. Scientific American, 230(5):108–119, May 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n5/pdf/scientificamerican0574-108.pdf.

Gardner:1974:MGe

[886] Martin Gardner. Mathematical games. Scientific American, 230(5):120– 125, May 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n5/pdf/scientificamerican0574-120.pdf.

Stong:1974:ASe

[887] C. L. Stong. The amateur scientist. Scientific American, 230(5):126– 131, May 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n5/pdf/scientificamerican0574-126.pdf.

Morrison:1974:Be

[888] Philip Morrison. Books. Scientific American, 230(5):133–143, May 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n5/pdf/scientificamerican0574-133.pdf.

Anonymous:1974:Be

[889] Anonymous. Bibliography. Scientific American, 230(5):144, May 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n5/pdf/scientificamerican0574-144.pdf.

Anonymous:1974:Lf

[890] Anonymous. Letters. Scientific American, 230(6):8, June 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n6/pdf/scientificamerican0674-8.pdf.

Anonymous:1974:YAf

[891] Anonymous. 50 and 100 years ago. Scientific American, 230(6):12, June 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n6/pdf/scientificamerican0674-12.pdf.

Anonymous:1974:Af

[892] Anonymous. The authors. Scientific American, 230(6):16–17, June 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n6/pdf/scientificamerican0674-16.pdf.

Bazelon:1974:PAP

[893] David L. Bazelon. Psychiatrists and the adversary process. Scientific American, 230(6):18–23, June 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n6/pdf/scientificamerican0674-
18.pdf.

Emmett:1974:FPL

[894] John L. Emmett, John Nuckolls, and Lowell Wood. Fusion power by laser implosion. Scientific American, 230(6):24–37, June 1974.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n6/pdf/scientificamerican0674-24.pdf.

Myers:1974:PCR

[895] Judith H. Myers and Charles J. Krebs. Population cycles in rodents. Scientific American, 230(6):38–46, June 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n6/pdf/scientificamerican0674-38.pdf.

Anonymous:1974:SCf

[896] Anonymous. Science and the citizen. Scientific American, 230(6):48–56, June 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n6/pdf/scientificamerican0674-48.pdf.

Axelrod:1974:N

[897] Julius Axelrod. Neurotransmitters. Scientific American, 230(6):58–71, June 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n6/pdf/scientificamerican0674-58.pdf.

Siever:1974:SSE

[898] Raymond Siever. The steady state of the Earth’s crust, atmosphere and oceans. Scientific American, 230(6):72–79, June 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n6/pdf/scientificamerican0674-72.pdf.

Winfree:1974:RCR

[899] Arthur T. Winfree. Rotating chemical reactions. Scientific American, 230(6):82–95, June 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n6/pdf/scientificamerican0674-
82.pdf.

Klein:1974:IAH

[900] Richard G. Klein. Ice-Age hunters of the Ukraine. Scientific American, 230(6):96–105, June 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n6/pdf/scientificamerican0674-
96.pdf.

Wagner:1974:WGV

[901] Philip Wagner. Wines, grape vines and climate. Scientific American, 230(6):106–115, June 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v230/n6/pdf/scientificamerican0674-
106.pdf.

Gardner:1974:MGf

[902] Martin Gardner. Mathematical games. Scientific American, 230(6):116– 121, June 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n6/pdf/scientificamerican0674-116.pdf.

Stong:1974:ASf

[903] C. L. Stong. The amateur scientist. Scientific American, 230(6):122– 127, June 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v230/n6/pdf/scientificamerican0674-122.pdf.

Morrison:1974:Bf

[904] Philip Morrison. Books. Scientific American, 230(6):129–135, June 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n6/pdf/scientificamerican0674-129.pdf.

Anonymous:1974:Bf

[905] Anonymous. Bibliography. Scientific American, 230(6):136, June 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v230/
n6/pdf/scientificamerican0674-136.pdf.

Anonymous:1974:Lg

[906] Anonymous. Letters. Scientific American, 231(1):10, July 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n1/pdf/scientificamerican0774-10.pdf.

Anonymous:1974:YAg

[907] Anonymous. 50 and 100 years ago. Scientific American, 231(1):12, July 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n1/pdf/scientificamerican0774-12.pdf.

Anonymous:1974:Ag

[908] Anonymous. The authors. Scientific American, 231(1):16, July 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n1/pdf/scientificamerican0774-16.pdf.

Emmons:1974:FFP

[909] Howard W. Emmons. Fire and fire protection. Scientific American, 231(1):21–27, July 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n1/pdf/scientificamerican0774-
21.pdf.

Schneider:1974:SAR

[910] Dietrich Schneider. The sex-attractant receptor of moths. Scientific American, 231(1):28–35, July 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n1/pdf/scientificamerican0774-
28.pdf.

Ruddle:1974:HCH

[911] Frank H. Ruddle and Raju S. Kucherlapati. Hybrid cells and human genes. Scientific American, 231(1):36–44, July 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n1/pdf/scientificamerican0774-36.pdf.

Anonymous:1974:SCg

[912] Anonymous. Science and the citizen. Scientific American, 231(1):46– 49, July 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n1/pdf/scientificamerican0774-46.pdf.

Weinberg:1974:UTE

[913] Steven Weinberg. Unified theories of elementary-particle interaction. Scientific American, 231(1):50–59, July 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v231/n1/pdf/scientificamerican0774-
50.pdf.

Bryant:1974:G

[914] Howard C. Bryant and Nelson Jarmie. The glory. Scientific American, 231(1):60–67, 70–71, July 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n1/pdf/scientificamerican0774-
60.pdf.

Stroud:1974:FPC

[915] Robert M. Stroud. A family of protein-cutting proteins. Scientific American, 231(1):74–88, July 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n1/pdf/scientificamerican0774-
74.pdf.

Teuber:1974:SAP

[916] Marianne L. Teuber. Sources of ambiguity in the prints of Maurits C. Escher. Scientific American, 231(1):90–104, July 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n1/pdf/scientificamerican0774-90.pdf.

Holloway:1974:CFH

[917] Ralph L. Holloway. The casts of fossil hominid brains. Scientific American, 231(1):106–115, July 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n1/pdf/scientificamerican0774-
106.pdf.

Gardner:1974:MGg

[918] Martin Gardner. Mathematical games. Scientific American, 231(1):116– 120, July 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n1/pdf/scientificamerican0774-116.pdf.

Stong:1974:ASg

[919] C. L. Stong. The amateur scientist. Scientific American, 231(1):122– 126, July 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n1/pdf/scientificamerican0774-122.pdf.

Morrison:1974:Bg

[920] Philip Morrison. Books. Scientific American, 231(1):129–135, July 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n1/pdf/scientificamerican0774-129.pdf.

Anonymous:1974:Bg

[921] Anonymous. Bibliography. Scientific American, 231(1):136, July 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n1/pdf/scientificamerican0774-136.pdf.

Anonymous:1974:Lh

[922] Anonymous. Letters. Scientific American, 231(2):10–11, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n2/pdf/scientificamerican0874-10.pdf.

Anonymous:1974:YAh

[923] Anonymous. 50 and 100 years ago. Scientific American, 231(2):12, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n2/pdf/scientificamerican0874-12.pdf.

Anonymous:1974:Ah

[924] Anonymous. The authors. Scientific American, 231(2):14, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n2/pdf/scientificamerican0874-14.pdf.

Bascom:1974:DWO

[925] Willard Bascom. The disposal of waste in the ocean. Scientific American, 231(2):16–25, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n2/pdf/scientificamerican0874-
16.pdf.

Webster:1974:CBR

[926] Adrian Webster. The cosmic background radiation. Scientific American, 231(2):26–33, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n2/pdf/scientificamerican0874-
26.pdf.

Bentley:1974:NCS

[927] David Bentley and Ronald R. Hoy. The neurobiology of Cricket song. Scientific American, 231(2):34–44, August 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n2/pdf/scientificamerican0874-34.pdf.

Anonymous:1974:SCh

[928] Anonymous. Science and the citizen. Scientific American, 231(2):46–50, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n2/pdf/scientificamerican0874-46.pdf.

Bronfenbrenner:1974:OA

[929] Urie Bronfenbrenner. The origins of alienation. Scientific American, 231(2):53–61, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n2/pdf/scientificamerican0874-
53.pdf.

Bundy:1974:SM

[930] Francis P. Bundy. Superhard materials. Scientific American, 231(2):62– 70, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n2/pdf/scientificamerican0874-62.pdf.

Bundy:1974:SMC

[931] F. P. Bundy. Superhard materials (crystal binding). Scientific American, 231(2):62–70, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hulse:1974:T

[932] Joseph H. Hulse and David Spurgeon. Triticale. Scientific American, 231(2):72–80, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n2/pdf/scientificamerican0874-
72.pdf.

Sobell:1974:HAB

[933] Henry M. Sobell. How actinomycin binds to DNA. Scientific American, 231(2):82–90, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n2/pdf/scientificamerican0874-
82.pdf.

Harris:1974:RCT

[934] John R. Harris. The rise of coal technology. Scientific American, 231(2):92–97, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n2/pdf/scientificamerican0874-
92.pdf.

Gardner:1974:MGh

[935] Martin Gardner. Mathematical games. Scientific American, 231 (2):98–103, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n2/pdf/scientificamerican0874-
98.pdf.

Stong:1974:ASh

[936] C. L. Stong. The amateur scientist. Scientific American, 231(2):104–109, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n2/pdf/scientificamerican0874-104.pdf.

Morrison:1974:Bh

[937] Philip Morrison. Books. Scientific American, 231(2):111–115, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n2/pdf/scientificamerican0874-111.pdf.

Anonymous:1974:Bh

[938] Anonymous. Bibliography. Scientific American, 231(2):116, August 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n2/pdf/scientificamerican0874-116.pdf.

Anonymous:1974:Li

[939] Anonymous. Letters. Scientific American, 231(3):8–12, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n3/pdf/scientificamerican0974-8.pdf.

Anonymous:1974:YAi

[940] Anonymous. 50 and 100 years ago. Scientific American, 231(3):16–20, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n3/pdf/scientificamerican0974-16.pdf.

Anonymous:1974:Ai

[941] Anonymous. The authors. Scientific American, 231(3):22–26, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n3/pdf/scientificamerican0974-22.pdf.

Freedman:1974:HP

[942] Ronald Freedman and Bernard Berelson. The human population. Scientific American, 231(3):30–39, September 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n3/pdf/scientificamerican0974-30.pdf.

Coale:1974:HHP

[943] Ansley J. Coale. The history of the human population. Scientific American, 231(3):40–51, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n3/pdf/scientificamerican0974-
40.pdf.

Segal:1974:PHR

[944] Sheldon J. Segal. The physiology of human reproduction. Scientific American, 231(3):52–62, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v231/n3/pdf/scientificamerican0974-
52.pdf.

Anonymous:1974:SCi

[945] Anonymous. Science and the citizen. Scientific American, 231(3):64–76, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n3/pdf/scientificamerican0974-64.pdf.

Cavalli-Sforza:1974:GHP

[946] L. L. Cavalli-Sforza. The genetics of human populations. Scientific American, 231(3):80–89, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n3/pdf/scientificamerican0974-
80.pdf.

Davis:1974:MHP

[947] Kingsley Davis. The migrations of human populations. Scientific American, 231(3):92–105, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n3/pdf/scientificamerican0974-
92.pdf.

Westoff:1974:PDC

[948] Charles F. Westoff. The populations of the developed countries. Scientific American, 231(3):108–120, September 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n3/pdf/scientificamerican0974-108.pdf.

Ryder:1974:FDC

[949] Norman B. Ryder. The family in developed countries. Scientific American, 231(3):122–132, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n3/pdf/scientificamerican0974-
122.pdf.

Blake:1974:CSW

[950] Judith Blake. The changing status of women in developed countries. Scientific American, 231(3):136–147, September 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n3/pdf/scientificamerican0974-136.pdf.

Demeny:1974:PUC

[951] Paul Demeny. The populations of the underdeveloped countries. Scientific American, 231(3):148–159, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v231/n3/pdf/scientificamerican0974-
148.pdf.

Revelle:1974:FP

[952] Roger Revelle. Food and population. Scientific American, 231 (3):160–170, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n3/pdf/scientificamerican0974-
160.pdf.

Myrdal:1974:TTU

[953] Gunnar Myrdal. The transfer of technology to underdeveloped countries. Scientific American, 231(3):172–182, September 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n3/pdf/scientificamerican0974-172.pdf.

Gardner:1974:MGi

[954] Martin Gardner and W. W. Rouse Ball. Mathematical games. Scientific American, 231(3):187–191, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v231/n3/pdf/scientificamerican0974-
187.pdf.

Stong:1974:ASi

[955] C. L. Stong. The amateur scientist. Scientific American, 231(3):192–198, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n3/pdf/scientificamerican0974-192.pdf.

Morrison:1974:Bi

[956] Philip Morrison. Books. Scientific American, 231(3):201–206, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n3/pdf/scientificamerican0974-201.pdf.

Anonymous:1974:Bi

[957] Anonymous. Bibliography. Scientific American, 231(3):210, September 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n3/pdf/scientificamerican0974-210.pdf.

Anonymous:1974:Lj

[958] Anonymous. Letters. Scientific American, 231(4):6–10, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n4/pdf/scientificamerican1074-6.pdf.

Anonymous:1974:YAj

[959] Anonymous. 50 and 100 years ago. Scientific American, 231(4):13, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n4/pdf/scientificamerican1074-13.pdf.

Anonymous:1974:Aj

[960] Anonymous. The authors. Scientific American, 231(4):14–16, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n4/pdf/scientificamerican1074-14.pdf.

Myrdal:1974:ICD

[961] Alva Myrdal. The international control of disarmament. Scientific American, 231(4):21–33, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n4/pdf/scientificamerican1074-
21.pdf.

Miller:1974:SEN

[962] Joseph S. Miller. The structure of emission nebulas. Scientific American, 231(4):34–43, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n4/pdf/scientificamerican1074-
34.pdf.

Satir:1974:HCM

[963] Peter Satir. How cilia move. Scientific American, 231(4):44–52, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n4/pdf/scientificamerican1074-44.pdf.

Anonymous:1974:SCj

[964] Anonymous. Science and the citizen. Scientific American, 231(4):55–63, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n4/pdf/scientificamerican1074-55.pdf.

Safrany:1974:NF

[965] David R. Safrany. Nitrogen fixation. Scientific American, 231(4):64–80, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n4/pdf/scientificamerican1074-64.pdf.

Fitch:1974:DS

[966] James Marston Fitch, John Templer, and Paul Corcoran. The dimensions of stairs. Scientific American, 231(4):82–90, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n4/pdf/scientificamerican1074-82.pdf.

Nicolai:1974:MPB

[967] Jürgen Nicolai. Mimicry in parasitic birds. Scientific American, 231(4):92–98, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n4/pdf/scientificamerican1074-
92.pdf.

Bizzi:1974:CEH

[968] Emilio Bizzi. The coordination of eye-head movements. Scientific American, 231(4):100–106, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n4/pdf/scientificamerican1074-
100.pdf.

Jameson:1974:EDG

[969] Michael H. Jameson. The excavation of a drowned Greek Temple. Scientific American, 231(4):110–119, October 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n4/pdf/scientificamerican1074-110.pdf.

Gardner:1974:MGj

[970] Martin Gardner. Mathematical games. Scientific American, 231 (4):120–125, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n4/pdf/scientificamerican1074-
120.pdf.

Stong:1974:ASj

[971] C. L. Stong. The amateur scientist. Scientific American, 231(4):126–130, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n4/pdf/scientificamerican1074-126.pdf.

Morrison:1974:Bj

[972] Philip Morrison. Books. Scientific American, 231(4):135–141, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n4/pdf/scientificamerican1074-135.pdf.

Anonymous:1974:Bj

[973] Anonymous. Bibliography. Scientific American, 231(4):142, October 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n4/pdf/scientificamerican1074-142.pdf.

Anonymous:1974:Lk

[974] Anonymous. Letters. Scientific American, 231(5):8–9, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n5/pdf/scientificamerican1174-8.pdf.

Anonymous:1974:YAk

[975] Anonymous. 50 and 100 years ago. Scientific American, 231(5):10, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n5/pdf/scientificamerican1174-10.pdf.

Anonymous:1974:Ak

[976] Anonymous. The authors. Scientific American, 231(5):12–14, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n5/pdf/scientificamerican1174-12.pdf.

Bok:1974:EGP

[977] Sissela Bok. The ethics of giving placebos. Scientific American, 231(5):17–23, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n5/pdf/scientificamerican1174-
17.pdf.

Will:1974:GT

[978] Clifford M. Will. Gravitation theory. Scientific American, 231(5):24–33, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n5/pdf/scientificamerican1174-24.pdf.

Glavitsch:1974:CCE

[979] Hans Glavitsch. Computer control of electric-power systems. Scientific American, 231(5):34–44, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v231/n5/pdf/scientificamerican1174-
34.pdf.

Anonymous:1974:SCk

[980] Anonymous. Science and the citizen. Scientific American, 231(5):49–54, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n5/pdf/scientificamerican1174-49.pdf.

Cooper:1974:DIS

[981] Max D. Cooper and Alexander R. Lawton III. The development of the immune system. Scientific American, 231(5):58–72, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n5/pdf/scientificamerican1174-58.pdf.

Patterson:1974:MD

[982] Blake Patterson. Musical dynamics. Scientific American, 231(5):78–95, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n5/pdf/scientificamerican1174-78.pdf.

Warren:1974:PG

[983] James V. Warren. The physiology of the giraffe. Scientific American, 231(5):96–105, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n5/pdf/scientificamerican1174-
96.pdf.

Campbell:1974:CSF

[984] Fergus W. Campbell and Lamberto Maffei. Contrast and spatial frequency. Scientific American, 231(5):106–114, November 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n5/pdf/scientificamerican1174-106.pdf.

Vanek:1974:TSH

[985] Joann Vanek. Time spent in housework. Scientific American, 231(5):116–120, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n5/pdf/scientificamerican1174-
116.pdf.

Gardner:1974:MGk

[986] Martin Gardner. Mathematical games. Scientific American, 231 (5):122–124, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n5/pdf/scientificamerican1174-
122.pdf.

Stong:1974:ASk

[987] C. L. Stong. The amateur scientist. Scientific American, 231(5):126–134, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n5/pdf/scientificamerican1174-126.pdf.

Morrison:1974:Bk

[988] Philip Morrison. Books. Scientific American, 231(5):137–145, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n5/pdf/scientificamerican1174-137.pdf.

Anonymous:1974:Bk

[989] Anonymous. Bibliography. Scientific American, 231(5):146, November 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n5/pdf/scientificamerican1174-146.pdf.

Anonymous:1974:Ll

[990] Anonymous. Letters. Scientific American, 231(6):9–11, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n6/pdf/scientificamerican1274-9.pdf.

Anonymous:1974:YAl

[991] Anonymous. 50 and 100 years ago. Scientific American, 231(6):14, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n6/pdf/scientificamerican1274-14.pdf.

Anonymous:1974:Al

[992] Anonymous. The authors. Scientific American, 231(6):19, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n6/pdf/scientificamerican1274-19.pdf.

Buckhout:1974:ET

[993] Robert Buckhout. Eyewitness testimony. Scientific American, 231(6):23–31, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n6/pdf/scientificamerican1274-
23.pdf.

Thorne:1974:SBH

[994] Kip S. Thorne. The search for black holes. Scientific American, 231(6):32–43, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n6/pdf/scientificamerican1274-
32.pdf.

Gierer:1974:HMD

[995] Alfred Gierer. Hydra as a model for the development of biological form. Scientific American, 231(6):44–54, December 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n6/pdf/scientificamerican1274-44.pdf.

Anonymous:1974:SCl

[996] Anonymous. Science and the citizen. Scientific American, 231(6):56–66, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n6/pdf/scientificamerican1274-56.pdf.

Govindjee:1974:ALP

[997] Govindjee and Rajni Govindjee. The absorption of light in photosynthesis. Scientific American, 231(6):68–82, December 1974. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v231/
n6/pdf/scientificamerican1274-68.pdf.

Flemings:1974:SC

[998] Merton C. Flemings. The solidification of castings. Scientific American, 231(6):88–95, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n6/pdf/scientificamerican1274-
88.pdf.

Keeton:1974:MPH

[999] William T. Keeton. The mystery of pigeon homing. Scientific American, 231(6):96–107, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n6/pdf/scientificamerican1274-
96.pdf.

Cline:1974:DNW

[1000] David B. Cline, Alfred K. Mann, and Carlo Rubbia. The detection of neutral weak currents. Scientific American, 231(6):108–109, 112–119, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n6/pdf/scientificamerican1274-108.pdf.

Boon:1974:CRB

[1001] George C. Boon. Counterfeiting in Roman Britain. Scientific American, 231(6):120–130, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n6/pdf/scientificamerican1274-
120.pdf.

Gardner:1974:MGl

[1002] Martin Gardner. Mathematical games. Scientific American, 231 (6):132–136, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v231/n6/pdf/scientificamerican1274-
132.pdf.

Stong:1974:ASl

[1003] C. L. Stong. The amateur scientist. Scientific American, 231(6):138–143, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n6/pdf/scientificamerican1274-138.pdf.

Philip:1974:B

[1004] Philip and Phylis Morrison. Books. Scientific American, 231(6):144–161, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n6/pdf/scientificamerican1274-144.pdf.

Anonymous:1974:AIa

[1005] Anonymous. Annual index. Scientific American, 231(6):162–165, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n6/pdf/scientificamerican1274-162.pdf.

Anonymous:1974:Bl

[1006] Anonymous. Bibliography. Scientific American, 231(6):166–167, December 1974. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v231/n6/pdf/scientificamerican1274-166.pdf.

Anonymous:1975:La

[1007] Anonymous. Letters. Scientific American, 232(1):8–11, January 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n1/pdf/scientificamerican0175-8.pdf.

Anonymous:1975:YAa

[1008] Anonymous. 50 and 100 years ago. Scientific American, 232(1):12, January 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n1/pdf/scientificamerican0175-12.pdf.

Anonymous:1975:Aa

[1009] Anonymous. The authors. Scientific American, 232(1):14–15, January 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n1/pdf/scientificamerican0175-14B.pdf.

Moore:1975:AEI

[1010] Geoffrey H. Moore. The analysis of economic indicators. Scientific American, 232(1):17–23, January 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n1/pdf/scientificamerican0175-
17.pdf.

Chapman:1975:NA

[1011] Clark R. Chapman. The nature of asteroids. Scientific American, 232(1):24–33, January 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n1/pdf/scientificamerican0175-
24.pdf.

Pierce:1975:FCA

[1012] John R. Pierce. The fuel consumption of automobiles. Scientific American, 232(1):34–44, January 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n1/pdf/scientificamerican0175-
34.pdf.

Anonymous:1975:SCa

[1013] Anonymous. Science and the citizen. Scientific American, 232(1):48–55, January 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n1/pdf/scientificamerican0175-48.pdf.

Llin000E1s:1975:CC

[1014] Rodolfo R. Llinás. The cortex of the cerebellum. Scientific American, 232(1):56–71, January 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n1/pdf/scientificamerican0175-
56.pdf.

Lemmon:1975:HER

[1015] Richard M. Lemmon and Wallace R. Erwin. High-energy reactions of carbon. Scientific American, 232(1):72–79, January 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n1/pdf/scientificamerican0175-72.pdf.

Strobel:1975:MDR

[1016] Gary A. Strobel. A mechanism of disease resistance in plants. Scientific American, 232(1):80–88, January 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n1/pdf/scientificamerican0175-
80.pdf.

Bishop:1975:MMC

[1017] J. A. Bishop and Laurence M. Cook. Moths, melanism and clean air. Scientific American, 232(1):90–99, January 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n1/pdf/scientificamerican0175-90.pdf.

Bryant:1975:CM

[1018] Vaughn M. Bryant, Jr. and Glenna Williams-Dean. The coprolites of Man. Scientific American, 232(1):100–109, January 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n1/pdf/scientificamerican0175-100.pdf.

Gardner:1975:MGa

[1019] Martin Gardner. Mathematical games. Scientific American, 232 (1):110–117, January 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n1/pdf/scientificamerican0175-
110.pdf.

Stong:1975:ASa

[1020] C. L. Stong. The amateur scientist. Scientific American, 232(1):118–124, January 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n1/pdf/scientificamerican0175-118.pdf.

Morrison:1975:Ba

[1021] Philip Morrison. Books. Scientific American, 232(1):126–133, January 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n1/pdf/scientificamerican0175-126.pdf.

Anonymous:1975:Ba

[1022] Anonymous. Bibliography. Scientific American, 232(1):134–136, January 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n1/pdf/scientificamerican0175-134.pdf.

Anonymous:1975:Lb

[1023] Anonymous. Letters. Scientific American, 232(2):6–7, February 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n2/pdf/scientificamerican0275-6.pdf.

Anonymous:1975:YAb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n2/pdf/scientificamerican0275-8.pdf.

Anonymous:1975:Ab

[1025] Anonymous. The authors. Scientific American, 232(2):12, February 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n2/pdf/scientificamerican0275-12.pdf.

Mick:1975:FMG

[1026] Stephen S. Mick. The foreign medical graduate. Scientific American, 232(2):14–21, February 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n2/pdf/scientificamerican0275-
14.pdf.

Cook:1975:CMP

[1027] Nathan H. Cook. Computer-managed parts manufacture. Scientific American, 232(2):22–29, February 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n2/pdf/scientificamerican0275-
22.pdf.

Cook:1975:CPM

[1028] N. H. Cook. Computer-managed parts manufacture. Scientific American, 232(2):22–29, February 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Grossman:1975:MPO

[1029] Lawrence Grossman. The most primitive objects in the solar system. Scientific American, 232(2):30–39, February 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n2/pdf/scientificamerican0275-30.pdf.

Anonymous:1975:SCb

[1030] Anonymous. Science and the citizen. Scientific American, 232(2):40–43, February 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n2/pdf/scientificamerican0275-40.pdf.

Stein:1975:CPG

[1031] Gary S. Stein, Janet Swinehart Stein, and Lewis J. Kleinsmith. Chromosomal proteins and gene regulation. Scientific American, 232(2):46–60, February 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n2/pdf/scientificamerican0275-46.pdf.

Schwarz:1975:DME

[1032] J. H. Schwarz. Dual-resonance models of elementary particles. Scientific American, 232(2):61–67, February 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schwarz:1975:DRM

[1033] John H. Schwarz. Dual-resonance models of elementary particles. Scientific American, 232(2):61–69, February 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n2/pdf/scientificamerican0275-61.pdf.

Palmer:1975:BCT

[1034] John D. Palmer. Biological clocks of the tidal zone. Scientific American, 232(2):70–79, February 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n2/pdf/scientificamerican0275-
70.pdf.

Moscati:1975:CFS

[1035] Sabatino Moscati. A Carthaginian fortress in Sardinia. Scientific American, 232(2):80–87, February 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n2/pdf/scientificamerican0275-
80.pdf.

Hallam:1975:AWH

[1036] A. Hallam. Alfred Wegener and the hypothesis of continental drift. Scientific American, 232(2):88–97, February 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n2/pdf/scientificamerican0275-88.pdf.

Gardner:1975:MGM

[1037] Martin Gardner. Mathemtical games: Mathematical games. Scientific American, 232(2):98–103, February 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v232/n2/pdf/scientificamerican0275-
98.pdf.

Stong:1975:ASb

[1038] C. L. Stong. The amateur scientist. Scientific American, 232(2):104–108, February 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n2/pdf/scientificamerican0275-104.pdf.

Morrison:1975:Bb

[1039] Philip Morrison. Books. Scientific American, 232(2):109–113, February 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n2/pdf/scientificamerican0275-109.pdf.

Anonymous:1975:Bb

[1040] Anonymous. Bibliography. Scientific American, 232(2):114–115, February 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n2/pdf/scientificamerican0275-114.pdf.

Anonymous:1975:Lc

[1041] Anonymous. Letters. Scientific American, 232(3):6–7, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n3/pdf/scientificamerican0375-6.pdf.

Anonymous:1975:YAc

[1042] Anonymous. 50 and 100 years ago. Scientific American, 232(3):8–11, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n3/pdf/scientificamerican0375-8.pdf.

Anonymous:1975:A

[1043] Anonymous. The authors. Scientific American, 232(3):12–14, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n3/pdf/scientificamerican0375-12.pdf.

Bjork:1975:EWS

[1044] Lars E. Björk. An experiment in work satisfaction. Scientific American, 232(3):17–23, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n3/pdf/scientificamerican0375-
17.pdf.

Gursky:1975:XDS

[1045] H. Gursky and E. P. J. van den Heuvel. X-ray-emitting double stars. Scientific American, 232(3):24–35, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gursky:1975:XRE

[1046] Herbert Gursky and Edward P. J. van den Heuvel. X-ray-emitting double stars. Scientific American, 232(3):24–35, March 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n3/pdf/scientificamerican0375-24.pdf.

Chapanis:1975:IHC

[1047] Alphonse Chapanis. Interactive human communication. Scientific American, 232(3):36–42, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n3/pdf/scientificamerican0375-
36.pdf.

Anonymous:1975:SCc

[1048] Anonymous. Science and the citizen. Scientific American, 232(3):47–49, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n3/pdf/scientificamerican0375-47.pdf.

Wyllie:1975:EM

[1049] Peter J. Wyllie. The Earth’s mantle. Scientific American, 232 (3):50–57, 60–63, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n3/pdf/scientificamerican0375-
50.pdf.

Rushton:1975:VPC

[1050] W. A. H. Rushton. Visual pigments and color blindness. Scientific American, 232(3):64–65, 68–74, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n3/pdf/scientificamerican0375-
64.pdf.

Benson:1975:RWO

[1051] Andrew A. Benson and Richard F. Lee. The role of wax in oceanic food chains. Scientific American, 232(3):76–86, March 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n3/pdf/scientificamerican0375-76.pdf.

Litten:1975:MPM

[1052] Walter Litten. The most poisonous mushrooms. Scientific American, 232(3):90–101, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n3/pdf/scientificamerican0375-
90.pdf.

Drake:1975:GDP

[1053] Stillman Drake and James MacLachlan. Galileo’s discovery of the parabolic trajectory. Scientific American, 232(3):102–111, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n3/pdf/scientificamerican0375-102.pdf.

Gardner:1975:MGb

[1054] Martin Gardner. Mathematical games. Scientific American, 232(3):112– 116, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n3/pdf/scientificamerican0375-112.pdf.

Stong:1975:ASc

[1055] C. L. Stong. The amateur scientist. Scientific American, 232(3):118–124, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n3/pdf/scientificamerican0375-118.pdf.

Morrison:1975:Bc

[1056] Philip Morrison. Books. Scientific American, 232(3):125–129, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n3/pdf/scientificamerican0375-125.pdf.

Anonymous:1975:Bc

[1057] Anonymous. Bibliography. Scientific American, 232(3):130–131, March 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n3/pdf/scientificamerican0375-130.pdf.

Anonymous:1975:Ld

[1058] Anonymous. Letters. Scientific American, 232(4):8, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n4/pdf/scientificamerican0475-8.pdf.

Anonymous:1975:YAd

[1059] Anonymous. 50 and 100 years ago. Scientific American, 232(4):13–15, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n4/pdf/scientificamerican0475-13.pdf.

Anonymous:1975:Ad

[1060] Anonymous. The authors. Scientific American, 232(4):15, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n4/pdf/scientificamerican0475-15A.pdf.

Epstein:1975:PNW

[1061] William Epstein. The proliferation of nuclear weapons. Scientific American, 232(4):18–33, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n4/pdf/scientificamerican0475-
18.pdf.

Julesz:1975:EVP

[1062] Bela Julesz. Experiments in the visual perception of texture. Scientific American, 232(4):34–43, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n4/pdf/scientificamerican0475-
34.pdf.

Cerami:1975:CSC

[1063] Anthony Cerami and Charles M. Peterson. Cyanate and sicklecell disease. Scientific American, 232(4):44–50, April 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n4/pdf/scientificamerican0475-44.pdf.

Anonymous:1975:SCd

[1064] Anonymous. Science and the citizen. Scientific American, 232(4):53–57, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n4/pdf/scientificamerican0475-53.pdf.

Bakker:1975:DR

[1065] Robert T. Bakker. Dinosaur renaissance. Scientific American, 232(4):58– 79, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n4/pdf/scientificamerican0475-58.pdf.

Albersheim:1975:WGP

[1066] Peter Albersheim. The walls of growing plant cells. Scientific American, 232(4):80–95, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n4/pdf/scientificamerican0475-
80.pdf.

Yonge:1975:GC

[1067] C. M. Yonge. Giant clams. Scientific American, 232(4):96–105, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n4/pdf/scientificamerican0475-96.pdf.

Howard:1975:RS

[1068] Robert Howard. The rotation of the Sun. Scientific American, 232(4):106–115, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n4/pdf/scientificamerican0475-
106.pdf.

McQueen:1975:DMA

[1069] H. J. McQueen and W. J. McGregor Tegart. The deformation of metals at high temperatures. Scientific American, 232(4):116–125, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

McQueen:1975:DMH

[1070] Hugh J. McQueen and W. J. McGregor Tegart. The deformation of metals at high temperatures. Scientific American, 232(4):116–125, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n4/pdf/scientificamerican0475-116.pdf.

Gardner:1975:MGc

[1071] Martin Gardner. Mathematical games. Scientific American, 232(4):126– 133, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n4/pdf/scientificamerican0475-126.pdf.

Stong:1975:ASd

[1072] C. L. Stong. The amateur scientist. Scientific American, 232(4):134– 142, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n4/pdf/scientificamerican0475-134.pdf.

Morrison:1975:Bd

[1073] Philip Morrison. Books. Scientific American, 232(4):143–147, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n4/pdf/scientificamerican0475-143.pdf.

Anonymous:1975:Bd

[1074] Anonymous. Bibliography. Scientific American, 232(4):148, April 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n4/pdf/scientificamerican0475-148.pdf.

Anonymous:1975:Le

[1075] Anonymous. Letters. Scientific American, 232(5):6–8, May 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n5/pdf/scientificamerican0575-6.pdf.

Anonymous:1975:YAe

[1076] Anonymous. 50 and 100 years ago. Scientific American, 232(5):10, May 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n5/pdf/scientificamerican0575-10.pdf.

Anonymous:1975:Ae

[1077] Anonymous. The authors. Scientific American, 232(5):11, May 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n5/pdf/scientificamerican0575-11.pdf.

Press:1975:EP

[1078] Frank Press. Earthquake prediction. Scientific American, 232(5):14– 23, May 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n5/pdf/scientificamerican0575-14.pdf.

Spector:1975:MBP

[1079] Deborah H. Spector and David Baltimore. The molecular biology of poliovirus. Scientific American, 232(5):24–31, May 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n5/pdf/scientificamerican0575-24.pdf.

Vacroux:1975:M

[1080] André G. Vacroux. Microcomputers. Scientific American, 232(5):32– 40, May 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n5/pdf/scientificamerican0575-32.pdf.

Anonymous:1975:SCe

[1081] Anonymous. Science and the citizen. Scientific American, 232(5):42– 45, May 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n5/pdf/scientificamerican0575-42.pdf.

Chaitin:1975:RMP

[1082] Gregory J. Chaitin. Randomness and mathematical proof. Scientific American, 232(5):47–52, May 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n5/pdf/scientificamerican0575-
47.pdf.

Bertram:1975:SSL

[1083] Brian C. R. Bertram. The social system of Lions. Scientific American, 232(5):54–65, May 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n5/pdf/scientificamerican0575-
54.pdf.

Plattner:1975:RMN

[1084] Stuart Plattner. Rural market networks. Scientific American, 232(5):66– 79, May 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n5/pdf/scientificamerican0575-66.pdf.

Sagan:1975:SEI

[1085] Carl Sagan and Frank Drake. The search for extraterrestrial intelligence. Scientific American, 232(5):80–89, May 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n5/pdf/scientificamerican0575-80.pdf.

Horn:1975:FS

[1086] Henry S. Horn. Forest succession. Scientific American, 232(5):90–98, May 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n5/pdf/scientificamerican0575-90.pdf.

Gardner:1975:MGd

[1087] Martin Gardner. Mathematical games. Scientific American, 232(5):102– 107, May 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n5/pdf/scientificamerican0575-102.pdf.

Stong:1975:ASe

[1088] C. L. Stong. The amateur scientist. Scientific American, 232(5):109– 115, May 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n5/pdf/scientificamerican0575-109.pdf.

Morrison:1975:Be

[1089] Philip Morrison. Books. Scientific American, 232(5):117–121, May 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n5/pdf/scientificamerican0575-117.pdf.

Anonymous:1975:Be

[1090] Anonymous. Bibliography. Scientific American, 232(5):122, May 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n5/pdf/scientificamerican0575-122.pdf.

Anonymous:1975:Lf

[1091] Anonymous. Letters. Scientific American, 232(6):6, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n6/pdf/scientificamerican0675-6.pdf.

Anonymous:1975:YAf

[1092] Anonymous. 50 and 100 years ago. Scientific American, 232(6):8, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n6/pdf/scientificamerican0675-8.pdf.

Anonymous:1975:Af

[1093] Anonymous. The authors. Scientific American, 232(6):10, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n6/pdf/scientificamerican0675-10.pdf.

Wortman:1975:AC

[1094] Sterling Wortman. Agriculture in China. Scientific American, 232(6):13– 21, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n6/pdf/scientificamerican0675-13.pdf.

Kappas:1975:HLM

[1095] Attallah Kappas and Alvito P. Alvares. How the liver metabolizes foreign substances. Scientific American, 232(6):22–31, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n6/pdf/scientificamerican0675-22.pdf.

Wilson:1975:SA

[1096] Edward O. Wilson. Slavery in ants. Scientific American, 232(6):32–36, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n6/pdf/scientificamerican0675-32.pdf.

Anonymous:1975:SCf

[1097] Anonymous. Science and the citizen. Scientific American, 232(6):41–49, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n6/pdf/scientificamerican0675-41.pdf.

Drell:1975:EAN

[1098] S. D. Drell. Electron-positron annihilation and the new particles. Scientific American, 232(6):50–62, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Drell:1975:EPA

[1099] Sidney D. Drell. Electron-positron annihilation and the new particles. Scientific American, 232(6):50–62, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v232/n6/pdf/scientificamerican0675-
50.pdf.

Percy:1975:PS

[1100] John R. Percy. Pulsating stars. Scientific American, 232(6):66–75, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n6/pdf/scientificamerican0675-66.pdf.

Johansson:1975:VMP

[1101] Gunnar Johansson. Visual motion perception. Scientific American, 232(6):76–88, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n6/pdf/scientificamerican0675-
76.pdf.

Butler:1975:PT

[1102] James N. Butler. Pelagic tar. Scientific American, 232(6):90–97, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n6/pdf/scientificamerican0675-90.pdf.

Drake:1975:RMG

[1103] Stillman Drake. The role of music in Galileo’s experiments. Scientific American, 232(6):98–104, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v232/n6/pdf/scientificamerican0675-
98.pdf.

Gardner:1975:MGe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n6/pdf/scientificamerican0675-106.pdf.

Stong:1975:ASf

[1105] C. L. Stong. The amateur scientist. Scientific American, 232(6):112– 120, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v232/n6/pdf/scientificamerican0675-112.pdf.

Morrison:1975:Bf

[1106] Philip Morrison. Books. Scientific American, 232(6):123–127, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n6/pdf/scientificamerican0675-123.pdf.

Anonymous:1975:Bf

[1107] Anonymous. Bibliography. Scientific American, 232(6):128, June 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v232/
n6/pdf/scientificamerican0675-128.pdf.

Anonymous:1975:Lg

[1108] Anonymous. Letters. Scientific American, 233(1):8, July 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n1/pdf/scientificamerican0775-8.pdf.

Anonymous:1975:YAg

[1109] Anonymous. 50 and 100 years ago. Scientific American, 233(1):10, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n1/pdf/scientificamerican0775-10B.pdf.

Anonymous:1975:Ag

[1110] Anonymous. The authors. Scientific American, 233(1):12, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n1/pdf/scientificamerican0775-12.pdf.

Tsipis:1975:ASM

[1111] Kosta Tsipis. The accuracy of strategic missiles. Scientific American, 233(1):14–23, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n1/pdf/scientificamerican0775-
14.pdf.

Cohen:1975:MG

[1112] Stanley N. Cohen. The manipulation of genes. Scientific American, 233(1):24–33, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n1/pdf/scientificamerican0775-
24.pdf.

Brandt:1975:PPS

[1113] Werner Brandt. Positrons as a probe of the solid state. Scientific American, 233(1):34–42, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n1/pdf/scientificamerican0775-
34.pdf.

Anonymous:1975:SCg

[1114] Anonymous. Science and the citizen. Scientific American, 233(1):45– 48, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n1/pdf/scientificamerican0775-45.pdf.

Bierlein:1975:JB

[1115] John C. Bierlein. The journal bearing. Scientific American, 233(1):50– 64, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n1/pdf/scientificamerican0775-50.pdf.

Wurtman:1975:ELH

[1116] Richard J. Wurtman. The effects of light on the human body. Scientific American, 233(1):68–77, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n1/pdf/scientificamerican0775-
68.pdf.

Few:1975:T

[1117] Arthur A. Few. Thunder. Scientific American, 233(1):80–90, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n1/pdf/scientificamerican0775-80.pdf.

McMahon:1975:MDT

[1118] Thomas A. McMahon. The mechanical design of trees. Scientific American, 233(1):92–102, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n1/pdf/scientificamerican0775-
92.pdf.

Wright:1975:WMR

[1119] R. H. Wright. Why mosquito repellents repel. Scientific American, 233(1):104–111, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n1/pdf/scientificamerican0775-
104.pdf.

Gardner:1975:MGf

[1120] Martin Gardner. Mathematical games. Scientific American, 233(1):112– 117, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n1/pdf/scientificamerican0775-112.pdf.

Stong:1975:ASg

[1121] C. L. Stong. The amateur scientist. Scientific American, 233(1):120– 125, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n1/pdf/scientificamerican0775-120.pdf.

Layzer:1975:B

[1122] David Layzer. Books. Scientific American, 233(1):126–131, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n1/pdf/scientificamerican0775-126.pdf.

Anonymous:1975:Bg

[1123] Anonymous. Bibliography. Scientific American, 233(1):132, July 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n1/pdf/scientificamerican0775-132.pdf.

Anonymous:1975:Lh

[1124] Anonymous. Letters. Scientific American, 233(2):9, August 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n2/pdf/scientificamerican0875-9.pdf.

Anonymous:1975:YAh

[1125] Anonymous. 50 and 100 years ago. Scientific American, 233(2):12, August 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n2/pdf/scientificamerican0875-12.pdf.

Anonymous:1975:Ah

[1126] Anonymous. The authors. Scientific American, 233(2):13, August 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n2/pdf/scientificamerican0875-13.pdf.

White:1975:ICM

[1127] Kerr L. White. International comparisons of medical care. Scientific American, 233(2):17–25, August 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n2/pdf/scientificamerican0875-
17.pdf.

Strom:1975:GRG

[1128] Richard G. Strom, George K. Miley, and Jan Oort. Giant radio galaxies. Scientific American, 233(2):26–35, August 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n2/pdf/scientificamerican0875-26.pdf.

Berg:1975:HBS

[1129] Howard C. Berg. How bacteria swim. Scientific American, 233(2):36–44, August 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n2/pdf/scientificamerican0875-36.pdf.

Anonymous:1975:SCh

[1130] Anonymous. Science and the citizen. Scientific American, 233(2):46–49, August 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n2/pdf/scientificamerican0875-46.pdf.

Clarke:1975:CBD

[1131] Bryan Clarke. The causes of biological diversity. Scientific American, 233(2):50–60, August 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n2/pdf/scientificamerican0875-
50.pdf.

Beck:1975:PSC

[1132] Jacob Beck. The perception of surface color. Scientific American, 233(2):62–75, August 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n2/pdf/scientificamerican0875-
62.pdf.

Heirtzler:1975:FMA

[1133] J. R. Heirtzler and W. B. Bryan. The floor of the Mid-Atlantic Rift. Scientific American, 233(2):78–90, August 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v233/n2/pdf/scientificamerican0875-
78.pdf.

Heirtzler:1975:FMR

[1134] J. R. Heirtzler and W. B. Bryan. The floor of the Mid-Atlantic Rift. Scientific American, 233(2):79–90, August 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Fowler:1975:PCU

[1135] Melvin L. Fowler. A pre-Columbian urban center on the Mississippi. Scientific American, 233(2):92–101, August 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n2/pdf/scientificamerican0875-92.pdf.

Emlen:1975:SOS

[1136] Stephen T. Emlen. The stellar-orientation system of a migratory bird. Scientific American, 233(2):102–111, August 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n2/pdf/scientificamerican0875-102.pdf.

Gardner:1975:MGg

[1137] Martin Gardner. Mathematical games. Scientific American, 233 (2):112–115, August 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n2/pdf/scientificamerican0875-
112.pdf.

Stong:1975:ASh

[1138] C. L. Stong. The amateur scientist. Scientific American, 233(2):116–121, August 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n2/pdf/scientificamerican0875-116.pdf.

Morrison:1975:Bg

[1139] Philip Morrison. Books. Scientific American, 233(2):123–127, August 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n2/pdf/scientificamerican0875-123.pdf.

Anonymous:1975:Bh

[1140] Anonymous. Bibliography. Scientific American, 233(2):128, August 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n2/pdf/scientificamerican0875-128.pdf.

Anonymous:1975:Li

[1141] Anonymous. Letters. Scientific American, 233(3):8–10, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n3/pdf/scientificamerican0975-8.pdf.

Anonymous:1975:YAi

[1142] Anonymous. 50 and 100 years ago. Scientific American, 233(3):12, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n3/pdf/scientificamerican0975-12.pdf.

Anonymous:1975:Ai

[1143] Anonymous. The authors. Scientific American, 233(3):16–20, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n3/pdf/scientificamerican0975-16.pdf.

Sagan:1975:SS

[1144] Carl Sagan. The solar system. Scientific American, 233(3):22–31, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n3/pdf/scientificamerican0975-22.pdf.

Cameron:1975:OES

[1145] A. G. W. Cameron. The origin and evolution of the solar system. Scientific American, 233(3):32–41, September 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n3/pdf/scientificamerican0975-32.pdf.

Parker:1975:S

[1146] E. N. Parker. The Sun. Scientific American, 233(3):42–50, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n3/pdf/scientificamerican0975-42.pdf.

Anonymous:1975:SCi

[1147] Anonymous. Science and the citizen. Scientific American, 233(3):53–61, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n3/pdf/scientificamerican0975-53.pdf.

Murray:1975:M

[1148] Bruce C. Murray. Mercury. Scientific American, 233(3):58–68, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n3/pdf/scientificamerican0975-58.pdf.

Andrew:1975:V

[1149] Andrew and Louise Young. Venus. Scientific American, 233(3):70–78, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n3/pdf/scientificamerican0975-70.pdf.

Young:1975:V

[1150] Andrew Young and Louise Young. Venus. Scientific American, 233(3): 70–78, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Siever:1975:E

[1151] Raymond Siever. The Earth. Scientific American, 233(3):82–90, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n3/pdf/scientificamerican0975-82.pdf.

Wood:1975:M

[1152] John A. Wood. The Moon. Scientific American, 233(3):92–102, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n3/pdf/scientificamerican0975-92.pdf.

Pollack:1975:M

[1153] James B. Pollack. Mars. Scientific American, 233(3):106–117, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n3/pdf/scientificamerican0975-106.pdf.

Wolfe:1975:J

[1154] John H. Wolfe. Jupiter. Scientific American, 233(3):118–126, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n3/pdf/scientificamerican0975-118.pdf.

Hunten:1975:OP

[1155] Donald M. Hunten. The outer planets. Scientific American, 233(3):130– 132, 134, 136, 138, 140, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n3/pdf/scientificamerican0975-
130.pdf.

Hartmann:1975:SBS

[1156] William K. Hartmann. The smaller bodies of the solar system. Scientific American, 233(3):142–147, 149–154, 156, 159, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n3/pdf/scientificamerican0975-142.pdf.

VanAllen:1975:IPF

[1157] James A. Van Allen. Interplanetary particles and fields. Scientific American, 233(3):160–162, 164, 166, 169–173, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n3/pdf/scientificamerican0975-160.pdf.

Gardner:1975:MGh

[1158] Martin Gardner. Mathematical games. Scientific American, 233 (3):174–180, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n3/pdf/scientificamerican0975-
174.pdf.

Stong:1975:ASi

[1159] C. L. Stong. The amateur scientist. Scientific American, 233(3):182–188, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n3/pdf/scientificamerican0975-182.pdf.

Morrison:1975:Bh

[1160] Philip Morrison. Books. Scientific American, 233(3):190–202, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n3/pdf/scientificamerican0975-190.pdf.

Anonymous:1975:Bi

[1161] Anonymous. Bibliography. Scientific American, 233(3):204–206, September 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n3/pdf/scientificamerican0975-204.pdf.

Anonymous:1975:Lj

[1162] Anonymous. Letters. Scientific American, 233(4):6–8, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n4/pdf/scientificamerican1075-6.pdf.

Anonymous:1975:YAj

[1163] Anonymous. 50 and 100 years ago. Scientific American, 233(4):10, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n4/pdf/scientificamerican1075-10.pdf.

Anonymous:1975:Aj

[1164] Anonymous. The authors. Scientific American, 233(4):12, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n4/pdf/scientificamerican1075-12.pdf.

McIntyre:1975:NHR

[1165] H. C. McIntyre. Natural-uranium heavy-water reactors. Scientific American, 233(4):17–27, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

McIntyre:1975:NUH

[1166] Hugh C. McIntyre. Natural-uranium heavy-water reactors. Scientific American, 233(4):17–27, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n4/pdf/scientificamerican1075-
17.pdf.

Satir:1975:FSS

[1167] Birgit Satir. The final steps in secretion. Scientific American, 233(4):28– 37, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n4/pdf/scientificamerican1075-28.pdf.

Glashow:1975:QCF

[1168] Sheldon Lee Glashow. Quarks with color and flavor. Scientific American, 233(4):38–50, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n4/pdf/scientificamerican1075-
38.pdf.

Anonymous:1975:SCj

[1169] Anonymous. Science and the citizen. Scientific American, 233(4):52–54, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n4/pdf/scientificamerican1075-52.pdf.

Gordon:1975:IRP

[1170] Richard Gordon, Gabor T. Herman, and Steven A. Johnson. Image reconstruction from projections. Scientific American, 233(4):56–61, 64–65, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n4/pdf/scientificamerican1075-56.pdf.

Sabloff:1975:RMM

[1171] Jeremy A. Sabloff and William L. Rathje. The rise of a Maya merchant class. Scientific American, 233(4):72–82, October 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n4/pdf/scientificamerican1075-72.pdf.

Isaacs:1975:AAD

[1172] John D. Isaacs and Richard A. Schwartzlose. Active animals of the deep-sea floor. Scientific American, 233(4):84–91, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n4/pdf/scientificamerican1075-84.pdf.

Deutsch:1975:MI

[1173] Diana Deutsch. Musical illusions. Scientific American, 233(4):92–104, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n4/pdf/scientificamerican1075-92.pdf.

York:1975:DHB

[1174] Herbert F. York. The debate over the hydrogen bomb. Scientific American, 233(4):106–113, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n4/pdf/scientificamerican1075-
106.pdf.

Gardner:1975:MGi

[1175] Martin Gardner. Mathematical games. Scientific American, 233 (4):114–118, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n4/pdf/scientificamerican1075-
114.pdf.

Stong:1975:ASj

[1176] C. L. Stong. The amateur scientist. Scientific American, 233(4):120–126, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n4/pdf/scientificamerican1075-120.pdf.

Bonner:1975:B

[1177] John Tyler Bonner. Books. Scientific American, 233(4):129–137, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n4/pdf/scientificamerican1075-129.pdf.

Anonymous:1975:Bj

[1178] Anonymous. Bibliography. Scientific American, 233(4):138–140, October 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n4/pdf/scientificamerican1075-138.pdf.

Anonymous:1975:Lk

[1179] Anonymous. Letters. Scientific American, 233(5):8–13, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n5/pdf/scientificamerican1175-8.pdf.

Anonymous:1975:YAk

[1180] Anonymous. 50 and 100 years ago. Scientific American, 233(5):14–16, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n5/pdf/scientificamerican1175-14.pdf.

Anonymous:1975:Ak

[1181] Anonymous. The authors. Scientific American, 233(5):18–20, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n5/pdf/scientificamerican1175-18.pdf.

Epstein:1975:NFZ

[1182] William Epstein. Nuclear-free zones. Scientific American, 233(5):25–35, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n5/pdf/scientificamerican1175-25.pdf.

Cohen:1975:PSM

[1183] Carolyn Cohen. The protein switch of muscle contraction. Scientific American, 233(5):36–45, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v233/n5/pdf/scientificamerican1175-
36.pdf.

Kolm:1975:HGM

[1184] Henry Kolm, John Oberteuffer, and David Kelland. High-gradient magnetic separation. Scientific American, 233(5):46–54, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n5/pdf/scientificamerican1175-46.pdf.

Kolm:1975:HMS

[1185] H. Kolm, J. Oberteuffer, and D. Kelland. High-gradient magnetic separation. Scientific American, 233(5):46–54, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1975:SCk

[1186] Anonymous. Science and the citizen. Scientific American, 233(5):56–60, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n5/pdf/scientificamerican1175-56.pdf.

Cairns:1975:CP

[1187] John Cairns. The cancer problem. Scientific American, 233(5):64–78, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n5/pdf/scientificamerican1175-64.pdf.

Weis-Fogh:1975:UMG

[1188] Torkel Weis-Fogh. Unusual mechanisms for the generation of lift in flying animals. Scientific American, 233(5):80–87, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n5/pdf/scientificamerican1175-80.pdf.

Toksoz:1975:SL

[1189] M. Nafi Toksöz. The subduction of the lithosphere. Scientific American, 233(5):88–98, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n5/pdf/scientificamerican1175-
88.pdf.

Derjaguin:1975:SDA

[1190] B. V. Derjaguin and D. B. Fedoseev. The synthesis of diamond at low pressure. Scientific American, 233(5):102–109, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Derjaguin:1975:SDL

[1191] B. V. Derjaguin and D. B. Fedoseev. The synthesis of diamond at low pressure. Scientific American, 233(5):102–109, November 1975.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n5/pdf/scientificamerican1175-102.pdf.

Hess:1975:RPS

[1192] Eckhard H. Hess. The role of pupil size in communication. Scientific American, 233(5):110–119, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v233/n5/pdf/scientificamerican1175-
110.pdf.

Gardner:1975:MGj

[1193] Martin Gardner. Mathematical games. Scientific American, 233 (5):120–125, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n5/pdf/scientificamerican1175-
120.pdf.

Stong:1975:ASk

[1194] C. L. Stong. The amateur scientist. Scientific American, 233(5):128–134, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n5/pdf/scientificamerican1175-128.pdf.

Morrison:1975:Bi

[1195] Philip Morrison. Books. Scientific American, 233(5):136–144, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n5/pdf/scientificamerican1175-136.pdf.

Anonymous:1975:Bk

[1196] Anonymous. Bibliography. Scientific American, 233(5):148, November 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n5/pdf/scientificamerican1175-148.pdf.

Anonymous:1975:Ll

[1197] Anonymous. Letters. Scientific American, 233(6):8–10, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n6/pdf/scientificamerican1275-8.pdf.

Anonymous:1975:YAl

[1198] Anonymous. 50 and 100 years ago. Scientific American, 233(6):15–17, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n6/pdf/scientificamerican1275-15.pdf.

Anonymous:1975:Al

[1199] Anonymous. The authors. Scientific American, 233(6):21, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n6/pdf/scientificamerican1275-21.pdf.

Atwood:1975:SMW

[1200] Genevieve Atwood. The strip-mining of western coal. Scientific American, 233(6):23–29, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n6/pdf/scientificamerican1275-
23.pdf.

Luria:1975:CEC

[1201] Salvador E. Luria. Colicins and the energetics of cell membranes. Scientific American, 233(6):30–37, December 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n6/pdf/scientificamerican1275-30.pdf.

Charles:1975:XRS

[1202] Philip A. Charles and J. Leonard Culhane. X rays from supernova remnants. Scientific American, 233(6):38–46, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n6/pdf/scientificamerican1275-38.pdf.

Charles:1975:XSR

[1203] P. A. Charles and J. L. Culhane. X-rays from supernova remnants. Scientific American, 233(6):38–46, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1975:SCl

[1204] Anonymous. Science and the citizen. Scientific American, 233(6):48–55, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n6/pdf/scientificamerican1275-48.pdf.

Layzer:1975:AT

[1205] David Layzer. The arrow of time. Scientific American, 233(6):56–69, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n6/pdf/scientificamerican1275-56.pdf.

vanHeyningen:1975:WHH

[1206] Ruth van Heyningen. What happens to the human lens in cataract. Scientific American, 233(6):70–81, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v233/n6/pdf/scientificamerican1275-
70.pdf.

James:1975:SEM

[1207] Wendy James. Sister-exchange marriage. Scientific American, 233(6):84–94, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n6/pdf/scientificamerican1275-
84.pdf.

Uhlmann:1975:MPM

[1208] D. R. Uhlmann and A. G. Kolbeck. The microstructure of polymeric materials. Scientific American, 233(6):96–106, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n6/pdf/scientificamerican1275-96.pdf.

Evans:1975:SWA

[1209] Howard E. Evans and Robert W. Matthews. The sand wasps of Australia. Scientific American, 233(6):108–115, December 1975. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v233/
n6/pdf/scientificamerican1275-108.pdf.

Gardner:1975:MGk

[1210] Martin Gardner. Mathematical games. Scientific American, 233 (6):116–119, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v233/n6/pdf/scientificamerican1275-
116.pdf.

Stong:1975:ASl

[1211] C. L. Stong. The amateur scientist. Scientific American, 233(6):120–125, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n6/pdf/scientificamerican1275-120.pdf.

Philip:1975:B

[1212] Philip and Phylis Morrison. Books. Scientific American, 233(6):127–142, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n6/pdf/scientificamerican1275-127.pdf.

Anonymous:1975:AIa

[1213] Anonymous. Annual index. Scientific American, 233(6):144–146, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n6/pdf/scientificamerican1275-144.pdf.

Anonymous:1975:Bl

[1214] Anonymous. Bibliography. Scientific American, 233(6):148, December 1975. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v233/n6/pdf/scientificamerican1275-148.pdf.

Anonymous:1976:La

[1215] Anonymous. Letters. Scientific American, 234(1):6–9, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n1/pdf/scientificamerican0176-6.pdf.

Anonymous:1976:Lg

[1216] Anonymous. Letters. Scientific American, 234(1):6–10, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n1/pdf/scientificamerican0776-6.pdf.

Anonymous:1976:YAa

[1217] Anonymous. 50 and 100 years ago. Scientific American, 234(1):10–15, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n1/pdf/scientificamerican0176-10.pdf.

Anonymous:1976:YAg

[1218] Anonymous. 50 and 100 years ago. Scientific American, 234(1):11– 16, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n1/pdf/scientificamerican0776-11.pdf.

Anonymous:1976:Aa

[1219] Anonymous. The authors. Scientific American, 234(1):16–17, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n1/pdf/scientificamerican0176-16.pdf.

Anonymous:1976:Ag

[1220] Anonymous. The authors. Scientific American, 234(1):18–25, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n1/pdf/scientificamerican0776-18.pdf.

Bethe:1976:NFP

[1221] H. A. Bethe. The necessity of fission power. Scientific American, 234(1):21–31, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n1/pdf/scientificamerican0176-
21.pdf.

Keyfitz:1976:WRW

[1222] Nathan Keyfitz. World resources and the world middle class. Scientific American, 234(1):28–35, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n1/pdf/scientificamerican0776-
28.pdf.

Carr:1976:VM

[1223] Michael H. Carr. The volcanoes of Mars. Scientific American, 234(1):32– 43, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n1/pdf/scientificamerican0176-32.pdf.

Cline:1976:SNF

[1224] David B. Cline, Alfred K. Mann, and Carlo Rubbia. The search for new families of elementary particles. Scientific American, 234(1):44–55, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n1/pdf/scientificamerican0176-44.pdf.

McEwen:1976:IBH

[1225] Bruce S. McEwen. Interactions between hormones and nerve tissue. Scientific American, 234(1):48–58, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v235/n1/pdf/scientificamerican0776-
48.pdf.

Anonymous:1976:SCa

[1226] Anonymous. Science and the citizen. Scientific American, 234(1):56–63, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n1/pdf/scientificamerican0176-56.pdf.

Anonymous:1976:SCg

[1227] Anonymous. Science and the citizen. Scientific American, 234(1):60– 67, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n1/pdf/scientificamerican0776-60.pdf.

Maniatis:1976:DOR

[1228] Tom Maniatis and Mark Ptashne. A DNA operator-repressor system. Scientific American, 234(1):64–76, January 1976. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n1/pdf/scientificamerican0176-64.pdf.

Miller:1976:DRI

[1229] Jack Robert Miller. The direct reduction of iron Ore. Scientific American, 234(1):68–80, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n1/pdf/scientificamerican0776-
68.pdf.

Caldwell:1976:S

[1230] Roy L. Caldwell and Hugh Dingle. Stomatopods. Scientific American, 234(1):80–89, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n1/pdf/scientificamerican0176-
80.pdf.

Almgren:1976:GSF

[1231] Frederick J. Almgren, Jr. and Jean E. Taylor. The geometry of soap films and soap bubbles. Scientific American, 234(1):82–93, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n1/pdf/scientificamerican0776-82.pdf.

Jerison:1976:PEM

[1232] Harry J. Jerison. Paleoneurology and the evolution of mind. Scientific American, 234(1):90–101, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n1/pdf/scientificamerican0176-
90.pdf.

Alldredge:1976:A

[1233] Alice Alldredge. Appendicularians. Scientific American, 234(1):94–105, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n1/pdf/scientificamerican0776-94.pdf.

Fraser:1976:M

[1234] Alistair B. Fraser and William H. Mach. Mirages. Scientific American, 234(1):102–111, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n1/pdf/scientificamerican0176-
102.pdf.

Wehner:1976:PLN

[1235] Rüdiger Wehner. Polarized-light navigation by insects. Scientific American, 234(1):106–115, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n1/pdf/scientificamerican0776-
106.pdf.

Langer:1976:IAS

[1236] William L. Langer. Immunization against smallpox before Jenner. Scientific American, 234(1):112–117, January 1976. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n1/pdf/scientificamerican0176-112.pdf.

McNown:1976:CA

[1237] John S. McNown. Canals in America. Scientific American, 234(1):116– 125, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n1/pdf/scientificamerican0776-116.pdf.

Gardner:1976:MGa

[1238] Martin Gardner. Mathematical games. Scientific American, 234 (1):118–123, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n1/pdf/scientificamerican0176-
118.pdf.

Stong:1976:ASa

[1239] C. L. Stong. The amateur scientist. Scientific American, 234(1):124–128, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n1/pdf/scientificamerican0176-124.pdf.

Gardner:1976:MGg

[1240] Martin Gardner. Mathematical games. Scientific American, 234(1):126– 131, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n1/pdf/scientificamerican0776-126.pdf.

Morrison:1976:Ba

[1241] Philip Morrison. Books. Scientific American, 234(1):130–135, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n1/pdf/scientificamerican0176-130.pdf.

Morrison:1976:Bg

[1242] Philip Morrison. Books. Scientific American, 234(1):132–137, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n1/pdf/scientificamerican0776-132.pdf.

Anonymous:1976:Ba

[1243] Anonymous. Bibliography. Scientific American, 234(1):136, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n1/pdf/scientificamerican0176-136.pdf.

Anonymous:1976:Bg

[1244] Anonymous. Bibliography. Scientific American, 234(1):138, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n1/pdf/scientificamerican0776-138.pdf.

Anonymous:1976:Lb

[1245] Anonymous. Letters. Scientific American, 234(2):8–15, February 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n2/pdf/scientificamerican0276-8.pdf.

Anonymous:1976:YAb

[1246] Anonymous. 50 and 100 years ago. Scientific American, 234(2):16–19, February 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n2/pdf/scientificamerican0276-16.pdf.

Anonymous:1976:Ab

[1247] Anonymous. The authors. Scientific American, 234(2):20, February 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n2/pdf/scientificamerican0276-20.pdf.

Barber:1976:EEH

[1248] Bernard Barber. The ethics of experimentation with human subjects. Scientific American, 234(2):25–31, February 1976. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n2/pdf/scientificamerican0276-25.pdf.

OMalley:1976:RSH

[1249] Bert W. O’Malley and William T. Schrader. The receptors of steroid hormones. Scientific American, 234(2):32–43, February 1976. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n2/pdf/scientificamerican0276-32.pdf.

Flandern:1976:GGW

[1250] Thomas C. Van Flandern. Is gravity getting weaker? Scientific American, 234(2):44–53, February 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n2/pdf/scientificamerican0276-
44.pdf.

VanFlandern:1976:GGW

[1251] T. C. Van Flandern. Is gravity getting weaker? Scientific American, 234 (2):44–59, February 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1976:SCb

[1252] Anonymous. Science and the citizen. Scientific American, 234(2):54–59, February 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n2/pdf/scientificamerican0276-54.pdf.

Hudgson:1976:FC

[1253] Harlow J. Hudgson. Forage crops. Scientific American, 234(2):60–75, February 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n2/pdf/scientificamerican0276-60.pdf.

Albus:1976:RS

[1254] James S. Albus and John M. Evans, Jr. Robot systems. Scientific American, 234(2):76–86, February 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n2/pdf/scientificamerican0276-
76.pdf.

Schild:1976:FPS

[1255] Romuald Schild. The final paleolithic settlements of the European plain. Scientific American, 234(2):88–99, February 1976. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n2/pdf/scientificamerican0276-88.pdf.

Jones:1976:C

[1256] Maitland Jones, Jr. Carbenes. Scientific American, 234(2):101–113, February 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n2/pdf/scientificamerican0276-101.pdf.

Saunders:1976:BCI

[1257] D. S. Saunders. The biological clock of insects. Scientific American, 234(2):114–121, February 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n2/pdf/scientificamerican0276-
114.pdf.

Gardner:1976:MGb

[1258] Martin Gardner. Mathematical games. Scientific American, 234 (2):122–127, February 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n2/pdf/scientificamerican0276-
122.pdf.

Stong:1976:ASb

[1259] C. L. Stong. The amateur scientist. Scientific American, 234(2):128–133, February 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n2/pdf/scientificamerican0276-128.pdf.

Morrison:1976:Bb

[1260] Philip Morrison. Books. Scientific American, 234(2):134–137, February 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n2/pdf/scientificamerican0276-134.pdf.

Anonymous:1976:Bb

[1261] Anonymous. Bibliography. Scientific American, 234(2):140, February 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n2/pdf/scientificamerican0276-140.pdf.

Anonymous:1976:Lc

[1262] Anonymous. Letters. Scientific American, 234(3):7–13, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n3/pdf/scientificamerican0376-7.pdf.

Anonymous:1976:YAc

[1263] Anonymous. 50 and 100 years ago. Scientific American, 234(3):14–19, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n3/pdf/scientificamerican0376-14.pdf.

Anonymous:1976:Ac

[1264] Anonymous. The authors. Scientific American, 234(3):20–21, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n3/pdf/scientificamerican0376-20.pdf.

Lieber:1976:MA

[1265] Charles S. Lieber. The metabolism of alcohol. Scientific American, 234(3):25–32, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n3/pdf/scientificamerican0376-
25.pdf.

Blumer:1976:PAC

[1266] Max Blumer. Polycyclic aromatic compounds in nature. Scientific American, 234(3):34–45, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n3/pdf/scientificamerican0376-
34.pdf.

Ingersoll:1976:MJ

[1267] Andrew P. Ingersoll. The meteorology of Jupiter. Scientific American, 234(3):46–57, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n3/pdf/scientificamerican0376-
46.pdf.

Anonymous:1976:SCc

[1268] Anonymous. Science and the citizen. Scientific American, 234(3):60–61, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n3/pdf/scientificamerican0376-60A.pdf.

Gott:1976:WUE

[1269] J. Richard Gott III, James E. Gunn, David N. Schramm, and Beatrice M. Tinsley. Will the Universe expand forever? Scientific American, 234(3):62–79, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n3/pdf/scientificamerican0376-
62.pdf.

Ross:1976:RBP

[1270] John Ross. The resources of binocular perception. Scientific American, 234(3):80–86, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n3/pdf/scientificamerican0376-
80.pdf.

McWhorter:1976:SEC

[1271] Eugene W. McWhorter. The small electronic calculator. Scientific American, 234(3):88–96, 98, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n3/pdf/scientificamerican0376-
88.pdf.

Burgess:1976:SS

[1272] J. Wesley Burgess. Social spiders. Scientific American, 234(3):100–106, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n3/pdf/scientificamerican0376-100.pdf.

Herzenberg:1976:FCS

[1273] L. A. Herzenberg, R. G. Sweet, and A. Herzenberg. Fluorescenceactivated cell sorting. Scientific American, 234(3):108–116, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Herzenberg:1976:FAC

[1274] Leonard A. Herzenberg, Richard G. Sweet, and Leonore A. Herzenberg. Fluorescence-activated cell sorting. Scientific American, 234(3):108–117, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n3/pdf/scientificamerican0376-108.pdf.

Gardner:1976:MGc

[1275] Martin Gardner. Mathematical games. Scientific American, 234(3):119– 124, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n3/pdf/scientificamerican0376-119.pdf.

Morrison:1976:Bc

[1276] Philip Morrison. Books. Scientific American, 234(3):126–129, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n3/pdf/scientificamerican0376-126.pdf.

Anonymous:1976:Bc

[1277] Anonymous. Bibliography. Scientific American, 234(3):130–131, March 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n3/pdf/scientificamerican0376-130.pdf.

Anonymous:1976:Ld

[1278] Anonymous. Letters. Scientific American, 234(4):8–21, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n4/pdf/scientificamerican0476-8.pdf.

Anonymous:1976:YAd

[1279] Anonymous. 50 and 100 years ago. Scientific American, 234(4):22–29, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n4/pdf/scientificamerican0476-22.pdf.

Anonymous:1976:Ad

[1280] Anonymous. The authors. Scientific American, 234(4):30, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n4/pdf/scientificamerican0476-30.pdf.

Nelkin:1976:STC

[1281] Dorothy Nelkin. The science-textbook controversies. Scientific American, 234(4):33–39, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n4/pdf/scientificamerican0476-
33.pdf.

Adler:1976:SCB

[1282] Julius Adler. The sensing of chemicals by bacteria. Scientific American, 234(4):40–47, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n4/pdf/scientificamerican0476-
40.pdf.

Kanizsa:1976:SC

[1283] Gaetano K. Kanizsa. Subjective contours. Scientific American, 234(4):48–52, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n4/pdf/scientificamerican0476-
48.pdf.

Anonymous:1976:SCd

[1284] Anonymous. Science and the citizen. Scientific American, 234(4):54–64, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n4/pdf/scientificamerican0476-54.pdf.

Zeeman:1976:CT

[1285] E. C. Zeeman. Catastrophe theory. Scientific American, 234(4):65–83, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n4/pdf/scientificamerican0476-65.pdf.

Darragh:1976:O

[1286] P. J. Darragh, A. J. Gaskin, and J. V. Sanders. Opals. Scientific American, 234(4):84–95, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n4/pdf/scientificamerican0476-
84.pdf.

Stern:1976:AMX

[1287] Edward A. Stern. The analysis of materials by X-ray absorption. Scientific American, 234(4):96–103, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v234/n4/pdf/scientificamerican0476-
96.pdf.

Drake:1976:GFM

[1288] Stillman Drake. Galileo and the first mechanical computing device. Scientific American, 234(4):104–113, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v234/n4/pdf/scientificamerican0476-
104.pdf.

Pooley:1976:NC

[1289] Anthony C. Pooley and Carl Gans. The Nile crocodile. Scientific American, 234(4):114–125, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n4/pdf/scientificamerican0476-
114.pdf.

Gardner:1976:MGd

[1290] Martin Gardner. Mathematical games. Scientific American, 234(4):126– 130, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n4/pdf/scientificamerican0476-126.pdf.

Morrison:1976:Bd

[1291] Philip Morrison. Books. Scientific American, 234(4):132–137, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n4/pdf/scientificamerican0476-132.pdf.

Anonymous:1976:Bd

[1292] Anonymous. Bibliography. Scientific American, 234(4):138, April 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n4/pdf/scientificamerican0476-138.pdf.

Anonymous:1976:Le

[1293] Anonymous. Letters. Scientific American, 234(5):7–11, May 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n5/pdf/scientificamerican0576-7.pdf.

Anonymous:1976:YAe

[1294] Anonymous. 50 and 100 years ago. Scientific American, 234(5):13– 16, May 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n5/pdf/scientificamerican0576-13.pdf.

Anonymous:1976:Ae

[1295] Anonymous. The authors. Scientific American, 234(5):20, May 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n5/pdf/scientificamerican0576-20.pdf.

Cochran:1976:OGC

[1296] Neal P. Cochran. Oil and gas from coal. Scientific American, 234(5):24– 29, May 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n5/pdf/scientificamerican0576-24.pdf.

Raff:1976:CSI

[1297] Martin C. Raff. Cell-surface immunology. Scientific American, 234(5): 30–39, May 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n5/pdf/scientificamerican0576-30.pdf.

Benjamin:1976:MA

[1298] J. S. Benjamin. Mechanical alloying. Scientific American, 234(5):40– 49, May 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n5/pdf/scientificamerican0576-40.pdf.

Anonymous:1976:SCe

[1299] Anonymous. Science and the citizen. Scientific American, 234(5):50– 57, May 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n5/pdf/scientificamerican0576-50.pdf.

Folkman:1976:VT

[1300] Judah Folkman. The vascularization of tumors. Scientific American, 234(5):58–73, May 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n5/pdf/scientificamerican0576-
58.pdf.

John:1976:SF

[1301] John and Elisabeth Buck. Synchronous fireflies. Scientific American, 234(5):74–85, May 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n5/pdf/scientificamerican0576-
74.pdf.

Goldhaber:1976:MP

[1302] Alfred Scharff Goldhaber and Michael Martin Nieto. The mass of the photon. Scientific American, 234(5):86–91, 94, May 1976. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n5/pdf/scientificamerican0576-86.pdf.

Cohen:1976:SH

[1303] I. Bernard Cohen. Stephen hales. Scientific American, 234(5):98–107, May 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n5/pdf/scientificamerican0576-98.pdf.

Cruikshank:1976:GSJ

[1304] Dale P. Cruikshank and David Morrison. The Galilean satellites of Jupiter. Scientific American, 234(5):108–116 (Intl. ed. 40–48), May 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n5/pdf/scientificamerican0576-108.pdf.

Gardner:1976:MGe

[1305] Martin Gardner. Mathematical games. Scientific American, 234(5):118– 123, May 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n5/pdf/scientificamerican0576-118.pdf.

Morrison:1976:Be

[1306] Philip Morrison. Books. Scientific American, 234(5):124–130, May 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n5/pdf/scientificamerican0576-124.pdf.

Anonymous:1976:Be

[1307] Anonymous. Bibliography. Scientific American, 234(5):132, May 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n5/pdf/scientificamerican0576-132.pdf.

Anonymous:1976:Lf

[1308] Anonymous. Letters. Scientific American, 234(6):7–13, June 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n6/pdf/scientificamerican0676-7.pdf.

Anonymous:1976:YAf

[1309] Anonymous. 50 and 100 years ago. Scientific American, 234(6):14, June 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n6/pdf/scientificamerican0676-14.pdf.

Anonymous:1976:Af

[1310] Anonymous. The authors. Scientific American, 234(6):16–17, June 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n6/pdf/scientificamerican0676-16.pdf.

Niemi:1976:CVS

[1311] Richard G. Niemi and William H. Riker. The choice of voting systems. Scientific American, 234(6):21–27, June 1976. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n6/pdf/scientificamerican0676-21.pdf.

Thomas:1976:EHL

[1312] Gordon A. Thomas. An electron-hole liquid. Scientific American, 234(6):28–37, June 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n6/pdf/scientificamerican0676-
28.pdf.

Thomas:1976:EL

[1313] G. A. Thomas. An electron-hole liquid. Scientific American, 234(6):28– 37, June 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stoeckenius:1976:PMS

[1314] Walther Stoeckenius. The purple membrane of salt-loving bacteria. Scientific American, 234(6):38–47, June 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v234/n6/pdf/scientificamerican0676-
38.pdf.

Anonymous:1976:SCf

[1315] Anonymous. Science and the citizen. Scientific American, 234(6):48–55, June 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n6/pdf/scientificamerican0676-48.pdf.

Melbourne:1976:NBP

[1316] William G. Melbourne. Navigation between the planets. Scientific American, 234(6):58–64, 68–74, June 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n6/pdf/scientificamerican0676-
58.pdf.

Jacobsen:1976:YGP

[1317] Thomas W. Jacobsen. 17,000 years of Greek prehistory. Scientific American, 234(6):76–89, June 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n6/pdf/scientificamerican0676-
76.pdf.

Splinter:1976:CPI

[1318] William E. Splinter. Center-pivot irrigation. Scientific American, 234(6):90–99, June 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v234/n6/pdf/scientificamerican0676-
90.pdf.

Stephenson:1976:HS

[1319] F. Richard Stephenson and David H. Clark. Historical supernovas. Scientific American, 234(6):100–107, June 1976. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n6/pdf/scientificamerican0676-100.pdf.

Ryder:1976:FPF

[1320] Henry W. Ryder, Harry Jay Carr, and Paul Herget. Future performance in footracing. Scientific American, 234(6):109–119, June 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n6/pdf/scientificamerican0676-109.pdf.

Gardner:1976:MGf

[1321] Martin Gardner. Mathematical games. Scientific American, 234(6):120– 125, June 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v234/n6/pdf/scientificamerican0676-120.pdf.

Morrison:1976:Bf

[1322] Philip Morrison. Books. Scientific American, 234(6):126–131, June 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n6/pdf/scientificamerican0676-126.pdf.

Anonymous:1976:Bf

[1323] Anonymous. Bibliography. Scientific American, 234(6):132, June 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v234/
n6/pdf/scientificamerican0676-132.pdf.

Cowan:1976:NFR

[1324] George A. Cowan. A natural fission reactor. Scientific American, 235(1): 36–47, July 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n1/pdf/scientificamerican0776-36.pdf.

Anonymous:1976:Lh

[1325] Anonymous. Letters. Scientific American, 235(2):9–10, August 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n2/pdf/scientificamerican0876-9.pdf.

Anonymous:1976:YAh

[1326] Anonymous. 50 and 100 years ago. Scientific American, 235(2):14, August 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n2/pdf/scientificamerican0876-14.pdf.

Anonymous:1976:Ah

[1327] Anonymous. The authors. Scientific American, 235(2):16, August 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n2/pdf/scientificamerican0876-16.pdf.

Rubsamen:1976:MM

[1328] David S. Rubsamen. Medical malpractice. Scientific American, 235(2):18–23, August 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n2/pdf/scientificamerican0876-
18.pdf.

Pappenheimer:1976:SF

[1329] John R. Pappenheimer. The sleep factor. Scientific American, 235(2):24– 29, August 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n2/pdf/scientificamerican0876-24.pdf.

Smith:1976:SAM

[1330] Philip E. L. Smith. Stone-age man on the nile. Scientific American, 235(2):30–41, August 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n2/pdf/scientificamerican0876-
30.pdf.

Anonymous:1976:SCh

[1331] Anonymous. Science and the citizen. Scientific American, 235(2):42–45, August 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n2/pdf/scientificamerican0876-42.pdf.

Burke:1976:HSE

[1332] Kevin C. Burke and J. Tuzo Wilson. Hot spots on the Earth’s surface. Scientific American, 235(2):46–59, August 1976. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n2/pdf/scientificamerican0876-46.pdf.

Lane:1976:RHF

[1333] Charles Lane. Rabbit hemoglobin from frog eggs. Scientific American, 235(2):60–71, August 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n2/pdf/scientificamerican0876-
60.pdf.

Price:1976:PL

[1334] William H. Price. The photographic lens. Scientific American, 235(2):72– 83, August 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n2/pdf/scientificamerican0876-72.pdf.

Milne:1976:SBB

[1335] Lorus J. Milne and Margery Milne. The social behavior of burying beetles. Scientific American, 235(2):84–89, August 1976. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n2/pdf/scientificamerican0876-84.pdf.

Callahan:1976:CSF

[1336] J. J. Callahan. The curvature of space in a finite Universe. Scientific American, 235(2):90–100, August 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n2/pdf/scientificamerican0876-
90.pdf.

Gardner:1976:MGh

[1337] Martin Gardner. Mathematical games. Scientific American, 235 (2):102–107, August 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n2/pdf/scientificamerican0876-
102.pdf.

Morrison:1976:Bh

[1338] Philip Morrison. Books. Scientific American, 235(2):110–115, August 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n2/pdf/scientificamerican0876-110.pdf.

Anonymous:1976:Bh

[1339] Anonymous. Bibliography. Scientific American, 235(2):116, August 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n2/pdf/scientificamerican0876-116.pdf.

Anonymous:1976:Li

[1340] Anonymous. Letters. Scientific American, 235(3):8–11, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n3/pdf/scientificamerican0976-8.pdf.

Anonymous:1976:YAi

[1341] Anonymous. 50 and 100 years ago. Scientific American, 235(3):12–17, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n3/pdf/scientificamerican0976-12.pdf.

Anonymous:1976:Ai

[1342] Anonymous. The authors. Scientific American, 235(3):18–26, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n3/pdf/scientificamerican0976-18.pdf.

Wortman:1976:FA

[1343] Sterling Wortman. Food and agriculture. Scientific American, 235(3):30–39, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n3/pdf/scientificamerican0976-
30.pdf.

Mayer:1976:DHH

[1344] Jean Mayer. The dimensions of human hunger. Scientific American, 235(3):40–49, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n3/pdf/scientificamerican0976-
40.pdf.

Scrimshaw:1976:RHN

[1345] Nevin S. Scrimshaw and Vernon R. Young. The requirements of human nutrition. Scientific American, 235(3):50–65, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n3/pdf/scientificamerican0976-50.pdf.

Anonymous:1976:SCi

[1346] Anonymous. Science and the citizen. Scientific American, 235(3):66–70, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n3/pdf/scientificamerican0976-66.pdf.

Janick:1976:CPA

[1347] Jules Janick, Carl H. Noller, and Charles L. Rhykerd. The cycles of plant and animal nutrition. Scientific American, 235(3):74–86, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n3/pdf/scientificamerican0976-74.pdf.

Harlan:1976:PAN

[1348] Jack R. Harlan. The plants and animals that nourish man. Scientific American, 235(3):88–97, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v235/n3/pdf/scientificamerican0976-
88.pdf.

Loomis:1976:AS

[1349] Robert S. Loomis. Agricultural systems. Scientific American, 235 (3):98–105, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n3/pdf/scientificamerican0976-
98.pdf.

Heady:1976:AU

[1350] Earl O. Heady. The agriculture of the U.S. Scientific American, 235(3):106–127, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n3/pdf/scientificamerican0976-
106.pdf.

Wellhausen:1976:AM

[1351] Edwin J. Wellhausen. The agriculture of Mexico. Scientific American, 235(3):128–150, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n3/pdf/scientificamerican0976-
128.pdf.

Mellor:1976:AI

[1352] John W. Mellor. The agriculture of India. Scientific American, 235(3):154–163, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n3/pdf/scientificamerican0976-
154.pdf.

Revelle:1976:RAA

[1353] Roger Revelle. The resources available for agriculture. Scientific American, 235(3):164–178, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n3/pdf/scientificamerican0976-
164.pdf.

Jennings:1976:AAP

[1354] Peter R. Jennings. The amplification of agricultural production. Scientific American, 235(3):180–194, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v235/n3/pdf/scientificamerican0976-
180.pdf.

Hopper:1976:DAD

[1355] W. David Hopper. The development of agriculture in developing countries. Scientific American, 235(3):196–205, September 1976. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n3/pdf/scientificamerican0976-196.pdf.

Gardner:1976:MGi

[1356] Martin Gardner. Mathematical games. Scientific American, 235 (3):206–211, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n3/pdf/scientificamerican0976-
206.pdf.

Morrison:1976:Bi

[1357] Philip Morrison. Books. Scientific American, 235(3):212–218, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n3/pdf/scientificamerican0976-212.pdf.

Anonymous:1976:Bi

[1358] Anonymous. Bibliography. Scientific American, 235(3):220–222, September 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n3/pdf/scientificamerican0976-220.pdf.

Anonymous:1976:Lj

[1359] Anonymous. Letters. Scientific American, 235(4):7–12, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n4/pdf/scientificamerican1076-7.pdf.

Anonymous:1976:YAj

[1360] Anonymous. 50 and 100 years ago. Scientific American, 235(4):14–17, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n4/pdf/scientificamerican1076-14.pdf.

Anonymous:1976:Aj

[1361] Anonymous. The authors. Scientific American, 235(4):20–21, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n4/pdf/scientificamerican1076-20.pdf.

Henderson:1976:ES

[1362] Donald A. Henderson. The eradication of smallpox. Scientific American, 235(4):25–33, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n4/pdf/scientificamerican1076-
25.pdf.

Chalmers:1976:PGE

[1363] Bruce Chalmers. The photovoltaic generation of electricity. Scientific American, 235(4):34–43, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n4/pdf/scientificamerican1076-
34.pdf.

Engelman:1976:NSS

[1364] Donald M. Engelman and Peter B. Moore. Neutron-scattering studies of the ribosome. Scientific American, 235(4):44–54, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n4/pdf/scientificamerican1076-44.pdf.

Anonymous:1976:SCj

[1365] Anonymous. Science and the citizen. Scientific American, 235(4):57–65, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n4/pdf/scientificamerican1076-57.pdf.

Strong:1976:CGB

[1366] I. B. Strong and R. W. Klebesadel. Cosmic gamma-ray bursts. Scientific American, 235(4):66–70, 75–90, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Strong:1976:CGR

[1367] Ian B. Strong and Ray W. Klebesadel. Cosmic gamma-ray bursts. Scientific American, 235(4):66–79, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v235/n4/pdf/scientificamerican1076-
66.pdf.

Leith:1976:WH

[1368] E. N. Leith. White-light holograms. Scientific American, 235(4):80– 88, 92–95, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Leith:1976:WLH

[1369] Emmett N. Leith. White-light holograms. Scientific American, 235(4):80–95, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n4/pdf/scientificamerican1076-
80.pdf.

Eaton:1976:SOJ

[1370] G. Gray Eaton. The social order of Japanese macaques. Scientific American, 235(4):96–106, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n4/pdf/scientificamerican1076-
96.pdf.

Idso:1976:DS

[1371] Sherwood B. Idso. Dust storms. Scientific American, 235(4):108– 111, 113–114, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n4/pdf/scientificamerican1076-
108.pdf.

Anonymous:1976:Bj

[1372] Anonymous. Bibliography. Scientific American, 235(4):114, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n4/pdf/scientificamerican1076-114.pdf.

Beresford:1976:DMV

[1373] Maurice Beresford. A deserted Medieval village in England. Scientific American, 235(4):116–129, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v235/n4/pdf/scientificamerican1076-
116.pdf.

Gardner:1976:MGj

[1374] Martin Gardner. Mathematical games. Scientific American, 235 (4):131–137, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n4/pdf/scientificamerican1076-
131.pdf.

Morrison:1976:Bj

[1375] Philip Morrison. Books. Scientific American, 235(4):138–143, October 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n4/pdf/scientificamerican1076-138.pdf.

Anonymous:1976:Lk

[1376] Anonymous. Letters. Scientific American, 235(5):7–11, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n5/pdf/scientificamerican1176-7.pdf.

Anonymous:1976:YAk

[1377] Anonymous. 50 and 100 years ago. Scientific American, 235(5):14–20, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n5/pdf/scientificamerican1176-14.pdf.

Anonymous:1976:Ak

[1378] Anonymous. The authors. Scientific American, 235(5):22–25, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n5/pdf/scientificamerican1176-22.pdf.

Drell:1976:LNW

[1379] Sidney D. Drell and Frank von Hippel. Limited nuclear war. Scientific American, 235(5):27–37, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v235/n5/pdf/scientificamerican1176-
27.pdf.

Bower:1976:RPC

[1380] T. G. R. Bower. Repetitive processes in child development. Scientific American, 235(5):38–47, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v235/n5/pdf/scientificamerican1176-
38.pdf.

Nambu:1976:CQ

[1381] Yoichiro Nambu. The confinement of quarks. Scientific American, 235(5):48–61, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n5/pdf/scientificamerican1176-
48.pdf.

Anonymous:1976:SCk

[1382] Anonymous. Science and the citizen. Scientific American, 235(5):64–70, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n5/pdf/scientificamerican1176-64.pdf.

McKenzie:1976:CCE

[1383] D. P. McKenzie and Frank Richter. Convection currents in the Earth’s mantle. Scientific American, 235(5):72–74, 79–84, 89, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n5/pdf/scientificamerican1176-72.pdf.

Glickstein:1976:VCP

[1384] Mitchell Glickstein and Alan R. Gibson. Visual cells in the pons of the brain. Scientific American, 235(5):90–98, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n5/pdf/scientificamerican1176-90.pdf.

Hecker:1976:FSM

[1385] S. S. Hecker and A. K. Ghosh. The forming of sheet metal. Scientific American, 235(5):100–108, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v235/n5/pdf/scientificamerican1176-
100.pdf.

Elias:1976:UT

[1386] Thomas S. Elias and Howard S. Irwin. Urban trees. Scientific American, 235(5):110–121, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n5/pdf/scientificamerican1176-
110.pdf.

Tuck:1976:AIB

[1387] James A. Tuck and Robert J. McGhee. An archaic Indian burial mound in Labrador. Scientific American, 235(5):122–129, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n5/pdf/scientificamerican1176-122.pdf.

Gardner:1976:MGk

[1388] Martin Gardner. Mathematical games. Scientific American, 235 (5):132–137, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n5/pdf/scientificamerican1176-
132.pdf.

Morrison:1976:Bk

[1389] Philip Morrison. Books. Scientific American, 235(5):138–149, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n5/pdf/scientificamerican1176-138.pdf.

Anonymous:1976:Bk

[1390] Anonymous. Bibliography. Scientific American, 235(5):150, November 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n5/pdf/scientificamerican1176-150.pdf.

Anonymous:1976:Ll

[1391] Anonymous. Letters. Scientific American, 235(6):8–11, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n6/pdf/scientificamerican1276-8.pdf.

Anonymous:1976:YAl

[1392] Anonymous. 50 and 100 years ago. Scientific American, 235(6):14–17, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n6/pdf/scientificamerican1276-14.pdf.

Anonymous:1976:Al

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n6/pdf/scientificamerican1276-20.pdf.

Ginzberg:1976:PEU

[1394] Eli Ginzberg. The pluralistic economy of the U.S. Scientific American, 235(6):25–29, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n6/pdf/scientificamerican1276-
25.pdf.

Bebbington:1976:RNF

[1395] William P. Bebbington. The reprocessing of nuclear fuels. Scientific American, 235(6):30–41, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v235/n6/pdf/scientificamerican1276-
30.pdf.

Favreau:1976:NAV

[1396] Olga Eizner Favreau and Michael C. Corballis. Negative aftereffects in visual perception. Scientific American, 235(6):42–49, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n6/pdf/scientificamerican1276-42.pdf.

Anonymous:1976:SCl

[1397] Anonymous. Science and the citizen. Scientific American, 235(6):50–54, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n6/pdf/scientificamerican1276-50.pdf.

Mermin:1976:SH

[1398] N. David Mermin and David M. Lee. Superfluid helium 3. Scientific American, 235(6):56–60, 62, 64, 67–68, 70–71, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n6/pdf/scientificamerican1276-56.pdf.

Pearson:1976:CW

[1399] Keir Pearson. The control of walking. Scientific American, 235(6):72–86, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n6/pdf/scientificamerican1276-72.pdf.

Kirshner:1976:SOG

[1400] Rober P. Kirshner. Supernovas in other galaxies. Scientific American, 235(6):88–95, 97–98, 100–101, December 1976. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n6/pdf/scientificamerican1276-88.pdf.

Campbell:1976:HVI

[1401] Allan M. Campbell. How viruses insert their DNA into the DNA of the host cell. Scientific American, 235(6):102–113, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v235/
n6/pdf/scientificamerican1276-102.pdf.

Macdougall:1976:FD

[1402] J. D. Macdougall. Fission-track dating. Scientific American, 235(6): 114–122, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Macdougall:1976:FTD

[1403] J. D. Macdougall. Fission-track dating. Scientific American, 235 (6):114–123, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n6/pdf/scientificamerican1276-
114.pdf.

Gardner:1976:MGl

[1404] Martin Gardner. Mathematical games. Scientific American, 235 (6):124–133, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v235/n6/pdf/scientificamerican1276-
124.pdf.

Philip:1976:B

[1405] Philip and Phylis Morrison. Books. Scientific American, 235(6):134–147, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n6/pdf/scientificamerican1276-134.pdf.

Anonymous:1976:AIa

[1406] Anonymous. Annual index. Scientific American, 235(6):148–151, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n6/pdf/scientificamerican1276-148.pdf.

Anonymous:1976:Bl

[1407] Anonymous. Bibliography. Scientific American, 235(6):152, December 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v235/n6/pdf/scientificamerican1276-152.pdf.

Anonymous:1977:La

[1408] Anonymous. Letters. Scientific American, 236(1):8–13, January 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n1/pdf/scientificamerican0177-8.pdf.

Anonymous:1977:YAa

[1409] Anonymous. 50 and 100 years ago. Scientific American, 236(1):14, January 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n1/pdf/scientificamerican0177-14.pdf.

Anonymous:1977:Aa

[1410] Anonymous. The authors. Scientific American, 236(1):16, January 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n1/pdf/scientificamerican0177-16.pdf.

Tietze:1977:LA

[1411] Christopher Tietze and Sarah Lewit. Legal abortion. Scientific American, 236(1):21–27, January 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n1/pdf/scientificamerican0177-
21.pdf.

Triplett:1977:AT

[1412] Glover B. Triplett, Jr. and David M. Van Doren, Jr. Agriculture without tillage. Scientific American, 236(1):28–33, January 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n1/pdf/scientificamerican0177-28.pdf.

Hawking:1977:QMB

[1413] S. W. Hawking. The quantum mechanics of black holes. Scientific American, 236(1):34–41, January 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n1/pdf/scientificamerican0177-
34.pdf.

Anonymous:1977:SCa

[1414] Anonymous. Science and the citizen. Scientific American, 236(1):43–49, January 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n1/pdf/scientificamerican0177-43.pdf.

Capra:1977:ACS

[1415] J. Donald Capra and Allen B. Edmundson. The antibody combining site. Scientific American, 236(1):50–59, January 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n1/pdf/scientificamerican0177-50.pdf.

Anonymous:1977:PMT

[1416] Anonymous. The perception of moving targets. Scientific American, 236 (1):60–64, 70–73, January 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Sekuler:1977:PMT

[1417] Robert Sekuler and Eugene Levinson. The perception of moving targets. Scientific American, 236(1):60–73, January 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n1/pdf/scientificamerican0177-60.pdf.

Rabinowicz:1977:E

[1418] Ernest Rabinowicz. Exoelectrons. Scientific American, 236(1):74–82, January 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n1/pdf/scientificamerican0177-74.pdf.

Hartmann:1977:CSS

[1419] William K. Hartmann. Cratering in the solar system. Scientific American, 236(1):84–86, 89–99, January 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n1/pdf/scientificamerican0177-
84.pdf.

Talmon:1977:S

[1420] Shemaryahu Talmon. The Samaritans. Scientific American, 236 (1):100–108, January 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n1/pdf/scientificamerican0177-
100.pdf.

Gardner:1977:MGa

[1421] Martin Gardner. Mathematical games. Scientific American, 236 (1):110–221, January 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n1/pdf/scientificamerican0177-
110.pdf.

Morrison:1977:Ba

[1422] Philip Morrison. Books. Scientific American, 236(1):122–131, January 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n1/pdf/scientificamerican0177-122.pdf.

Anonymous:1977:Ba

[1423] Anonymous. Bibliography. Scientific American, 236(1):132, January 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n1/pdf/scientificamerican0177-132.pdf.

Anonymous:1977:Lb

[1424] Anonymous. Letters. Scientific American, 236(2):8–10, February 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n2/pdf/scientificamerican0277-8.pdf.

Anonymous:1977:YAb

[1425] Anonymous. 50 and 100 years ago. Scientific American, 236(2):14, February 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n2/pdf/scientificamerican0277-14.pdf.

Anonymous:1977:Ab

[1426] Anonymous. The authors. Scientific American, 236(2):16, February 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n2/pdf/scientificamerican0277-16.pdf.

Tsipis:1977:CM

[1427] Kosta Tsipis. Cruise missiles. Scientific American, 236(2):20–29, February 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n2/pdf/scientificamerican0277-20.pdf.

Veverka:1977:PD

[1428] Joseph Veverka. Phobos and Deimos. Scientific American, 236(2):30–37, February 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n2/pdf/scientificamerican0277-30.pdf.

Birley:1977:FPR

[1429] Robin Birley. A frontier post in Roman Britain. Scientific American, 236(2):38–46, February 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n2/pdf/scientificamerican0277-
38.pdf.

Anonymous:1977:SCb

[1430] Anonymous. Science and the citizen. Scientific American, 236(2):48–56, February 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n2/pdf/scientificamerican0277-48.pdf.

Edelson:1977:GSC

[1431] Burton I. Edelson. Global satellite communications. Scientific American, 236(2):58–68, 73, February 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n2/pdf/scientificamerican0277-
58.pdf.

Benditt:1977:OA

[1432] Earl P. Benditt. The origin of atherosclerosis. Scientific American, 236(2):74–84, February 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n2/pdf/scientificamerican0277-
74.pdf.

Zare:1977:LSI

[1433] Richard N. Zare. Laser separation of isotopes. Scientific American, 236(2):86, 90–98, February 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n2/pdf/scientificamerican0277-
86.pdf.

Krauss:1977:SNS

[1434] Robert M. Krauss and Sam Glucksberg. Social and nonsocial speech. Scientific American, 236(2):100–105, February 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n2/pdf/scientificamerican0277-100.pdf.

Lester:1977:RA

[1435] Henry A. Lester. The response to acetylcholine. Scientific American, 236(2):106–116, 118, February 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n2/pdf/scientificamerican0277-
106.pdf.

Gardner:1977:MGb

[1436] Martin Gardner. Mathematical games. Scientific American, 236 (2):121–125, February 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n2/pdf/scientificamerican0277-
121.pdf.

Morrison:1977:Bb

[1437] Philip Morrison. Books. Scientific American, 236(2):128–136, February 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n2/pdf/scientificamerican0277-128.pdf.

Anonymous:1977:Bb

[1438] Anonymous. Bibliography. Scientific American, 236(2):138, February 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n2/pdf/scientificamerican0277-138.pdf.

Anonymous:1977:Lc

[1439] Anonymous. Letters. Scientific American, 236(3):8–10, March 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n3/pdf/scientificamerican0377-8.pdf.

Anonymous:1977:YAc

[1440] Anonymous. 50 and 100 years ago. Scientific American, 236(3):14–18, March 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n3/pdf/scientificamerican0377-14.pdf.

Anonymous:1977:Ac

[1441] Anonymous. The authors. Scientific American, 236(3):22–24, March 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n3/pdf/scientificamerican0377-22.pdf.

Vendryes:1977:SFB

[1442] G. A. Vendryes. Superphenix: a full-scale breeder reactor. Scientific American, 236(3):26–35, March 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Vendryes:1977:SFS

[1443] Georges A. Vendryes. Superphénix: A full-scale breeder reactor. Scientific American, 236(3):26–35, March 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v236/n3/pdf/scientificamerican0377-
26.pdf.

Gosling:1977:WSW

[1444] J. T. Gosling and A. J. Hundhausen. Waves in the solar wind. Scientific American, 236(3):36–43, March 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n3/pdf/scientificamerican0377-
36.pdf.

Snyder:1977:ORI

[1445] Solomon H. Snyder. Opiate receptors and internal opiates. Scientific American, 236(3):44–57, March 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n3/pdf/scientificamerican0377-
44.pdf.

Anonymous:1977:SCc

[1446] Anonymous. Science and the citizen. Scientific American, 236(3):58–64, March 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n3/pdf/scientificamerican0377-58.pdf.

Brill:1977:BNF

[1447] Winston J. Brill. Biological nitrogen fixation. Scientific American, 236(3):68–81, March 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n3/pdf/scientificamerican0377-
68.pdf.

Sundberg:1977:ASV

[1448] Johan Sundberg. The acoustics of the singing voice. Scientific American, 236(3):82–84, 86, 88–91, March 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n3/pdf/scientificamerican0377-
82.pdf.

Moorbath:1977:ORG

[1449] Stephen Moorbath. The oldest rocks and the growth of continents. Scientific American, 236(3):92–104, March 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n3/pdf/scientificamerican0377-92.pdf.

McCosker:1977:FF

[1450] John E. McCosker. Flashlight fishes. Scientific American, 236(3):106– 114, March 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n3/pdf/scientificamerican0377-106.pdf.

Hammond:1977:EM

[1451] Norman Hammond. The earliest Maya. Scientific American, 236(3):116– 133, March 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n3/pdf/scientificamerican0377-116.pdf.

Gardner:1977:MGc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n3/pdf/scientificamerican0377-134.pdf.

Morrison:1977:Bc

[1453] Philip Morrison. Books. Scientific American, 236(3):142–148, March 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n3/pdf/scientificamerican0377-142.pdf.

Anonymous:1977:Bc

[1454] Anonymous. Bibliography. Scientific American, 236(3):150, March 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n3/pdf/scientificamerican0377-150.pdf.

Anonymous:1977:Ld

[1455] Anonymous. Letters. Scientific American, 236(4):8, April 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n4/pdf/scientificamerican0477-8.pdf.

Anonymous:1977:YAd

[1456] Anonymous. 50 and 100 years ago. Scientific American, 236(4):10–14, April 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n4/pdf/scientificamerican0477-10.pdf.

Anonymous:1977:Ad

[1457] Anonymous. The authors. Scientific American, 236(4):20, April 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n4/pdf/scientificamerican0477-20.pdf.

Drake:1977:ILN

[1458] Elisabeth Drake and Robert C. Reid. The importation of liquefied natural gas. Scientific American, 236(4):22–29, April 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n4/pdf/scientificamerican0477-22.pdf.

Molnar:1977:CBI

[1459] Peter Molnar and Paul Tapponnier. The collision between India and Eurasia. Scientific American, 236(4):30–41, April 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n4/pdf/scientificamerican0477-30.pdf.

Burke:1977:SI

[1460] Derek C. Burke. The status of Interferon. Scientific American, 236(4):42– 50, April 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n4/pdf/scientificamerican0477-42.pdf.

Anonymous:1977:SCd

[1461] Anonymous. Science and the citizen. Scientific American, 236(4):52–60, April 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n4/pdf/scientificamerican0477-52.pdf.

Knuth:1977:A

[1462] Donald E. Knuth. Algorithms. Scientific American, 236(4):63–66, 69–72, 79–78, 80, April 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n4/pdf/scientificamerican0477-
63.pdf.

Inoue:1977:B

[1463] Shinya Inoué and Kayo Okazaki. Biocrystals. Scientific American, 236(4):82–92, April 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n4/pdf/scientificamerican0477-
82.pdf.

Abt:1977:CSS

[1464] Helmut A. Abt. The companions of sunlike stars. Scientific American, 236(4):96–104, April 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n4/pdf/scientificamerican0477-
96.pdf.

Phillipson:1977:SBL

[1465] D. W. Phillipson. The spread of the Bantu language. Scientific American, 236(4):106–114, April 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n4/pdf/scientificamerican0477-
106.pdf.

Nussenzveig:1977:TR

[1466] H. Moysés Nussenzveig. The theory of the rainbow. Scientific American, 236(4):116–127, April 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n4/pdf/scientificamerican0477-
116.pdf.

Gardner:1977:MGd

[1467] Martin Gardner. Mathematical games. Scientific American, 236(4):129– 136, April 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n4/pdf/scientificamerican0477-129.pdf.

Morrison:1977:Bd

[1468] Philip Morrison. Books. Scientific American, 236(4):140–146, April 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n4/pdf/scientificamerican0477-140.pdf.

Anonymous:1977:Bd

[1469] Anonymous. Bibliography. Scientific American, 236(4):148, April 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n4/pdf/scientificamerican0477-148.pdf.

Anonymous:1977:Le

[1470] Anonymous. Letters. Scientific American, 236(5):8–10, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n5/pdf/scientificamerican0577-8.pdf.

Anonymous:1977:YAe

[1471] Anonymous. 50 and 100 years ago. Scientific American, 236(5):14, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n5/pdf/scientificamerican0577-14.pdf.

Anonymous:1977:Ae

[1472] Anonymous. The authors. Scientific American, 236(5):15–18, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n5/pdf/scientificamerican0577-15.pdf.

Ambroggi:1977:URC

[1473] Robert P. Ambroggi. Underground reservoirs to control the water cycle. Scientific American, 236(5):21–27, May 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n5/pdf/scientificamerican0577-21.pdf.

Simons:1977:R

[1474] Elwyn L. Simons. Ramapithecus. Scientific American, 236(5):28–35, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n5/pdf/scientificamerican0577-28.pdf.

Adler:1977:AD

[1475] D. Adler. Amorphous-semiconductor devices. Scientific American, 236 (5):36–48, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Adler:1977:ASD

[1476] David Adler. Amorphous-semiconductor devices. Scientific American, 236(5):36–48, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n5/pdf/scientificamerican0577-
36.pdf.

Anonymous:1977:SCe

[1477] Anonymous. Science and the citizen. Scientific American, 236(5):50– 61, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n5/pdf/scientificamerican0577-50.pdf.

Old:1977:CI

[1478] Lloyd J. Old. Cancer immunology. Scientific American, 236(5):62– 79, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n5/pdf/scientificamerican0577-62.pdf.

Eddy:1977:CMS

[1479] John A. Eddy. The case of the missing sunspots. Scientific American, 236(5):80–92, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n5/pdf/scientificamerican0577-
80.pdf.

Altschul:1977:EH

[1480] Siri von Reis Altschul. Exploring the herbarium. Scientific American, 236(5):96–104, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n5/pdf/scientificamerican0577-
96.pdf.

Lore:1977:RS

[1481] Richard Lore and Kevin Flannelly. Rat societies. Scientific American, 236(5):106–116, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n5/pdf/scientificamerican0577-
106.pdf.

Efron:1977:SPS

[1482] Bradley Efron and Carl Morris. Stein’s paradox in statistics. Scientific American, 236(5):119–127, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n5/pdf/scientificamerican0577-
119.pdf.

Gardner:1977:PST

[1483] M. Gardner. The “jump proof” and its similarity to the toppling of a row of dominoes. Scientific American, 236(5):128, 131–132, 134–135, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gardner:1977:MGe

[1484] Martin Gardner. Mathematical games. Scientific American, 236(5):128– 135, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n5/pdf/scientificamerican0577-128.pdf.

Morrison:1977:Be

[1485] Philip Morrison. Books. Scientific American, 236(5):139–145, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n5/pdf/scientificamerican0577-139.pdf.

Anonymous:1977:Be

[1486] Anonymous. Bibliography. Scientific American, 236(5):146, May 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n5/pdf/scientificamerican0577-146.pdf.

Anonymous:1977:Lf

[1487] Anonymous. Letters. Scientific American, 236(6):10, June 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n6/pdf/scientificamerican0677-10.pdf.

Anonymous:1977:YAf

[1488] Anonymous. 50 and 100 years ago. Scientific American, 236(6):16, June 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n6/pdf/scientificamerican0677-16.pdf.

Anonymous:1977:Af

[1489] Anonymous. The authors. Scientific American, 236(6):18, June 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n6/pdf/scientificamerican0677-18.pdf.

Cohen:1977:DRW

[1490] Bernard L. Cohen. The disposal of radioactive wastes from fission reactors. Scientific American, 236(6):21–31, June 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n6/pdf/scientificamerican0677-21.pdf.

Rowe:1977:USR

[1491] Ednor M. Rowe and John H. Weaver. The uses of synchrotron radiation. Scientific American, 236(6):32–41, June 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n6/pdf/scientificamerican0677-32.pdf.

Jannasch:1977:MLD

[1492] Holger W. Jannasch and Carl O. Wirsen. Microbial life in the deep sea. Scientific American, 236(6):42–52, June 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n6/pdf/scientificamerican0677-42.pdf.

Anonymous:1977:SCf

[1493] Anonymous. Science and the citizen. Scientific American, 236(6):54–64, June 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n6/pdf/scientificamerican0677-54.pdf.

Dickman:1977:BG

[1494] Robert L. Dickman. Bok globules. Scientific American, 236(6):66–81, June 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n6/pdf/scientificamerican0677-66.pdf.

Olton:1977:SM

[1495] David S. Olton. Spatial memory. Scientific American, 236(6):82–98, June 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n6/pdf/scientificamerican0677-82.pdf.

Silverman:1977:LRF

[1496] William A. Silverman. The lesson of retrolental fibroplasia. Scientific American, 236(6):100–107, June 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v236/n6/pdf/scientificamerican0677-
100.pdf.

Sharon:1977:L

[1497] Nathan Sharon. Lectins. Scientific American, 236(6):108–119, June 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n6/pdf/scientificamerican0677-108.pdf.

Byard:1977:PRC

[1498] Margaret M. Byard. Poetic responses to the Copernican Revolution. Scientific American, 236(6):120–129, June 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n6/pdf/scientificamerican0677-120.pdf.

Gardner:1977:MGf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v236/n6/pdf/scientificamerican0677-131.pdf.

Morrison:1977:Bf

[1500] Philip Morrison. Books. Scientific American, 236(6):136–141, June 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n6/pdf/scientificamerican0677-136.pdf.

Anonymous:1977:Bf

[1501] Anonymous. Bibliography. Scientific American, 236(6):142, June 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v236/
n6/pdf/scientificamerican0677-142.pdf.

Anonymous:1977:Lg

[1502] Anonymous. Letters. Scientific American, 237(1):8–9, July 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n1/pdf/scientificamerican0777-8.pdf.

Anonymous:1977:YAg

[1503] Anonymous. 50 and 100 years ago. Scientific American, 237(1):10– 14, July 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n1/pdf/scientificamerican0777-10.pdf.

Anonymous:1977:Ag

[1504] Anonymous. The authors. Scientific American, 237(1):18–20, July 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n1/pdf/scientificamerican0777-18.pdf.

Grobstein:1977:RDD

[1505] Clifford Grobstein. The recombinant-DNA debate. Scientific American, 237(1):22–33, July 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n1/pdf/scientificamerican0777-
22.pdf.

Leovy:1977:AM

[1506] Conway B. Leovy. The atmosphere of Mars. Scientific American, 237(1): 34–43, July 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n1/pdf/scientificamerican0777-34.pdf.

Melnick:1977:VH

[1507] Joseph L. Melnick, Gordon R. Dreesman, and F. Blaine Hollinger. Viral hepatitis. Scientific American, 237(1):44–52, July 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n1/pdf/scientificamerican0777-44.pdf.

Anonymous:1977:SCg

[1508] Anonymous. Science and the citizen. Scientific American, 237(1):56– 62, July 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n1/pdf/scientificamerican0777-56.pdf.

Bryant:1977:BRP

[1509] Peter J. Bryant, Susan V. Bryant, and Vernon French. Biological regeneration and pattern formation. Scientific American, 237(1):66–81, July 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n1/pdf/scientificamerican0777-66.pdf. Discusses

the role of cellular interactions in development.

Double:1977:SC

[1510] D. D. Double and A. Hellawell. The solidification of cement. Scientific American, 237(1):82–90, July 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n1/pdf/scientificamerican0777-
82.pdf.

Dye:1977:AAM

[1511] James L. Dye. Anions of the alkali metals. Scientific American, 237(1):92, 96, 98, 100, 102, 104–105, July 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n1/pdf/scientificamerican0777-92.pdf.

Horridge:1977:CEI

[1512] G. Adrian Horridge. The compound eye of insects. Scientific American, 237(1):108–120, July 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n1/pdf/scientificamerican0777-
108.pdf.

Stewart:1977:G

[1513] Ian Stewart. Gauss. Scientific American, 237(1):122–131, July 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n1/pdf/scientificamerican0777-122.pdf.

Gardner:1977:MGg

[1514] Martin Gardner. Mathematical games. Scientific American, 237(1):132– 137, July 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n1/pdf/scientificamerican0777-132.pdf.

Walker:1977:ASa

[1515] Jearl Walker. The amateur scientist. Scientific American, 237(1):138– 144, July 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n1/pdf/scientificamerican0777-138.pdf.

Morrison:1977:Bg

[1516] Philip Morrison. Books. Scientific American, 237(1):146–153, July 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n1/pdf/scientificamerican0777-146.pdf.

Anonymous:1977:Bg

[1517] Anonymous. Bibliography. Scientific American, 237(1):154, July 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n1/pdf/scientificamerican0777-154.pdf.

Anonymous:1977:YAh

[1518] Anonymous. 50 and 100 years ago. Scientific American, 237(2):18, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n2/pdf/scientificamerican0877-18.pdf.

Anonymous:1977:Ah

[1519] Anonymous. The authors. Scientific American, 237(2):20, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n2/pdf/scientificamerican0877-20.pdf.

Scoville:1977:SN

[1520] Herbert Scoville, Jr. The SALT negotiations. Scientific American, 237(2):24–31, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n2/pdf/scientificamerican0877-
24.pdf.

Disney:1977:BLO

[1521] Michael J. Disney and Philippe Véron. BL Lacertae objects. Scientific American, 237(2):32–39, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n2/pdf/scientificamerican0877-
32.pdf.

Boyle:1977:LC

[1522] W. S. Boyle. Light-wave communications. Scientific American, 237(2): 40–48, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Boyle:1977:LWC

[1523] W. S. Boyle. Light-wave communications. Scientific American, 237(2):40–48, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n2/pdf/scientificamerican0877-
40.pdf.

Anonymous:1977:SCh

[1524] Anonymous. Science and the citizen. Scientific American, 237(2):52–58, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n2/pdf/scientificamerican0877-52.pdf.

Pollack:1977:FHE

[1525] Henry N. Pollack and David S. Chapman. The flow of heat from the Earth’s interior. Scientific American, 237(2):60–68, 73–76, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n2/pdf/scientificamerican0877-60.pdf.

Dawson:1977:K

[1526] T. J. Dawson. Kangaroos. Scientific American, 237(2):78–89, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n2/pdf/scientificamerican0877-78.pdf.

Walsby:1977:GVB

[1527] A. E. Walsby. The gas vacuoles of blue-green algae. Scientific American, 237(2):90–97, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n2/pdf/scientificamerican0877-
90.pdf.

Irwin:1977:HAC

[1528] Howard S. Irwin. The history of the airflow car. Scientific American, 237(2):98–106, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n2/pdf/scientificamerican0877-
98.pdf.

Nathanson:1977:MB

[1529] J. A. Nathanson and P. Greengard. ‘Second messengers’ in the brain. Scientific American, 237(2):108–119, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Nathanson:1977:SMB

[1530] James A. Nathanson and Paul Greengard. “Second messengers” in the brain. Scientific American, 237(2):108–119, August 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n2/pdf/scientificamerican0877-108.pdf.

Gardner:1977:MGN

[1531] Martin Gardner. Mathematical games: A new kind of cipher that would take millions of years to break. Scientific American, 237(2):120–124, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n2/pdf/scientificamerican0877-120.pdf.

Walker:1977:ASb

[1532] Jearl Walker. The amateur scientist. Scientific American, 237 (2):126–131, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n2/pdf/scientificamerican0877-
126.pdf.

Morrison:1977:Bh

[1533] Philip Morrison. Books. Scientific American, 237(2):132–138, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n2/pdf/scientificamerican0877-132.pdf.

Anonymous:1977:Bh

[1534] Anonymous. Bibliography. Scientific American, 237(2):140, August 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n2/pdf/scientificamerican0877-140.pdf.

Anonymous:1977:Lh

[1535] Anonymous. Letters. Scientific American, 237(3):8–9, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n3/pdf/scientificamerican0977-8.pdf.

Anonymous:1977:YAi

[1536] Anonymous. 50 and 100 years ago. Scientific American, 237(3):13, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n3/pdf/scientificamerican0977-13.pdf.

Anonymous:1977:Ai

[1537] Anonymous. The authors. Scientific American, 237(3):14–23, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n3/pdf/scientificamerican0977-14.pdf.

Gardner:1977:MGi

[1538] Martin Gardner. Mathematical games. Scientific American, 237 (3):24–42, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n3/pdf/scientificamerican0977-
24.pdf.

Morrison:1977:Bi

[1539] Philip Morrison. Books. Scientific American, 237(3):46–61, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n3/pdf/scientificamerican0977-46.pdf.

Noyce:1977:M

[1540] Robert N. Noyce. Microelectronics. Scientific American, 237(3):62– 69 (or 63–69??), September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n3/pdf/scientificamerican0977-
62.pdf.

Various:1977:SIM

[1541] Various authors. Special issue on microelectronics. Scientific American, 237(3):62–245, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Meindl:1977:MCE

[1542] James D. Meindl. Microelectronic circuit elements. Scientific American, 237(3):70–81, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n3/pdf/scientificamerican0977-
70.pdf.

Holton:1977:LIM

[1543] W. C. Holton. The large-scale integration of microelectronic circuits. Scientific American, 237(3):82–94, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holton:1977:LSI

[1544] William C. Holton. The large-scale integration of microelectronic circuits. Scientific American, 237(3):82–94, September 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n3/pdf/scientificamerican0977-82.pdf.

Anonymous:1977:SCi

[1545] Anonymous. Science and the citizen. Scientific American, 237(3):96–106, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n3/pdf/scientificamerican0977-96.pdf.

Oldham:1977:FMC

[1546] William G. Oldham. The fabrication of microelectronic circuits. Scientific American, 237(3):110–128, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v237/n3/pdf/scientificamerican0977-
110.pdf.

Hodges:1977:MM

[1547] David A. Hodges. Microelectronic memories. Scientific American, 237(3):130–131, 135–136, 138–140, 142, 144–145, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n3/pdf/scientificamerican0977-130.pdf.

Toong:1977:M

[1548] Hoo-Min D. Toong. Microprocessors. Scientific American, 237(3):146– 147, 151–156, 158, 160–161, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v237/n3/pdf/scientificamerican0977-
146.pdf.

Terman:1977:RMD

[1549] Lewis M. Terman. The role of microelectronics in data processing. Scientific American, 237(3):162–177, September 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n3/pdf/scientificamerican0977-162.pdf.

Oliver:1977:RMI

[1550] Bernard M. Oliver. The role of microelectronics in instrumentation and control. Scientific American, 237(3):180–181, 183–184, 186–188, 190, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n3/pdf/scientificamerican0977-180.pdf.

Mayo:1977:RMC

[1551] John S. Mayo. The role of microelectronics in communication. Scientific American, 237(3):192–193, 196–198, 200, 202, 205–206, 208–209, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n3/pdf/scientificamerican0977-192.pdf.

Sutherland:1977:MCS

[1552] Ivan E. Sutherland and Carver A. Mead. Micro-electronics and computer science. Scientific American, 237(3):210–212, 214, 218–220, 222, 224, 226, 228, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n3/pdf/scientificamerican0977-
210.pdf.

Kay:1977:MPC

[1553] Alan C. Kay. Microelectronics and the personal computer. Scientific American, 237(3):230–244, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v237/n3/pdf/scientificamerican0977-
230.pdf.

Walker:1977:ASc

[1554] Jearl Walker. The amateur scientist. Scientific American, 237 (3):246–257, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n3/pdf/scientificamerican0977-
246.pdf.

Anonymous:1977:Bi

[1555] Anonymous. Bibliography. Scientific American, 237(3):258–262, September 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n3/pdf/scientificamerican0977-258.pdf.

Anonymous:1977:Li

[1556] Anonymous. Letters. Scientific American, 237(4):7–12, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n4/pdf/scientificamerican1077-7.pdf.

Anonymous:1977:YAj

[1557] Anonymous. 50 and 100 years ago. Scientific American, 237(4):14, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n4/pdf/scientificamerican1077-14.pdf.

Anonymous:1977:Aj

[1558] Anonymous. The authors. Scientific American, 237(4):15, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n4/pdf/scientificamerican1077-15.pdf.

Gardner:1977:MGj

[1559] Martin Gardner. Mathematical games. Scientific American, 237(4):18– 25, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n4/pdf/scientificamerican1077-18.pdf.

Morrison:1977:Bj

[1560] Philip Morrison. Books. Scientific American, 237(4):26–32, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n4/pdf/scientificamerican1077-26.pdf.

Cole:1977:PRS

[1561] Stephen Cole, Leonard Rubin, and Jonathan R. Cole. Peer review and the support of science. Scientific American, 237(4):34–41, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n4/pdf/scientificamerican1077-34.pdf.

Clark:1977:XRS

[1562] George W. Clark. X-ray stars in globular clusters. Scientific American, 237(4):42–55, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n4/pdf/scientificamerican1077-
42.pdf.

Clark:1977:XSG

[1563] G. W. Clark. X-ray stars in globular clusters. Scientific American, 237 (4):42–55, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schwitters:1977:FPC

[1564] Roy F. Schwitters. Fundamental particles with Charm. Scientific American, 237(4):56–70, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n4/pdf/scientificamerican1077-
56.pdf.

Anonymous:1977:SCj

[1565] Anonymous. Science and the citizen. Scientific American, 237(4):74–82, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n4/pdf/scientificamerican1077-74.pdf.

Jensen:1977:SAR

[1566] H. Jensen, L. C. Graham, L. J. Porcello, and E. N. Leith. Side-looking airborne radar. Scientific American, 237(4):84–95, October 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Jensen:1977:SLA

[1567] Homer Jensen, L. C. Graham, Leonard J. Porcello, and Emmett N. Leith. Side-looking airborne radar. Scientific American, 237(4):84–95, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n4/pdf/scientificamerican1077-84.pdf.

Cunningham:1977:SFH

[1568] Bruce A. Cunningham. The structure and function of histocompatibility antigens. Scientific American, 237(4):96–107, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n4/pdf/scientificamerican1077-96.pdf.

Appel:1977:SFC

[1569] Kenneth Appel and Wolfgang Haken. The solution of the four-color-map problem. Scientific American, 237(4):108–121, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n4/pdf/scientificamerican1077-108.pdf.

Maddin:1977:HIA

[1570] Robert Maddin, James D. Muhly, and Tamara S. Wheeler. How the iron age began. Scientific American, 237(4):122–131, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n4/pdf/scientificamerican1077-122.pdf.

Siegel:1977:H

[1571] Ronald K. Siegel. Hallucinations. Scientific American, 237(4):132–139, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n4/pdf/scientificamerican1077-132.pdf.

Walker:1977:ASd

[1572] Jearl Walker. The amateur scientist. Scientific American, 237 (4):142–150, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n4/pdf/scientificamerican1077-
142.pdf.

Anonymous:1977:Bj

[1573] Anonymous. Bibliography. Scientific American, 237(4):152, October 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n4/pdf/scientificamerican1077-152.pdf.

Anonymous:1977:Lj

[1574] Anonymous. Letters. Scientific American, 237(5):6–10, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n5/pdf/scientificamerican1177-6.pdf.

Anonymous:1977:YAk

[1575] Anonymous. 50 and 100 years ago. Scientific American, 237(5):12–14, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n5/pdf/scientificamerican1177-12.pdf.

Anonymous:1977:Ak

[1576] Anonymous. The authors. Scientific American, 237(5):15, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n5/pdf/scientificamerican1177-15.pdf.

Gardner:1977:MGk

[1577] Martin Gardner. Mathematical games. Scientific American, 237 (5):18–28, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n5/pdf/scientificamerican1177-
18.pdf.

Morrison:1977:Bk

[1578] Philip Morrison. Books. Scientific American, 237(5):30–38, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n5/pdf/scientificamerican1177-30.pdf.

Ginzberg:1977:JP

[1579] Eli Ginzberg. The job problem. Scientific American, 237(5):43–51, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n5/pdf/scientificamerican1177-43.pdf.

Horowitz:1977:SLM

[1580] Norman H. Horowitz. The search for life on Mars. Scientific American, 237(5):52–61, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n5/pdf/scientificamerican1177-
52.pdf.

Shoji:1977:DI

[1581] Kobe Shoji. Drip irrigation. Scientific American, 237(5):62–68, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n5/pdf/scientificamerican1177-62.pdf.

Anonymous:1977:SCk

[1582] Anonymous. Science and the citizen. Scientific American, 237(5):70–75, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n5/pdf/scientificamerican1177-70.pdf.

Groth:1977:CG

[1583] Edward J. Groth, P. James E. Peebles, Michael Seldner, and Raymond M. Soneira. The clustering of Galaxies. Scientific American, 237(5):76, 84, 87–90, 95–98, November 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n5/pdf/scientificamerican1177-76.pdf.

Todd:1977:CC

[1584] Neil B. Todd. Cats and commerce. Scientific American, 237(5):100–107, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n5/pdf/scientificamerican1177-100.pdf.

Keeley:1977:FPF

[1585] Lawrence H. Keeley. The functions of paleolithic flint tools. Scientific American, 237(5):108–126, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v237/n5/pdf/scientificamerican1177-
108.pdf.

Epel:1977:PF

[1586] David Epel. The program of fertilization. Scientific American, 237(5):128–138, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n5/pdf/scientificamerican1177-
128.pdf.

Nef:1977:EEC

[1587] John U. Nef. An early energy crisis and its consequences. Scientific American, 237(5):140–151, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n5/pdf/scientificamerican1177-
140.pdf.

Walker:1977:ASe

[1588] Jearl Walker. The amateur scientist. Scientific American, 237 (5):152–160, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n5/pdf/scientificamerican1177-
152.pdf.

Anonymous:1977:Bk

[1589] Anonymous. Bibliography. Scientific American, 237(5):163–164, November 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n5/pdf/scientificamerican1177-163.pdf.

Anonymous:1977:Lk

[1590] Anonymous. Letters. Scientific American, 237(6):6–7, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n6/pdf/scientificamerican1277-6.pdf.

Anonymous:1977:YAl

[1591] Anonymous. 50 and 100 years ago. Scientific American, 237(6):8–9, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n6/pdf/scientificamerican1277-8.pdf.

Anonymous:1977:Al

[1592] Anonymous. The authors. Scientific American, 237(6):10–15, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n6/pdf/scientificamerican1277-10.pdf.

Gardner:1977:MGl

[1593] Martin Gardner. Mathematical games. Scientific American, 237 (6):17–24, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n6/pdf/scientificamerican1277-
17.pdf.

Philip:1977:B

[1594] Philip and Phylis Morrison. Books. Scientific American, 237(6):26–42, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n6/pdf/scientificamerican1277-26.pdf.

Montgomery:1977:PTC

[1595] G. Franklin Montgomery. Product technology and the consumer. Scientific American, 237(6):47–53, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v237/n6/pdf/scientificamerican1277-
47.pdf.

Fiddes:1977:NSV

[1596] John C. Fiddes. The nucleotide sequence of a viral DNA. Scientific American, 237(6):54–67, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n6/pdf/scientificamerican1277-
54.pdf.

Boore:1977:MGE

[1597] David M. Boore. The motion of the ground in earthquakes. Scientific American, 237(6):68–78, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v237/n6/pdf/scientificamerican1277-
68.pdf.

Anonymous:1977:SCl

[1598] Anonymous. Science and the citizen. Scientific American, 237(6):82–87, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n6/pdf/scientificamerican1277-82.pdf.

Kaplan:1977:EI

[1599] Martin M. Kaplan and Robert G. Webster. The epidemiology of influenza. Scientific American, 237(6):88–106, December 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n6/pdf/scientificamerican1277-88.pdf.

Land:1977:RTC

[1600] Edwin H. Land. The Retinex theory of color vision. Scientific American, 237(6):108–120, 122–123, 126, 128, December 1977. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v237/
n6/pdf/scientificamerican1277-108.pdf.

Harris:1977:D

[1601] William F. Harris. Disclinations. Scientific American, 237(6):130–136, 138–142, 144–145, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n6/pdf/scientificamerican1277-
130.pdf.

Holldobler:1977:WA

[1602] Berthold K. Hölldobler and Edward O. Wilson. Weaver ants. Scientific American, 237(6):146–154, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v237/n6/pdf/scientificamerican1277-
146.pdf.

Wainwright:1977:CFS

[1603] Geoffrey Wainwright. A Celtic farmstead in Southern Britain. Scientific American, 237(6):156–169, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v237/n6/pdf/scientificamerican1277-
156.pdf.

Walker:1977:ASf

[1604] Jearl Walker. The amateur scientist. Scientific American, 237 (6):172–180, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v237/n6/pdf/scientificamerican1277-
172.pdf.

Anonymous:1977:AIa

[1605] Anonymous. Annual index. Scientific American, 237(6):183–189, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n6/pdf/scientificamerican1277-183.pdf.

Anonymous:1977:Bl

[1606] Anonymous. Bibliography. Scientific American, 237(6):190, December 1977. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v237/n6/pdf/scientificamerican1277-190.pdf.

Anonymous:1978:La

[1607] Anonymous. Letters. Scientific American, 238(1):8, January 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n1/pdf/scientificamerican0178-8.pdf.

Anonymous:1978:YAa

[1608] Anonymous. 50 and 100 years ago. Scientific American, 238(1):10, January 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n1/pdf/scientificamerican0178-10.pdf.

Anonymous:1978:Aa

[1609] Anonymous. The authors. Scientific American, 238(1):12, January 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n1/pdf/scientificamerican0178-12.pdf.

Gardner:1978:MGa

[1610] Martin Gardner. Mathematical games. Scientific American, 238(1):14– 27, January 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n1/pdf/scientificamerican0178-14.pdf.

Morrison:1978:Ba

[1611] Philip Morrison. Books. Scientific American, 238(1):28–31, January 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n1/pdf/scientificamerican0178-28.pdf.

Woodwell:1978:CDQ

[1612] George M. Woodwell. The carbon dioxide question. Scientific American, 238(1):34–43, January 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n1/pdf/scientificamerican0178-
34.pdf.

Sonstegard:1978:SRH

[1613] David A. Sonstegard, Larry S. Matthews, and Herbert Kaufer. The surgical replacement of the human knee joint. Scientific American, 238(1):44– 51, January 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n1/pdf/scientificamerican0178-44.pdf.

Rich:1978:TDS

[1614] Alexander Rich and Sung Hou Kim. The three-dimensional structure of transfer RNA. Scientific American, 238(1):52–62, January 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n1/pdf/scientificamerican0178-52.pdf.

Anonymous:1978:SCa

[1615] Anonymous. Science and the citizen. Scientific American, 238(1):64–69, January 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n1/pdf/scientificamerican0178-64.pdf.

Heiles:1978:SIM

[1616] Carl Heiles. The structure of the interstellar medium. Scientific American, 238(1):74–84, January 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n1/pdf/scientificamerican0178-
74.pdf.

Costerton:1978:HBS

[1617] J. W. Costerton, G. G. Geesey, and K.-J. Cheng. How bacteria stick. Scientific American, 238(1):86–95, January 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n1/pdf/scientificamerican0178-86.pdf.

Lewis:1978:EA

[1618] Harry R. Lewis and Christos H. Papadimitriou. The efficiency of algorithms. Scientific American, 238(1):96–109, January 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n1/pdf/scientificamerican0178-96.pdf.

Humphrey:1978:RC

[1619] John H. Humphrey and John Griffiths Pedley. Roman Carthage. Scientific American, 238(1):110–120, January 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v238/n1/pdf/scientificamerican0178-
110.pdf.

Dunn-Rankin:1978:VCW

[1620] Peter Dunn-Rankin. The visual characteristics of words. Scientific American, 238(1):122–130, January 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n1/pdf/scientificamerican0178-
122.pdf.

Walker:1978:ASa

[1621] Jearl Walker. The amateur scientist. Scientific American, 238 (1):132–136, January 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n1/pdf/scientificamerican0178-
132.pdf.

Anonymous:1978:Ba

[1622] Anonymous. Bibliography. Scientific American, 238(1):138–139, January 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n1/pdf/scientificamerican0178-138.pdf.

Anonymous:1978:Lb

[1623] Anonymous. Letters. Scientific American, 238(2):8–9, February 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n2/pdf/scientificamerican0278-8.pdf.

Anonymous:1978:YAb

[1624] Anonymous. 50 and 100 years ago. Scientific American, 238(2):12–14, February 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n2/pdf/scientificamerican0278-12.pdf.

Anonymous:1978:Ab

[1625] Anonymous. The authors. Scientific American, 238(2):16, February 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n2/pdf/scientificamerican0278-16.pdf.

Gardner:1978:MGb

[1626] Martin Gardner. Mathematical games. Scientific American, 238 (2):19–32, February 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n2/pdf/scientificamerican0278-
19.pdf.

Morrison:1978:Bb

[1627] Philip Morrison. Books. Scientific American, 238(2):34–44, February 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n2/pdf/scientificamerican0278-34.pdf.

Bassuk:1978:DMH

[1628] Ellen L. Bassuk and Samuel Gerson. Deinstitutionalization and mental health services. Scientific American, 238(2):46–53, February 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n2/pdf/scientificamerican0278-46.pdf.

Bonatii:1978:OMD

[1629] E. Bonatii. The origin of metal deposits in the oceanic lithosphere. Scientific American, 238(2):54–61, February 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bonatti:1978:OMD

[1630] Enrico Bonatti. The origin of metal deposits in the oceanic lithosphere. Scientific American, 238(2):54–61, February 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n2/pdf/scientificamerican0278-54.pdf.

Nevins:1978:CA

[1631] J. L. Nevins and D. E. Whitney. Computer-controlled assembly. Scientific American, 238(2):62–74, February 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Nevins:1978:CCA

[1632] James L. Nevins and Daniel E. Whitney. Computer-controlled assembly. Scientific American, 238(2):62–74, February 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n2/pdf/scientificamerican0278-62.pdf.

Anonymous:1978:SCb

[1633] Anonymous. Science and the citizen. Scientific American, 238(2):76–90, February 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n2/pdf/scientificamerican0278-76.pdf.

Shepherd:1978:MNS

[1634] Gordon M. Shepherd. Microcircuits in the nervous system. Scientific American, 238(2):92–98, 100, 102–103, February 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n2/pdf/scientificamerican0278-92.pdf.

Heslop-Harrison:1978:CP

[1635] Yolande Heslop-Harrison. Carnivorous plants. Scientific American, 238(2):104–114, February 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n2/pdf/scientificamerican0278-
104.pdf.

Croce:1978:GHC

[1636] Carlo M. Croce and Hilary Koprowski. The genetics of human cancer. Scientific American, 238(2):117–125, February 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n2/pdf/scientificamerican0278-117.pdf.

Freedman:1978:SUL

[1637] Daniel Z. Freedman and Peter van Nieuwenhuizen. Supergravity and the unification of the laws of Physics. Scientific American, 238(2):126, 128–129, 131–139, 141–143, February 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n2/pdf/scientificamerican0278-126.pdf.

Bahadori:1978:PCS

[1638] Mehdi N. Bahadori. Passive cooling systems in Iranian architecture. Scientific American, 238(2):144–154, February 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n2/pdf/scientificamerican0278-144.pdf.

Walker:1978:ASb

[1639] Jearl Walker. The amateur scientist. Scientific American, 238 (2):156–161, February 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n2/pdf/scientificamerican0278-
156.pdf.

Anonymous:1978:Bb

[1640] Anonymous. Bibliography. Scientific American, 238(2):162, February 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n2/pdf/scientificamerican0278-162.pdf.

Anonymous:1978:Lc

[1641] Anonymous. Letters. Scientific American, 238(3):10–15, March 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n3/pdf/scientificamerican0378-10.pdf.

Anonymous:1978:YAc

[1642] Anonymous. 50 and 100 years ago. Scientific American, 238(3):18–20, March 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n3/pdf/scientificamerican0378-18.pdf.

Anonymous:1978:Ac

[1643] Anonymous. The authors. Scientific American, 238(3):22–23, March 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n3/pdf/scientificamerican0378-22.pdf.

Gardner:1978:MGc

[1644] Martin Gardner. Mathematical games. Scientific American, 238(3):25– 29, March 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n3/pdf/scientificamerican0378-25.pdf.

Morrison:1978:Bc

[1645] Philip Morrison. Books. Scientific American, 238(3):30–41, March 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n3/pdf/scientificamerican0378-30.pdf.

Flower:1978:WOP

[1646] Andrew R. Flower. World oil production. Scientific American, 238(3):42– 49, March 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n3/pdf/scientificamerican0378-42.pdf.

Perl:1978:HL

[1647] Martin L. Perl and William T. Kirk. Heavy leptons. Scientific American, 238(3):50–57, March 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n3/pdf/scientificamerican0378-
50.pdf.

Luff:1978:ET

[1648] Peter P. Luff. The electronic telephone. Scientific American, 238(3):58– 64, March 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n3/pdf/scientificamerican0378-58.pdf.

Anonymous:1978:SCc

[1649] Anonymous. Science and the citizen. Scientific American, 238(3):69–74, March 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n3/pdf/scientificamerican0378-69.pdf.

Arvidson:1978:SM

[1650] Raymond E. Arvidson, Alan B. Binder, and Kenneth L. Jones. The surface of Mars. Scientific American, 238(3):76–89, March 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n3/pdf/scientificamerican0378-76.pdf.

Gosz:1978:FEF

[1651] James R. Gosz, Richard T. Holmes, Gene E. Likens, and F. Herbert Bormann. The flow of energy in a forest ecosystem. Scientific American, 238(3):92–102, March 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n3/pdf/scientificamerican0378-
92.pdf.

Hinkle:1978:HCM

[1652] Peter C. Hinkle and Richard E. McCarty. How cells make ATP. Scientific American, 238(3):104–117, 121–123, March 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n3/pdf/scientificamerican0378-104.pdf.

Graham:1978:CMS

[1653] Ronald L. Graham. The combinatorial mathematics of scheduling. Scientific American, 238(3):124–132, March 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n3/pdf/scientificamerican0378-124.pdf.

Klein:1978:PBE

[1654] H. Arthur Klein. Pieter bruegel the elder as a guide to 16th-century technology. Scientific American, 238(3):134–140, March 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n3/pdf/scientificamerican0378-134.pdf.

Walker:1978:ASc

[1655] Jearl Walker. The amateur scientist. Scientific American, 238(3):142– 153, March 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n3/pdf/scientificamerican0378-142.pdf.

Anonymous:1978:Bc

[1656] Anonymous. Bibliography. Scientific American, 238(3):154, March 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n3/pdf/scientificamerican0378-154.pdf.

Anonymous:1978:Ld

[1657] Anonymous. Letters. Scientific American, 238(4):8, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n4/pdf/scientificamerican0478-8.pdf.

Anonymous:1978:YAd

[1658] Anonymous. 50 and 100 years ago. Scientific American, 238(4):11–12, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n4/pdf/scientificamerican0478-11.pdf.

Anonymous:1978:Ad

[1659] Anonymous. The authors. Scientific American, 238(4):14, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n4/pdf/scientificamerican0478-14.pdf.

Gardner:1978:MGd

[1660] Martin Gardner. Mathematical games. Scientific American, 238(4):16– 32, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n4/pdf/scientificamerican0478-16.pdf.

Gardner:1978:MGW

[1661] M. Gardner. Mathematical games: White and brown music, fractal curves and 1/f fluctuations. Scientific American, 238(4):16–32, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1978:Bd

[1662] Philip Morrison. Books. Scientific American, 238(4):34–40, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n4/pdf/scientificamerican0478-34.pdf.

Rose:1978:NPN

[1663] David J. Rose and Richard K. Lester. Nuclear power, nuclear weapons and international stability. Scientific American, 238(4):45–57, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n4/pdf/scientificamerican0478-45.pdf.

Fein:1978:MSS

[1664] Jack M. Fein. Microvascular surgery for stroke. Scientific American, 238(4):58–67, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n4/pdf/scientificamerican0478-
58.pdf.

Albrecht-Buehler:1978:TMC

[1665] Guenter Albrecht-Buehler. The tracks of moving cells. Scientific American, 238(4):68–76, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n4/pdf/scientificamerican0478-

68.pdf. Description of similarities in the behavior of the daughter cells resulting from a cell-division.

Anonymous:1978:SCd

[1666] Anonymous. Science and the citizen. Scientific American, 238(4):78–88, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n4/pdf/scientificamerican0478-78.pdf.

Isaac:1978:FSB

[1667] Glynn Isaac. The food-sharing behavior of protohuman hominids. Scientific American, 238(4):90–108, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v238/n4/pdf/scientificamerican0478-
90.pdf.

Zeilik:1978:BMS

[1668] Michael Zeilik. The birth of massive stars. Scientific American, 238(4):110–118, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n4/pdf/scientificamerican0478-
110.pdf.

Cox:1978:KP

[1669] Keith G. Cox. Kimberlite pipes. Scientific American, 238(4):120– 123, 125–128, 130, 132, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n4/pdf/scientificamerican0478-
120.pdf.

Milne:1978:IWS

[1670] Lorus J. Milne and Margery Milne. Insects of the water surface. Scientific American, 238(4):134–142, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n4/pdf/scientificamerican0478-
134.pdf.

Lynch:1978:AH

[1671] David K. Lynch. Atmospheric halos. Scientific American, 238(4):144– 152, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n4/pdf/scientificamerican0478-144.pdf.

Walker:1978:ASd

[1672] Jearl Walker. The amateur scientist. Scientific American, 238(4):154– 162, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n4/pdf/scientificamerican0478-154.pdf.

Anonymous:1978:Bd

[1673] Anonymous. Bibliography. Scientific American, 238(4):164, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n4/pdf/scientificamerican0478-164.pdf.

Anonymous:1978:Le

[1674] Anonymous. Letters. Scientific American, 238(5):10–12, May 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n5/pdf/scientificamerican0578-10.pdf.

Anonymous:1978:YAe

[1675] Anonymous. 50 and 100 years ago. Scientific American, 238(5):14– 15, May 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n5/pdf/scientificamerican0578-14.pdf.

Anonymous:1978:Ae

[1676] Anonymous. The authors. Scientific American, 238(5):18–23, May 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n5/pdf/scientificamerican0578-18.pdf.

Gardner:1978:MGe

[1677] Martin Gardner. Mathematical games. Scientific American, 238(5):24– 30, May 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n5/pdf/scientificamerican0578-24.pdf.

Morrison:1978:Be

[1678] Philip Morrison. Books. Scientific American, 238(5):33–43, May 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n5/pdf/scientificamerican0578-33.pdf.

Kaplan:1978:ERW

[1679] Fred M. Kaplan. Enhanced-radiation weapons. Scientific American, 238(5):44–51, May 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n5/pdf/scientificamerican0578-
44.pdf.

Hsu:1978:WBS

[1680] Kenneth J. Hsü. When the Black Sea was drained. Scientific American, 238(5):52–63, May 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n5/pdf/scientificamerican0578-
52.pdf.

Muller:1978:CBR

[1681] Richard A. Muller. The cosmic background radiation and the new aether drift. Scientific American, 238(5):64–74, May 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n5/pdf/scientificamerican0578-64.pdf.

Anonymous:1978:SCe

[1682] Anonymous. Science and the citizen. Scientific American, 238(5):81– 96, May 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n5/pdf/scientificamerican0578-81.pdf.

Devey:1978:UMD

[1683] Gilbert B. Devey and Peter N. T. Wells. Ultrasound in medical diagnosis. Scientific American, 238(5):98–104, 106, 108, 110, 112, May 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n5/pdf/scientificamerican0578-98.pdf.

Wiley:1978:LMS

[1684] R. Haven Wiley, Jr. The Lek mating system of the sage grouse. Scientific American, 238(5):114–125, May 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n5/pdf/scientificamerican0578-
114.pdf.

Gogel:1978:APV

[1685] Walter C. Gogel. The adjacency principle in visual perception. Scientific American, 238(5):126–139, May 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n5/pdf/scientificamerican0578-
126.pdf.

Staehelin:1978:JBL

[1686] L. Andrew Staehelin and Barbara E. Hull. Junctions between living cells. Scientific American, 238(5):140–152, May 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n5/pdf/scientificamerican0578-140.pdf.

Smith:1978:RHT

[1687] Norman Smith. Roman hydraulic technology. Scientific American, 238(5):154–161, May 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n5/pdf/scientificamerican0578-
154.pdf.

Walker:1978:ASe

[1688] Jearl Walker. The amateur scientist. Scientific American, 238(5):162– 171, May 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n5/pdf/scientificamerican0578-162.pdf.

Anonymous:1978:Be

[1689] Anonymous. Bibliography. Scientific American, 238(5):172, May 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n5/pdf/scientificamerican0578-172.pdf.

Anonymous:1978:Lf

[1690] Anonymous. Letters. Scientific American, 238(6):8–11, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n6/pdf/scientificamerican0678-8.pdf.

Anonymous:1978:YAf

[1691] Anonymous. 50 and 100 years ago. Scientific American, 238(6):14, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n6/pdf/scientificamerican0678-14.pdf.

Anonymous:1978:Af

[1692] Anonymous. The authors. Scientific American, 238(6):17, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n6/pdf/scientificamerican0678-17.pdf.

Gardner:1978:MGM

[1693] M. Gardner. Mathematical games: a mathematical zoo of astounding critters, imaginary and otherwise. Scientific American, 238(6):18–28, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). Organisms with mathematically interesting structure.

Gardner:1978:MGf

[1694] Martin Gardner. Mathematical games. Scientific American, 238(6):18– 28, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n6/pdf/scientificamerican0678-18.pdf.

Morrison:1978:Bf

[1695] Philip Morrison. Books. Scientific American, 238(6):30–41, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n6/pdf/scientificamerican0678-30.pdf.

Taylor:1978:ATR

[1696] D. Garth Taylor, Paul B. Sheatsley, and Andrew M. Greeley. Attitudes toward racial integration. Scientific American, 238(6):42–49, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n6/pdf/scientificamerican0678-42.pdf.

Schmandt-Besserat:1978:EPW

[1697] Denise Schmandt-Besserat. The earliest precursor of writing. Scientific American, 238(6):50–59, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n6/pdf/scientificamerican0678-
50.pdf.

Cerny:1978:ELN

[1698] Joseph Cerny and Arthur M. Poskanzer. Exotic light nuclei. Scientific American, 238(6):60–72, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n6/pdf/scientificamerican0678-
60.pdf.

March:1978:SC

[1699] Long March. Science and the citizen. Scientific American, 238(6):74–86, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n6/pdf/scientificamerican0678-74.pdf.

aPippenger:1978:CT

[1700] N. Pippenger. Complexity theory. Scientific American, 238(6), June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Dickinson:1978:CM

[1701] Dale F. Dickinson. Cosmic masers. Scientific American, 238(6):90–105, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n6/pdf/scientificamerican0678-90.pdf.

Gordon:1978:STE

[1702] Richard Gordon and Antone G. Jacobson. The shaping of tissues in embryos. Scientific American, 238(6):106–113, June 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n6/pdf/scientificamerican0678-106.pdf.

Pippenger:1978:CT

[1703] Nicholas Pippenger. Complexity theory. Scientific American, 238(6):114– 124, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n6/pdf/scientificamerican0678-114.pdf.

Gauri:1978:PS

[1704] K. Lal Gauri. The preservation of Stone. Scientific American, 238(6):126–136, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n6/pdf/scientificamerican0678-
126.pdf.

Jones:1978:FBM

[1705] Jack Colvard Jones. The feeding behavior of mosquitoes. Scientific American, 238(6):138–148, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v238/n6/pdf/scientificamerican0678-
138.pdf.

Walker:1978:ASf

[1706] Jearl Walker. The amateur scientist. Scientific American, 238(6):151– 158, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v238/n6/pdf/scientificamerican0678-151.pdf.

Anonymous:1978:Bf

[1707] Anonymous. Bibliography. Scientific American, 238(6):160, June 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v238/
n6/pdf/scientificamerican0678-160.pdf.

Anonymous:1978:Lg

[1708] Anonymous. Letters. Scientific American, 239(1):8–10, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n1/pdf/scientificamerican0778-8.pdf.

Anonymous:1978:YAg

[1709] Anonymous. 50 and 100 years ago. Scientific American, 239(1):15, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n1/pdf/scientificamerican0778-15.pdf.

Anonymous:1978:Ag

[1710] Anonymous. The authors. Scientific American, 239(1):16–17, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n1/pdf/scientificamerican0778-16.pdf.

Gardner:1978:MGg

[1711] Martin Gardner. Mathematical games. Scientific American, 239(1):18– 26, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n1/pdf/scientificamerican0778-18.pdf.

Morrison:1978:Bg

[1712] Philip Morrison. Books. Scientific American, 239(1):28–36, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n1/pdf/scientificamerican0778-28.pdf.

Wilson:1978:AAE

[1713] David Gordon Wilson. Alternative automobile engines. Scientific American, 239(1):39–49, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n1/pdf/scientificamerican0778-
39.pdf.

Linsley:1978:HCR

[1714] J. Linsley. The highest-energy cosmic rays. Scientific American, 239(1): 48–58, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Patterson:1978:CDN

[1715] Paul H. Patterson, David D. Potter, and Edwin J. Furshpan. The chemical differentiation of nerve cells. Scientific American, 239(1):50–59, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n1/pdf/scientificamerican0778-50.pdf.

Linsley:1978:HEC

[1716] John Linsley. The highest-energy cosmic rays. Scientific American, 239(1):60–70, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n1/pdf/scientificamerican0778-
60.pdf.

Anonymous:1978:SCf

[1717] Anonymous. Science and the citizen. Scientific American, 239(1):72– 85, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n1/pdf/scientificamerican0778-72.pdf.

Croft:1978:NIE

[1718] Thomas A. Croft. Nighttime images of the Earth from Space. Scientific American, 239(1):86–98, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n1/pdf/scientificamerican0778-
86.pdf.

Menzel:1978:LMB

[1719] Randolf Menzel and Jochen Erber. Learning and memory in bees. Scientific American, 239(1):102–111, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v239/n1/pdf/scientificamerican0778-
102.pdf.

Gregor:1978:SMT

[1720] Harry P. Gregor and Charles D. Gregor. Synthetic-membrane technology. Scientific American, 239(1):112–128, July 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n1/pdf/scientificamerican0778-112.pdf.

Calder:1978:K

[1721] William A. Calder III. The kiwi. Scientific American, 239(1):132–142, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n1/pdf/scientificamerican0778-132.pdf.

Findler:1978:CP

[1722] Nicholas V. Findler. Computer poker. Scientific American, 239(1):144– 151, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n1/pdf/scientificamerican0778-144.pdf.

Walker:1978:ASg

[1723] Jearl Walker. The amateur scientist. Scientific American, 239(1):152– 160, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n1/pdf/scientificamerican0778-152.pdf.

Anonymous:1978:Bg

[1724] Anonymous. Bibliography. Scientific American, 239(1):162, July 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n1/pdf/scientificamerican0778-162.pdf.

Anonymous:1978:Lh

[1725] Anonymous. Letters. Scientific American, 239(2):8–10, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n2/pdf/scientificamerican0878-8.pdf.

Anonymous:1978:YAh

[1726] Anonymous. 50 and 100 years ago. Scientific American, 239(2):15, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n2/pdf/scientificamerican0878-15.pdf.

Anonymous:1978:Ah

[1727] Anonymous. The authors. Scientific American, 239(2):16–17, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n2/pdf/scientificamerican0878-16.pdf.

Gardner:1978:MGh

[1728] Martin Gardner. Mathematical games. Scientific American, 239(2):18– 26, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n2/pdf/scientificamerican0878-18.pdf.

Morrison:1978:Bh

[1729] Philip Morrison. Books. Scientific American, 239(2):27–33, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n2/pdf/scientificamerican0878-27.pdf.

Olander:1978:GC

[1730] Donald R. Olander. The gas centrifuge. Scientific American, 239(2):37– 43, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n2/pdf/scientificamerican0878-37.pdf.

Weymann:1978:SW

[1731] Ray J. Weymann. Stellar winds. Scientific American, 239(2):44–53, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n2/pdf/scientificamerican0878-44.pdf.

Walker:1978:HET

[1732] Alan Walker and Richard E. F. Leakey. The hominids of East Turkana. Scientific American, 239(2):54–67, August 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n2/pdf/scientificamerican0878-54.pdf.

Anonymous:1978:SCg

[1733] Anonymous. Science and the citizen. Scientific American, 239(2):68–72, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n2/pdf/scientificamerican0878-68.pdf.

Rick:1978:T

[1734] Charles M. Rick. The tomato. Scientific American, 239(2):76–87, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n2/pdf/scientificamerican0878-76.pdf.

Proctor:1978:NAT

[1735] Warren G. Proctor. Negative absolute temperatures. Scientific American, 239(2):90–99, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n2/pdf/scientificamerican0878-
90.pdf.

Heller:1978:TVA

[1736] H. Craig Heller, Larry I. Crawshaw, and Harold T. Hammel. The thermostat of vertebrate animals. Scientific American, 239(2):102–113, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n2/pdf/scientificamerican0878-102.pdf.

OKeefe:1978:TP

[1737] John A. O’Keefe. The tektite problem. Scientific American, 239(2):116–125, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n2/pdf/scientificamerican0878-
116.pdf.

Vogel:1978:OCC

[1738] Steven Vogel. Organisms that capture currents. Scientific American, 239(2):128–139, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n2/pdf/scientificamerican0878-
128.pdf.

Walker:1978:ASh

[1739] Jearl Walker. The amateur scientist. Scientific American, 239 (2):140–146, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n2/pdf/scientificamerican0878-
140.pdf.

Anonymous:1978:Bh

[1740] Anonymous. Bibliography. Scientific American, 239(2):148, August 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n2/pdf/scientificamerican0878-148.pdf.

Anonymous:1978:Li

[1741] Anonymous. Letters. Scientific American, 239(3):8–13, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n3/pdf/scientificamerican0978-8.pdf.

Anonymous:1978:YAi

[1742] Anonymous. 50 and 100 years ago. Scientific American, 239(3):14, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n3/pdf/scientificamerican0978-14.pdf.

Anonymous:1978:Ai

[1743] Anonymous. The authors. Scientific American, 239(3):16–18, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n3/pdf/scientificamerican0978-16.pdf.

Gardner:1978:MGi

[1744] Martin Gardner. Mathematical games. Scientific American, 239 (3):20–30, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n3/pdf/scientificamerican0978-
20.pdf.

Morrison:1978:Bi

[1745] Philip Morrison. Books. Scientific American, 239(3):32–44, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n3/pdf/scientificamerican0978-32.pdf.

Mayr:1978:E

[1746] Ernst Mayr. Evolution. Scientific American, 239(3):46–55, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n3/pdf/scientificamerican0978-46.pdf.

Ayala:1978:ME

[1747] Francisco J. Ayala. The mechanisms of evolution. Scientific American, 239(3):56–69, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n3/pdf/scientificamerican0978-
56.pdf.

Dickerson:1978:CEO

[1748] Richard E. Dickerson. Chemical evolution and the origin of life. Scientific American, 239(3):70–87, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v239/n3/pdf/scientificamerican0978-
70.pdf.

Anonymous:1978:SCh

[1749] Anonymous. Science and the citizen. Scientific American, 239(3):88–108, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n3/pdf/scientificamerican0978-88.pdf.

Schopf:1978:EEC

[1750] J. William Schopf. The evolution of the earliest cells. Scientific American, 239(3):110–138, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n3/pdf/scientificamerican0978-
110.pdf.

Valentine:1978:EMP

[1751] James W. Valentine. The evolution of multicellular plants and animals. Scientific American, 239(3):140–158, September 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n3/pdf/scientificamerican0978-140.pdf.

May:1978:EES

[1752] Robert M. May. The evolution of ecological systems. Scientific American, 239(3):160–175, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n3/pdf/scientificamerican0978-
160.pdf.

Smith:1978:EB

[1753] John Maynard Smith. The evolution of behavior. Scientific American, 239(3):176–192, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n3/pdf/scientificamerican0978-
176.pdf.

Washburn:1978:EM

[1754] Sherwood L. Washburn. The evolution of Man. Scientific American, 239(3):194–206, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n3/pdf/scientificamerican0978-
194.pdf.

Lewontin:1978:A

[1755] Richard C. Lewontin. Adaptation. Scientific American, 239(3):212–230, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n3/pdf/scientificamerican0978-212.pdf.

Walker:1978:ASi

[1756] Jearl Walker. The amateur scientist. Scientific American, 239 (3):232–236, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n3/pdf/scientificamerican0978-
232.pdf.

Anonymous:1978:Bi

[1757] Anonymous. Bibliography. Scientific American, 239(3):242, September 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n3/pdf/scientificamerican0978-242.pdf.

Anonymous:1978:Lj

[1758] Anonymous. Letters. Scientific American, 239(4):10–13, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n4/pdf/scientificamerican1078-10.pdf.

Anonymous:1978:YAj

[1759] Anonymous. 50 and 100 years ago. Scientific American, 239(4):16, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n4/pdf/scientificamerican1078-16.pdf.

Anonymous:1978:Aj

[1760] Anonymous. The authors. Scientific American, 239(4):20, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n4/pdf/scientificamerican1078-20.pdf.

Gardner:1978:MGj

[1761] Martin Gardner. Mathematical games. Scientific American, 239(4):23– 30, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n4/pdf/scientificamerican1078-23.pdf.

Morrison:1978:Bj

[1762] Philip Morrison. Books. Scientific American, 239(4):32–47, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n4/pdf/scientificamerican1078-32.pdf.

Morrison:1978:NSM

[1763] Philip Morrison and Paul F. Walker. A new strategy for military spending. Scientific American, 239(4):48–61, October 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n4/pdf/scientificamerican1078-48.pdf.

Lassen:1978:BFB

[1764] Niels A. Lassen, David H. Ingvar, and Erik Skinhøj. Brain function and blood flow. Scientific American, 239(4):62–71, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n4/pdf/scientificamerican1078-62.pdf.

Lederman:1978:UP

[1765] Leon M. Lederman. The upsilon particle. Scientific American, 239(4):72– 81, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n4/pdf/scientificamerican1078-72.pdf.

Anonymous:1978:SCi

[1766] Anonymous. Science and the citizen. Scientific American, 239(4):85–102, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n4/pdf/scientificamerican1078-85.pdf.

Edwards:1978:FLT

[1767] Harold M. Edwards. Fermat’s last theorem. Scientific American, 239(4):104–122, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n4/pdf/scientificamerican1078-
104.pdf.

Schramm:1978:DST

[1768] David N. Schramm and Robert N. Clayton. Did a supernova trigger the formation of the solar system? Scientific American, 239(4):124–139, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n4/pdf/scientificamerican1078-124.pdf.

Isbell:1978:PGD

[1769] William H. Isbell. The prehistoric ground drawings of Peru. Scientific American, 239(4):140–153, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v239/n4/pdf/scientificamerican1078-
140.pdf.

Wolpert:1978:PFB

[1770] Lewis Wolpert. Pattern formation in biological development. Scientific American, 239(4):154–164, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v239/n4/pdf/scientificamerican1078-

154.pdf. This and the preceding paper discuss of the manner in which the context in which cells find themselves affects their development.

Williams:1978:OMM

[1771] Timothy C. Williams and Janet M. Williams. An oceanic mass migration of land birds. Scientific American, 239(4):166–176, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n4/pdf/scientificamerican1078-166.pdf.

Walker:1978:ASj

[1772] Jearl Walker. The amateur scientist. Scientific American, 239 (4):179–187, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n4/pdf/scientificamerican1078-
179.pdf.

Anonymous:1978:Bj

[1773] Anonymous. Bibliography. Scientific American, 239(4):188, October 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n4/pdf/scientificamerican1078-188.pdf.

Anonymous:1978:Lk

[1774] Anonymous. Letters. Scientific American, 239(5):10–15, November 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n5/pdf/scientificamerican1178-10.pdf.

Anonymous:1978:YAk

[1775] Anonymous. 50 and 100 years ago. Scientific American, 239(5):16, November 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n5/pdf/scientificamerican1178-16.pdf.

Anonymous:1978:Ak

[1776] Anonymous. The authors. Scientific American, 239(5):18–20, November 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n5/pdf/scientificamerican1178-18.pdf.

Gardner:1978:MGk

[1777] Martin Gardner. Mathematical games. Scientific American, 239 (5):22–30, November 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n5/pdf/scientificamerican1178-
22.pdf.

Morrison:1978:Bk

[1778] Philip Morrison. Books. Scientific American, 239(5):35–47, November 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n5/pdf/scientificamerican1178-35.pdf.

Yonas:1978:FPP

[1779] Gerold Yonas. Fusion power with particle beams. Scientific American, 239(5):50–61, November 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n5/pdf/scientificamerican1178-
50.pdf.

Jonathan:1978:AV

[1780] P. Jonathan, G. Butler, and Aaron Klug. The assembly of a virus. Scientific American, 239(5):62–69, November 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n5/pdf/scientificamerican1178-62.pdf.

Spiller:1978:OLW

[1781] Eberhard Spiller and Ralph Feder. The optics of long-wavelength X rays. Scientific American, 239(5):70–78, November 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n5/pdf/scientificamerican1178-70.pdf.

Anonymous:1978:SCj

[1782] Anonymous. Science and the citizen. Scientific American, 239(5):80–91, November 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n5/pdf/scientificamerican1178-80.pdf.

Moskowitz:1978:AL

[1783] Breyne Arlene Moskowitz. The acquisition of language. Scientific American, 239(5):92–108, November 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n5/pdf/scientificamerican1178-
92.pdf.

Gorenstein:1978:RCG

[1784] Paul Gorenstein and Wallace Tucker. Rich clusters of galaxies. Scientific American, 239(5):110–128, November 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n5/pdf/scientificamerican1178-110.pdf.

Samuels:1978:MAM

[1785] Leonard E. Samuels. The mechanisms of abrasive machining. Scientific American, 239(5):132–152, November 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v239/n5/pdf/scientificamerican1178-
132.pdf.

Routtenberg:1978:RSB

[1786] Aryeh Routtenberg. The reward system of the brain. Scientific American, 239(5):154–164, November 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n5/pdf/scientificamerican1178-
154.pdf.

Hooven:1978:WBF

[1787] Frederick J. Hooven. The Wright brothers’ flight-control system. Scientific American, 239(5):166–184, November 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n5/pdf/scientificamerican1178-166.pdf.

Walker:1978:ASk

[1788] Jearl Walker. The amateur scientist. Scientific American, 239 (5):186–196, November 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n5/pdf/scientificamerican1178-
186.pdf.

Anonymous:1978:Bk

[1789] Anonymous. Bibliography. Scientific American, 239(5):198, November 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n5/pdf/scientificamerican1178-198.pdf.

Anonymous:1978:Ll

[1790] Anonymous. Letters. Scientific American, 239(6):8–10, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n6/pdf/scientificamerican1278-8.pdf.

Anonymous:1978:YAl

[1791] Anonymous. 50 and 100 years ago. Scientific American, 239(6):12, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n6/pdf/scientificamerican1278-12.pdf.

Anonymous:1978:Al

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n6/pdf/scientificamerican1278-15.pdf.

Gardner:1978:MGl

[1793] Martin Gardner. Mathematical games. Scientific American, 239 (6):18–26, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n6/pdf/scientificamerican1278-
18.pdf.

Philip:1978:B

[1794] Philip and Phylis Morrison. Books. Scientific American, 239(6):28–48, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n6/pdf/scientificamerican1278-28.pdf.

Westoff:1978:MFD

[1795] Charles F. Westoff. Marriage and fertility in the developed countries. Scientific American, 239(6):51–57, December 1978. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v239/
n6/pdf/scientificamerican1278-51.pdf.

Fickett:1978:FPP

[1796] A. P. Fickett. Fuel-cell power plants. Scientific American, 239(6):54–60, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bromley:1978:NM

[1797] D. Allan Bromley. Nuclear molecules. Scientific American, 239(6):58–68, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n6/pdf/scientificamerican1278-58.pdf.

Fickett:1978:FCP

[1798] Arnold P. Fickett. Fuel-cell power plants. Scientific American, 239(6):70–76, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n6/pdf/scientificamerican1278-
70.pdf.

Anonymous:1978:SCk

[1799] Anonymous. Science and the citizen. Scientific American, 239(6):78–84, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n6/pdf/scientificamerican1278-78.pdf.

Perutz:1978:HSR

[1800] M. F. Perutz. Hemoglobin structure and respiratory transport. Scientific American, 239(6):92–125, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v239/n6/pdf/scientificamerican1278-
92.pdf.

Land:1978:AEM

[1801] Michael F. Land. Animal eyes with mirror optics. Scientific American, 239(6):126–134, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n6/pdf/scientificamerican1278-
126.pdf.

Redford:1978:RTA

[1802] Donald B. Redford. The razed Temple of Akhenaten. Scientific American, 239(6):136–147, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n6/pdf/scientificamerican1278-
136.pdf.

McMahon:1978:FRT

[1803] Thomas A. McMahon and Peter R. Greene. Fast running tracks. Scientific American, 239(6):148–163, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v239/n6/pdf/scientificamerican1278-
148.pdf.

Chaisson:1978:GN

[1804] Eric J. Chaisson. Gaseous nebulas. Scientific American, 239(6):164–180, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n6/pdf/scientificamerican1278-164.pdf.

Walker:1978:ASl

[1805] Jearl Walker. The amateur scientist. Scientific American, 239 (6):182–190, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v239/n6/pdf/scientificamerican1278-
182.pdf.

Anonymous:1978:AIa

[1806] Anonymous. Annual index. Scientific American, 239(6):192–197, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n6/pdf/scientificamerican1278-192.pdf.

Anonymous:1978:Bl

[1807] Anonymous. Bibliography. Scientific American, 239(6):198, December 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v239/n6/pdf/scientificamerican1278-198.pdf.

Anonymous:1979:La

[1808] Anonymous. Letters. Scientific American, 240(1):8–10, January 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n1/pdf/scientificamerican0179-8.pdf.

Anonymous:1979:YAa

[1809] Anonymous. 50 and 100 years ago. Scientific American, 240(1):15, January 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n1/pdf/scientificamerican0179-15.pdf.

Anonymous:1979:Aa

[1810] Anonymous. The authors. Scientific American, 240(1):16, January 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n1/pdf/scientificamerican0179-16.pdf.

Gardner:1979:MGa

[1811] Martin Gardner. Mathematical games. Scientific American, 240(1):18– 24, January 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n1/pdf/scientificamerican0179-18.pdf.

Morrison:1979:Ba

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n1/pdf/scientificamerican0179-33.pdf.

Griffith:1979:WCP

[1813] Edward D. Griffith and Alan W. Clarke. World coal production. Scientific American, 240(1):38–47, January 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n1/pdf/scientificamerican0179-38.pdf.

Lodish:1979:ACM

[1814] Harvey F. Lodish and James E. Rothman. The assembly of cell membranes. Scientific American, 240(1):48–63, January 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n1/pdf/scientificamerican0179-48.pdf. Details of low-level self-

assembly of cell membranes.

Yariv:1979:GO

[1815] A. Yariv. Guided-wave optics. Scientific American, 240(1):54–62, January 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Yariv:1979:GWO

[1816] Amnon Yariv. Guided-wave optics. Scientific American, 240(1):64–73, January 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n1/pdf/scientificamerican0179-64.pdf.

Anonymous:1979:SCa

[1817] Anonymous. Science and the citizen. Scientific American, 240(1):74–87, January 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n1/pdf/scientificamerican0179-74.pdf.

Jordan:1979:DSC

[1818] Thomas H. Jordan. The deep structure of the continents. Scientific American, 240(1):92–107, January 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n1/pdf/scientificamerican0179-
92.pdf.

Bahill:1979:TSE

[1819] A. Terry Bahill and Lawrence Stark. The trajectories of saccadic eye movements. Scientific American, 240(1):108–117, January 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n1/pdf/scientificamerican0179-108.pdf.

Weinreich:1979:CMP

[1820] Gabriel Weinreich. The coupled motions of piano strings. Scientific American, 240(1):118–127, January 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v240/n1/pdf/scientificamerican0179-
118.pdf.

Clarke:1979:HSW

[1821] Malcolm R. Clarke. The head of the sperm whale. Scientific American, 240(1):128–141, January 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n1/pdf/scientificamerican0179-
128.pdf.

Kevles:1979:RM

[1822] Daniel J. Kevles. Robert A. Millikan. Scientific American, 240 (1):142–151, January 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n1/pdf/scientificamerican0179-
142.pdf.

Walker:1979:ASa

[1823] Jearl Walker. The amateur scientist. Scientific American, 240 (1):152–163, January 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n1/pdf/scientificamerican0179-
152.pdf.

Anonymous:1979:Ba

[1824] Anonymous. Bibliography. Scientific American, 240(1):164, January 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n1/pdf/scientificamerican0179-164.pdf.

Anonymous:1979:Lb

[1825] Anonymous. Letters. Scientific American, 240(2):11, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n2/pdf/scientificamerican0279-11.pdf.

Anonymous:1979:YAb

[1826] Anonymous. 50 and 100 years ago. Scientific American, 240(2):12–13, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n2/pdf/scientificamerican0279-12.pdf.

Anonymous:1979:Ab

[1827] Anonymous. The authors. Scientific American, 240(2):14–15, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n2/pdf/scientificamerican0279-14.pdf.

Gardner:1979:MGb

[1828] Martin Gardner. Mathematical games. Scientific American, 240 (2):16–27, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n2/pdf/scientificamerican0279-
16.pdf.

Morrison:1979:Bb

[1829] Philip Morrison. Books. Scientific American, 240(2):28–34, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n2/pdf/scientificamerican0279-28.pdf.

Aspin:1979:VSI

[1830] Les Aspin. The verification of the SALT II Agreement. Scientific American, 240(2):38–45, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n2/pdf/scientificamerican0279-
38.pdf.

Rahn:1979:HBE

[1831] Hermann Rahn, Amos Ar, and Charles V. Paganelli. How bird eggs breathe. Scientific American, 240(2):46–55, February 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n2/pdf/scientificamerican0279-46.pdf.

Bocquet:1979:LBA

[1832] Aimé Bocquet. Lake-bottom archaeology. Scientific American, 240(2):56–64, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n2/pdf/scientificamerican0279-
56.pdf.

Anonymous:1979:SCb

[1833] Anonymous. Science and the citizen. Scientific American, 240(2):67–75, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n2/pdf/scientificamerican0279-67.pdf.

Spurr:1979:S

[1834] Stephen H. Spurr. Silviculture. Scientific American, 240(2):76–91, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n2/pdf/scientificamerican0279-76.pdf.

Rebbi:1979:S

[1835] Claudio Rebbi. Solitons. Scientific American, 240(2):92–117, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n2/pdf/scientificamerican0279-92.pdf.

Carrigan:1979:SEM

[1836] Charles R. Carrigan and David Gubbins. The source of the Earth’s magnetic field. Scientific American, 240(2):118–130, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n2/pdf/scientificamerican0279-118.pdf.

Crow:1979:GVM

[1837] James F. Crow. Genes that violate Mendel’s rules. Scientific American, 240(2):134–147, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n2/pdf/scientificamerican0279-
134.pdf.

Reiser:1979:MIS

[1838] Stanley Joel Reiser. The medical influence of the stethoscope. Scientific American, 240(2):148–157, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v240/n2/pdf/scientificamerican0279-
148.pdf.

Walker:1979:ASb

[1839] Jearl Walker. The amateur scientist. Scientific American, 240 (2):158–166, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n2/pdf/scientificamerican0279-
158.pdf.

Anonymous:1979:Bb

[1840] Anonymous. Bibliography. Scientific American, 240(2):168, February 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n2/pdf/scientificamerican0279-168.pdf.

Anonymous:1979:Lc

[1841] Anonymous. Letters. Scientific American, 240(3):10–15, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n3/pdf/scientificamerican0379-10.pdf.

Anonymous:1979:YAc

[1842] Anonymous. 50 and 100 years ago. Scientific American, 240(3):16–17, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n3/pdf/scientificamerican0379-16.pdf.

Anonymous:1979:Ac

[1843] Anonymous. The authors. Scientific American, 240(3):18, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n3/pdf/scientificamerican0379-18.pdf.

Gardner:1979:MGc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n3/pdf/scientificamerican0379-21.pdf.

Morrison:1979:Bc

[1845] Philip Morrison. Books. Scientific American, 240(3):32–47, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n3/pdf/scientificamerican0379-32.pdf.

Ginzberg:1979:PUL

[1846] Eli Ginzberg. The professionalization of the U.S. labor force. Scientific American, 240(3):48–53, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n3/pdf/scientificamerican0379-
48.pdf.

Wetherill:1979:AO

[1847] George W. Wetherill. Apollo objects. Scientific American, 240(3):54–65, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n3/pdf/scientificamerican0379-54.pdf.

Nicolson:1979:CM

[1848] Garth L. Nicolson. Cancer metastasis. Scientific American, 240(3):66– 77, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n3/pdf/scientificamerican0379-66.pdf.

Anonymous:1979:SCc

[1849] Anonymous. Science and the citizen. Scientific American, 240(3):78–93, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n3/pdf/scientificamerican0379-78.pdf.

Hansch:1979:SAH

[1850] Theodor W. Hänsch, Arthur L. Schawlow, and George W. Series. The spectrum of atomic hydrogen. Scientific American, 240(3):94–111, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n3/pdf/scientificamerican0379-94.pdf.

Gilchrist:1979:PSB

[1851] Alan L. Gilchrist. The perception of surface blacks and whites. Scientific American, 240(3):112–125, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n3/pdf/scientificamerican0379-
112.pdf.

Keynes:1979:ICN

[1852] Richard D. Keynes. Ion channels in the nerve-cell membrane. Scientific American, 240(3):126–135, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n3/pdf/scientificamerican0379-
126.pdf.

Wursig:1979:D

[1853] Bernd Würsig. Dolphins. Scientific American, 240(3):136–149, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n3/pdf/scientificamerican0379-136.pdf.

Soedel:1979:AC

[1854] Werner Soedel and Vernard Foley. Ancient catapults. Scientific American, 240(3):150–161, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n3/pdf/scientificamerican0379-
150.pdf.

Walker:1979:ASc

[1855] Jearl Walker. The amateur scientist. Scientific American, 240(3):162– 173, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n3/pdf/scientificamerican0379-162.pdf.

Anonymous:1979:Bc

[1856] Anonymous. Bibliography. Scientific American, 240(3):174, March 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n3/pdf/scientificamerican0379-174.pdf.

Anonymous:1979:Ld

[1857] Anonymous. Letters. Scientific American, 240(4):10, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n4/pdf/scientificamerican0479-10.pdf.

Anonymous:1979:YAd

[1858] Anonymous. 50 and 100 years ago. Scientific American, 240(4):12–15, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n4/pdf/scientificamerican0479-12.pdf.

Anonymous:1979:Ad

[1859] Anonymous. The authors. Scientific American, 240(4):16–17, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n4/pdf/scientificamerican0479-16.pdf.

Gardner:1979:MGd

[1860] Martin Gardner. Mathematical games. Scientific American, 240(4):18– 28, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n4/pdf/scientificamerican0479-18.pdf.

Morrison:1979:Bd

[1861] Philip Morrison. Books. Scientific American, 240(4):37–53, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n4/pdf/scientificamerican0479-37.pdf.

Parmentola:1979:PBW

[1862] John Parmentola and Kosta Tsipis. Particle-beam weapons. Scientific American, 240(4):54–65, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n4/pdf/scientificamerican0479-
54.pdf.

Binkley:1979:TEP

[1863] Sue Binkley. A timekeeping enzyme in the pineal gland. Scientific American, 240(4):66–71, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n4/pdf/scientificamerican0479-
66.pdf.

Strom:1979:EDG

[1864] Stephen E. Strom and Karen M. Strom. The evolution of disk galaxies. Scientific American, 240(4):72–83, April 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n4/pdf/scientificamerican0479-72.pdf.

Anonymous:1979:SCd

[1865] Anonymous. Science and the citizen. Scientific American, 240(4):84– 105, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n4/pdf/scientificamerican0479-84.pdf.

Millot:1979:C

[1866] Georges Millot. Clay. Scientific American, 240(4):108–118, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n4/pdf/scientificamerican0479-108.pdf.

Kuhn:1979:IP

[1867] L. Kuhn and R. A. Myers. Ink-jet printing. Scientific American, 240 (4):120–128, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Illmensee:1979:TC

[1868] Karl Illmensee and Leroy C. Stevens. Teratomas and chimeras. Scientific American, 240(4):120–132, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n4/pdf/scientificamerican0479-
120.pdf.

Chou:1979:COB

[1869] Hung-Hsiang Chou. Chinese oracle bones. Scientific American, 240(4):134–148, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n4/pdf/scientificamerican0479-
134.pdf.

Feld:1979:PK

[1870] Michael S. Feld, Ronald E. McNair, and Stephen R. Wilk. The physics of karate. Scientific American, 240(4):150–158, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n4/pdf/scientificamerican0479-150.pdf.

Kuhn:1979:IJP

[1871] Larry Kuhn and Robert A. Myers. Ink-jet printing. Scientific American, 240(4):162–178, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n4/pdf/scientificamerican0479-
162.pdf.

Walker:1979:ASd

[1872] Jearl Walker. The amateur scientist. Scientific American, 240(4):180– 189, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n4/pdf/scientificamerican0479-180.pdf.

Anonymous:1979:Bd

[1873] Anonymous. Bibliography. Scientific American, 240(4):190, April 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n4/pdf/scientificamerican0479-190.pdf.

Anonymous:1979:Le

[1874] Anonymous. Letters. Scientific American, 240(5):11, May 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n5/pdf/scientificamerican0579-11.pdf.

Anonymous:1979:YAe

[1875] Anonymous. 50 and 100 years ago. Scientific American, 240(5):13, May 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n5/pdf/scientificamerican0579-13.pdf.

Anonymous:1979:Ae

[1876] Anonymous. The authors. Scientific American, 240(5):14–17, May 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n5/pdf/scientificamerican0579-14.pdf.

Gardner:1979:MGe

[1877] Martin Gardner. Mathematical games. Scientific American, 240(5):18– 25, May 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n5/pdf/scientificamerican0579-18.pdf.

Morrison:1979:Be

[1878] Philip Morrison. Books. Scientific American, 240(5):26–45, May 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n5/pdf/scientificamerican0579-26.pdf.

Bartky:1979:SDS

[1879] Ian R. Bartky and Elizabeth Harrison. Standard and daylightsaving time. Scientific American, 240(5):46–53, May 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n5/pdf/scientificamerican0579-46.pdf.

Gordon:1979:CMG

[1880] M. A. Gordon and W. B. Burton. Carbon monoxide in the Galaxy. Scientific American, 240(5):54–67, May 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n5/pdf/scientificamerican0579-54.pdf.

Krisch:1979:SP

[1881] Alan D. Krisch. The spin of the proton. Scientific American, 240(5):68– 81, May 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n5/pdf/scientificamerican0579-68.pdf.

Anonymous:1979:SCe

[1882] Anonymous. Science and the citizen. Scientific American, 240(5):82– 98, May 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n5/pdf/scientificamerican0579-82.pdf.

Lazarides:1979:MBC

[1883] Elias Lazarides and Jean Paul Revel. The molecular basis of cell movement. Scientific American, 240(5):100–113, May 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n5/pdf/scientificamerican0579-100.pdf.

Baker:1979:BSM

[1884] M. A. Baker. A brain-cooling system in mammals. Scientific American, 240(5):114–122, May 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ronn:1979:LC

[1885] Avigdor M. Ronn. Laser chemistry. Scientific American, 240(5):114– 129, May 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n5/pdf/scientificamerican0579-114.pdf.

Baker:1979:BCS

[1886] Mary Ann Baker. A brain-cooling system in mammals. Scientific American, 240(5):130–139, May 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n5/pdf/scientificamerican0579-
130.pdf.

Stockmeyer:1979:IDP

[1887] Larry J. Stockmeyer and Ashok K. Chandra. Intrinsically difficult problems. Scientific American, 240(5):140–159, May 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n5/pdf/scientificamerican0579-140.pdf.

Olson:1979:GPH

[1888] Roberta J. M. Olson. Giotto’s portrait of Halley’s comet. Scientific American, 240(5):160–170, May 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n5/pdf/scientificamerican0579-
160.pdf.

Walker:1979:ASe

[1889] Jearl Walker. The amateur scientist. Scientific American, 240(5):172– 183, May 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n5/pdf/scientificamerican0579-172.pdf.

Anonymous:1979:Be

[1890] Anonymous. Bibliography. Scientific American, 240(5):184, May 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n5/pdf/scientificamerican0579-184.pdf.

Anonymous:1979:Lf

[1891] Anonymous. Letters. Scientific American, 240(6):10–13, June 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n6/pdf/scientificamerican0679-10.pdf.

Anonymous:1979:YAf

[1892] Anonymous. 50 and 100 years ago. Scientific American, 240(6):14–15, June 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n6/pdf/scientificamerican0679-14.pdf.

Anonymous:1979:Af

[1893] Anonymous. The authors. Scientific American, 240(6):16–18, June 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n6/pdf/scientificamerican0679-16.pdf.

Gardner:1979:MGf

[1894] Martin Gardner. Mathematical games. Scientific American, 240(6):20– 35, June 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n6/pdf/scientificamerican0679-20.pdf.

Morrison:1979:Bf

[1895] Philip Morrison. Books. Scientific American, 240(6):36–54, June 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n6/pdf/scientificamerican0679-36.pdf.

Grobstein:1979:EHF

[1896] Clifford Grobstein. External human fertilization. Scientific American, 240(6):57–67, June 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v240/n6/pdf/scientificamerican0679-
57.pdf.

Levi-Montalcini:1979:NGF

[1897] Rita Levi-Montalcini and Pietro Calissano. The nerve-growth factor. Scientific American, 240(6):68–77, June 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v240/n6/pdf/scientificamerican0679-
68.pdf.

Kahne:1979:ACD

[1898] Stephen Kahne, Irving Lefkowitz, and Charles Rose. Automatic control by distributed intelligence. Scientific American, 240(6):78–90, June 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n6/pdf/scientificamerican0679-78.pdf.

Anonymous:1979:SCf

[1899] Anonymous. Science and the citizen. Scientific American, 240(6):93– 109, June 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n6/pdf/scientificamerican0679-93.pdf.

Neuwirth:1979:TK

[1900] Lee Neuwirth. The theory of knots. Scientific American, 240(6):110– 124, June 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n6/pdf/scientificamerican0679-110.pdf.

Bosch:1979:NFM

[1901] Peter W. Bosch. A Neolithic flint mine. Scientific American, 240(6):126– 133, June 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n6/pdf/scientificamerican0679-126.pdf.

Golub:1979:UN

[1902] R. Golub, W. Mampe, J. M. Pendlebury, and P. Ageron. Ultracold neutrons. Scientific American, 240(6):134–154, June 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n6/pdf/scientificamerican0679-134.pdf.

Sclater:1979:HA

[1903] John G. Sclater and Christopher Tapscott. The history of the Atlantic. Scientific American, 240(6):156–175, June 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n6/pdf/scientificamerican0679-156.pdf.

Stommel:1979:YS

[1904] Henry Stommel and Elizabeth Stommel. The year without a Summer. Scientific American, 240(6):176–186, June 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n6/pdf/scientificamerican0679-176.pdf.

Walker:1979:ASf

[1905] Jearl Walker. The amateur scientist. Scientific American, 240(6):188– 200, June 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v240/n6/pdf/scientificamerican0679-188.pdf.

Anonymous:1979:Bf

[1906] Anonymous. Bibliography. Scientific American, 240(6):202, June 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v240/
n6/pdf/scientificamerican0679-202.pdf.

Anonymous:1979:Lg

[1907] Anonymous. Letters. Scientific American, 241(1):8–9, July 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n1/pdf/scientificamerican0779-8.pdf.

Anonymous:1979:YAg

[1908] Anonymous. 50 and 100 years ago. Scientific American, 241(1):13, July 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n1/pdf/scientificamerican0779-13.pdf.

Anonymous:1979:Ag

[1909] Anonymous. The authors. Scientific American, 241(1):14–15, July 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n1/pdf/scientificamerican0779-14.pdf.

Gardner:1979:MGg

[1910] Martin Gardner. Mathematical games. Scientific American, 241(1):16– 24, July 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n1/pdf/scientificamerican0779-16.pdf.

Morrison:1979:Bg

[1911] Philip Morrison. Books. Scientific American, 241(1):29–32, July 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n1/pdf/scientificamerican0779-29.pdf.

Lewis:1979:PDE

[1912] Kevin N. Lewis. The prompt and delayed effects of nuclear war. Scientific American, 241(1):35–47, July 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n1/pdf/scientificamerican0779-
35.pdf.

Henle:1979:EBV

[1913] Warner Henle, Gertrude Henle, and Evelyne T. Lennette. The Epstein–Barr virus. Scientific American, 241(1):48–59, July 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n1/pdf/scientificamerican0779-48.pdf.

Geballe:1979:CPG

[1914] Thomas R. Geballe. The central parsec of the Galaxy. Scientific American, 241(1):60–71, July 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n1/pdf/scientificamerican0779-
60.pdf.

Anonymous:1979:SCg

[1915] Anonymous. Science and the citizen. Scientific American, 241(1):72– 85B, July 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n1/pdf/scientificamerican0779-72.pdf.

Culick:1979:OFP

[1916] F. E. C. Culick. The origins of the first powered, man-carrying airplane. Scientific American, 241(1):86–101, July 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n1/pdf/scientificamerican0779-86.pdf.

Garcia-Bellido:1979:CAD

[1917] Antonio García-Bellido, Peter A. Lawrence, and Gines Morata. Compartments in animal development. Scientific American, 241(1):102–111, July 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n1/pdf/scientificamerican0779-102.pdf.

Johnson:1979:BMQ

[1918] Kenneth A. Johnson. The bag model of quark confinement. Scientific American, 241(1):112–121, July 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n1/pdf/scientificamerican0779-
112.pdf.

Morris:1979:ABS

[1919] Simon Conway Morris and H. B. Whittington. The animals of the Burgess Shale. Scientific American, 241(1):122–135, July 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n1/pdf/scientificamerican0779-122.pdf.

Regan:1979:VPM

[1920] David Regan, Kenneth Beverley, and Max Cynader. The visual perception of motion in depth. Scientific American, 241(1):136–151, July 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n1/pdf/scientificamerican0779-136.pdf.

Walker:1979:ASg

[1921] Jearl Walker. The amateur scientist. Scientific American, 241(1):152– 161, July 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n1/pdf/scientificamerican0779-152.pdf.

Anonymous:1979:Bg

[1922] Anonymous. Bibliography. Scientific American, 241(1):162, July 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n1/pdf/scientificamerican0779-162.pdf.

Anonymous:1979:Lh

[1923] Anonymous. Letters. Scientific American, 241(2):10, August 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n2/pdf/scientificamerican0879-10.pdf.

Anonymous:1979:YAh

[1924] Anonymous. 50 and 100 years ago. Scientific American, 241(2):15, August 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n2/pdf/scientificamerican0879-15.pdf.

Anonymous:1979:Ah

[1925] Anonymous. The authors. Scientific American, 241(2):16, August 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n2/pdf/scientificamerican0879-16.pdf.

Gardner:1979:MGh

[1926] Martin Gardner. Mathematical games. Scientific American, 241(2):18– 24, August 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n2/pdf/scientificamerican0879-18.pdf.

Morrison:1979:Bh

[1927] Philip Morrison. Books. Scientific American, 241(2):28–39, August 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n2/pdf/scientificamerican0879-28.pdf.

Devoret:1979:BTP

[1928] Raymond Devoret. Bacterial tests for potential carcinogens. Scientific American, 241(2):40–49, August 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n2/pdf/scientificamerican0879-
40.pdf.

Furth:1979:PTT

[1929] Harold P. Furth. Progress toward a Tokamak fusion reactor. Scientific American, 241(2):50–61, August 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n2/pdf/scientificamerican0879-
50.pdf.

Moore:1979:PNF

[1930] Andrew M. T. Moore. A pre-Neolithic farmers’ village on the Euphrates. Scientific American, 241(2):62–71, August 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n2/pdf/scientificamerican0879-62.pdf.

Jove:1979:SC

[1931] Jove. Science and the citizen. Scientific American, 241(2):72–80, August 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n2/pdf/scientificamerican0879-72.pdf.

Goreau:1979:CCR

[1932] Thomas F. Goreau, Nora I. Goreau, and Thomas J. Goreau. Corals and coral reefs. Scientific American, 241(2):124–136, August 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n2/pdf/scientificamerican0879-124.pdf.

Herbst:1979:SSF

[1933] William Herbst and George E. Assousa. Supernovas and star formation. Scientific American, 241(2):138–145, August 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n2/pdf/scientificamerican0879-138.pdf.

Hellman:1979:MPK

[1934] Martin E. Hellman. The mathematics of public-key cryptography. Scientific American, 241(2):146–157, August 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n2/pdf/scientificamerican0879-146.pdf.

Wilson:1979:PPM

[1935] Kenneth G. Wilson. Problems in physics with many scales of length. Scientific American, 241(2):158–179, August 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n2/pdf/scientificamerican0879-158.pdf.

Crews:1979:HCB

[1936] David Crews. The hormonal control of behavior in a lizard. Scientific American, 241(2):180–187, August 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n2/pdf/scientificamerican0879-
180.pdf.

Walker:1979:ASh

[1937] Jearl Walker. The amateur scientist. Scientific American, 241 (2):188–197, August 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n2/pdf/scientificamerican0879-
188.pdf.

Anonymous:1979:Bh

[1938] Anonymous. Bibliography. Scientific American, 241(2):198, August 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n2/pdf/scientificamerican0879-198.pdf.

Anonymous:1979:Li

[1939] Anonymous. Letters. Scientific American, 241(3):8, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n3/pdf/scientificamerican0979-8.pdf.

Anonymous:1979:YAi

[1940] Anonymous. 50 and 100 years ago. Scientific American, 241(3):10, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n3/pdf/scientificamerican0979-10.pdf.

Anonymous:1979:Ai

[1941] Anonymous. The authors. Scientific American, 241(3):14–20, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n3/pdf/scientificamerican0979-14.pdf.

Gardner:1979:MGi

[1942] Martin Gardner. Mathematical games. Scientific American, 241 (3):22–32, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n3/pdf/scientificamerican0979-
22.pdf.

Morrison:1979:Bi

[1943] Philip Morrison. Books. Scientific American, 241(3):34–42, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n3/pdf/scientificamerican0979-34.pdf.

Hubel:1979:B

[1944] David H. Hubel. The brain. Scientific American, 241(3):44–52, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n3/pdf/scientificamerican0979-44.pdf.

Stevens:1979:N

[1945] Charles F. Stevens. The neuron. Scientific American, 241(3):54–65, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n3/pdf/scientificamerican0979-54.pdf.

Kandel:1979:SSN

[1946] Eric R. Kandel. Small systems of neurons. Scientific American, 241(3):66–76, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n3/pdf/scientificamerican0979-
66.pdf.

Anonymous:1979:SCh

[1947] Anonymous. Science and the citizen. Scientific American, 241(3):78–84, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n3/pdf/scientificamerican0979-78.pdf.

Nauta:1979:OB

[1948] Walle J. H. Nauta and Michael Feirtag. The organization of the brain. Scientific American, 241(3):88–111, September 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n3/pdf/scientificamerican0979-88.pdf.

Cowan:1979:DB

[1949] W. Maxwell Cowan. The development of the brain. Scientific American, 241(3):112–133, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n3/pdf/scientificamerican0979-
112.pdf.

Iversen:1979:CB

[1950] Leslie L. Iversen. The chemistry of the brain. Scientific American, 241(3):134–149, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n3/pdf/scientificamerican0979-
134.pdf.

Hubel:1979:BMV

[1951] David H. Hubel and Torsten N. Wiesel. Brain mechanisms of vision. Scientific American, 241(3):150–162, September 1979. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n3/pdf/scientificamerican0979-150.pdf.

Evarts:1979:BMM

[1952] Edward V. Evarts. Brain mechanisms of movement. Scientific American, 241(3):164–179, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n3/pdf/scientificamerican0979-
164.pdf.

Geschwind:1979:SHB

[1953] Norman Geschwind. Specializations of the human brain. Scientific American, 241(3):180–199, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n3/pdf/scientificamerican0979-
180.pdf.

Kety:1979:DHB

[1954] Seymour S. Kety. Disorders of the human brain. Scientific American, 241(3):202–214, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n3/pdf/scientificamerican0979-
202.pdf.

Crick:1979:TAB

[1955] F. H. C. Crick. Thinking about the brain. Scientific American, 241(3):219–232, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n3/pdf/scientificamerican0979-
219.pdf.

Walker:1979:ASi

[1956] Jearl Walker. The amateur scientist. Scientific American, 241 (3):234–246, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n3/pdf/scientificamerican0979-
234.pdf.

Anonymous:1979:Bi

[1957] Anonymous. Bibliography. Scientific American, 241(3):245–250, September 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n3/pdf/scientificamerican0979-245.pdf.

Anonymous:1979:Lj

[1958] Anonymous. Letters. Scientific American, 241(4):10, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n4/pdf/scientificamerican1079-10.pdf.

Anonymous:1979:YAj

[1959] Anonymous. 50 and 100 years ago. Scientific American, 241(4):15, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n4/pdf/scientificamerican1079-15.pdf.

Anonymous:1979:Aj

[1960] Anonymous. The authors. Scientific American, 241(4):16, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n4/pdf/scientificamerican1079-16.pdf.

Gardner:1979:MGj

[1961] Martin Gardner. Mathematical games. Scientific American, 241(4):18– 26, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n4/pdf/scientificamerican1079-18.pdf.

Morrison:1979:Bj

[1962] Philip Morrison. Books. Scientific American, 241(4):28–39, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n4/pdf/scientificamerican1079-28.pdf.

Likens:1979:AR

[1963] Gene E. Likens, Richard F. Wright, James N. Galloway, and Thomas J. Butler. Acid rain. Scientific American, 241(4):43–51, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n4/pdf/scientificamerican1079-43.pdf.

Epstein:1979:LC

[1964] Arthur J. Epstein and Joel S. Miller. Linear-chain conductors. Scientific American, 241(4):52–61, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n4/pdf/scientificamerican1079-
52.pdf.

Quate:1979:AM

[1965] Calvin F. Quate. The acoustic microscope. Scientific American, 241(4):62–70, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n4/pdf/scientificamerican1079-
62.pdf.

Anonymous:1979:SCi

[1966] Anonymous. Science and the citizen. Scientific American, 241(4):72–78, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n4/pdf/scientificamerican1079-72.pdf.

Fraser:1979:L

[1967] David W. Fraser and Joseph E. McDade. Legionellosis. Scientific American, 241(4):82–99, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n4/pdf/scientificamerican1079-
82.pdf.

Miller:1979:PM

[1968] Kenneth R. Miller. The photosynthetic membrane. Scientific American, 241(4):102–113, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n4/pdf/scientificamerican1079-
102.pdf.

Peck:1979:LLK

[1969] Dallas L. Peck, Thomas L. Wright, and Robert W. Decker. The lava lakes of Kilauea. Scientific American, 241(4):114–117, 119–122B, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n4/pdf/scientificamerican1079-114.pdf.

Buffetaut:1979:EC

[1970] Eric Buffetaut. The evolution of the crocodilians. Scientific American, 241(4):130–144, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n4/pdf/scientificamerican1079-
130.pdf.

Calder:1979:CM

[1971] Allan Calder. Constructive mathematics. Scientific American, 241(4):146–171, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n4/pdf/scientificamerican1079-
146.pdf.

Walker:1979:ASj

[1972] Jearl Walker. The amateur scientist. Scientific American, 241 (4):172–184, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n4/pdf/scientificamerican1079-
172.pdf.

Anonymous:1979:Bj

[1973] Anonymous. Bibliography. Scientific American, 241(4):186, October 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n4/pdf/scientificamerican1079-186.pdf.

Anonymous:1979:Lk

[1974] Anonymous. Letters. Scientific American, 241(5):8–10, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n5/pdf/scientificamerican1179-8.pdf.

Anonymous:1979:YAk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n5/pdf/scientificamerican1179-12.pdf.

Anonymous:1979:Ak

[1976] Anonymous. The authors. Scientific American, 241(5):17–18, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n5/pdf/scientificamerican1179-17.pdf.

Gardner:1979:MGk

[1977] Martin Gardner. Mathematical games. Scientific American, 241 (5):20–34, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n5/pdf/scientificamerican1179-
20.pdf.

Morrison:1979:Bk

[1978] Philip Morrison. Books. Scientific American, 241(5):36–49, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n5/pdf/scientificamerican1179-36.pdf.

Feld:1979:LBI

[1979] Bernard T. Feld and Kosta Tsipis. Land-based intercontinental ballistic missiles. Scientific American, 241(5):50–60, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n5/pdf/scientificamerican1179-50.pdf.

Notkins:1979:CD

[1980] Abner Louis Notkins. The causes of diabetes. Scientific American, 241(5):62–73, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n5/pdf/scientificamerican1179-
62.pdf.

Schetky:1979:SA

[1981] L. M. Schetky. Shape-memory alloys. Scientific American, 241(5):68–76, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schetky:1979:SMA

[1982] L. McDonald Schetky. Shape-memory alloys. Scientific American, 241(5):74–82, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n5/pdf/scientificamerican1179-
74.pdf.

Glue:1979:SC

[1983] Quark Glue. Science and the citizen. Scientific American, 241(5):84–93, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n5/pdf/scientificamerican1179-84.pdf.

Kimura:1979:NTM

[1984] Motoo Kimura. The neutral theory of molecular evolution. Scientific American, 241(5):98–126, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v241/n5/pdf/scientificamerican1179-
98.pdf.

Meier:1979:PG

[1985] David L. Meier and Rashid A. Sunyaev. Primeval galaxies. Scientific American, 241(5):130–144, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v241/n5/pdf/scientificamerican1179-
130.pdf.

Heinrich:1979:EAD

[1986] Bernd Heinrich and George A. Bartholomew. The ecology of the African dung beetle. Scientific American, 241(5):146–156, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n5/pdf/scientificamerican1179-146.pdf.

dEspagnat:1979:QTR

[1987] Bernard d’Espagnat. The quantum theory and reality. Scientific American, 241(5):158–181, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n5/pdf/scientificamerican1179-
158.pdf.

Dixon:1979:NIA

[1988] P. W. Dixon. A Neolithic and Iron Age site on a hilltop in Southern England. Scientific American, 241(5):182–190, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n5/pdf/scientificamerican1179-182.pdf.

Walker:1979:ASk

[1989] Jearl Walker. The amateur scientist. Scientific American, 241 (5):192–204, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n5/pdf/scientificamerican1179-
192.pdf.

Anonymous:1979:Bk

[1990] Anonymous. Bibliography. Scientific American, 241(5):206, November 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n5/pdf/scientificamerican1179-206.pdf.

Anonymous:1979:Ll

[1991] Anonymous. Letters. Scientific American, 241(6):13–14, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n6/pdf/scientificamerican1279-13.pdf.

Anonymous:1979:YAl

[1992] Anonymous. 50 and 100 years ago. Scientific American, 241(6):16, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n6/pdf/scientificamerican1279-16.pdf.

Anonymous:1979:Al

[1993] Anonymous. The authors. Scientific American, 241(6):19, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n6/pdf/scientificamerican1279-19.pdf.

Gardner:1979:MGl

[1994] Martin Gardner. Mathematical games. Scientific American, 241 (6):22–30, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n6/pdf/scientificamerican1279-
22.pdf.

Philip:1979:B

[1995] Philip and Phylis Morrison. Books. Scientific American, 241(6):33–53, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n6/pdf/scientificamerican1279-33.pdf.

Kalhammer:1979:ES

[1996] F. R. Kalhammer. Energy-storage systems. Scientific American, 241 (6):42–51, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kalhammer:1979:ESS

[1997] Fritz R. Kalhammer. Energy-storage systems. Scientific American, 241(6):56–65, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n6/pdf/scientificamerican1279-
56.pdf.

Blackshear:1979:IDD

[1998] Perry J. Blackshear. Implantable drug-delivery systems. Scientific American, 241(6):66–73, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n6/pdf/scientificamerican1279-
66.pdf.

Robertis:1979:GTA

[1999] E. M. De Robertis and J. B. Gurdon. Gene transplantation and the analysis of development. Scientific American, 241(6):74–82, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n6/pdf/scientificamerican1279-74.pdf.

Anonymous:1979:SCj

[2000] Anonymous. Science and the citizen. Scientific American, 241(6):84–92, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n6/pdf/scientificamerican1279-84.pdf.

Feldman:1979:PL

[2001] Jerome A. Feldman. Programming languages. Scientific American, 241(6):94–116, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n6/pdf/scientificamerican1279-
94.pdf.

Trinkaus:1979:N

[2002] Erik Trinkaus and William W. Howells. The Neanderthals. Scientific American, 241(6):118–133, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v241/n6/pdf/scientificamerican1279-
118.pdf.

Regan:1979:ERE

[2003] David Regan. Electrical responses evoked from the human brain. Scientific American, 241(6):134–146, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v241/n6/pdf/scientificamerican1279-
134.pdf.

Fulcher:1979:DV

[2004] Lewis P. Fulcher, Johann Rafelski, and Abraham Klein. The decay of the vacuum. Scientific American, 241(6):150–159, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v241/
n6/pdf/scientificamerican1279-150.pdf.

Kittleman:1979:T

[2005] Laurence R. Kittleman. Tephra. Scientific American, 241(6):160–177, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n6/pdf/scientificamerican1279-160.pdf.

Walker:1979:ASl

[2006] Jearl Walker. The amateur scientist. Scientific American, 241 (6):178–190, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v241/n6/pdf/scientificamerican1279-
178.pdf.

Anonymous:1979:AIa

[2007] Anonymous. Annual index. Scientific American, 241(6):192–200, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n6/pdf/scientificamerican1279-192.pdf.

Anonymous:1979:Bl

[2008] Anonymous. Bibliography. Scientific American, 241(6):202, December 1979. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v241/n6/pdf/scientificamerican1279-202.pdf.

SCIAM:1845:SA

[2009] Scientific American, 1845. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). Scientific American, Inc., 415 Madison Avenue, New York, NY 10017-1111, USA.

Beeler:1973:PW

[2010] M. Beeler. Patterson’s worm. AI Memo 290, Massachusetts Institute of Technology, A. I. Lab., Cambridge, MA, USA, 1973. ???? pp. Description of simple system for exploring optimal exploration pattern for a simple organism “feeding” on a lattice. See also [786].

Berlekamp:1982:WWY

[2011] E. Berlekamp, J. H. Conway, and R. Guy. Winning Ways for Your Mathematical Plays. Academic Press, New York, NY, USA, 1982. ISBN 0-12-091150-7 (vol. 1), 0-12-091152-3 (vol. 2). various pp. LCCN QA95 .B446 1982. In two volumes. Volume II contains the proof that the game of LIFE is computation universal and describes the construction of a self-reproducing configuration. See [2012].

Bays:1987:CGL

[2012] C. Bays. Candidates for the game of Life in three dimensions. Complex Systems, 1(3):373–400, ???? 1987. ISSN 0891-2513. 3D CA rules with dynamics similar to Conway’s game of Life [165, 2011].See also [2013].

Dewdney:1987:CRG

[2013] A. K. Dewdney. Computer recreations: The game of Life acquires some successors in three dimensions. Scientific American, 256(2): 16–17, 20–22, 24, February 1987. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v256/n2/pdf/scientificamerican0287-

REFERENCES 324 16.pdf. Description of Carter Bays’ 3D extensions of Conway’s game of LIFE. See also [2012].